A Bibliography of Publications about the Mathematica Symbolic Algebra Language

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

12 May 2026 Version 2.159

Title word cross-reference

$145.00 [Ano98a]. $19.95 [Ano98b]. 2 [Ozb17]. 2.0 [FKS20]. $29.95 [Ano97d, Ano00b]. 3 [Arc20, IIS18, Ozb17]. 3 + 1 [Tor20]. $42.95 [Ano95b]. $45.00 [Ano99a]. $49.95 [Ano98c, Ano00a]. $54.95 [MT96]. $59.95 [Ano97d]. $695 [Ano90a]. $74.95 [Ano00b]. $75.00 [Rue07]. $79.97 [McC02]. $84.95 [Ano98d]. $99.00 [Ano97b]. ^{1} [WW93a]. ^{(}R) [Rav07, Ish16]. _{2} [Gro01, KLO96]. _{p}F_{p−1}, F_{1}, F_{2}, F_{3}, F_{4} [BKK13]. α [Jon92]. $Apart [Fen12]. E_{6} [Dep17]. ℓ_{1} [Lor08]. F [Che94]. F_{c} [BK16]. F_{D} [BKM14]. F_{S} [BKM14]. γ [JL93]. GL(nR) [Bro09]. K [Bzo21, Che94]. n [Che94]. N = 3 [EM93]. P [Alc97]. π [AW97]. q [AU21, Kra93, Kra95]. QR [BQO98]. R [Str93a]. ⋆ [Tos10]. τ [Kei96]. → [WW93a]. x [GK92]. z [EMM90b, Gra04]. Z(ζ) [ADG94].

-6 [Ano96d]. -adic [Alc97]. -algebra [JL93]. -binomial [Kra93, Kra95]. -Dirichlet [Kei96]. -function [Jon92]. -out-of- [Che94]. -penalized [Lor08]. -product [Tos10]. -range [Str93a]. -series [AU21]. -strategy [Arc20, IIS18]. -transforms [EMM90b, Gra04].

/sup [Maj96].

0 [Tay99]. 0-387-94274-2 [MT96].

0-387-95234-9 [McC02]. 0-521-84150-X [Rue07]. 0-8493-2016-X [Tay99].

1.4 [Wie15]. 1/D [WW93b]. 1/P [WW93b].

1/S [WW93b]. 1/sup [WW93b]. 10th [SBA94]. 11th [Kap92]. 13/C [Maj96]. 14 [MP10]. 15 [Fos11]. 16th [IEE95a]. 17-Eck [BH08a]. 17-gon [BH08a]. 1995 [PMDA96]. 1ères [Ziz97b].

2 [BCxx, CW05b, Fen16, Tam95a, Tho91]. 2-D [Tho91]. 2.0 [Pat17]. 2/sup [WW93b]. 2017 [Ano17]. 20th [IEE95d]. 21st [BH95, Jef08]. 22nd [PL92]. 25th [BH95, Ras91]. 28S [Tha89]. 29th [IEE90d]. 2èmes [Ziz97b]. 2nd [FMRP16, MT96, OT94].

3 [Cof92a, GG97, Oln98, Poi98, Sfe98, Ver97, WW93a, Ree98, SMC98]. 3-540-62743-X [Oln98]. 3-540-94047-2 [Pie96]. 3-7643-2832-0 [Ano94c]. 3-D [Cof92a]. 3.0 [Ano91b, Bra96, Coo98a, Gal98, Wol96c]. 3.0.1 [Ant98, Shi97]. 3/sup [WW93b]. 32-bit [Tam95a]. 32.50 [Ano99a]. 34th [Ano95h, IEE95e]. 386-based [Wol88b, WW88b, WW89a]. 3M [CW05d]. 3Ms [CW05b, CW05c, CW05a]. 3rd [Ano96a]. 3s [IIS18].

4 [Bra96, GG00, McC00, Wol99b, Ano00b]. 4.0 [BHJ^{+}00, Wol99b]. 4.0/Buchtipp [BHJ^{+}00]. 4.2 [Sta03]. 48.00 [Ano97a]. 4860HD [Ano92d]. 4th [Ano95d, DN93, HY14].

5.2 [Hil06]. 54 [Ano94c]. 5th [Car96, RCM03].

6 [Ano96d, HLO^{+}09]. 6th [GT94, SBA94].

7-10 [IEE91c]. 75th [Ano95d]. 782 [BQO98]. 78th [Ano94b].

8 [Fos11]. 80 [Pie96]. 85 [Stu95c].

9.3 [SMO20]. ’90 [IEE90c, Mio90, WN90, Tho97a]. 90’s [IEE90a]. ’91 [IEE91c, PMD92]. 910 [Kor11]. ’92 [DN93]. 923 [Wim12]. ’93 [Mio93]. ’94 [ACM94, Ano94g, GT94, IEE94a]. ’95 [Ano95h, IEE95b, Lev95]. ’96 [Lak96].

A-Train [Ano92d]. Aachen [Ano94g]. ABJ [Pre18]. Abotec [Shu93]. Abstract [SBQ14, Ano99a, Fat96, HL99, JM08, KSS07, MS08, Pet08]. Abstraction [Hay92b]. abstracts [IEE95c]. accelerated [Hug93]. accelerator [Eva93b, Ryn93]. Accessing [KG90]. Accompany [App94, GEJHH94, KN94]. Accord [Sta19]. Accumulated [DPU94g]. accuracy [McC00]. Accurate [ABF14]. achievements [Pet89]. Acoustics [IEE90b, IEE91b, IEE92b, IEE93, IEE95b]. Acrobat [BHJ^{+}00]. Action [O’R92, Wag91, Wagxxa, Wag99, Wag10, Sol92, Tay92, Ano00a]. ActiveX [Kro98]. Ada [Kro98]. Adaptive [PGB94]. Add [Wol96c, Ano97d, Vin03, Wol99b]. Add-on [Wol96c, Ano97d, Vin03, Wol99b]. adding [Ive92]. Addison [Ano95b]. Addison-Wesley [Ano95b]. addition [Ano91b]. additional [Wol99b]. Addressing [DPU94a]. Adds [Stu96]. adiabatic [Xav22]. adic [Alc97]. Adobe [BHJ^{+}00]. Adroit [Was05]. Advanced [Bha06, Cra96, Har00, Kou09, KN02, MM97, MM98b, Pid96a, SLR^{+}95, ZT94, Har15, Ish16, KN94]. Advances [Fat90a]. adventures [Cra91a]. Advertisers [Ano95f]. Advice [MB94]. affine [Naz12]. Affine.m [Naz12]. Again [Kob94a]. Agent [Pat93]. Aggregation [Gay92b]. AI [RSL^{+}26]. Aid [MG94, GH93, LJW91, Sav95]. aide [JLP99]. Aided [Ano96d, MGC94, Mil90, PMDA96, SYS^{+}90, EJ94, FM93, Ost94, PMD92, RMC96, RF96, YZPD92]. Aktion [Wagxxb]. Akutsu [EM93]. al. [GEJHH94]. Alamos [Ryn93]. Albert [BHJ^{+}00]. Albuquerque [IEE90b]. Alfred [Ano98d]. Algebra [AM94, Asl96, Bag95, BG01, Cro93, DFPA95, DPV14, EMB92, FSC95a, FEV93, HWH91, Hog07, HM92, Joh94, KWW92, LM90c, LM93, Mat89a, Mor93b, Rei93, Sob92, Sza15, AHR94, Ano99a, Ben98a, Che91, CLO92, CLO07, Cro91, DT00, DF94, EE96, Fat15, FP95, GKW03, HPR96, HPS05, HL99, Ioa92f, JL93, Jon92, KSS07, LW99, LM90b, Mak97, Mat89b, Mat89c, Mat91c, Mat91g, MT94b, MT94c, Mun96, Ohm94, PF95, Rad95, RF95, Sob95, Ste94, SH10, Sza00, TT99, TT09, Wes99, Wic96, Car96, Ree00]. Algebra-System [KWW92]. Algebraic [ACM94, BFZ14, BG92, Bro93, Gut04, Ioa92c, Lak96, Lev95, MCG13, RLRML04, RF96, SE11, SYS^{+}90, Stu95b, Stu95a, Tra00, WN90, Ziz92, vdHvH12, CLO92, CLO07, Fat90a, Fat90c, Fat15, Mit89, Pro96, San95, SBP96, WW93a, WW93b, MBD91, Stu95c]. Algebras [FKS20, Bd93a, Eks11, FK15, Mak97, Naz12]. Algèbres [Mak97]. Algorithm [GGK96, Bro09, EMEH96, GMP09, HBN95, Hen96, KA93, Luc95, PS95, PS06, BQO98, Kor11, Wim12]. Algorithmic [Koe93]. Algorithms [APL92, Eva93a, Lew94, Sob92, AGW95, BMM90, Can03, CLO92, CLO07, Sha98, dT93]. Aligning [Bar10]. Allen [Ano99a]. allows [Tod92]. Alone [Ano92d]. Alternating [PW94a, Ant98]. Alternative [ES10, Fat15, Sim91, Ste08, Ste13]. AMBRE [GKR07]. American [Ano91c, NEH90, SBA94]. AMFlow [LM23]. Amplitude [Sko94]. amplitudes [MBD91, Wes93]. Anaheim [IEE94b]. analog [CN94]. Analogy [Kro98]. Analysis [All94, Bau97, Bha05, Car92, CP90, CN94, DHE92, EM92, Eva93a, Fos93, Kro98, Lip12, LU95, Mil90, Mor98b, NR96, Ost94, Pao99, PC94, Rao98, Rav07, Rid90c, Rue07, Sch98b, Sti95, SS14, Vor97, AJ95, AC91, AdSR96, All92, Ano94h, Ano96f, Ano97c, Ant98, BM94, BEL^{+}95a, BGZ96, Bau98, Bau99, Bau00, Bel95b, BB95, Bha06, Che91, CP93a, Fol02, Gla90, Gre05, GI98, HFO94, Hen94, IOA^{+}95, JKP96, KZK10, Kul91, LBSA96, MTWW94, MGS96, NS94, NA90, NA92, Nov17, Noy95, Ozb17, Pao01, Rao93, Rid90a, Rid90b, RL19, Sav95, Sei92, Sha06, Skr93, Sol96, Tay99, TFS94, Var96, YK95, YZPD92, ter03, Kro98, Str97]. Analyst [Was05]. Analytic [Bah92, BK13, AC15, AC18, Pat15, Pat17, Vos99, Vos00]. analytica [Kor93, CZ92]. Analytical [Maj96, Bes98, Dub03, FMRP16, NA92, Var97, WW06]. analyze [Arc20, IIS18]. Analyzing [San95]. Anatomical [GDD^{+}96]. Andrew [GEJHH94]. Andrews [Tra00]. Anil [Ski90]. Animating [PW94b, PW97, ST90, Ano97a]. Animation [Ker96, Mur92, Swe91, MW95, Rom94]. Animationen [Ano95g]. Animations [Swe92, Cal04, Pid96a, Pid96b]. Annasoft [Kro98]. années [Ziz97b]. Anniversary [Ras91]. announcement [AC18]. Annual [Ano91c, Ano94b, Ano94i, Ano95d, Ano95h, BH95, LT94, Lum93, Lum94, NEH90, RPWW96, Jef08, VM90]. annular [KAG96]. anQCD [AC15, AC18]. antenna [CP93b]. antennas [IEE90a]. Anwendung [Ben98a]. Anwendungsbeispielen [Kau92, Ste93]. Anymore [Sim92d, Sim92a]. ao [Nor94]. APart [Fen16, Fen12]. aplicaciones [DFPA94, Dom98, Ram97]. appearing [Ioa92e]. Application [BH93, CN92, DHE92, Ioa92a, Ioa92d, Ioa92e, KMRC90, KYB94, NA90, Sch02, SBP96, Ano96b, Cau90, Fat15, FK15, FKS20, HFO94, Kau92, MC93, NG93, PDP00, Sch98a]. Applications [Bha05, EMM90a, Fre94a, GWWJS95, Gra94c, Gra97b, IEE91d, KSS07, MT96, Tro95b, Van93, WJ94, AU21, Alb06, Bha00, CLR98, DT00, Dot90, EE96, GKW03, Gra98b, Gue07, Har14, Ioa93, JKP96, KVF99, Kou09, Lyn07, Nam94, RLM06, SS95a, SC10, Tro98, ter03, Pie96]. applicazioni [Fal93]. Applied [ABD94, Ano96h, Gra04, Ham94, RD94, ST94, Var94, WA96, DCN96, FFVH94, HT13, Har94, Mar00, Mat91g, Rad95, Rus99]. Applying [Bes98, Chi92]. appraisal [Bel89]. Approach [Bla92a, Bla92b, Cal95a, FKM95, Lip12, Mac92, Mat89a, Ost94, Rav07, Rue07, SB92, SYS^{+}90, Squ92, ABMBR15, Aza13, BW99, Cal95b, Gra04, GMP97, Gre05, Haz10, Kou09, SLC07, Shi09, Tho90, WG92, Wic96, WW06, Xav22]. approaches [Cof92a]. Approche [Sfe93]. Approximate [CSA94, WW06]. Approximating [Mar94]. approximation [Ioa92b]. Approximations [CD92, GK92, DPU94c]. apps [Tod92]. Apr [IEE91c]. April [Ano94b, HRV92, IEE90b, IEE93, IEE96, Mio90, PMD92]. APS [GT94]. Arbeiten [KWW92, KP94]. Arbeitsbuch [HJ93]. Arbing [Mic92a]. arbitrarily [vdG94]. Arbitrary [BH08b, TF94, FRW17, JL93, Nam94]. Archimedean [GGV94]. architecture [Sam94]. Arclength [Coh94a]. arguments [dT93]. arising [HL08]. Arithmetic [BH07, BH08c, APU97, Kei93]. Arizona [MGC94]. ARKEIA [Ano97g]. Arles [Van95]. arrays [Bar90]. article [Bar92, Dev92a]. articles [Fos91b]. Artificial [Ano94f, Fre94a, IEE91d]. Artlandia [Kro98]. Asks [Lew94]. Aspects [SÜ18]. assembly [RF96]. assist [Gla90]. associated [NR96]. Associative [Mak97]. associatives [Mak97]. associativity [KA93]. astronomical [Rom09]. Astrophysics [KM13, Ish16]. asymmetric [NG93]. asymptotic [Sal89, Sal91]. asymptotics [TIM96]. Atlas [Tho97b, Tho97a, Ano97b, Ano98a]. Atmospheric [PSMM95]. Atom [WSW94, Wan96, YJ92]. atomic [WAW96]. atoms [Gui06]. attitudes [Gli95]. attraction [RM96]. Audio [PSW93]. Aug [GT94]. August [Ano94f, DN93, HY14, Tra00, Van95, WN90]. Austria [Mio93, PMD92, PMDA96, Jef08]. autocorrelation [Byk19]. Automata [Ada96, Ano97f, GN96]. Automated [AC13, Ano96b, BBK14, Kap92, Nai22, Nis94]. Automatic [ABMBR15, AdSR96, AHH93, Kor97, Kro98, MAC08, PS93, Pat93, Ros16, Sal89, Sal91, Shu93, Wie15]. Automation [IEE96]. autoparametric [BBD93]. aux [Ziz97b]. auxiliary [LM23]. Available [Kro98, Sch95]. avanzadas [MM98a]. Average [BYA94]. averaging [BBD93, TSC93]. Award [Ano17]. AXIOM [Ben98a, Ben98a, Ken01, Grä96].

B [Ano95b, Ree00, Röh99, Wee96]. back [Shu93]. back-of-the-envelope [Shu93]. Background [Str93c]. Baden [Joh95]. Baden-Baden [Joh95]. Bahder [Ano95b, Wee96]. Baker [KL02]. Banach [Wag93]. Banach-Tarski [Wag93]. Banded [Wim12]. Barnes [GKR07]. base [IMAS95]. Based [BG01, Eva93a, FH96, JS93b, ON92, BGH96, BKK13, BKM14, BK15, BK16, De 99, Gli94, Gli95, GMP09, IOA^{+}95, JS93a, JS93c, KDM^{+}92, KB93, MAL92, RB93, SKS95, Wol88b, WW88b, WW89a]. Basel [Ano94c]. Bases [RLRML04, Buc92, EMM90b]. Basic [Bel90, UOT94, BHH91, Kra93, Kra95]. Basics [BPU90b]. Basins [RM96]. Baumann [Ano97e, CTMC97]. Bayesian [APL92, CB95, Gre05, Kor93, Lip12, Rav07, Rue07]. Bd [Str97]. Beach [IEE91d]. Beam [Ano95a, AGW95, Auc96, IOA^{+}95, Kni95, Tho90, Yar95]. beams [Ano96f, Bha96, NA90, NA92]. Beginner [GG92, GGxx, Ree98, Sak95, SMC98, Ano00b, GG97, GG00]. beginners [MT06]. Behavior [McP92, DWY98, GD98]. Beispielen [KP94]. beliefs [Gli95]. Benchmark [Jon93, FMRP16, Sha98]. BenderWu [SÜ18]. Benefits [Rid90a]. Berlin [Pie96]. Bessel [Ada95, GM93, YM03, dT93]. better [Ive92]. Between [JS93b, LW22, AC13, JS93a, JS93c]. Bhansali [Ski90]. Bi [LW22]. Bi-Directional [LW22]. bias [SC09]. bimetric [Tor20]. bimEX [Tor20]. Binding [Kro98, Jac19]. Binomial [CZ09, Kra93, Kra95, PS95]. Biochemical [Alb06]. Biofilm [Zho97]. Biografie [BHJ^{+}00]. Biografie/Kuhn/Försterling [BHJ^{+}00]. Biological [Ano96a, GW95, GWWJS95, MT96, TMC96]. biomedical [Pid96a, Pid96b]. Birkhäuser [Ano94c]. bit [Tam95a]. Blachman [Mac92, Squ92]. Black [Kro98, Mil90]. blackboard [Swa91]. Blood [XKM93, GDD^{+}96, HKM93]. Boca [Ano98d]. Bodies [Mur92]. body [Xav22, XMLC16]. Book [Ano94d, Ano94c, Ano95b, Ano96c, Ano97a, Ano97b, Ano97c, Ano97e, Ano97f, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano00b, Ano00a, Ben98b, Che25, CAMM94, CTMC97, Dar94, Har19, Lev94, Lip12, McC02, MT96, Mor98b, Oln98, Pie96, Rao98, Rav07, Ree98, Ree00, Röh99, Rue07, Sha96, SMC98, Tay92, Tay99, TMC96, Wee96, WB94, WW95b, McC09, Tro98, Wol96a, Wol96b, Wol99a, Wol03]. Books [Smi93]. boosts [Var91c]. Boston [MCG13]. boundaries [Bis95, RCM03]. Boundary [Das90, EP96b, Asm05, KPS03, PH95]. Bounded [Mat92a]. brackets [XMLC16]. Bragg [Tho90]. Branch [DF94]. Bravais [MM93]. Breaking [GMM95]. Brenner [FMRP16]. Brian [Ano98b, Ish16, Sha96]. Bristol [Tay99]. broadened [NG93]. broadening [NG93]. Broker [Kro98]. BRST [PTL04]. BSP [GMP09]. BSP-based [GMP09]. Bucharest [Ano94j]. Buchtipp [BHJ^{+}00]. Buckling [KM95b, DF96, Ozb17]. Buckminsterfullerenes [MM93]. Buffering [KC94]. bug [Sim89]. bug-ridden [Sim89]. bugs [Old92]. Build [Per03]. building [Gui06]. Built [Sof93b].

Built-In [Sof93b]. bundig [Ano98e]. Burgers [GMM95, ZM93]. Butterworth [TF93]. Buyers [Gal91]. BYTE [Yag92].

C [Ano97b, Ano97g, Ano99a, Dev92a,

Kus97, Rav07, Rue07, Tay99, KR98, Kor11, Maj96, Oln98, PK10, Ste94, SH10, Tho97b, Zac98, Ano98c]. C-XSC [PK10]. CA [SE11, AKLG95]. CAD [BGH96, KSA92, OM92, OFM94, Rid90b, SHH93]. CADE [Kap92]. CADE-11 [Kap92]. CAI [Kro98]. Calc [MBD91]. CalcLabs [Bar98b, B^{+}96, Hol98, Hol03a]. Calcul [Sfe93, BESP94, Mak97]. Calculating [EM93, EMEH96, Gro01]. Calculation [Koe93, WK17, WKB20, XKM93, ABF14, AGW95, Aza13, Bd93a, HKM93, KZK10, Kul91, MBD91, Pat15, Pat17, TIM96, Var97, XMLC16]. Calculations [ADG94, Ben93, BA96, Kor11, AC15, AC18, BK13, Fer07, Fon21, HY92, Jac19, Luc95, Pid96a, Pid96b, Shr91, WL11, Zit11]. calculator [Shu93]. Calculo [AF97b, BLM97, FSC95b, Mor98a, Pér95]. Calculus [ACCM93a, ACCM93b, App94, Bar98b, B^{+}92, BPU90a, BPU90b, BPU90c, BPU91, BUP91, Bro91, BD91, BD92, BDxx, Coh94a, Cro93, CR91, DPU94c, DPU94e, DPU94g, DPU94f, DPU94b, DPU94d, DPU94a, Ede91, EN92, FL92b, KN93, Kir94, KSJE95, LS95, Mor94, PW94b, Ree00, SDB93b, SDB93a, Str92b, Str93b, Str93c, Str94, Str98, Uhl95, Ano97a, BSB94, CJL94, CLR98, Cro91, De 99, GLMG12, GEJHH94, Hol98, Ken93, Ken01, May94a, PW97, Par93, SBQ14, Str99, TT99, TT09, UK11, GEJHH94, Höf92, Hol03a]. calculus-based [De 99]. Calendar [FF93]. California [IEE92b, IEE94b, LT94, Rog92]. Calling [Was05]. Cambridge [Ano97d, Ano00b, Rue07]. Campbell [Har15]. campus [PL92]. Can [Lew94, Ste08, Ive92, DPV14]. Canada [IEE91b, Lev95, OT94]. Canonical [GGK96]. Cantabria [Gut04]. cantilever [Auc96]. Cantor [Hen96]. Canty [BHJ^{+}00]. capabilities [Sim91]. Capital [Ben93, Kor94]. Capri [Mio90]. careful [AHR94]. Carl [BH08a]. CARTAN [Sol96]. CAS [BGS93]. cascade [WW93a, WW93b]. Case [ST90, ABFP24, FP90, Pat93, TIM96]. Castigliano [CP93a]. Catastrophes [Gay92a]. Cauchy [CKP13]. CD [Ano97b, Ano98a, Ano00a, MT96, ABGA96, GAM95, Tro98]. CD-ROM [Ano98a, Ano00a, MT96, ABGA96, GAM95, Tro98]. CDs [McC02]. CE [Ano95f, Kro98]. celestial [CD02]. Cellular [Ada96, Ano97f, GN96]. Center [Aga90, Ano02, IEE90b, IEE93, IEE94b, PL92, Jon93]. centered [Haz10]. Central [CZ09]. Centre [IEE91b]. Century [BH95]. CERN [Van95]. Certain [ZD02]. chains [KA99]. challenge [Tro97, Wat90b]. Challenges [Ano91c]. Challenging [RCM03]. Champaign [Ano97d, Pie96, Tho94]. change [Str98, Str99]. Changing [Ano91c]. channel [MN91]. channeled [Aza13]. channeling [Aza13]. Chaos [KS97, GI98, Nis94]. chaotic [DWY98]. characteristics [Rao93]. characterized [AF92]. Characters [Sko94]. charge [YK95]. charged [Hug93]. Chebyshev [LRW95]. Chemical [Ano95j, Ano96a, Fol02, TM14]. Chemistry [BHJ^{+}00, Cro98]. Chemistry/Wiser [BHJ^{+}00]. chess [RL19]. ChessY [RL19]. Chicago [Ano02, Bel91]. CHIPR [Xav22]. chromatic [MAL92]. Cinematic [DF91]. circles [Råd93c, Rad93b]. Circuit [Rid90c, Ano94h, Ant98, Rid90b, TFS94]. circuits [CN94, Doe95, FP90]. Circular [AB92a, BA96, KAG96]. Circulation [SGMM95]. circumference [BP95]. Class [Cro93, Cro91]. classes [CKP13, Ziz97a, Ziz97b]. Classical [CD02, CTMC97, Ano97e, Bau96, Luc95]. Classroom [GAM94b, Tay91]. Clebsch [Dil91c]. Click [SS94, Shu94]. Closed [GK92, SNLR92, Wel93]. Closed-form [SNLR92]. cloth [MT96]. Clouds [PSMM95]. cluster [Var97]. clusters [Bit93]. CMA [Ioa93]. CMU [Ras91]. CO [IOA^{+}95]. coal [KL92]. code [Byk19, Hah08, JM93, Kor97, Lew00]. codes [FMRP16, HHL06, BQO98]. Coding [TW06, JM08]. coefficient [PS95]. Coefficients [CZ09, Dil91c, MdR13, NG93, ÖAÖ24, ÖAÖ25]. Coincidences [WSW94]. Colin [McC02]. Collection [Tre91]. collective [Arc20]. College [Lum94]. Collegiate [Lum93, Lum94]. Collision [Lev92a, Bro94]. Collisions [WSW94]. Colloquium [Ano96g]. Color [MAL92]. Colorado [Wag94]. Column [DW92]. Combination [Tre91]. combinatorial [LEF91]. Combinatorics [Ski90, Skixx, PS03]. Combined [SBA94, Var97]. Combining [CH95, Bar98a]. Combustion [RB95]. combustors [KL92]. Coming [CW05e]. Commemorative [Ras91]. Commun [Stu95c, Ziz97b]. Commun. [Stu95a]. Communicating [PK10]. commutative [CLO92, CLO07]. compact [Cof92a]. companion [Hol03b]. Company [Ano95b]. Comparative [Lip12, Rav07, Rue07, Gre05, Par93]. compared [BFS93]. Comparison [CDS^{+}92, LW99, TS00, KA99, Yar95, Zot07]. compatible [Tod92]. Compensators [OM92]. competitor [Ano90a]. compilation [SLR^{+}95]. compiler [SLR^{+}95]. Complete [Ede92, GSOW93, APU97]. Complex [Asl96, Gay92a, GWWJS95, KW92, MT96, Sha06, TMC96, Gal91, Nov17, Sei92, dT93]. complexity [DWY98]. Component [Kro98]. composite [Bha96]. Composites [Kaw93]. Composition [FEV93]. compressible [RP92]. Compression [Wat94]. Comput [Stu95c]. Comput.

[Stu95a]. computable [IIS18]. Computation [ACM94, Ano96g, Bee17, BG92, Bro93, Cra94, Cra96, IEE91a, Ish16, JOS93, Jef08, Lak96, Lev95, MCG13, SE11, Str93a, Tra00, Van93, Var94, Var95, WN90, Wim12, vdHvH12, BBP97, Bar93, BP94, BB95, BSB94, CN92, GSR90, Har94, HS90, Hug93, Jac01, KM13, LM23, LS96, Mio90, Mio93, NA90, PS93, PTL04, RCM03, Sen95, SS05, TS00, Wat90b, XW95, XX95, dT93, Gut04]. Computational [Ano96a, Ano98c, BFZ14, Han04, PS03, PGB94, Ryn93, Sch98b, Sto03, Var91a, Varxx, Var96, WJ02, CLO92, CLO07, Gra04, Mor95, YK95, Zac98, Ano97c, Mor98b, Rao98]. computationalist [Mad91]. Computations [Bha05, BMM90, Bha06, Gal91, MdR13, Mit89, Naz12, NEH90, Rid90a, Sei92, Tor20, Wie15]. compute [HHP^{+}14, Pro96]. Computer [All94, Ano96d, AH90, Asl96, Bag95, BFS93, Bd93a, Buc93, BD91, BD92, BDxx, Car92, CP90, Cra90, Cro93, DPV14, EJ94, FEV93, FM93, GW95, GWWJS95, GKW03, HWH91, HPR96, HSW94, HSW98, HM92, IEE90e, IEE95d, Iss96, JDBT06, KN94, KWW92, LM90c, LM93, Mae93b, Mae00, Mat89a, Mat89b, MGC94, MT96, Mil90, Oln98, Ost94, PC94, PMD92, PMDA96, Ras91, Sal94, Sch92, Sob92, SYS^{+}90, SR91, Web92, Wes92, Wes99, Wol88a, Wol91a, Wolxxa, YZPD92, AHR94, All92, Ano95i, Bag05, Ben98a, Car96, Che91, Cro91, Cro98, DF94, EE96, Fat15, FP95, Gro01, Hay93, Ioa92f, Jon92, JDBT09, KR98, KN02, LW99, LM90b, Mae90a, Mak97, MZ92, Mat89c, Mat91c, Mat91g, MBD91, Mit89, Ohm94, Pet89, PF95, Rad95, RF96, SH98]. computer [Sob95, Ste94, SH10, TJD11, Tab16, WW95a, WW88a, WRR90, Wol91e, Wol91b, Wol91c, Wol91f, Wol93a, Wol93c, Wol93f, Wol93d, Wol93e, Wol93g, Wol94b, ter03, PMDA96, TMC96, Wolxxb, Kro89].

Computer-Aided [MGC94, Mil90, SYS^{+}90, FM93, Ost94, RF96]. Computer-driven [Sch92]. computer/math [Ano95i]. Computeralgebra [Ben98a]. Computeralgebra-Systemen [Ben98a]. Computerized [Ano95c]. Computers [Ano93a, BHJ^{+}00, Dev92b, DW92, Dev94a, Dev94b, GG91, RRM90, Var94]. Computers/Canty [BHJ^{+}00]. Computing [Agu90, Ano94g, BKR^{+}91, BCDH94, Cav93, CKP13, EP96a, EP96b, HSH07, HH98, HH03, KM95b, LP13, NP95, PL92, Ric91, RdG96, SK93, SK95, Van95, YM03, APL92, Ano97b, Ano98a, BS12, DN93, EP03, GT94, Gui06, Har14, HPR96, Hid21, HRV92, LP05, LR18a, LR18b, MR01, Möl07, RM96, RMC96, RLM06, Tha89, Thi91, Tho97b, Tho97a, WW93a, WW93b, WC98, Chu14]. con [AF97a, AF97b, BLM97, DFPA95, FSC95a, FSC95b, MM98a, Mor98a, Pér95, Ram97]. Concept [Cro93, Cro91, Hug93]. Concepts [Bha05, Tro95b, Bar98b, Hol98, Lew91]. Concerning [GG91]. Conditioned [KMRC90]. Conference [AKLG95, Ano91c, Ano92c, Ano94a, Ano94f, Ano95h, Ano96a, Ano96j, Ano02, AH90, BH95, EJ94, GT94, Gug95, Ham94, HY14, HRV92, IEE91a, IEE94a, IEE90b, IEE90d, IEE91b, IEE91d, IEE92a, IEE92b, IEE93, IEE95b, IEE95d, IEE95e, IEE96, IBM^{+}95, Joh95, Kap92, LT94, Lum93, Lum94, RRM90, RdG96, RPWW96, SBA94, Wol91d, Ano90c, DN93, HW92, IEE94b, IEE95c, OT94, VM90]. Confidence [MR94]. configurations [GHIL09]. confinement [SS95a]. conflicts [Can00, Can03]. conformal [Bzo21, LM91, Sch95]. Congress [Ano94g, SBA94]. Congressi [GT94]. conic [LM90b]. Conjecture [Ede92, De 94]. Connection [Ano99b, MdR13]. connects [Coh94b]. consecutive [Che94]. consecutive- [Che94]. Constants [RSL^{+}26, BBP97, LM91]. constraining [Auc96]. constraint [LEF91, RF96]. constraints [LM90a]. Construct [Bar91]. Constructible [Slo93]. Constructing [GGK96, IA93, Xav22]. Construction [BSB94, Fat90a, Fat90c, GKR07]. Consultants [Ano95f]. contained [KN02]. contexts [Bar98b, Hol98]. Continued [Ada92, Hen96]. continuous [Sah96, TIM96]. Continuum [RLM06]. contracted [AC13]. contractility [LBSA96]. contractility-digital [LBSA96]. contractions [Bd93a]. contrast [MAL92]. Control [Ano95a, Ano96g, BGZ96, Chr94, Fre94a, Gug95, IEE94a, IEE90d, IEE95e, IBM^{+}95, Kro98, MGC94, Ono92, PGB94, BB95, BGH96, Chr93, FM93, GI98, KSA92, Mun96, OFM94, SBP96, Str92a]. controller [LS96]. Controllers [Cal95a, Cal95b, RMC96]. Controlling [Fos01]. Controls [Kro98]. Convention [Aga90, IEE90b, IEE93, IEE94b]. Conventional [Sit97]. convergence [ABFP24, Ano96e]. Conversion [BHJ^{+}00]. converters [BYA94]. Convolution [Knu92, WM93]. convolutions [HHP^{+}14]. Cookbook [Man10]. cooling [JY93, YJ92]. Coombes [Ano98b, Röh99, Sha96]. cooperation [IEE90e]. CORBA [Kro98]. Corrections [Ano95f, SC09]. corrective [MS08]. correlation [Bzo21]. Cosine [Wat94]. Cost [Ben93, Bha96]. Coulomb [Jon93]. Counting [DK93]. coupled [Bit93, Xav22]. coupling [CKS00, ÖAÖ25, Pre18]. Course [Hol94, Sit97, Gli94, Gli95, Kau99, LP05, Noy95, Par93, Wag94]. Courses [SDB93a, Ano95c]. Courseware [BPU90a, BEL^{+}95a, Mae93a]. Coventry [IEE94a]. crack [Ioa92e]. cracks [Ioa92f]. Crandall [Tay92]. crash [Kau99]. CRC [Ano98d, Tay99, von93]. Create [ES10, Ste08]. Creating [Dil91a, Ste13]. Crichton [BHJ^{+}00]. Critical [KS93, Bel89, DK93, Kei96, MC93]. cross [Bro94]. cross-collision [Bro94]. crossing [Cha91]. CRTs [Yar95]. Cryptography [TW06, McA08]. crystal [Aza13]. Crystals [Bah92, Yar95]. Cubes [Hay92b]. Cumulative [BH07, BH08b, BH08c]. Cumulus [PSMM95]. Current [Abb90, RdG96, Ant98]. currents [Apt91]. Curriculum [FW91, Doe95, SW93]. curvas [CFGxx]. curvature [TPH96]. curve [Mat92a]. curved [NA90, NA92, Rao93]. Curves [Gra97a, von93, Ano98d, BSB94, GMP09, Gra93, Gra98a, GAS06, PL92, Sen95]. curvilinear [Ioa92f]. Cut [Yag92]. cuts [DF94, HW92]. Cybernetics [IEE92a]. cycles [Ziz97a]. Cylindrical [KM95b, RP92]. Czechoslovakia [DN93, Str92a].

D [Ano17, WW93a, Cof92a, Ozb17, Tho91, WW93b]. Dallas [Aga90, Ano95d, IEE90a, IEE90e, NEH90]. Darboux [Ohm94]. Darstellung [Gra94a]. Data [Kro98, Lip12, Rav07, Rue07, Sti95, Tre91, Wit90, Gre05, HD91, Ive92, MTWW94]. data-driven [Ive92]. databases [IA93]. Daubechies [RA95]. David [Har15, Ish16]. dBASE [Ano92d]. DC [BYA94]. DC-DC [BYA94]. Debye [Zyp93]. Debye-Scherrer [Zyp93]. decade [IEE90c]. decay [KA99, Wan96]. December [AKLG95, IEE90d, IEE90e, IEE95a, IEE95e]. Deceptive [Wag92]. Decision [Chi92, IEE90d, IEE95e, Kor93, Kor95]. Decisions [Kor94]. decoder [Sob95]. Decompositions [EGM94]. decoupling [CKS00]. Deduction [Kap92]. DeFinetti [Arc20]. Definite [Wag92, Lic11]. Definition [WK17, WKB20]. Degrees [ZD02]. Deguchi [EM93]. Deliver [WJ02]. demands [Ano91a]. demon [Ano95e]. demonstrations [MdR13]. demystified [Hos09]. Dense [Wim12, BQO98]. Density [vdG94, MN91, MC93]. Department [Lum93, Sch92]. dependence [MN91]. deployable [GGV94]. Derivation [Ioa92f, Bar90, PSW23, PSW23]. Derivative [Cro93, WGDZ04]. Derivatives [DPU94e, Sha98]. DERIVE [Ben98a, SC09, Ben98a, KWW92, Pet89]. Descartes [Mon93]. Descartes-Geometry [Mon93]. descendant [LM91]. Description [WSW94]. Design [BG92, Chr94, Eva93a, Fos93, MGC94, Mio93, Bha96, Chr93, CN94, Fat90a, FM93, GGV94, LS96, LTE01, Mae92, Mie95, MGS96, OFM94, Pat93, RMC96, Rom09, RF96, YZPD92, Zho97, Mio90]. designer [Mad91]. Desktop [Ano92d, Kul91, Swa91]. Detailed [Tro94, MK03]. detection [SKS95]. Determination [Dil91c]. determine [HBN95]. Determining [Bro94, BP95, GDD^{+}96]. deterministic [Hof91]. Detroit [IEE95b]. DEUG [Ziz97b]. Developer [Lew94]. Developers [Ano95f]. Development [Cro93, Kni95, Kro98, SHH93, Cro91]. Developments [JOS93]. devices [Yar95]. Diagnostics [Ano95a]. diagonalization [Dil91c]. diagrams [Bar10, Stu95b, Stu95a, Stu95c]. Dialogs [GG91]. diamond [Aza13]. diamond-structure [Aza13]. diatomic [Sen95]. Diego [AKLG95]. diferenciales [AF97a]. Difference [KS97, KM02, Mik94a]. differencial [CFGxx]. DIfferential [BKK13, BKM14, BK15, BK16, AB93, AB97a, AB04a, Ban01, Bau97, Che25, Coo98a, CHL^{+}95, EP96a, EP96b, EP03, GK11, Gra97a, HLO^{+}09, Mor94, PBA06, Ros95, Ros04, Sha96, Sit97, ST99, Tho91, Vor97, Vve93, Vve94, Ziz92, Ano98d, Apt91, Asm05, Bau98, Bau99, Bau00, Cal04, EP04, GV96, Gra93, GMP97, Gra98a, GAS06, Hol03b, KPS97, KPS03, MR01, Mat89b, Mat89c, Mat91c, PS93, SW97, SBP96, SH98, ST04, VMP94, ZL95, Ano98b, Che25]. differential-algebraic [SBP96]. Differentialgeometrie [Gra94a]. Differentialgleichungen [SG95]. differentiation [Gue07, PS93]. differentiel [BESP94]. DiffExp [Hid21]. diffraction [BA96, Gro01, Tho90, Zyp93]. diffractive [O’S96]. Diffusion [Gay92b]. Diffusion-Limited [Gay92b]. diffusions [SS93]. Digest [Ano96g, Aga90, IEE90a]. Digital [ZT94, LBSA96, PWW90, Rog92]. Digits [Hay92b]. Dill [Luc95]. dimension [Hen96]. Dimensional [WG94, CN92, Hid21, LM91, Naz12, WG92, WGDZ04, ZM93]. dimensionally [GMP09]. Dimensions [DPU94f, JL93]. d’initiation [GG93]. Direct [FP90, Ioa92a, Ioa92d]. Directional [LW22]. directions [Abb90]. Directly [WK17, WKB20]. Directories [Ano95f]. Directory [Ano95f]. Dirichlet [Kei96]. DISCO [Mio90, Mio93]. discover [May94a]. Discovering [B^{+}92, KSJE95]. Discrete [CD92, Hol94, KM02, Ski90, Skixx, Wat94, PS03, SW93, SRR94, XW95, XX95]. disk [BA96, Cha91]. diskette [Ano96c, Ano97c, Ano97e, Ano97f, Ano98c]. Disorder [BD93b]. Dispersion [Bah92, MN91, ADG94]. Dispersive [CT94]. displacement [Kni95]. display [MAL92, Rog92]. Distance [Uhl95]. distinctive [Cof92a]. Distributed [IEE90e, JS93b, LP13, Bar90, JS93a, JS93c, Sam94, SLR^{+}95]. distributed-memory [SLR^{+}95]. Distribution [BH07, BH08b, BH08c, CGO17, Str93a, Byk19, Tza09]. distributions [Sch95, SC09]. diversity [Mae92]. division [MP10]. DM [Ano94c, Ano96c, Ano97a, Ano97c, Ano97e, Ano97f, Ano99a, Ano00a, Pie96]. DNA [PC93]. do [Ben93, Gas98, Jan94]. document [IIS18]. Documents [Dil91b, Sla90, YF04]. does [GSR90, Lin91a]. Doing [Buc93, Kro89, PC94, SR91, Web92, Wes92, Wol88a, Wol91a, Wolxxa, Bel93, GLMG12, Hay93, Ken93, Luc95, WW95a, WW88a, Wol91e, Wol91b, Wol91c, Wol91f, Wol93a, Wol93c, Wol93f, Wol93d, Wol93e, Wol93g, Wol94b]. Domains [TF94, Jef93]. domes [GGV94]. Done [ST94]. Dot [BN91, Sch92]. Double [BS94, MC96]. Double-track [BS94]. Drainage [RRIR^{+}92]. Driven [Hay92b, SGMM95, Ive92, Sch92]. drudgery [Ano92a]. DSP [EMM90b, Kro98, Ono92]. Dual [CT94]. Dual-Frequency [CT94]. duct [Che91]. Duffing [WW06]. During [Mat91h]. Dynamic [Ano96a, OM92, Ano96f, BK96, CP93b, KB93, LS96, Rao93, Kro98]. Dynamical [Hol94, Lyn07, BPR99, BPR00, Cha91, KM02]. Dynamics [PC93, Arc20, BYA94, BK95, IIS18, KYB94, LJW91]. Dyson [BHJ^{+}00].

e-mails [Bra96]. E6Tensors [Dep17]. early [MAL92]. easy [Ano95e]. ECC [IBM^{+}95]. Eck [BH08a]. Econometrics [Bel90, Bel93]. Econometrics.m [Bel93, Bel90]. Economic [Dar94, Var92, Var93, MGS96]. Economics [Mor98b, Rao98, Sch98b, Ano97c, Var96]. economists [HC97]. economization [LRW95]. Ecuaciones [AF97a]. Ed [Ano97a, OT94]. ED-MEDIA [OT94]. EDA [Adi89]. eddy [Apt91]. eddy-currents [Apt91]. edge [Cof92a]. edges [TD95]. Edited [Ano97c, Ano97d]. Editing [CH95]. Edition [Ano00a, KN94, MT96, Web92, Wes92, Ano96c, Ano98d, Hol03a, KN02]. Editor [Ano95f, Stu96]. Education [Ano91c, Ano94i, Ano95d, BH95, Hay94b, RPWW96, Mat91b, Adi89, Mat94, Mat05, Fow96, SBP96, Swa91, Wat90b]. Educational [OT94, EMW91]. EEs [Ano95i]. efficiencies [VMP94]. Efficient [Gue07, Sof93a, Wim12]. EiffelBase [Kro98]. eigenproblems [HT13]. eigenseries [MC96]. eighth [KN02]. Einführung [Kau92, Kof93, KP94].

Einstein [BHJ^{+}00, Agu90, Har03]. elastic [KYB94]. elasticity [CN92, Ioa93, IA95]. Electric [Ost94, Wan96]. electrical [Ano95c, Rid90a]. electromagnetic [Hug93, Sch95]. Electromagnetics [BSHS93, IEE91a, CKS90, TS00]. Electron [TIM96, WSW94, Yar95]. Electron-Atom [WSW94]. Electron-beam [Yar95]. Electron-wave-function [TIM96]. Electronic [RD94, Ano90b, BEL^{+}95a, BP94, Rid90b]. electrons [Aza13]. Electrostatic [May94b, SS95a]. Electrotechnical [KZK10]. electroweak [HL08]. elegant [Var91b]. Element [Bha05, Mik94b, Bha06, Che91, CP93a, HT13, Ioa92d, Ioa93, Kni95, Kor97, Skr93]. Elementary [EP04, Ioa93, SK93, SK95, Wat90a, Wol15, FP95, Gli94, Gli95, IA93, PF95]. elements [AC13, Bit93, Bro94, FFVH94, HL08, O’S96, SNLR92, Var97]. elimination [LM90a]. Eliminations [Ede92]. elliptic [TSC93]. embedded [AHH93]. Emily [Ano97d]. emission [Yar95]. emphasizing [Bag05]. Employing [Mis13]. Empress [Ano97g]. encoder [Sob95]. encoder/decoder [Sob95]. End [LS95, Wol93g, ter03]. End-User [LS95]. Energy [BHJ^{+}00, RRIR^{+}92, HY92, Ioa92a, Sen95, TM14, Wie15, Xav22]. Enfoque [UB93]. Enführung [Mae93b]. Engineering [Ano91c, Ano96d, Bel95b, BH95, Fos92, Fos93, Fri94, Hay94b, KN94, Mae00, MB94, Pao99, Pao01, RD94, Rob95, Ano95c, Cap03, DT00, DWY98, EJ94, Fol02, Har14, HSW94, Joh95, KNJ93, KN02, MM97, MM98b, Mat91g, Pri93, Wat90b, Ben98a, Tay99]. Engineers [Bah94, Com96, Iss96, NEH90, Dub03, EM01, Gas98, GH93, Gra04, HS90, Sei92, Wee96, Ano95b]. England [ACM94]. English [Bur94a, Kau94]. Enhance [GMM95, PSMM95, SGMM95]. Enhanced [LaB94, Swe92]. enhancing [Doe95]. enough [AHR94]. enseignement [Ziz93]. enthusiasm [Buc92]. Enthusiasts [Was05]. Entrainment [PSMM95]. Enumeration [Slo93]. envelope [Shu93]. Environment [Eva93a, FEV93, Sof93a, Old92, Pri93]. Epidemic [GVV91]. EPS [GT94]. EPS-APS [GT94]. Equation [BH93, GMM95, Stu96, ST90, Apt91, Run89, Tho90, WGDZ04, WW06, ZM93]. Equations [AB93, Bau97, Che25, CHL^{+}95, EP96a, EP96b, GK11, Ioa92c, KS97, Lew94, PBA06, Ros95, Sit97, ST99, Tho91, Van93, Vor97, Vve93, Vve94, AB97a, AB04a, ABF14, AF92, Ano96b, Ano98b, Asm05, Ban01, Bau98, Bau99, Bau00, Bes98, Cal04, Coo98a, Dvo90, EP03, EP04, GV96, GMP97, Hol03b, HLO^{+}09, HHL06, Ioa92e, Ioa92f, KM02, KPS97, KPS03, MR01, Mat89b, Mat89c, Mat91c, Nam94, PSW23, Ros04, SW97, SBP96, Sha96, SH98, ST04, Tho90, ZL95]. Equilibria [DK91]. Equilibrium [Nog91]. equipment [MGS96]. equivalence [Rad93d]. Erratum [Stu95a, Stu95c]. error [BMM90]. error-free [BMM90]. Erste [Bur93, Bur94a]. Erwin [KN02]. erythrocyte [MK03]. ESCAPE [Ano96d]. ESCAPE-6 [Ano96d]. Essential [Sim93]. estimation [Cur95, RM96, Tza09]. estimator [vdG94]. EUFIT [Ano94g]. Euler [BK96, RSL^{+}26]. EUROCAST [PMDA96, PMD92]. European [Ano94g, IBM^{+}95, Ano96d]. Evaluate [BH07, BH08b, BH08c]. evaluating [Bzo21]. Evaluation [dT93, Ada95, AC13, Bro09, Sah96, Sim89]. evaluators [Fat96]. Event [LU95, XW95, XX95]. Event-Related [LU95]. evidence [KL02]. EvoDyn [IIS18]. EvoDyn-3s [IIS18]. Evolution [BHJ^{+}00, HHL06, Wan96]. evolutionary [IIS18, Jac01]. Exact [RLRML04, Tor20]. examined [Tho90]. Example [AB92b, AB94, ABxx, Coh94a, AB97b, AB04b, AB08, BB95, JM93, Kor93, Sti02]. Examples [CTMC97, Koe93, Sal89, Sal91, Tro95b, WK17, WKB20, Ano97e, Bau96, Kau92, Kau94, KLO96, Möl07, MK03]. Excel [Coh94b]. excellence [IEE90c]. exceptional [Neu11]. exchange [Bit93]. excited [Wan96]. Exercices [JLP99]. exercise [TPH96]. expanding [Ber24, HM06]. expansion [Pre18, Ros16]. Expansions [LP13, Tho92, Eks11, Hid21, Sal89, Sal91, Thi91]. Experimental [AGW95, Gli94, Gli95, HD91]. Expert [HRV92]. Expertensysteme [KWW92]. explicit [Bro09]. Exploration [Sco91, Nis94]. Explorations [GW95, GWWJS95, LM93, MT96, TMC96, WC98]. Exploring [Cha01, FL92b, FL92a, FL94, Fre94b, GG91, GG94, GEJHH94, HL99, LM90b, Mor95, Rei95, Sar97, Sol92, Vos99, Vos00, WG94, Wes96a, Ano99a]. Exponential [Van93, Byk19]. Exponents [MC93]. exported [Bar10]. expressing [Wes93]. Expressions [Chu14, GK92, Mez92, PK10, BBK14, Jef93, KLO96, Kor97, Maj96, YF04, Zit11]. Extended [HR11, Nam94, Fat96]. Extending [Sof93b]. Extensible [LW22]. extension [ST11, ST12]. extent [Jef93]. Exterior [Ede91, BSB94, CJL94]. External [KG90]. extinction [KLO96]. Extracting [YF04]. extraction [MGS96]. extreme [YM03].

F [Ree00]. facility [Grä96]. factorizations [BQO98]. factors [Ioa92b]. Faddeev [PS06]. False [Ede92]. Familiarizing [WK17, WKB20]. family [Neu11]. FAPT [BK13]. Fast [BMM90, Sob92, LRW95]. faster [Tam95a, Tod92]. fault [SKS95]. FDDI [AJ95]. Feagin [CAMM94, Tit94]. Feasible [Che96]. Featured [Car08]. Features [Shi97, SMO20]. Feb [Ryn93]. February [IEE91d, Rog92]. Fedosov [Tos10]. feedback [SBP96]. Fermi [ABF14].

Feyn [MBD91]. Feyn-Calc-computer-algebraic [MBD91]. FeynCalc [SMO20]. FeynGKZ [ABBD23]. Feynman [Stu95c, ABBD23, GKR07, Hid21, LM23, MBD91, Stu95b, Stu95a, Wes93]. FeynmanParameter [Wes93]. FFT [IOA^{+}95]. Fiber [CT94]. field [AC91, Sch95, ST11, Wan96, Yar95]. field-emission [Yar95]. Fields [Van93, Abe90, Has03, Hug93, LM91, ST12]. Fifth [Lum94, PMDA96, Hol03a]. Figure [Wig92]. Files [JS93b, JS93a, JS93c]. Filter [LTE01]. filters [TF93]. Finance [Mor98b, Rao98, Sch98b, Ano97c, Var96]. Finance-Modelling [Rao98]. Financial [Mil90, Var92, Var93, Dar94, Sha98, Sto03]. Find [GK92, MR94]. Finding [DK91, Mat91d, Gea91, Nam94]. finger [TPH96]. Finite [Bha05, Mik94a, Mik94b, Sut92, Van93, Bha06, Che91, CP93a, HT13, Ioa92d, Ioa93, Kni95, Kor97, Naz12, Nie03, SNLR92, Skr93]. finite-dimensional [Naz12]. finite-element [Kor97]. First [Bur94a, Bur94b, FW91, Hol94, KM95a, LP05, VM90, dT93]. First-Year [FW91]. fitting [Mat92a]. five [Sei92, XMLC16]. five-body [XMLC16]. fixed [Mar00]. FL [Ano98d]. flexible [BK95, BK96, KB93]. Flipbooks [Swe91, Swe92]. Florida [IEE91d]. Flow [Ano95j, CS96b, XKM93, Che91, GDD^{+}96, HKM93, LM23, MC96]. fluid [KVF99, MC96, MGS96]. fly [Mic92a]. focal [ZL95]. forget [Iss96]. Form [EGM94, GK92, SNLR92, ZL95, CMS17]. Formal [Koe93, Koe94, SDB93a, Bar90]. Formalism [GSOW93, MM08, PTL04]. Format [Sof93b]. formation [Bar93]. FormCalc [Hah08]. FormCalc-generated [Hah08]. formel [Mak97]. Forms [Ede91, Ziz92, Abe90, GH04, MC93]. FormTracer [CMS17]. Formula [AW97]. Formulae [CH95]. Formulas [Bar91, RSL^{+}26, HSW94, HSW98, KU94].

Försterling [BHJ^{+}00]. FORTRAN [Tay99, Ano98a, Pao99, Pao01, LP05, Tho97a, Var97]. forward [BK95]. Foundations [KVF99, GKW03]. Four [McP92, Yag92, Maj96]. Four-Layer [McP92]. four-spin [Maj96]. Fourier [Ano96e, Apt91, Asm05]. Fourth [Lum93]. Fractal [RRIR^{+}92, Bis95]. fractals [Bau96, GH04, CTMC97, Ano97e]. fraction [Hen96]. Fractional [BK13]. Fractions [Ada92, Mat90b]. fracture [IA93]. Framework [CH95, RB93, Kro98]. France [Van95, vdHvH12]. Francisco [IEE92b, LT94]. Free [Auc96, ES10, Kro98, BMM90, NR96, RP92, Ste08, Ste13]. French [Mak97]. Frequency [CT94, BA96, Cha91, Ozb17]. FRI [SS94]. FRI-S14 [SS94]. FRIL [MB94, BM94]. FRODO [AC13]. Front [ter03, Wol93g]. Front-end [ter03]. Frontiers [BH95]. fronts [EN96]. frustration [Buc92]. full [HW92]. fullerenes [MT94a]. Function [ABR94, Bee17, Kor11, Roa96, Str93a, Abe90, ADG94, Byk19, BK16, DCN96, Fen12, Fen16, Jon92, TIM96, Ada95]. Functional [Ano96e, PK10, Lor08, PSW23]. Functionality [Sne94]. Functionals [CD92]. Functions [Asl96, BH07, BH08b, BH08c, CGO17, MAC08, Ada95, ABFP24, Ano97b, Ano98a, Ber24, BKK13, BKM14, BK15, BK16, Bzo21, GM93, HM06, IA93, KW92, Kou13, LW96, ÖAÖ25, Tho97b, Tho97a, TSC93, Wol99b, Xav22, YM03, dT93]. Fundamental [Bha05, DPU94a]. fundamentals [RLM06]. fundamentos [DFPA94, Dom98]. Fusion [Ano94a]. future [Abb90, BS94, Mic92b]. Fuzzifying [BM94]. Fuzzy [Che94, Fre94a, RB93, SRR94].

G [Oln98, Röh99, Ada95]. G-function [Ada95]. gain [BS94]. Gained [Kor94]. galaxies [Cha91]. games [Can00, Can03, IIS18]. Gamma [Bah92, Tza09]. Gamma-point [Bah92]. Gamma1 [WSW94]. Gamma2 [WSW94]. Gammel [KL02]. Garrett [Ano98b, Sha96]. Gas [Ano95j]. Gaseous [RB95, Bro94]. gauge [RF95]. Gaunt [ÖAÖ24]. Gauß [BH08a, Edi90, BH08a]. Gaussian [Ede92]. Gaylord [Ano96c, Ano97f, MT96, TMC96, Ada96]. GBP [Ano99a, Ano00a]. gearing [Ano96b]. Gegenständen [GAM94a]. General [KU94, Nog91, Sal94, AC15, AC18, Har03, GLMG12]. Generalized [BH93, BS11a, BS11b, BKK13, BKM14, BK15, BK16, Fen12, Fen16, ÖAÖ25]. generate [HHL06]. generated [Hah08]. Generating [Mis12, Mis13]. Generatingfunctionology [Wil06]. Generation [MAC08, O’S96, Ano96b, Byk19, FMRP16, HHWT95, KDM^{+}92, Kor97, RL19]. generator [SKS95]. generators [MZ14, Mis13]. Generic [CH95]. Genome [BHJ^{+}00]. geodesic [AF92]. Geographical [Dau17]. Geomathematics [Ria91]. Geometria [CFGxx]. Geometric [Mat93, Rom09, GGV94, RF96]. geometrically [NA90]. geometries [AF92]. Geometry [BFZ14, Gra97a, Hay94a, Mon93, Raa94, RLRML04, Ano98d, CLR98, CLO92, CLO07, Gra93, Gra98a, GAS06, RF96, Vos99, Vos00]. Geopsychical [Sca94]. Geoscience [Ria91]. geosciences [Han04]. Gerd [Ano97e, CTMC97]. German [Can00, BH08a, Ben98a, Kau92]. Germany [Ano94g, Ano97e, Joh95]. Getting [Che98, Che05, ST94]. GFUN [HR11]. Give [BPU90c, MB94]. given [Byk19, Sch95]. Giving [AB92a]. GKZ [ABBD23]. Glacial [SYS^{+}90]. Glasgow [AH90]. Gleason [GEJHH94]. global [GSR90, Ozb17]. Glynn [Ano00b, Ree98, SMC98, Sol92]. Gmunden [Mio93]. Golden [Mon93]. Golf [Mur93]. gon [BH08a]. Gordan [Dil91c]. Gordian [Kro98]. Gould [MT96]. graded [Bd93a]. graduate [Ish16]. Grafik [Nol97, Sch94a]. Grand [Ano92d]. Graph [Ski90, Skixx, PS03]. Graphic [CH95]. Graphical [Car94a, PSW93, Tre91, Ban01, HS90, LP05, RM96, WA96]. Graphics [GWWJS95, GH04, HSH07, Kro98, Lev92b, Phi90, SB95, Stu96, WJ94, Cof92a, Fos92, Gue07, Pie96, Rus99, Rus04, Rus09, Sch95, Sim92b, Tro04a]. graphing [Ano95e, Sim91]. graphs [FRW17, RL19]. grating [VMP94]. Gravity [Har03]. Gray [Ano98d, Ano00b, SMC98, Sol92, Ree98]. great [Ano91b]. greater [Cof92b]. Green [DCN96]. GreenTree [Kro98]. Gregory [Rue07, Lip12, Rav07]. Grenoble [vdHvH12]. griffbereit [BCxx]. Gröbner [Buc92, RLRML04]. Ground [CS96b]. Ground-Water [CS96b]. group [Fon21, RF96, Kro98, Har19]. GroupMath [Fon21]. Growing [Mad91]. Growth [DPU94e, DPU94g, Mat92a]. guarantees [GMP09]. GUI [Nov17]. Guide [AR93, Ben98b, Boy91, Gal91, GG92, GG93, GGxx, HWH91, Ree98, Sak95, SMC98, Sza15, Tam97, Wol92, Wol93b, Wol93h, Wol95b, Ano97b, Ano98a, Ano00b, GG97, GG00, Har15, KNJ93, KN02, MT06, Tam08, Tho97a, Tro98, WW95a, Wes99, WW89a, Wol93f, Wol93e, Wol93g, Wol99b]. Guidebook [CR91, Miy94, SB95, Tro95b, Tro06a, Tro06b, Tho97b, Tro00, Tro04a, Tro04b, Tro05]. guided [SH98].

H [Mor98b, Rao98, Gro01]. Hagenberg [Jef08]. Hal [Ano97c, Dar94, Sch98b]. Hallett [GEJHH94]. Hamiltonian [FP90]. Handbook [AB92c, Bee17, Hog07, Ree00, GKW03, TT99, TT09]. Handling [FEV93]. Hands [HMM15, HMM16]. Hands-on [HMM15, HMM16]. Hard [Str93a]. Hard-Sphere [Str93a]. hardback [Ano97d, Ano00b]. Hardcover [Pie96, Rue07]. Hardware [Kro98]. Harmonic [ABR94, CZ09, PW94a, HL08, Maî06]. Harper [Lum94]. Harvey [MT96]. Hausdorff [Hen96]. Hawaii [IEE90d]. Hawaiian [IEE90d]. HDF [JS93a, JS93b, JS93c]. Healthcare [RdG96]. heart [GDD^{+}96]. Heat [ST90, Che91, MC96]. Heidelberg [Ano97e]. Heisenberg [Bes98]. held [PL92, Rog92, VM90]. helicity [MM08]. helium [WW93a, WW93b]. Help [Mat91a, Ive92, Sei92, Var90]. helps [Zyp93]. HEPMath [Wie15]. Herglotz [GGK96]. Herrgott [BHJ^{+}00]. heterogeneous [LZ96]. Hewlett [Tha89]. Hibbard [Ano99a]. Hickory [Ano02]. High [KMRC90, UOT94, Ano91a, Ano95e, EN96, HT13, HY92, Hug93, Pet89, Saf91, SW93, Wie15]. high-energy [HY92]. high-level [Ano95e]. High-Precision [KMRC90]. Higher [Ano96f, Sim90]. Hilbert [LM91]. Hilton [IEE90d, IEE95e]. HIP [HY92]. HiQ [Wes96a]. Höhere [Str97]. Hokkaido [Ano95h]. holomorphic [KW92]. Holonomic [Kou13, Kou09]. Home [Ano92d]. honest [ABK95]. Honolulu [IEE90d]. Horn [BKM14, BK15]. Horn-type [BKM14, BK15]. Hotel [IEE91b]. HotGAMS [Bra96]. Houches [Sta19]. Hours [Mat91h]. HPL [Maî06]. Hughes [GEJHH94]. Human [Lev92a, Rog92]. humvee [Sch92]. Hunt [Ano98b, Röh99, Sha96]. Hybrid [NA92]. hydrograph [Cav93]. HYP [Kra93, Kra95]. Hyperbolic [Wag93]. HyperCard [CKS90]. HYPERDIRE [BKM14, BKK13, BK15, BK16]. HYPERgeometric [BK15, Roa96, ABBD23, ABFP24, Ber24, BKK13, BKM14, BK16, HM06, Kra93, Kra95, ÖAÖ25, BKK13, BKM14, BK16]. hypermedia [OT94]. HypExp [HM06]. HYPQ [Kra93, Kra95]. HYTECH [HHWT95].

I/O [Shi97]. IASTED [Ham94]. IBM [Tam95a]. IC [Kro98]. ICASSP [IEE90b, IEE91b, IEE92b, IEE93, IEE95b]. ICASSP-92 [IEE92b]. ICMS [HY14]. Icosahedra [Mae94b]. Idea [IEE94b, Wol16]. Idea/microelectronics [IEE94b]. Ideals [CLO92, CLO07]. ideas [Wol16]. Identifying [HD91]. identities [Ada95, PS95]. IEE [Ano96g]. IEEE [Aga90, IEE90c, IEE90d, IEE90e, IEE91c, IEE92b, IEE93, LT94, MGC94]. IEEE-Dallas [IEE90e]. IEEE/IFAC [MGC94]. IFAC [Gug95, Joh95, MGC94, Str92a]. II [Abe94b, Mae96a, Tam91b, WW88c, WW88d]. III [Cha91, LW96, Rog92]. IL/Cambridge [Ano97d]. Ill [KMRC90]. Ill-Conditioned [KMRC90]. Illinois [Lum94]. Illustrated [GAM94b, GAM95, Sza15, Ano97b, Ano98a, Tho97b, Tho97a]. Illustrating [Jac01, Tho91]. Illustrierte [GAM94a]. IMACS [Ano96a, HRV92, JOS93]. Image [Wat94, ter03]. images [PL92]. Imaging [KMRC90]. immersed [Cro91]. impact [NEH90]. implement [SZ96]. Implementation [GSR90, GSOW93, Hof91, Mil90, SRR94, ABMBR15, BEL^{+}95a, HT13, Maî06, MM08, Mio90, Mio93, Stu95b, Stu95a, Stu95c]. Implementations [Kro98]. implemented [VMP94]. Implementing [BSS93, Dil91c, Har02, Nie03, Ski90, Skixx, SDB93a]. Implicit [Bar93, GMP09]. Imposing [EGM94]. Improved [NG93]. improvements [SMO20]. improves [Ano92a]. Inc. [Sim92c]. included [Ano96c, Ano97b, Ano97c, Ano97e, Ano97f, Ano98a, Ano98c, Ano00a]. includes [Bar92, Ste94]. Including [Mez92, KLO96, MK03, SH10]. indefinite [Lic11]. Index [Ano95f]. Industrial [Gil95]. inertial [SS95a]. inertial-electrostatic [SS95a]. Inference [CH95, Kor93]. Infinite [CZ09]. inflating [Rom94]. Informal [BD93b]. Informant [Bra96]. informatica [DFPA94, Dom98]. Informatics [Ham94, Pri93]. Informatik [Mae93b]. Information [Ano92c, RPWW96, Ano95f]. informatique [Wol95a, Ziz93]. infrastructure [Bar10]. Ingenieure [San99]. Ingenieurmathematik [Ben98a]. Inheritance [FEV93, Har94]. Initiation [BESP94]. Innovation [KMH97]. Innsbruck [PMDA96]. Insertion [Ros16]. insight [Ano92a, Doe95]. insights [Cal04]. Instabilities [SBA94]. institute [SW93]. Instruction [CW05b, CW05c, RRM90]. Instructional [IMAS95]. integer [HM06, dT93]. integer-order [dT93]. integer-valued [HM06]. integers [San95]. Integral [Mor94, Wag92, WK17, AGW95, BESP94, Ioa92e, Ioa92f, Jon92, Sah96, Tho90]. Integrals [BS11a, DPU94g, Ada95, ABBD23, BS11b, Bzo21, CKP13, GKR07, Hid21, Jon93, LM23, MS98, Pat15, Pat17, Stu95b, Stu95a, Stu95c, Wes93]. Integrand [BFZ14]. Integrated [FW91, Joh95, GMP97]. Integrating [Adi89, BEL^{+}95a]. Integration [Ano97g, Jef93, LT94, SW93, Fat15, KW92, Lic11, Mit89, Sam94]. Intelligence [Ano94f, Fre94a, IEE91d]. Intelligent [Ano94g, ABK95]. intensities [TD95]. intensity [Ioa92b, YJ92]. intensive [Bag05]. interactions [Arc20, Cro91, HL08]. Interactive [BSHS93, Dil91b, Har15, IMAS95, Pao99, Phi90, Sla90, Tay99, Uhl95, CMT12, Pao01, Wic96]. interaktiv [Nol97]. Interface [CW92, KG90, LW22, PL92, Phi90, SS94, Ano93c, Hah08, MZ14, Nov17, TM14]. interiors [Ma95]. intermediate [KAG96]. International [ACM94, Aga90, Ano94a, Ano95h, Ano95j, Ano96a, Ano02, AH90, Bro93, GT94, Gut04, Ham94, HY14, HRV92, IEE91a, IEE94a, IEE90b, IEE90a, IEE91b, IEE92a, IEE92b, IEE93, IEE95c, IEE95b, IEE96, JOS93, Jef08, Kap92, KM95a, KMH97, Lak96, LT94, Lev95, Lum93, Lum94, Mio90, Mio93, MCG13, PMD92, PMDA96, RCM03, RPWW96, SBA94, SE11, Tra00, WN90, vdHvH12, Ano94j, DN93]. Internet [BHJ^{+}00, Ano99b, Fos99]. interpersonal [Phi91]. Interpolation [PS06]. Interpretation [Wag93]. Interpreter [Mae94a]. Interval [Kei93, APU97]. Intervals [MR94]. Interview [SV93]. intra [LBSA96]. intra-uterine [LBSA96]. Introducão [Nor94]. introduccion [Ram97]. Introducing [WWH89]. Introduction [Abe94a, Abe94b, All92, All94, BGS93, Cul91, Cul93, EL90, EL91, Fol02, Fri94, GKW93, GKW94, GMP97, Has01, Has06, Has10, Hon94, Jae94, Kob94b, Mae93b, MT96, Mor93a, Ree00, Ros95, ST99, TW06, Tro94, Var95, Wol15, Zac98, Ano96c, BS12, CLO92, CLO07, GKW96, Har03, Kau92, Kau94, KN02, Ros04, SC10, ST04, Sza00, TT99, TT09, WGK05, Wel13, Ano98c]. Introductory [DWY98, KNJ93, Was05]. Inverse [Sca94, YK95, BK95, BK96, KYB94]. Inversion [CSA94, KMRC90]. Investigating [EMT93]. Investigation [HM92]. involving [ADG94, Cha91]. ion [JY93]. IONA [Kro98]. irradiance [Wei92]. irreducible [San95]. ISAAC [Gut04]. ISBN [Ano94c, McC02, MT96, Pie96, Rue07]. ISE [Kro98]. ISO [AGW95]. isolation [ST11, ST12]. isotropic [Maj96]. isovector [CH94]. ISSAC [Jef08, ACM94, Lak96, Lev95, MCG13, SE11, Tra00, Tro97, WN90, vdHvH12]. ISSAC’93 [Bro93]. Issues [Chr94, Lic11]. Itô [Ken01]. Italy [IEE95a, Mio90]. Iterative [Ioa92c, Ioa92e]. Itovsn3 [Ken93]. IV [Ano92d]. IZIC [FKM95].

J [Ano97b, Ano98a, Ano98b, Ree98, Röh99, Sha96, TMC96, KS93]. James [Tit94, CAMM94]. Jan [MT96]. January [Ano95d]. Japan [WN90]. Japanese [Ohm94, SB92]. Java [LP05]. JavaBeans [Kro98]. jeep [Sch92]. Jenks [Ano17]. Jerry [Ano00b, Sol92, SMC98]. John [Ano97b, Ano98a, Ano98b, Sha96]. Joint [GT94, MGC94, SBA94]. Jones [Pie96, EMEH96]. Jonesboro [Ano94b]. Jose [Rog92, SE11]. Joseph [Ano98c]. journal [Ano90b, Mat98, MAT90a]. Journey [Ano92d]. Joy [SS94, Shu94, SS95b]. Julia [GH04]. July [ACM94, Ano94f, Ano95h, Bro93, Gut04, Jef08, Lak96, Lev95, vdHvH12]. June [Ano94j, Ano02, IEE95c, Kap92, MCG13, OT94, SE11]. Just [Sim92d, Lew91, Sim92a].

K-S [BH07, BH08b, BH08c]. Kai [Ano97g]. Kamin [Ano96c]. Kaufmann [Ano94c, Wee96]. Kazume [Ano97f]. KdV [ZM93]. Kellogg [PL92]. Kenneth [Ano99a]. Kerr [SR90]. Kevin [Ano98b, Sha96]. Khoros [Ano97g]. Kiev [Bro93, SBA94]. kills [KL02]. kind [McC09, Wol02, YM03, dT93]. kinematics [KYB94, LJW91, PS93]. Kinetic [Bel91]. Kingdom [ACM94]. Kinzel [Oln98]. Kit [Dub90, Tod92]. KL [Kro98]. Klassische [Gra94a]. KLEIN [BG01]. Kloosterman [Bro09]. knot [EM93]. Knowledge [Eva93a, KDM^{+}92, ON92, EMM90b, RB93]. Knowledge-Based [Eva93a, ON92, KDM^{+}92, RB93]. Koberlein [Ish16]. Konfliktlösungen [BHJ^{+}00, Can00]. Korea [HY14]. Koster [KU94]. Kranc [HHL06]. Krems [PMD92]. Kreyszig [KN94]. Kuhn [BHJ^{+}00]. kurz [Ano98e].

L [Ano98b, Ano98c, MT96, Röh99, Sha96]. L1Packv2 [Lor08]. Lab [Yag92]. Labels [Lev92b]. Laboratories [Höf92, KN93, Rei93, Bag05]. Laboratory [DRS97, PW94b, SDB93a, PW97, Tre90, Wic96, Ano97a]. Labs [SDB93b, Saf91]. LabVIEW [Tre90, Tre91]. lacks [Cof91]. Lagrangian [TPH96]. Lambda [Eks11]. laminated [RP92]. LAMMPS [FMRP16]. langage [Sfe98]. Language [Wol15, EN96, Fat90b, HMM15, HMM16, LEF91, Mae92, MdR13, Str98, Str99, YK95]. LANs [AJ95]. Laplace [CSA94, Gra04, Mat91g]. Large [Str94, BBK14]. laser [IOA^{+}95, JY93, YJ92]. laser-cooling [JY93]. Lasers [Ano95j]. Latin [SBA94]. Lattice [PC93]. Lattices [MM93]. Laue [Tho90]. Laurent [Koe93]. Lauricella [BK16]. Layer [Bob93, McP92]. LC [TF93]. Leads [Shi97]. Leadville [Wag94]. Lean [LW22]. leaner [Tod92]. Learn [Shu94]. Learning [EKWM93, GMM95, Kob94a, Mor93a, Mor93b, PSMM95, SGMM95, EJ94, Möl07, Sti02]. Least [Mat91d, vdG94]. Lebesgue [WK17, WKB20]. Lecture [SS14]. lectures [CD02, Tro95a]. left [GDD^{+}96]. Lehrbuch [Ste93]. Leitfaden [Kof93]. Length [MR94]. LERG [Stu95b, Stu95a, Stu95c]. LERG-I [Stu95b, Stu95a, Stu95c]. L’essentiel [Ver97]. Lesson [Rad92, Mat92a]. Let [Win91]. Levasseur [Ano99a]. level [Ano95e, EN96, Sam94, Har15, Ish16]. Leverrier [PS06]. L’Hôpital [Mat91e]. Libraries [KG90]. Library [Bee17, BKR^{+}91, Kro98, ABGA96, TM14]. Lie [Bd93a, FK15, FKS20, Naz12]. LieART [FK15, FKS20]. lifetime [MC93]. Light [KA93]. like [MN91]. likelihood [Cur95, Tza09]. limit [YJ92]. limitations [BB95]. Limited [Gay92b]. Limits [Ada92]. Lindstedt [Mar00]. line [Kni95, NG93]. line-broadening [NG93]. Lineal [DFPA95, FSC95a]. Linear [AM94, BG92, Car94a, Car94b, CD92, DT00, EMB92, Hog07, Ioa92c, Joh94, Mor93b, Ngu93, Ree00, Rei93, Sch98a, Sof96, Sza00, Sza15, Wic96, Cav93, FP90, HSW98, HPS05, JKP96, Kni95, LM90a, LW96, LEF91, Pet08, RP92, SNLR92, TT99, TT09, vdG94]. linearization [FM93]. linearized [Bro94]. linearly [Auc96]. Lines [Mat91d, NG93]. linkable [Tod92]. linkages [Kou16]. Links [BKR^{+}91]. Linux [Ano95f, Gal98]. Lipsman [Ano98b, Röh99, Sha96]. Lisp [Fat90b, FH96]. Lisp-Based [FH96]. Lisp-language [Fat90b]. List [Shi97]. Literacy [BPU90b]. Little [Hay94a]. lives [Wol16]. loading [CP93b]. Local [IEE95d]. Location [IEE90c]. Log [BS11a, BS11b]. Log-sine [BS11a, BS11b]. Logic [Coh92, Mae94a, Wat90a, LEF91]. Logical [Lip12, Rav07, Rue07, Gre05]. Logistic [Mar94]. Look [BD93b, Lew91]. looks [Yag92]. loop [BFZ14, Hof91, MS98, Pat15, Pat17]. loops [Pre18]. Louisiana [IEE95e]. Low [YJ92, BA96, MN91]. low-density [MN91]. low-frequency [BA96]. Low-intensity [YJ92]. Löwdin [Jon92]. LRFD [Bha96]. Lugano [GT94]. luminance [MAL92].

M [Ano99a, CAMM94, GEJHH94, Har15, Pre18, Tit94, MM08, Xav22]. M-CHIPR [Xav22]. M. [Dev92a]. M31 [Str94]. M38 [ABD94]. MA [Ano95b, MCG13]. Mac [Way89]. machine [Cra90, FFVH94]. Machines [Ost94, Sut92]. Macintosh [AC91, Bos96, MZ92, WW88c, WW88d, Wol88d, Wol88e, Wol88c, WW88e, Wol89, WW89b, Wol93e]. Macintoshes [Swa91]. MACSYMA [Ben98a, KWW92, Ben98a, Grä96, PZ96]. Madison [IEE95c]. Magma [ES10, Ste08, Ste13]. magnetic [AC91, Hug93, KS93]. magnetofluid [Ma95]. mails [Bra96]. Mainframes [Sim92d, Sim92a]. majors [KNJ93]. makers [Wol16]. makes [Dvo90]. Making [Swe91, De 94]. Man [IEE92a, SV93].

Management [Kro98]. ManeParse [CGO17]. Manipulating [CDS^{+}92]. Manipulation [Chu14, CN94, Erl90, Fat90a, Fat90c, GGV94, HFO94, KYB94, Kra93, Kra95]. Manipulator [Kaw93]. manipulators [BK95, BK96, KB93, Sch94b]. Manual [KN94, WW89b]. many [PL92, Xav22]. MAPLE [Ben98a, Ben98a, ST04, Ste94, Tho94, Ano90a, Bag95, Bar92, BLM97, Bel95b, BFS93, Bra96, CW05b, CW05c, CW05d, CW05e, Dur98, ES10, Fos11, Grä96, Har14, JLP99, JKP96, KSS07, LP05, Möl07, MP10, PF95, PZ96, SLC07, Shi09, ST99, SH10, Ste08, Ste13, SC09, Wes95, Zot07]. Maple/Mathematica [Bel95b]. maps [Sch95]. Marbles [Rad95]. March [IEE94a, IEE92b, MGC94]. Marotta [Har15]. Marriott [IEE92b]. Martin [Ano97d]. mask [O’S96]. mass [Auc96, LM23, Ros16]. masses [CKS00, Was05]. MassToMI [Ros16]. Master [BH93]. Mastering [Gra94b, Gra94c, Gra97b, Gra98b]. matching [Fat15]. Matematicas [MM98a, Ram97, DFPA94, Dom98]. Materials [App94, Gup91]. Math [Ano90a, Ano92a, Bet90, Cof92a, Dvo90, Fos91c, Fos99, Kob94a, Lew91, Ria91, Saf91, Sim92d, Ano95e, Ano95i, Fos91d, Fos91b, Fos92, Gal91, Old92, Sei92, Sim90, Sim92a, Sim92b, Sta90, Way89, Way90, PC94]. MATHCAD [Ben98a, Ben98a, Ano92f, Bar92, BS94, Möl07, BFS93, Cof92a]. MathCW [Bee17]. MathemaTB [Jac19]. MATHEMATICA [Ben98a, Ree00, Ano92d, Ano95f, Ano95g, Ano97g, Ano98c, BHJ^{+}00, Bur94a, CMT12, Can00, HJ93, Ish16, Kau94, KMH97, KP94, KWW92, Mae93b, Oln98, Pao99, RCM03, Ree98, Röh99, Rue07, SMC98, Sol92, Str97, Stu95c, Tay99, TMC96, Ano00b, BH08a, Ben98a, BB95, Bis95, BKM14, CF96, CS96a, Cal95a, Das90, Gla90, Hay94b, Ioa92a, Ioa92e, Ioa92c, IA93, Ioa93, JL93, LJW91, Rom94, Skr93, Ste94, Tho94, Yu94, Mat95a, ABMBR15, Abb90, Abb92, Abe94a, Abe94b, AB92a, AB92b, AB92c, AB93, AB94, ABD94, ABGA96, AB97a, AB97b, ABR99, ABxx, AB04a, AB04b, AB08, AU21, AR93, Ada96, Adi89, AMR91, AF97a, APU97, APL92, Alb06, Alc97, AH91, All92, All94, ABBD23, ABFP24, ACCM93a, ACCM93b, AM94, AC13, Ano88a, Ano88b, Ano89, Ano90b, Ano90a]. Mathematica [Ano91a, Ano91b, Ano92b, Ano92f, Ano93b, Ano93c, Mat94, Ano94e, Ano94h, Ano95e, Ano96j, Ano96i, Ano98e, Ano99b, Mat05, Ant98, App94, Arc20, AF97b, AC15, AC18, Aza13, Bag95, Bag05, Bah94, BK13, BM94, Bar92, BGZ96, Bar98a, BC07, Bar10, BLM97, Bar98b, BH93, Bat93, BL93, Bau93, Bau96, Bau97, Bau98, Bau99, Bau00, Bau05, BESP94, BG01, BPR99, BPR00, Bel89, Bel93, Bel95b, BG92, BFS93, Ben93, Ber24, Bes98, Bha00, Bha05, Bha06, BM92, BSB94, Bla90, Bla92a, Bla92b, BC92, BC94, B^{+}96, BW99, Blaxx, BCxx, BA96, Boy91, B^{+}92, Bra96, BGH96, Bro94, BPU90a, BPU90b, BPU90c, BHH91, BPU91, BUP91, Bro91, BH07, BH08b, BH08c, Buc92, Buc93, BD91, BD92, BDxx, Bur93, Bur94b, Byk19, BKK13]. Mathematica [BK15, BK16, Bzo21, CW92, Cal04, Can00, Can03, Cap03, Car94b, Car04, Car08, CIG^{+}93, Cau90, CS96b, CKS00, Che98, Che05, Chi92, CW05b, CW05c, CW05d, CW05e, Chu14, CGO17, CZ92, Cof89, Cof91, Cof92a, Cof92b, Coh94b, CKS90, CB95, Coo98a, CLR98, Coo98b, CHL^{+}95, Cra91b, Cra91c, Cro91, Cro93, CR91, Cro98, Cul91, Cul93, Cur95, CMS17, Dau17, DPU94c, DPU94e, DPU94g, DPU94f, DPU94b, DPU94d, DPU94a, DT00, De 94, De 99, DF96, Dep17, Dev92a, DW92, DRS97, Dil91a, Dil91c, Dil91b, DFPA94, DFPA95, Dom98, Don01, Don09, Dot90, Dub03, DF91, Dur98, Eks11, EL90, EL91, EN92, EM01, ES10, EMW91, EMT93, EKWM93, Fal93, FRW17, Fat90b, Fat92, FF93, FMRP16, Fea94, FM02, FK15, FKS20, Fen12, Fen16, FSC95a, FSC95b, Fer07, FL92b, FL92a]. Mathematica [FL94, Fol02, Fon21, Fos91a, Fos01, Fos11, Fre92, Fre94b, Fri94, FM93, Gal98, GV96, GMM95, Gas98, GKW93, GKW94, GW95, GWWJS95, GKW96, GN96, GD98, Gea91, GH04, GM93, Gil95, Gli94, Gli95, GKR07, GG97, GG00, GK11, GLMG12, Grä96, Gra04, GG91, GG92, GK92, GG93, GG94, Gra94b, Gra94c, GMP97, Gra97a, Gra97b, Gra98a, Gra98b, GGxx, GAS06, GHIL09, GEJHH94, Gre05, Gro01, GSOW93, Gup91, GI98, HBN95, Hah08, HT13, Han04, Har00, Har14, Has03, Has01, Has06, Has10, HMM15, HMM16, HL08, Hay93, Haz10, HKM93, HSH07, Her88, HL99, Hid21, Hil06, HS90, Höf92, HH98, HH03, Hol98, Hol03a, Hol03b, Hon94, HHP^{+}14, Hos09, HC97, HM06, HLO^{+}09, HHL06, Ioa92d, Ioa92b, IA95, Iss96, IMAS95, IIS18, Jac01, Jac19, Jae94, JS93a, JS93b]. Mathematica [Jan94, JS93c, JLP99, JM93, Joh94, JDBT06, JDBT09, Jur92, KW92, KNJ93, Kau92, Kau94, Kau99, KA93, Kei93, Ken93, KM95a, KN93, KA99, KR98, Kir94, KVF99, KSA92, KC94, KS97, Kni95, KSJE95, KL02, Kob94b, KM13, Kof93, KZK10, Kor93, Kor95, Kou13, Kra93, Kra95, KN94, KN02, Kro98, Kro89, KM02, Kul91, Kus97, KM95b, KPS97, KPS03, LaB94, LP13, LP05, LR18a, LR18b, LB91, LS95, LC90, Lev92b, Lew94, Lew00, LW22, Lin91a, Lin91b, LM23, LU95, Lor08, LTE01, Lyn07, Mac92, MAC08, Mad91, Mae89, Mae90b, Mae91, Mae92, Mae93a, Mae94c, Mae94a, Mae96a, Mae96b, Mae97, Maexx, Mae00, Maî06, MM08, MO94, MM98a, MM97, MM98b, Man10, Mar00, MR01, MdR13, MZ14, MB94, Mat91a, Mat98, Mat90b]. Mathematica [Mat91d, Mat91e, Mat91f, Mat92c, Mat92b, Mat91h, May94a, May94b, McC09, McC00, MT06, MC93, MW95, MT96, MS98, Mie95, Mik94a, Mik94b, MTWW94, Mis12, Mis13, MA00, MA01, Miy94, Möl07, MR94, MG94, Mor95, Mor98a, Mor93a, Mor93b, Mor94, MK03, Nag07, Nai22, Naz12, NS94, Nie03, Nol97, Nor94, Nov17, O’R92, O’S96, Ohm94, OM92, OFM94, Old92, Ono92, Ost94, Ozb17, PW94b, PW97, Pao01, Par93, Pat15, Pat17, Pat93, PS95, PSW23, PS03, Pér95, Per03, Pet89, Phi90, PTL04, Pid96a, Pid96b, PH95, PSW93, Poi98, PK10, PZ96, Pre18, PSMM95, Pri93, PDP00, Råd93c, Raa94, Rad92, Rad93a, Rad93b, Ram97, Rav07, RB95, Rei95, Ria91, RD94, Rob95, RLM06, Rom09, RS02, Ros16, Ros95, Ros04, RA95, RL19, Run89, Rus99]. Mathematica [Rus04, Rus09, SB92, San95, SZ96, San99, Sar97, SC10, Sav95, Sch94a, Sch98a, SW93, SW97, Sch95, Sco91, Sei90, Sei93, Sen95, Sfe93, Sfe98, ST94, Sha98, Sha06, SLC07, Shi09, Shi97, Shi02, Shr91, SS94, Shu94, SS95b, Sim89, Sim90, Sim92b, Sim92c, Sit97, SK93, SK95, Ski90, Skixx, Slo93, Smi91, SB95, Sof93a, Soi95, Sol96, SDB93b, SDB93a, SRR94, Squ92, SBQ14, Rev93, Sta03, Sta19, ST99, ST04, SH10, SS93, Ste08, Ste13, Ste93, SR91, Sti93, Sti95, Sti02, Sto03, SC09, SG95, SGMM95, Str92b, Str93b, Str93c, Str94, Stu95b, Stu95a, Stu96, SS14, Sza00, Sza15, TJD11, Tab16, TM14, Tam91a, Tam91b, Tam97, Tam08, Tam95a, Tam95b, Tay91, TFS94, Tha89, Thi91, Tho97b, Tho97a, Tit94, Tod92, TT99, TT09]. Mathematica [Tor20, Tos10, Tre90, Tre91, Tro94, Tro95b, Tro95a, Tro97, Tro98, Tro00, Tro04a, Tro04b, Tro05, Tro06a, Tro06b, Tza09, UB93, UOT94, UK11, VMP94, Van93, Var91a, Varxx, Var97, Var90, Var92, Var93, Var96, Var91c, Ver97, Vin03, Vor97, Vos99, Vos00, Vve93, Vve94, Wag91, Wag94, Wag99, Wagxxb, Wagxxa, Wag10, WG92, WW95a, WGDZ04, WL11, Was05, Way89, Web92, Wee96, Mat91b, WGK05, Wel13, Wes92, WJ94, Wic96, Wie15, Wig92, Wit91a, Wit91b, WK17, WKB20, Wol88b, WW88b, WW88c, WW88d, Wol88d, Wol88e, Wol88c, Wol88a, WW88a, WW88e, WWH89, WW89a, Wol89, WW89b, WRR90, MAT90a, Wol91d, Wol91a, Wol91e, Wol91b, Wol91c, Wol91f, Wol92, Wol93b, Wol93a, Wol93c, Wol93f, Wol93d, Wol93e, Wol93g, Wol93i, WB94, Wol94b, Wol95b, WW95b, Wol95a]. Mathematica [Wol96a, Wol96b, Wol96c, WD97, Wol99a, Wol99b, Wolxxa, Wolxxb, Wol03, WW06, Xav22, Xem98, XMLC16, XKM93, YK95, Zac93, Zac98, ZD02, Zho97, ZO95, ZO02, Zit11, Ziz97a, Ziz97b, Zot07, dA93, ter03, vS89, von93, BS12, Ano97b, Ano98a, Ben98b, CTMC97, CKP13, Dar94, Har02, McC02, Mor98b, NP95, Rao98, Sch02, Sch98b, Sha96, Sol92, WJ02, Ano98d, Ano94c, Che25, Har19, Pie96, Tay92, Ano97d]. Mathematica-based [BKK13, BK15, BK16, Gli94, Gli95]. Mathematica’-charge [YK95]. Mathematica-to-Lisp [Fat90b]. Mathematical [Abe94a, Abe94b, Bee17, Cap03, Chu14, ES10, FKM95, Has03, HY14, Kan93, Mur93, RSL^{+}26, RS02, Soi95, Str93c, Zac93, Ano95c, Ano97b, Ano98a, Bag05, BS12, FFVH94, Hen94, KDM^{+}92, KLO96, MR01, Sha98, Sim91, Tho97b, Tho97a, YF04, McC02]. Mathematical-Function [Bee17]. Mathematica(R) [Ano97a, SÜ18, Ano96c, Ano97c, Ano97f, Ano99a, Ano00a, CAMM94, Ano97e, Ano95b]. Mathematica^{T M} [Har15, Lip12, Tam91d, Tam91c, Ano98b]. Mathematicians [DPV14, Mae93b]. Mathematics [ABD94, Buc93, Dev92b, Dev94a, Dev94b, FH96, FL92a, FL94, Fos93, GAM94b, GAM95, GG91, GG94, Har14, HC97, KM95a, KN94, Kro89, Lum93, Lum94, Mae90a, Mae00, Mor93a, Mor93b, Sei92, Ski90, Skixx, Sol92, SR91, UOT94, Var94, Web92, Wes92, Wol88a, WRR90, Wol91a, Wol93c, Wolxxa, Bat93, Ben98a, Fat96, Fow96, Gli94, Gli95, Has06, Hay93, KMH97, KS97, KN02, MM97, MM98b, Mat91g, MW95, PS03, Pet89, Rob95, Rus99, Rus04, Rus09, SW93, SLC07, Shi09, Sto03, Tha89, Tro00, WW95a, Way90, WW88a, Wol91e, Wol91b, Wol91c, Wol91f, Wol93a, Wol93f, Wol93d, Wol93e, Wol93g, Wol94b]. Mathematics^{T M} [Fos91c]. Mathematik [GAM94a, Str97, Wolxxb, San99]. Mathematiker [Mae93b]. Mathématiques [Xem98, Ziz93, Dur98, JLP99, Wol95a]. Mathematische [KWW92]. mathematischen [GAM94a]. MathHDF [JS93a, JS93b, JS93c]. MATHLAP [Ben98a, Ben98a]. MathLink [Coh94b, JS93a, JS93b, JS93c, MA00, MA01, Tod92, Wol93h]. MathLink-Based [JS93b, JS93a, JS93c]. MathML [Ano02, Har02]. MathQCDSR [WL11]. Maths [Sne94]. Maths-Phobics [Sne94]. mathscape [Bar98a]. MATHSCOUT [BC07]. MathScript [Ano99b]. Mathsource [Wol94a]. mathStatica [Vin03]. MathTensor [CP90, Sol96, TPH96]. MathUser [Mat95b]. MATLAB [Tay99, Bha05, Bha06, ES10, KSS07, LTE01, Pao99, Pao01, Ste13, CW05b, CW05c, CW05d, CW05e, MAC08, MM97, MM98b, Möl07, Ste08]. Matrices [EGM94, KMRC90, WG94, Wim12, BQO98, Kni95, PS06, SNLR92]. Matrix [Bit93, AC13, BMM90, Bro94, Dil91c, HL08, MAL92, Var97, WG92]. Maxima [LM90c, SH10]. Maximum [Cur95, Tza09, Jef93]. May [Aga90, IEE90a, IEE91b, IEE95b, PL92, PMDA96, VM90, HW92]. MD [FMRP16]. Mean [AC91]. Mean-field [AC91]. Meaning [AB92a]. Measure [WKB20, MS08]. measurements [KS93]. measures [Neu11]. Measuring [DPU94e, DPU94g, DPU94f, DPU94c]. mecanismelor [Ano94j]. Mechanical [DHE92, NEH90]. Mechanically [Bar91]. Mechanics [Ano99b, BPR99, BPR00, CTMC97, Fea94, Kaw93, Tit94, WG94, Ano97e, Bau96, CD02, Dil91c, IA93, KVF99, NEH90, RLM06, Ste94, SH10, SÜ18, WG92]. mechanics-related [IA93]. mechanisms [Ano94j]. mechano [Rid90a]. mechano-electrical [Rid90a]. Media [Ano97d, SR90, LZ96, OT94]. Media/Cambridge [Ano97d]. Medical [KMRC90]. Meeting [NEH90, Ano94b, Ano95d, Jef08, Rog92]. Meets [Tam95a, Tam95b]. Meisel [Ish16]. Mellin [GKR07]. Membrane [AB92a]. Memorial [Ano17]. memory [Ano91a, GSR90, SLR^{+}95]. Merges [Cha91]. merging [Cha91, IEE90a]. Message [Kro98]. messages [Wit91a]. metabolism [MK03]. metal [Bit93]. Method [BH93, Edi90, HSH07, Mik94b, ABMBR15, Bis95, BP95, DF96, HT13, Ioa92a, Ioa93, Mar00, Mik94a, Nam94, Skr93, VMP94, WGDZ04, YK95, dA93]. méthodes [Sfe98]. Methodist [IEE90e]. Methods [Ano96h, CAMM94, GK11, Gra94c, Gra97b, MT96, Pao99, AJ95, AGW95, Bag05, Ban01, Bha00, Cap03, CH94, Dub03, FM02, Gra98b, Has03, Lic11, LRW95, Pao01, Rus99, Tay99, WA96]. Mexico [IEE90b, Ryn93]. Miami [IEE91d]. Michigan [IEE95b, PL92]. microeconomics [Sti02]. microelectronics [IEE94b]. micrographs [Nov17]. microscopic [Var97]. microscopy [Sar97]. Microsoft [WW95a, Wol94b]. Microwave [Aga90, Rid90b]. Middleware [Kro98]. MIMO [SHH93]. MIMOQCAD [BGH96]. Mindlin [DCN96]. Minima [LM90c]. Minimal [BCDH94, MR94, CH94]. Minimax [Ioa92b]. minimizing [Lor08]. Minimum [RRIR^{+}92]. Minneapolis [IEE93, IEE95d, IEE96]. Minnesota [IEE93, IEE95d, IEE96]. Misfortunes [DPV14]. Mitte [BHJ^{+}00]. Mixed [Kro98]. Mixed-Signal [Kro98]. Mixing [SR90, Maj96]. mixtures [Bro94]. Modal [McP92, Wei92]. mode [BYA94, IOA^{+}95, LS96, Sch90]. Model [AH91, Ano92d, Bob93, Mar94, Mil90, Nog91, PC93, BM94, BK96, FP90, KL92, Ma95, Var97, vdG94]. Modeling [Ano97c, EP96a, EP96b, GN96, Kan93, Kor94, RB95, Sch98b, Van93, Var92, Var93, Zac93, Dar94, EP03, GD98, KDM^{+}92, KB93, SC10, Var96, Ano97f]. Modelling [Ada96, HW92, Mor98b, MK03, Rao98, Sha98, Chr93, FFVH94]. Models [BSS93, Dau17, MB94, Apt91, AC15, AC18, BYA94, GSR90, Gui06, Mor95, PL92, Sha98, SRR94, Sti93]. Modern [Gra93, Gra97a, Gra98a, GAS06, SC10, BS12, Noy95, Ano98d]. modernen [Kus97]. moderner [Gra94a]. modulation [Das90]. Molecular [DHE92, McC09]. molecule [Xav22]. molecules [NG93, Sen95]. moment [KS93]. Moments [BM92, Pro96]. Money [BHJ^{+}00]. Montreal [Lev95]. Moshinsky [XMLC16]. Moss [BHJ^{+}00]. Motion [Kou16, BM94, Mar00]. Mountain [Kro98]. Moving [Yag92]. MP [Ziz97b]. MPSI [Ziz97b]. MS [Fos91c, Wol88b, WW88b, WW89a]. MS-DOS [Fos91c, Wol88b, WW88b, WW89a]. MT [HHP^{+}14]. MTT [Aga90]. MTT-S [Aga90]. Multi [BFZ14, Kro98, TF93, KS93, WGDZ04, Xav22, ter03]. multi-derivative [WGDZ04]. multi-j [KS93]. Multi-loop [BFZ14]. Multi-ORB [Kro98]. multi-scale [ter03]. Multi-solutions [TF93]. multi-state [Xav22]. multicell [AdSR96]. multicomputers [SLR^{+}95]. Multidimensional [EMB92, Eva93a, EBM94, MC96]. MultiHypExp [Ber24]. Multimedia [MG94, De 94, GMP97, OT94, Phi91, PWW90]. Multiparadigm [SV93]. Multiple [Asl96, Kor11, Ber24, Iss96, ST12]. Multiple-Precision [Kor11]. Multiple-Valued [Asl96]. multiplication [MP10]. multiply [Bet90]. multipole [LRW95]. Multirate [EMB92, Eva93a, EBM94]. multistate [JY93, YJ92]. Multivariable [BG92, CLR98, ABFP24, BGH96, Hol98, Hol03a]. multivariate [Ber24, LR18a, LR18b, Pet08, vdG94]. MultivariateResidues [LR18a, LR18b]. MuPAD [Ben98a, Ben98a, Grä96, PZ96]. Murray [McC02]. Music [AH90].

NAG [BKR^{+}91]. Name [Lev92a]. Nancy [Squ92, Mac92]. nanoparticles [Nov17]. Nash [DK91]. National [Ano94f]. native [Ano95e]. Natural [WY93, GH04]. Nature [Ada96, Ano97f, GN96]. Naturwissenschaftler [Mae93b]. Navigation [IEE90c]. navigator [Rus99, Rus04, Rus09]. Near [Bah92, CW05e]. nearness [DPU94c]. NEC [Ano92d]. Nested [Ada92]. Nets [MM93]. NetTracker [Ano97g]. Network [BCDH94, Kro98, MA00, MA01, Rid90a]. Networks [Fre92, Fre94b, IEE95d, RRIR^{+}92, Cau90, Yu94]. Neural [Cau90, Fre92, Fre94b]. News [Bra96, Wee96]. Newton [Bis95, BK96, Edi90, Nam94]. Newton-Euler [BK96]. next [HHWT95, Cra90]. Nineties [BPU90a]. Nishidate [Ano97f, Ada96]. NIST [Bra96]. nitrogen [NG93]. NMR [GSOW93, Shr91]. no [Tam95a]. No.1996 [Ano96g]. No.1996/078 [Ano96g]. Nodal [Rid90c]. Non [Per03, RP92, Che96, Hof91, Neu11]. non-deterministic [Hof91]. Non-linear [RP92]. Non-Parametric [Per03]. non-uniform [Neu11]. nonhomogeneous [GMP09]. Nonlinear [Ano96a, CMT12, CT94, Edi90, EM01, OM92, PGB94, Vor97, BB95, Bes98, FM93, GI98, HD91, NA90, NA92, OFM94, Rao93, RMC96, SKS95, TSC93, WW06, Yu94, Har15]. nonlinearity [NR96]. Normal [Abe90, ZL95, vdG94]. Northcon [Ano90c].

Northern [Lum94]. Norton [Ano92d]. Norwegian [San95]. notable [Wol16]. notations [Har00]. Note [Sah96]. Notebook [Ano92d, Dub90, Mez92, CMT12, von93, Har15]. Notebooks [DF91, PW94b, Sit97, Win91, Ano97a, PW97, Smi93]. Notes [SS14, WW93c, Wag94, Wol93i]. Nothing [Kor94]. Nouvelle [Sfe93]. November [IEE91a, Lum93, Lum94, NEH90]. Nucci [Dev92a]. nuclear [HL08]. Nucleus [Bob93]. number [MT94a, Mis13]. Numbers [CZ09, Alc97, Lew91]. Numeric [BKR^{+}91, Sof93a, HS90, LP05]. Numerical [BSS93, Car92, Dub03, GV96, HSH07, HRV92, KG90, PGB94, SK93, SK95, SS14, Wim12, ABMBR15, AJ95, Ban01, BP95, Byk19, HHL06, Ioa92d, Ioa92c, KW92, KS97, Mit89, Noy95, PH95, Rid90b, SS05, WGDZ04, WA96]. numerico [AF97b, Pér95]. Numerics [Tro06a, HPR96, Tro05]. NY [Kap92].

O [Shi97]. Object [FFVH94, IA93]. Object-oriented [FFVH94]. ObjectMath [FEV93]. Observation [CF96, CS96a]. observer [SKS95]. observer-based [SKS95]. obtain [Jef93]. Occurrence [BHJ^{+}00]. Oceanographic [GMM95, SGMM95]. October [Ano90c, IEE95d]. ODE [BGS93, PZ96]. Office [Mat91h]. official [Wol99b]. old [Old92]. Oliveira [BHJ^{+}00]. Olsson.wl [ABFP24]. One [BH07, BH08b, BH08c, WG94, Fos91b, Hid21, Pat15, Pat17, WG92, WGDZ04]. One-Dimensional [WG94, Hid21, WG92, WGDZ04]. one-loop [Pat15, Pat17]. One-sample [BH07]. One-Sided [BH08b, BH07]. ongoing [Ken01]. online [Mis13]. only [JM08, MS08, Pet08]. Ontario [IEE91b]. Open [Ano97g, Lic11, McA08, Ste08, Ste13]. open-source [McA08]. OpenMath [PDP00]. operational [Ano91a]. operations [Ano95e, Has06]. Operator [GSOW93, Bes98, Eks11, Shr91, Thi91, Zit11]. operators [Bit93, Bro94, DT00]. opportunity [Wat90b]. Optical [CT94, O’S96, Rom09, Wei92]. Optics [Möl07, Rom09]. optimal [Sto03]. Optimization [Gil95, ST90, Bha00, Cal95b, Kor97, LEF91, Mat91f]. optimized [DF96]. Optimizing [Sam94]. Optimum [Bha96]. Option [BSS93, Lew00]. options [Sto03]. Optische [Ano95g]. ORB [Kro98]. orbital [Cha91]. orbitals [Jon93, WAW96]. Order [BD93b, Sof96, Ano96f, BBD93, HT13, Mat89b, Mat89c, Pro96, WGDZ04, dT93]. Ordering [MC96, AC91]. Ordinary [Apt91, Che25, GK11, GMP97, MR01, ZL95]. Oregon [Lum93]. organized [MC93]. oriented [FFVH94, LW99]. Orleans [IEE95e]. orthogonal [MdR13]. orthogonality [ÖAÖ24]. orthotropic [NR96, Rao93]. OS/2 [Tam95a]. Osborn [Ano98b, Röh99, Sha96]. oscillator [HL08]. other [BC07, Tod92]. Outer [WKB20]. outline [Don01]. output [BC07]. over-$400 [Gal91]. Overdamped [FP90]. overlap [Jon92]. own [Sch92, Win91]. Oxford [ACM94]. ozone [NG93].

P [Rav07, Rue07, WW93a, WW93b].

Package

[Ano95f, BH93, BG01, Bel90, CS96b, Chu14, Koe94, NP95, Ngu93, Wat90a, AU21, ABBD23, AC15, AC18, Aza13, BK13, BC07, Bel93, Ber24, BGH96, Bzo21, CKS00, CMS17, Dep17, Dvo90, Eks11, FRW17, Fer07, Fon21, GKR07, GLMG12, Gra04, GHIL09, GI98, Hid21, HHP^{+}14, HM06, HHL06, Jac19, JL93, KSA92, LR18a, LR18b, LM23, Lor08, Luc95, Mad91, McC00, Mie95, Nag07, Naz12, NS94, Old92, Pat15, Pat17, PSW23, PTL04, Pre18, Ros16, SHH93, Sch94b, SÜ18, Thi91, Tor20, Var97, WL11, Way90, Wie15, Zit11, Pat15, Pat17]. Package-X [Pat15, Pat17]. Packages [Coh92, Wol93b, Wol95b, Wol96c, AR93, ABFP24, BFS93, Bet90, Boy91, BKK13, BKM14, BK15, BK16, Cof92a, Dil91a, Gal91, Kra93, Kra95, Sim91, Sol96, Wit91b, Wol99b, Ano97d]. Packard [Tha89]. Packel [Ano97a]. pages [Ano95b, Ano96c, Ano97a, Ano97b, Ano97c, Ano97e, Ano97f, Ano97d, Ano98b, Ano98d, Ano99a, Ano00b, Ano00a, Rue07]. Pair [Str93a]. Pair-Distribution [Str93a]. Pais [BHJ^{+}00]. Palazzo [GT94]. panels [RP92, Rao93]. Pao [Tay99]. paper [HW92]. paperback [Ano97d, Ano00b]. Papers [PMDA96, Rog92, Ano95h, PMD92]. paradigm [Mae92, SLR^{+}95]. Paradigms [Swa93a, Swa93b]. Paradox [Wag93]. Parallel [IEE90e, JM93]. parameter [KLO96, SC09, Tza09, YM03]. Parameters [Mar94, HM06, KW92, Neu11, ÖAÖ25, PL92, Wes93]. Parametric [LM90a, Per03, TF94, BSB94]. Part [Abe94a, Abe94b, BPU90b, BPU90c, BPU91, CDS^{+}92, FP90, Iss96, TIM96, CW05b]. Partial [Asm05, Che25, KPS97, KPS03, Mat90b, ST99, ST04, Tho91, Tho92, Vve93, Vve94, GV96]. partially [Bha96]. ParticleRecognition [Nov17]. particles [Hug93, KLO96]. particulate [HW92]. partition [AU21]. Partitioning [Fat15, Xav22]. Parton [CGO17]. paso [Mor98a]. Path [LM93]. Paths [Wel93]. Patrick [Ben98b]. pattern [Wei92]. Patterns [PW94a, GH04, Gro01, O’S96, Sch92, Wei92]. Paul [Ano96c, MT96, TMC96]. Pawlowski [TPH96]. PC [EN96, GT94, IOA^{+}95, KB93, Ziz97b]. PC-based [KB93]. PCSI [Ziz97b]. penalized [Lor08]. penetration [HFO94]. people [Wol16]. PerfectBACKUP [Ano97g]. Performance [AJ95, All94, Hay92a, SS95a, Zot07, All92, Var91c]. Performances [BB95]. Performing [CP90]. period [Cha91]. periodic [LZ96]. person [Can00, Can03]. personal [Wol16]. Perspectives [RdG96, Wol16]. Perturbation [BK13, DHE92, SÜ18]. Pfaffian [Wim12]. Phase [Sko94]. Phenomena [Ano96a, Gay92c, May94b, WY93, DWY98]. phenomenology [GHIL09]. Phil [Lip12]. Phobics [Sne94]. Phone [Ano92d]. photonics [Har19]. Phys [Stu95a, Stu95c]. Physical [Ano96a, BHJ^{+}00, GW95, GWWJS95, Iss96, Lip12, MT96, Oln98, Rav07, Rue07, TMC96, Cro98, Gre05, Har14, KR98]. Physicist [Ben98b, Tam97, Tam08]. Physics [Abe94a, Abe94b, Ano96a, Bau96, CTMC97, Com96, DN93, KR98, RRM90, SBA94, Tam91c, Tam91a, Tam91b, Tam91d, ZO95, Ano97e, Bau05, Cap03, CLR98, Cra90, Cra91a, De 99, EE96, EM01, GT94, Har19, Has03, HY92, McC09, Ryn93, Wie15, ZO02, Oln98]. Physik [Kus97, Bau93]. PID [Cal95a, Cal95b, RMC96]. Pisa [IEE95a]. Pittsburgh [VM90]. Pivoting [Ede92]. plagued [Ano91a]. Planar [McP92, Aza13, GMP09, Kou16]. Plane [Cox93, Tho92, IA95]. planes [BP95]. planetary [Ma95]. planning [RF96]. Plans [IEE90c, Dot90, Mad91]. Planteamiento [DFPA95, AF97b]. Plasma [IEE95c, SBA94, Gla90]. Plasmas [SBA94]. plasmons [SC10]. plate [HFO94, KAG96, Skr93]. plates [DCN96]. Platforms [Fos91c]. platonic [GGV94]. plotting [ABK95]. Plug [Ano96i]. Plug-In [Ano96i]. plus [MT96, BS94]. Poincaré [Mar00]. Point [LS95, SS94, Shu94, Bah92, Mar00]. Point-and-Click [SS94, Shu94]. points [DK93, Råd93c, Rad93b]. Polarimetry [BG01]. polyhedra [GGV94]. polylogarithmic [BBP97]. polylogarithms [Ber24, Maî06]. polynomial [BMM90, Grä96, Hen96, LW99, MP10, PS06]. polynomial-oriented [LW99]. Polynomials [HM92, Knu92, Mat89a, ADG94, DK93, EM93, EMEH96, Kou16, MdR13]. Polywell [SS95a]. POMA [GSOW93]. population [Mat92a]. porous [LZ96]. Portable [Bee17, Kor11]. Portland [Lum93]. Position [IEE90c]. positional [RL19]. Post [Mac92, SYS^{+}90]. Post-Glacial [SYS^{+}90]. Post-Use [Mac92]. postgraduate [Har15]. postprint [Joh95]. postprocess [BC07]. PostScript [YF04]. Potential [LU95, Wat90b, FMRP16, HBN95, Hug93, Sen95, TIM96, TM14, Xav22, YK95]. POTLIB2Math [TM14]. pour [Wol95a, Ziz97a, Ziz97b]. Power [HSH07, Koe93, Koe94, Wag94, Fow96, MC93, Tre90, dA93]. powerful [Pri93]. PowerPC [Ano95e]. PowerPC-native [Ano95e]. pp [Ano94c, McC02, MT96, Pie96]. practica [Ano94j]. Practical [Bha00, Bla92a, Bla92b, Mac92, SB92, Squ92, BW99, KNJ93, Kau94, Wes99]. practicas [AF97a, BLM97, FSC95a, FSC95b]. Practice [Mae00, Ano94j, Mat90b]. Práctico [UB93]. practitioners [Ano97b, Ano98a, Tho97b, Tho97a]. Prague [DN93, Str92a]. Praktiker [Kof93]. Praktisches [KWW92]. pratique [Poi98]. Precalculus [Ree00, TT99, TT09]. precedes [Dev92a]. Precise [SS05]. Precision [BH08b, KMRC90, Kor11, Hug93]. premiers [Ziz97a]. Prepackaged [Fos91d, Fos91b]. préparatoires [Ziz97a, Ziz97b]. preprints [Str92a]. Presentation [Stu96, Tre91, Ano96e, Yag92]. Presentations [Yag92]. presented [NEH90]. Press [Ano97d, Ano98d, Ano00b, Rue07, Tay99]. pressure [LBSA96]. Pricing [BSS93]. prime [San95]. primer [Coo98b, Röh99]. primeras [Ram97]. primitive [BP95].

Principi [Fal93]. Principles

[BHJ^{+}00, SBP96]. Printed [Sko94]. Prix [Ano92d]. Pro [Ano92e, Ano97g]. probabilistic [Gui06]. Probability [Kob94b, Rad92, Rad93a, Has01, Has10, Rad95]. probe [Sar97]. Problem [Abb92, AB92a, FW91, SYS^{+}90, Ano98c, Cro91, GV96, Har19, Haz10, KVF99, LP05, SLC07, Shi09, YK95, Zac98]. problem-centered [Haz10]. Problemas [Mor98a, AF97b, DFPA95]. Problèmes [Dur98]. Problems [Car94a, EP96b, KC94, Oln98, Sca94, Ano91a, Asm05, DF96, Don01, EE96, Ioa92a, Ioa92d, Ioa92e, Ioa93, IA95, KR98, KPS03, MR01, MC96, PH95, Pri93, TS00]. Procedure [LP13]. Proceeding [Ano94f]. Proceedings [ACM94, Ano91c, Ano92c, Ano94g, AH90, BH95, Ham94, IEE90d, IEE90e, IEE91d, IEE95b, IEE95d, IEE95e, IEE96, IBM^{+}95, JOS93, Kap92, Lak96, Lev95, Lum93, Lum94, MCG13, PL92, PMDA96, RPWW96, SBA94, SE11, VM90, vdHvH12, AKLG95, Ano95h, Bro93, DN93, GT94, Gut04, HRV92, IEE91c, IEE95a, KM95a, KMH97, LT94, MGC94, Mio93, PMD92, RCM03, RRM90, Tra00, Van95, WN90, Car96, HY14, Jef08, Mio90]. Process [Ano96d, Byk19, SBP96, Zho97]. Processes [GMM95, PSMM95, SGMM95, HW92]. Processing [EMM90a, EMT93, Eva93a, IEE90b, IEE90e, IEE91b, IEE92b, IEE93, IEE95b, ON92, Wit90, ZT94, BEL^{+}95a, Chr93, EMW91, LTE01, PWW90, Rog92, Run89, Sch92, Sla90]. Product [Gal98, GSOW93, Eks11, Shr91, Thi91, Tos10]. Product-Operator [GSOW93, Shr91]. Production [BHJ^{+}00]. Products [Ano95f, Ano97g]. Profiles [XKM93, HKM93]. Program [DK91, Fos91c, Phi90, Rid90c, ZD02, AC13, Arc20, FMRP16, Gro01, JM93, KDM^{+}92, MS98, Rid90b, Skr93, Var97, WGDZ04, WW06, Xav22, XMLC16]. Programm [Vor97]. Programme [Ziz97b, BSB94]. Programmer [Mae94c, Mae94a, Mae96a]. programmiert [Ano95g]. Programmierung [Nol97]. Programming [Bee17, Car94a, GKW93, GKW94, Gra94c, Gra97b, Mae89, Mae90b, Mae91, Mae94a, Mae96b, Mae97, Maexx, Oln98, Swa93a, Swa93b, Wel13, Ziz92, Ano96c, Ano98c, Car94b, Cav93, CP93a, EN96, GKW96, Gra98b, HMM15, HMM16, KR98, LEF91, Mae92, MA00, MA01, Sch98a, Tro04b, Was05, WGK05, Zac98, Zyp93]. Programs [CDS^{+}92, Hay92a, Pao99, Yag92, Ano97b, Ano98a, Bar90, BC07, Cro98, Lew91, Pao01, Pet89, Sei92, Ski90, Ste94, SH10, Tay99, Tho97b, Tho97a, Var91b, Wes93, dT93]. Project [Kro98, ES10]. Projectile [LM93]. Projective [Cox93]. Projects [ACCM93b, AM94, Cra94, EP96a, Str93c, Str99, ACCM93a, Cal04, Kir94]. promising [Old92]. Proof [Slo93]. Proofs [CH95, Nai22]. Propagation [Ano95a, CT94, IEE90a]. propagator [MS98]. properties [KL92, KLO96]. proposed [AGW95]. Prospects [Chr93, Chr94]. protected [Sch90]. Prover [CZ92]. provides [Mae92]. proving [PS95]. PSI [Ziz97b]. PSpice [KZK10]. PT [Ziz97b]. PTSI [Ziz97b]. Publication [CW05e]. Publishing [Ano95b, BP94]. Puisieux [Koe93]. Pulsatile [XKM93, HKM93]. pulse [HFO94]. Pulses [CT94]. pulverized [KL92]. Purdue [HRV92]. Pure [Var94]. Python [Sta19].

QCD [AC15, AC18, BK13, WL11]. QCWAVE [TJD11, Tab16]. QDENSITY [JDBT06, JDBT09, Tab16]. QDENSITY/QCWAVE [Tab16]. QFT [SHH93]. qFunctions [AU21]. QMeS [PSW23]. QMeS-Derivation [PSW23]. QSWalk [FRW17]. quadratic [Kni95, SNLR92, TPH96]. quadrilateral [Kni95]. Qualitative [Nis94, Cro91, Rao93]. Quantenmechanik [Kus97]. Quantitative [Ano96h]. quantization [Zit11]. quantizer [Mie95]. Quantum [DHE92, Fea94, FM02, JDBT06, Ste94, SH10, Tit94, WG94, Aza13, FRW17, Fer07, JDBT09, Mis12, Mis13, SÜ18, TJD11, Tab16, WG92, WC98, CAMM94]. quark [CKS00]. quasicrystals [Gro01]. queries [LM90a]. question [Bar10]. Questions [DPU94a]. Queueing [AH91, FP95, PF95]. Quick [BC92, BC94]. QuickBasic [Tay99, Pao99, Pao01]. Quickly [Sch95]. QuickTime [Tod92]. QuickTime-compatible [Tod92]. quotient [Ano96f].

R [Ano97c, Ano98b, Dar94, MT96, Mor98b, Rao98, Röh99, Sch98b, Sha96, TMC96]. Raczka [TPH96]. Radar [BG01]. radiation [Aza13]. radiative [KL92]. Radicals [Ada92]. radioactive [KA99]. Raffiniert [BHJ^{+}00]. Rainey [Lum94]. Ramanujan [Kei96, KL02]. Random [BN91, Råd93c, Rad93b, Byk19, DWY98, Mis12, Mis13, Yu94]. Random-Dot [BN91]. randomness [Cha01]. range [Str93a]. rank [BQO98]. rank-revealing [BQO98]. Raphson [Bis95]. rapid [BBP97]. rate [HR11]. ratio [Ano92a]. Rational [BH07, BH08c, EGM94, Ohm94, Kro98]. Raton [Ano98d]. ray [HFO94]. Rayleigh [MN91, DF96]. Rayleigh-like [MN91]. RComp [NP95]. RDBMS [Ano97g]. reactor [Kul91]. reader [CGO17, Sta19]. reading [Ano95b]. Real [Cox93, IEE95a, KW92, DK93, ST11, ST12]. Real-Time [IEE95a]. Rearrangement [PW94a]. reasoning [Nis94]. Rebound [SYS^{+}90]. receives [Ano93c]. Recife [CD02]. reconstruction [JM93]. record [Ano90c, IEE90c, IEE94b, IEE95c]. recording [LBSA96]. Recreations [Var91a, Varxx]. Rectangles [Slo93]. rectangular [Che91]. recurrences [Pet08]. Recursive [BK95, NP95]. Recursively [Mat93]. Red [Ano95f]. Redirector [Kro98]. REDUCE [Ste94, Grä96]. REduction [BKM14, BFZ14, BKK13, BK15, BK16, MS98, Stu95b, Stu95a, Stu95c, BKK13, BK15, BK16]. redundant [Che96]. Reed [Sob95]. Reents [Oln98]. Reference [BC92, BC94, Wol92, Wol93h]. References [Ano95f, Bat93]. regimes [YM03]. regions [ABFP24]. Regression [Edi90]. regular [BP95]. reinforced [TD95]. Reise [BHJ^{+}00]. Related [LU95, Bar92, Has03, IA93]. relation [ADG94]. Relations [Bah92]. relationships [ÖAÖ24]. relativistic [Aza13]. Relativity [HPR96, Sal94, Har03, Tor20, GLMG12]. Release [Bra96, Wol93i]. Reliability [Rad93d, Che94, Sam94]. remediation [MGS96]. renormalization [RF95]. Replace [Lew94]. replicator [Arc20]. Report [Fos93]. Representation [FKS20, FK15, Naz12]. Representations [Roa96, TF94, ABFP24, Bd93a, GKR07]. Representing [GVV91]. Reprint [HW92]. Resampling [EGM94]. Research [Chr94, Pie96, Sim92c, Mat91b, Mat05, Bos96, Has06, Ano91b, Kro89]. researchers [Ive92]. reshaping [Sim92b]. residues [LR18a, LR18b]. resistor [Yu94]. resolucion [AF97b, DFPA95]. résolus [Dur98, JLP99]. Resolving [Can03, Can00]. Resources [Ano95f, BHJ^{+}00]. respectively [Kra93, Kra95]. Response [CW05a, Iss96, CP93b]. results [MT94b, MT94c]. revamped [Ano93c]. revealing [BQO98]. Review [Ano94c, Ano95b, Ano96c, Ano97a, Ano97b, Ano97c, Ano97e, Ano97f, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano00b, Ano00a, Bar92, Ben98b, Ber96, Che25, CAMM94, CTMC97, Fat92, Gal98, Har19, Her88, Lip12, Mac92, McC02, Mor98b, Oln98, Pie96, Rao98, Rav07, Ree98, Ree00, Röh99, Rue07, Sch94b, Sim91, SMC98, Tay99, TMC96, Vin03, Web92, Wes92, Bel89, BHH91, Car08, Dar94, MZ92, PZ96, Wes95, Wes96b]. Reviews [Ano93d, Ano94d, Bra96, CW05c, CW05e, Kro89, Lev94, MT96, Sch98b, Sha96, Sol92, Tay92, Wee96, Dev92a]. revised [Coo98a]. Revisited [Hay92a]. revolutionary [Sim89]. Revolutions [BHJ^{+}00]. Revolutions/Pais [BHJ^{+}00]. rewrite [Har94]. RHINE [Car96]. Ricci [Agu90]. Richard [Ano96c, Ano97f, Ano17, MT96, Tay92, TMC96]. ridden [Sim89]. Ridge [Ano02]. Rigid [Mur92]. ring [KAG96]. Riverside [IEE95e]. RKR [Sen95]. RLC [FP90]. robot [Che96, LJW91, NS94, PS93]. robotic [Sch94b]. Robotica [NS94, Sch94b]. Robotics [IEE96, IMAS95]. Robots [Gil95, KYB94]. Robust [SKS95]. ROC2.wl [ABFP24]. Role [Car92, Gli95]. ROM [Ano98a, Ano00a, MT96, ABGA96, GAM95, Tro98]. Ronald [Ano98b, Sha96]. Root [Gea91, GK92, ST11, ST12]. roots [Nam94]. Rose [McC02]. Rosemont [Lum94]. Rotating [Mur92]. Rotational [DHE92]. rotor [NG93]. Rule [Mat91e, Bar93, Har94]. Rules [EBM94, Sof93b, MC96, WL11]. RunDec [CKS00]. running [CKS00]. runs [Tam95a].

S [Aga90, BH07, BH08b, BH08c, CW92, MM08, WW93a, WW93b]. S115 [DPU94a]. S14 [SS94]. S94 [Raa94]. SAGE [JM08, ABMBR15, ES10, Ste13]. Sample [BH08b, BH08c, BH07, Ste94, SH10]. sampler [Var91b]. Sampling [BH08b, BH08c]. Samuel [Ano96c]. San [AKLG95, IEE92b, LT94, Rog92, SE11, Ioa92e]. Santander [Gut04]. Saratoga [Kap92]. SAT [Raa94, Str94]. SAT-M31 [Str94]. SAT-S94 [Raa94]. satisfaction [RF96]. saturation [EN96]. scalar [Stu95b, Stu95a, Stu95c]. Scale [LT94, MGS96, Sam94, ter03]. Scale-up [MGS96]. Scales [UK11]. scaling [Bis95]. scanning [Sar97, Yar95]. scans [Bra96]. scattering [KLO96]. Schaum [Don01]. Scherrer [Zyp93]. Scholes [Mil90]. School [UOT94, Van95, Saf91, SW93]. Schritte [Bur93, Bur94a]. Schrödinger [WGDZ04]. Schubert [ZD02]. Schwerpunkt [Kus97]. Science [Har15, IEE95c, Mae93b, Mae00, Ras91, CMT12, DWY98, Gas98, Gup91, Har14, KNJ93, PL92, Wol02]. Sciences [Ano92c, Ano96h, Cra91b, Cra91c, GW95, Lip12, Rav07, Rue07, Bat93, Gre05, IEE90c, Tay92]. Scientific [BHJ^{+}00, Cra94, Cra96, Doe95, MR01, Sim93, Str93c, Ano98c, BC07, HPR96, HRV92, Ive92, LP05, MZ92, RLM06, Zac98]. Scientifique [Sfe93]. scientifiques [Ziz97a, Ziz97b]. Scientists [Bah94, Iss96, Mae93b, Wee96, Ano95b, Dub03, EM01, Gas98, Gra04, Pid96a, Pid96b, Sei92]. Scope [Har15, Ish16]. Scotland [AH90, Tra00]. script [HL08]. ScriptEase [Ano97g]. SDK [Ano97g]. Seattle [Ano90c, Ano94f]. Sebastian [Har15]. Second [Ano94g, Ano00a, BBD93, HRV92, IEE90e, PMD92, Web92, Wes92, Ano96c, Ano98d, Mat89b, Mat89c, YM03, Zit11, dT93, KMH97]. second-quantization-operator [Zit11]. secondary [SW93]. Section [IEE90e]. sections [AdSR96, JM93, LM90b]. Sector [Nog91]. See [Ano92f, Bar92]. seeing [Ive92]. Selected [Ano97e, Bat93, CTMC97, WW93c, Adi89, Bau96]. Selection [PMDA96, PMD92]. Self [BD91, BD92, BDxx, KN02, MC93, Neu11]. self-contained [KN02]. self-organized [MC93]. self-similar [Neu11]. Self-Tutor [BD91, BD92, BDxx]. Semi [Ioa92c, Ioa92d, Wie15]. semi-automatic [Wie15]. Semi-numerical [Ioa92c, Ioa92d]. Semiclassical [JY93]. Semigroups [Sut92]. semileptonic [HL08]. sensitivity [MAL92]. Seoul [HY14]. September [Ano94g, EJ94, IEE94b, Joh95, Mio93, Str92a, Van95].

Sequences [Mat92c, Mat92b, NP95]. serial [JM93]. Series [CZ09, Ioa92c, Koe93, Koe94, Mat92c, Mat92b, Mat93, PW94a, ABMBR15, AU21, ABFP24, Ano96e, Asm05, Hid21, Kei96, Kra93, Kra95, Sti93]. Server [FKM95, Sch02, WJ02]. Services [WJ02]. Session [Ano95h]. Set [Coh92, RB93, SZ96]. sets [GH04, Hen96, Nie03]. Setting [Tre91]. seven [Fos91b]. SevenOperators [HL08]. Seventh [IEE91d, KN94, RPWW96]. Several [CDS^{+}92]. sFr [Ano97a, Ano97c, Ano97e, Ano97f, Ano99a, Ano00a]. Shape [LM93]. shaped [GGV94, Nov17]. Sheets [BPU90b]. Sheffield [EJ94]. Shell [KM95b, NR96, RP92, Wol93f]. Sheraton [IEE91b]. Sherwood [Ano94a]. shock [CP93b]. short [HW92]. Should [Iss96]. Show [Ano92e]. si [Ano94j]. SICE [Ano95h]. Sided [BH08b, BH08c, BH07]. Signal [EMM90a, EMT93, Eva93a, IEE90b, IEE91b, IEE92b, IEE93, IEE95b, Kro98, ON92, ZT94, BEL^{+}95a, EMW91, LTE01, PWW90, Sla90]. Signals [EM92, Wol97, EKWM93]. silence [Ive92]. Silentwriter [Ano92d]. simbólico [Pér95]. similar [Neu11]. Simple [AW97, Bar91, HKM93, KLO96, Wel93, XKM93]. simplification [BBK14, HSW94, HSW98, SBQ14]. Simplified [KL92, PC93]. Simpozion [Ano94j]. Simulating [GD98]. Simulation [AKLG95, BK96, CH94, MT96, Sob95, Bag05, BEL^{+}95a, Eva93b, JDBT06, JDBT09, KB93, MAL92, Mie95, RB93, Sav95, TJD11, Tab16, YK95, Yu94]. Simulations [GW95, GWWJS95, PSW93, TMC96, Ano97f, Pid96a, Pid96b]. Simulator [Kro98, Bel91]. Simultaneous [PS93, Kor97]. sine [BS11a, BS11b]. Single [Tre91, Aza13, Bar98b, Cha91, Mae92]. singular [CKP13, Ioa92e, Ioa92f]. singularity [Che96]. SITE [RPWW96]. six [ÖAÖ25]. Sixth [LT94]. size [KLO96]. size-parameter [KLO96]. skeletal [Nie03].

Skew [Wim12]. Skew-Symmetric [Wim12]. Skiena [Ski90]. SLAM [MZ14]. Slater [Jon93, KU94]. Slater-type [Jon93]. Slavnov [PTL04]. Slavnov-Taylor1.0 [PTL04]. sliding [LS96]. slightly [RP92]. slowly [TIM96]. small [KLO96]. Smith [McC02, EGM94]. smoothly [Ano96f]. smooths [Gal91]. SNAP [Nag07]. SNEG [Zit11]. Soar [Pat93]. Social [GW95]. Society [Ano91c, IEE90e, NEH90, RPWW96, GD98]. socioeconomic [GD98]. Soft [Ano94g, Kro98]. Software [Ano93a, Ano93d, Bar92, BH93, Bee17, BSHS93, Cal95a, Car92, Com96, ES10, Fos93, Fos01, Gal91, Hen94, HY14, Kan93, KC94, Kro98, Ngu93, Sim91, Sim92d, Wes95, Wes96b, Ano89, Ano92a, Bel89, BFS93, CKP13, Cra91a, Fos92, GSR90, Hug93, Ioa92f, Lew91, Luc95, Mad91, MZ92, Mat91c, McA08, Mie95, Pat93, Pro96, Ric91, Sch94b, Sim92a, Sim92b, vS89, Kro98]. software-math [Fos92]. soil [MGS96]. solid [Har19, Mar00]. Solitaire [Ano92d]. solitons [CF96, CS96a]. Solomon [Sob95]. Solution [Ano88a, Ioa92c, IA95, Van93, Bes98, BK95, DCN96, Ioa92a, Ioa92d, Ioa92e, Ioa93, KA99, MC96, Pri93, WGDZ04, WW06]. Solutions [Car94a, Tho91, ABF14, AF92, FP90, GV96, HW92, Mat89b, TF93, Tro97, ZM93]. Solve [Ano92f, LM90c, Bar92, Grä96, Sei92]. Solvers [Ber96, PZ96]. solveTruss [Ozb17]. Solving [AH91, Che25, FW91, GK11, KC94, Mat91c, Pet08, RLRML04, Abb92, ABBD23, Ano98c, Cro91, GV96, Har19, KVF99, LP05, PH95, SBP96, SLC07, Shi09, Zac98]. Some [Bar91, BKR^{+}91, EE96, JKP96, MT94b, MT94c, WKB20, CKP13, Wol16, dT93]. Sons [Ano97b, Ano98a, Ano98b]. Soon [CW05e]. soot [KLO96]. sophisticated [Old92]. Sound [Ano97g, Ive92, Sim92b, Nol97]. source [JM93, McA08, Ste08, Ste13]. South [HY14].

Southeastcon [IEE91c]. Southern [IEE90e]. Space [AF92, LM91, Nai22]. Space-time [AF92]. Spain [Gut04]. spanning [MT94a]. Sparse [MP10]. Spatial [FW91, PL92]. Special [BS11a, BS11b, Kor11]. Special-Function [Kor11]. Specifications [Rei95]. spectra [Aza13]. spectral [KL92, KLO96, MC93]. spectrum [MZ14, TIM96]. Speech [IEE90b, IEE91b, IEE92b, IEE93, IEE95b]. speed [Ano95e, Cof91]. Speeding [FH96]. Sphere [RB95, Str93a]. Spin [Ma95, Bit93, Maj96]. Spin-up [Ma95]. Spinor [BG01, GLMG12, MM08]. spinor-helicity [MM08]. Spinors [GLMG12]. Spins [Shi02]. Spirals [Mon93]. Sponsored [IEE90e]. Spread [GVV91]. Spreading [Gay92c]. Springer [Ano96c, Ano97a, Ano97c, Ano97e, Ano97f, Ano98c, Ano99a, Ano00a, McC02, MT96, Oln98, Pie96]. Springer-Verlag [Ano96c, Ano97a, Ano97c, Ano97e, Ano97f, Ano98c, Ano99a, Ano00a, McC02, MT96]. Springs [Kap92, Auc96]. Square [GK92]. Squares [Mat91d, vdG94]. SSP [Ano95i]. St [Tra00]. Stability [Sof96, Ano96f, Cof91, Gla90]. stable [Cof92a]. Stack [Mat91a, Var90]. Stan [Ano97a, Ano00a, Tay92, O’R92, Sol92]. Standard [Ano97d, Wol93b, Wol96c, von93, AR93, AGW95, Boy91, Run89, WW88e, Wol99b, BS94]. Standards [Ano95a]. Stars [Sof96]. start [HMM15, HMM16]. Started [ST94, Che98, Che05]. State [Lum93, PL92, Sut92, FP90, Har19, Xav22]. states [Mis12, Mis13]. Static [Ozb17, Sch95]. Statistical [Per03, McC00]. Statistics [ABR99, CB95, Kob94b, McC02, Bag05, DWY98, PL92, Pro96, RS02, Rus04, Rus09]. status [Abb90]. Stauffer [BHJ^{+}00]. steam [SKS95]. steel [Bha96]. Stellated [Mae94b]. step [WGDZ04]. Stephan [Ano94c, Wee96]. Stephen [Web92, Wes92, SV93]. Steps [Bur94a, Bur94b]. Stereograms [BN91]. Steven [Mad91, Ski90]. Stewart [Bar98b, Hol98, Hol03a]. stiffeners [TD95]. stiffness [Kni95, SNLR92]. still [Old92]. Stochastic [LP13, Raa94, TSC93, FRW17, Ken93, Lew00]. stocks [Sto03]. story [Ken01]. strain [SNLR92]. Strained [Bah92]. strategies [Kul91]. strategy [Arc20, IIS18]. stratified [Ma95]. Stress [TD95, HFO94, Ioa92b]. string [GHIL09, Ker96]. STRINGVACUA [GHIL09]. strip [Kei96, TD95]. strong [CKS00]. Structure [EGM94, Gug95, RRIR^{+}92, Aza13, LM91, SRR94, Str92a]. Structures [EBM94, Bha06, GDD^{+}96, Har94, Nie03]. Structuring [AMR91]. Stuck [Ano98b, Röh99, Sha96]. Student [Ree00, WB94, Doe95, Gli95, TT99, TT09]. Students [WK17, WKB20, Has03, Ish16, Win91]. Study [KM95b, ST90, WW93a, BBD93, Cha91, Cro91, Hug93, Par93, Pat93, Sch92, TSC93, WW93b]. Studying [LJW91, ZL95, BYA94, GHIL09, Mat91f]. Subclass [Dub90]. subexpressions [Fat15]. subjected [HFO94]. Subscription [Ano95f]. suffers [Old92]. Suitable [Com96]. sum [WL11]. summer [SW93]. Summing [Mat93]. Sums [Hay92b, Van93, Bro09, Kra93, Kra95]. Sun [BHJ^{+}00, Wol93i, ABD94, DPU94a]. SUN-M38 [ABD94]. SUN-S115 [DPU94a]. sup [Maj96, WW93b]. Super [Ngu93]. supercritical [MGS96]. superfices [CFGxx]. superfield [Fer07]. supersonic [Erl90]. supersymmetries [KS93]. supersymmetry [Luc95]. supplement [Ano90b]. Support [Chi92, Kor95, Gre05, KAG96, KM13, Tre90, Ish16, Lip12, Rav07, Rue07]. supports [Cra91a]. Surface [JM93, MM93, Ano96b, SC10, TM14]. Surfaces [BCDH94, FKM95, Gra97a, TF94, von93, Ano98d, CH94, Gra93, Gra98a, GAS06]. surprise [Tam95a]. surprising [Mae92]. surprisingly [Ano95e]. Survey [Wit90]. Survival [Sza15, Rad93d]. SUSY [MZ14]. SusyMath [Fer07]. Swing [Mur93]. switch [BYA94]. Switzerland [GT94, Lak96]. Symbol [CDS^{+}92, CN94]. Symbol-Manipulating [CDS^{+}92]. Symbolic [ACM94, Ano94h, Ano96g, BP94, BG92, Ber96, Bro93, BKR^{+}91, Cal95a, Cal95b, Che91, Cra91a, Erl90, EMM90a, EMM90b, EMB92, EM92, FH96, Fat96, Fos93, Fow96, Gut04, JOS93, Jef08, Kaw93, Ken01, KS97, Lak96, Lev95, Lic11, LS96, MCG13, Mon93, Mun96, NEH90, ON92, PGB94, RMC96, RF95, SE11, Shi97, Sim92d, Sof93a, Str93a, TFS94, Tra00, Var94, Var95, WN90, Way89, WM93, XW95, XX95, Yu94, Zyp93, vdHvH12, AHH93, BBK14, Bar93, BB95, BSB94, CJL94, CN92, Chr93, CP93a, Fat15, GGV94, GSR90, Gue07, Gui06, Har94, Har14, HFO94, HS90, HY92, KA99, KB93, KYB94, LP05, Luc95, MdR13, Mat89b, Mat91g, MW95, Mio90, Mio93, MT94b, MT94c, NA90, PSW23, RCM03, Rid90a, Rid90b, RM96, Sim92a, TS00, Wat90b, Way90]. symbolic [YK95, Zit11, Chu14]. Symbolic-Numeric [Sof93a]. SymbolicC [SH10]. Symbolics [Chr94, Tro06b]. SYMMAN [Vor97]. Symmetric [Wim12]. symmetrically [RP92]. Symmetry [Bau97, Bau98, Bau99, Bau00, McC09, Vor97]. Symposium [ACM94, Aga90, Ano95d, Ano95j, Ano96d, Bro93, Gut04, IEE90c, IEE90e, IEE95a, JOS93, Jef08, KM95a, Lak96, Lev95, MGC94, Mio90, Mio93, MCG13, PL92, SE11, Tra00, WN90, vdHvH12, IEE90a, KMH97, RCM03]. Syntactic [Sut92]. syntax [Cal04]. Synthesis [Kan93, Ano96b]. System [Ano97g, Buc93, CP90, Cro93, Gug95, HM92, IMAS95, Kor11, Kro98, Kro89, KM95b, KWW92, LM90c, LM93, MGC94, OM92, PC94, Str92a, Web92, Wes92, Wol88a, Wol91a, Wolxxa, Wolxxb, BBD93, BGZ96, BGH96, Che94, Chr93, Cro91, Dev92a, Fat96, FP95, FM93, Grä96, HFO94, Hay93, KSA92, LM90b, Mat89c, Nam94, Ohm94, OFM94, Pat93, Phi91, PF95, PWW90, Tro97, WW95a, WW88a, Wol91e, Wol91b, Wol91c, Wol91f, Wol93a, Wol93c, Wol93f, Wol93d, Wol93e, Wol93g, Wol94b, Kro98]. systematic [Fat90c]. systematisches [Ste93]. système [Wol95a]. Systemen [Ben98a]. Systems [Ano92c, Ano96a, Bag95, BG92, Chi92, DPV14, EMB92, EM92, Fre94a, Gay92a, GWWJS95, HWH91, Hol94, IEE92a, IEE95a, Kor95, Mat89a, Ngu93, PMDA96, RLRML04, Str93a, AHH93, ABBD23, BPR99, BPR00, Ben98a, EKWM93, Fat90a, Fat90c, Fat15, FM93, GKW03, GI98, HBN95, HSW94, HSW98, HRV92, HD91, Joh95, JKP96, KZK10, Kou09, KM02, LW99, Lyn07, Maj96, Mat91c, MC93, Mio90, Mio93, PMD92, Rom09, Sah96, Sch92, SBP96, Sob95, TM14, TSC93, Wes99, Wol88b, WW88b, WW89a, Wol97, XW95, XX95, YZPD92, MT96, TMC96]. systolic [Bar90].

T [Pie96, Ree98]. Tables [Per03, KU94]. Takagi [Tho90]. take [Cof92a]. Tam [Ben98b]. Tandy [Ano92d]. Tangent [HM92]. Tangrams [Slo93]. Tanury [Har15]. Tape [CW05d]. tapered [Ano96f, Auc96]. TARCER [MS98]. target [BM94]. Tarski [Wag93]. Tauschung [Ano95g]. Taylor [ABMBR15, Mat89a]. Taylor1.0 [PTL04]. Teach [Iss96, Adi89, Mat89c, Mat91g, PF95, Zyp93, GM93]. Teacher [Ano94i, RPWW96]. teachers [Gli94, Gli95, SW93]. Teaching [BL93, Car92, FP95, LB91, Mat89a, McA08, MG94, Noy95, Ria91, BEL^{+}95a, GH93, Mun96]. Technical [Bra96, Sim91]. technique [CN94, NA92, SZ96]. Techniques [Ano94g, GWWJS95, WJ94, Adi89, Apt91, Pie96, RM96, SLR^{+}95, TS00]. technologies [IEE90a, Kro98]. Technology [Ano94i, DPU94a, Gal91, Kro98, Lum93, Lum94, RPWW96, Adi89, SH98]. Telescoping [LRW95]. Tempra [Ano92e]. Tensor [BG01, CP90, PC94, CJL94, Sol96, PC94]. tensorial [HHL06]. Tensors [Agu90, Dep17]. Tenth [Ano95j]. Teoria [Ano94j]. Terminator [Sch92]. terms [Ber24, Hid21]. test [KA93, KS93]. Testing [FMRP16, AGW95]. tetrahedron [SNLR92]. tetrameric [Bit93]. TEX/Mathematica [Dil91a, Dil91b]. Texas [Aga90, IEE90e, NEH90]. Text [App94, Bar10, Was05]. textbook [Ish16]. Their [RLRML04, ABFP24, Ben98a, Bd93a, Gro01, NEH90, Win91]. Them [DPV14, Win91]. Theodore [Ano00b, SMC98, Sol92]. Theorem [CZ92]. theoretic [RB93]. Theoretical [Bau96, CTMC97, WSW94, Ano97e, Bau05, EE96]. Theoretischen [Bau93, Kus97]. Theorie [Gra94a]. theories [LM91, RF95]. Theorist [Wes96b]. Theory [Ano94a, Ano94j, ABR94, Coh92, DHE92, EMT93, FKS20, Mae00, PMDA96, SBA94, Ski90, Skixx, Sof96, TW06, AU21, Don01, FK15, FP95, Fon21, Har19, HFO94, HSW98, JY93, JM08, Kor93, Mat90b, McC09, Naz12, PS03, PMD92, PF95, Pre18, Rom09, RF96, SC10, SÜ18, ZL95, ter03, BK13]. Thermal [Kro98, EN96, Yar95]. Thermath [MAC08]. Thermodynamics [BL93, Alb06, GH93, LB91]. they’re [Ive92]. thin [AdSR96, Skr93]. thin-walled [AdSR96]. Thinking [DF91]. Third [IBM^{+}95]. Thomas [Ano95b, Har15, ABF14, Wee96]. Thompson [Ano97b, Ano98a]. thousand [Wol99b]. Three [DPU94f, Tro95a, BP95, BKM14, BK16, Maj96, Tho90, Tza09]. three- [Maj96]. three-beam [Tho90]. three-parameter [Tza09]. ties [Tod92]. tight [Jac19]. tight-binding [Jac19]. Time [IEE95a, LZ96, Sti93, Wan96, AF92, Cha91, Hen96, YZPD92, UK11]. Time-periodic [LZ96]. time-varying [YZPD92]. Timeline [BHJ^{+}00]. Timoshenko [Ano96f]. TiO [KLO96]. tip [Auc96]. Tobochnik [MT96]. Tokyo [WN90]. Tome [Ziz97b]. Too [Com96]. Tool [CS96b, Dub90, GVV91, Kro98, Ano95i, De 94, EMW91, Kau92, Kau94, Mat91f, MW95, MTWW94, Wee96]. toolbox [Ano90a, RL19]. Toolkit [Sim93, GD98]. Tools [BHJ^{+}00, CH95, Fos93, Fos11, Kro98, PGB94, Ano95c, BEL^{+}95a, BYA94, Iss96, Mun96]. Topics [ABD94, Cra96, Bha06, SW93]. topological [GMP09]. topologies [Maj96]. topos [Nie03]. Toronto [IEE91b]. Torrence [Ree00, Ree00]. torsion [Ioa92a]. total [Sah96]. Tour [SR91, SH98]. Towers [IEE91b]. Toys [Fos11]. TPC [Ziz97b]. trace [Wes93]. TRACER [JL93]. tracing [CMS17]. track [BS94]. Tracking [EN96]. Trader [Mic92b]. trading [Sto03]. Traditional [App94, Par93]. traffic [Sch92]. Train [Ano92d]. Trajectories [Gil95, Che96]. Transfer [WG94, Che91, WG92]. transfer-matrix [WG92]. Transform [CSA94, Wat94]. Transformation [HSH07]. Transformations [GGK96, ABFP24, Ohm94]. Transforms [EMM90a, EMM90b, Gra04, Mat91g]. Transient [CP93b, HFO94, SR90]. transition [Erl90]. translation [SB92]. translator [Fat90b]. transport [LZ96]. Transportation [Sch92]. Transputer [BCDH94]. transverse [CP93b]. trapped [JY93]. Travelling [ZM93]. trees [MT94a]. Trends [JOS93, Fat90a]. triangle [Kni95]. triatomic [Xav22]. Tricoloring [MS08]. tries [Old92]. Trio [DPV14]. triple [Bzo21, MC96]. triple- [Bzo21]. TripleK [Bzo21]. tritangential [BP95]. truly [Mis12, Mis13]. truncation [AGW95]. trusses [Ozb17]. Trust [DPV14]. Try [BPU90c]. TSI [Ziz97b]. Tsunami [Bos96]. TTC [CJL94]. tube [SKS95]. Tucson [MGC94]. Tuning [Cal95a, Cal95b]. Tutor [BD91, BD92, BDxx]. Tutorial [EL90, EL91, Mae92, WW93c, Adi89, Ant98, McC09]. Tutorials [AMR91, BPU90b, Pid96b, Pid96a]. tutoring [PWW90]. Twelfth [Ano94f, Ham94, WGDZ04]. twelfth-order [WGDZ04]. twenty [VM90]. twenty-first [VM90]. Two [BH08c, DPU94f, Nog91, SR90, Slo93, ABFP24, BK15, Can00, Can03, CN92, Dev92a, DK93, Jon93, LM91, MS98, Pet89, Sam94, SC09, WGDZ04, ZM93]. two-center [Jon93]. two-dimensional [CN92, LM91, ZM93]. two-level [Sam94]. two-loop [MS98]. two-person [Can00, Can03]. Two-Sector [Nog91]. Two-Sided [BH08c]. two-step [WGDZ04]. Two-Wave [SR90]. TX [Ano95d, IEE90a]. type [BKM14, BK15, CKP13, Jon93, PS06]. types [IA93]. Typesetting [Soi95].

U [SKS95]. U-tube [SKS95]. U.K

[Ano97d, Ano00b]. UK [IEE94a]. Ukraine [Bro93]. unbounded [Sch90]. Uncertainty [Kor94, DWY98]. undamped [WW06]. Undergraduate [BSHS93, GH93, Har15, LB91]. undergraduates [Ish16]. understand [Ive92]. Unification [UK11]. uniform [Neu11]. Unifying [ES10, RSL^{+}26]. unit [Cav93]. United [ACM94]. Univariate [ST11, ST12]. Universities [Str94]. University [Ano95h, Ano97d, Ano00b, Gut04, HRV92, IEE94a, IEE90e, Lum93, Lum94, PL92, Rue07, Tra00, VM90]. Unix [Wol93f]. Unlimited [Ano92d]. Update [Gal91, Fos91a, TJD11, Tab16]. Updated [Fos91c, HLO^{+}09]. Upgrading [Wit91b]. upon [MAL92]. urns [Rad95]. USA [Pie96, Ano02, HRV92, IEE90b, IEE91c, IEE95c, IEE95e, Kap92, LT94, MCG13, SE11]. Usage [Kro98]. Use [Bar91, BG92, BHJ^{+}00, GAM94b, Hug93, Kaw93, Mac92, Shu94, Skr93, Ano95e, Ben98a, Iss96, Pat93, Swa91, Tza09, Wol99b]. Use/Moss [BHJ^{+}00]. User [LS95, Wol93e, Wol93g, WW95a, WW89a, WW89b, Wol93f]. users [BS94]. uses [Sch90, Sch92]. Using [AMR91, ACCM93b, AM94, Bee17, Ben93, BD91, BD92, BDxx, CB95, Cro93, DPV14, GGV94, GMM95, GM93, Gil95, Gla90, GK92, Gui06, Gup91, HSH07, Höf92, KM95b, LM90c, LU95, MAC08, MB94, Mat89c, Mat91g, May94a, May94b, Mik94b, MR94, OM92, Oln98, Per03, Phi90, PSW93, PF95, PSMM95, Pro96, RB95, RLRML04, Sav95, Slo93, Sof93a, Sti95, SC09, SGMM95, Str92b, Str94, Tre90, Tre91, Wat94, WK17, WKB20, XKM93, Abb92, Ada96, Adi89, AHR94, AJ95, ABBD23, ACCM93a, Ano94h, Ano98c, Apt91, BM94, BB95, Bha96, BSB94, BK96, Bro94, Che91, CKS90, CN94, CMS17, Dub03, EMM90b, Fol02, FM93, GV96, Gas98, GN96, Gra04, Has03, HKM93, IA93, JM93, KW92, KA93, KA99, KR98, KVF99, Kni95, Kor93, Kul91, LP05]. using [LRW95, LTE01, Lyn07, Mie95, Mik94a, Mis12, Möl07, MT94b, MT94c, Nai22, O’S96, OFM94, Pid96a, Pid96b, RLM06, Rom09, SZ96, SW93, Sch95, Sen95, SH98, Sob95, SDB93b, SDB93a, Ste94, SH10, Sto03, Str93b, Str93c, Sza00, TFS94, Tho90, TSC93, WG92, Xav22, Zac98, ZL95, Zho97]. Uterine [LBSA96]. utilizing [Ohm94].

V [Bra96, Bar92, Wes95]. v.1 [Vin03]. v.1.2 [Cau90]. v1.0 [Ozb17]. v3.8 [Ano97g]. VA [IEE91c]. vacuum [GHIL09]. valid [Jef93]. valuation [Lew00]. Value [EP96b, Asm05, KPS03, PH95, ZL95]. Valued [Asl96, HM06]. Values [BS11a, BS11b, Jon93]. Vancouver [OT94]. variable [LM90a, Bar98b]. variables [ABFP24, BKM14, BK15, BK16, DK93, vdG94]. Varian [Ano97c, Dar94, Mor98b, Rao98, Sch98b]. Variant [FEV93]. variational [FP90]. Varieties [ZD02, CLO92, CLO07]. various [BBP97]. varying [TIM96, YZPD92]. Vector [DPU94f, Hug93, Mie95, Nai22, SBQ14]. vectorial [BLM97]. Vectors [Ede91]. Velocity [XKM93, HKM93]. ventricle [GDD^{+}96]. Venture [Kor94]. Ventured [Kor94]. verification [AHH93]. verifications [MdR13]. Verlag [Ano94c, Ano96c, Ano97a, Ano97c, Ano97e, Ano97f, Ano98c, Ano99a, Ano00a, McC02, MT96]. Versatile [Chu14]. Version [Ano88b, Ano94e, Ano97g, BC92, BC94, BCxx, Cof89, GG92, GGxx, LaB94, LC90, MO94, Ree98, Sei90, Sim90, SMC98, Wol93b, Wol93h, Wol95b, WW95b, WD97, AR93, AC13, Ano95e, Ano00b, AC18, Cof92b, Dev92a, Gal98, GG97, GG00, JL93, PS95, SHH93, Tam95a, WW95a, Wit91b, WW88c, WW88d, WW88e, Wol93f, Wol93e, Wol93g, Wol93i, Ziz97b]. versions [BS94]. vertex [Eks11]. Very [KMRC90, Hug93]. Vessel [XKM93, HKM93]. VEST [SBQ14]. via [Ada95, Cal95a, Cav93, LM23, Nie03, Smi93]. Viable [ES10, Ste08, Ste13]. vibrating [Ker96]. Vibration [KAG96, NR96]. Vibrational [DHE92]. Vibrational-Rotational [DHE92]. vibrations [Auc96, DF96, MN91, RP92]. Video [Yag92]. video-presentation [Yag92]. View [LS95]. Village [IEE90d]. vision [KM95a, MAL92, Rog92, Sch92, ter03]. Visual [CKS90, HPS05, Dvo90, Rog92]. VisualDSolve [SW97]. Visualisierung [GAM94a]. Visualization [FKM95, GVV91, JS93b, Wei92, WJ02, WY93, Doe95, HPR96, JS93a, JS93c, MZ92, Phi91, RL19, Wag94, WA96, JS93b]. Visualizations [GAM94b, Ive92]. Visualizing [May94b, Sta90, Tho92, WAW96, SW97]. volatility [Lew00]. volume [Joh95]. vs [Par93, Tha89].

W [Ree98, Sol92]. W. [Oln98]. Wadati [EM93]. Wafer [LT94, Sam94]. Wagon [Ano97a, Ano00a, O’R92, Sol92, Tay92]. walks [FRW17]. walled [AdSR96]. War [BHJ^{+}00]. warning [Wit91a]. Warp [Tam95a, Tam95b]. Warwick [IEE94a]. Washington [Ano90c, Ano94f]. Water [CS96b]. Waterloo [Tho94]. Wave [SR90, Tho92, HFO94, TIM96, ZM93]. wavefunctions [AC13]. Waveguide [McP92]. Wavelet [Ano96i]. Wavelets [Sak95, RA95]. Waves [GMM95, SBA94]. Way [Shu94]. Ways [Bar91]. weak [Pre18, Wan96]. Web [Bra96, Ano99b, Bra96, Shi02, WJ02]. wedges [Pet08]. Wellin [Ano96c, MT96, TMC96]. Werkzeug [Kau92, Kau94, Ano94c]. Wescon [IEE94b]. Wescon/94 [IEE94b]. Wesley [Ano95b]. whip [CP93b]. White [Kro98]. Whole [Str93a]. Wickham [Pie96]. Wickham-Jones [Pie96]. widths [AGW95]. WiLE [Pre18]. Wiley [Ano97b, Ano98a, Ano98b]. will [BS94]. William [Ano97b, Ano98a, Lum94]. Williamsburg [IEE91c]. Wills [KL02]. Wilson [Iss96, Pre18]. Wind [SGMM95]. Wind-Driven [SGMM95]. Window [Ano92d]. Windows [Ano92b, Ano92d, Nor94, Sim92c, Lin91a, Ano91b, Ano92f, Ano93c, Ano94e, Bar92, Cof92b, Eva93b, Kro98, LaB94, Lin91b, MO94, Old92, Sim92b, Smi91, Rev93, Tam95a, WW95a, Wol94b]. Winter [AKLG95, NEH90]. Wirtschaftswissenschaftler [San99]. Wisconsin [IEE95c]. Wiser [BHJ^{+}00]. wissenschaftliche [BHJ^{+}00]. without [CW05d]. Wizardry [Shi02]. Wolfram [Ano97d, Pie96, Ano91b, Bra96, HMM15, HMM16, Kro89, Mad91, Old92, Sim89, Sim92c, SV93, Var91c, Web92, Wes92, Wol15]. wonderful [Sim89]. work [KS97]. Working [Bla90]. Workshop [PMD92, PMDA96, SBA94, Car96, Str92a]. workstation [IOA^{+}95]. workstations [Wol93i]. World [Ano91c, Mat95a, OT94, Sim90]. Worldwide [Ano96j]. wormhole [Rom94]. worthy [Ano90a]. Wow [Win91]. Writing [Mae93a]. written [Gro01, ter03]. Wstep [Jan94].

X [Rue07, Eva93b, Oln98, Pat15, Pat17, Tay99, Wol93g]. X-Windows [Eva93b]. xiii [McC02]. Xorbit [Eva93b]. XSC [PK10]. xSLHA [Sta19].

Y. [Tay99]. Year [FW91]. York

[Ano96c, Ano97a, Ano97b, Ano97c, Ano97f, Ano98a, Ano98b, Ano98c, Ano99a, Ano00a, McC02, MT96, Oln98].

Zachary [Ano98c]. Zeilberger [PS95]. Zeit [BHJ^{+}00]. Zeit/Dyson [BHJ^{+}00]. zen [Way90]. zero [SZ96]. zeros [KW92, Kei96]. Zincblende [Bah92]. ZnS [Yar95]. Zurich [Lak96]. Zweipersonenspiele [BHJ^{+}00, Can00]. Zweipersonenspiele/Adobe [BHJ^{+}00].

References

Abell:1992:GMC

[AB92a] M. Abell and J. Braselton. Giving meaning to the circular membrane problem with Mathematica. Mathematica in Education, 1(3):11–16, Spring 1992. ISSN 1065-2965.

Abell:1992:ME

[AB92b] Martha L. Abell and James P.

Braselton. Mathematica by Example. Academic Press, New York, NY, USA, 1992. ISBN 0-12-041540-2. xiii + 654 pp. LCCN QA76.95 .A214 1994.

Abell:1992:MH

[AB92c] Martha L. Abell and James P.

Braselton. The Mathematica Handbook. Academic Press, New York, NY, USA, 1992. ISBN 0-12-041535-6, 0-12-041536-4. xvi + 789 pp. LCCN QA76.95 .A22 1992. US$39.95.

Abell:1993:DEM

[AB93] Martha L. Abell and James P.

Braselton. Differential Equations with Mathematica. Academic Press, New York, NY, USA, 1993. ISBN 0-12-041538-0, 0-12-041539-9. viii + 631 pp. LCCN QA371.5.D37 A24 1993. US$44.95.

Abell:1994:ME

[AB94] Martha L. Abell and James P.

Braselton. Mathematica by Example. Academic Press, New York, NY, USA, revised edition, 1994. ISBN 0-12-041530-5. xii + 523 pp. LCCN QA76.95 .A214 1994. US$39.95. See [ABxx].

Abell:1997:DEM

[AB97a] Martha L. Abell and James P.

Braselton. Differential equations with Mathematica. Academic Press, New York, NY, USA, second edition, 1997. ISBN 0-12-041550-X. xii + 807 pp. LCCN QA371.5.D37A24 1997. US$44.95.

Abell:1997:ME

[AB97b] Martha L. Abell and James P. Braselton. Mathematica by example. Academic Press, New York, NY, USA, second edition, 1997. ISBN 0-12-041552-6. xii + 603 pp. LCCN QA76.95.A214 1997. US$39.95.

Abell:19xx:ME

[ABxx] Martha L. Abell and James P. Braselton. Mathematica by Example. Academic Press and Fuji Software, New York, NY, USA, 19xx. ISBN 4-938455-35-8. ???? pp. LCCN ???? Japanese translation of [AB94].

Abell:2004:DEM

[AB04a] Martha L. Abell and James P. Braselton. Differential equations with Mathematica. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2004. ISBN 0-12-041562-3. xvi + 876 pp. LCCN QA371.5.D37 A24 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els041/2003061664.
                                        html;   http://www.loc.gov/
                                        catdir/toc/els041/2003061664.
                                        html.

Abell:2004:ME

[AB04b] Martha L. Abell and James P. Braselton. Mathematica by example. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2004. ISBN 0-12-041563-1. xii + 571 pp. LCCN QA76.95 .A214 2004. Includes CD-ROM.

Abell:2008:ME

[AB08] Martha L. Abell and James P.

Braselton. Mathematica by example. Academic Press, New York, NY, USA, fourth edition, 2008. ISBN 0-12-374318-4 (paperback). 564 pp. LCCN QA76.95 .A214 2009.

Abbott:1990:CSF

[Abb90] Paul C. Abbott. Current status and future directions of Mathematica. In Noor et al. [NEH90], pages 29–41. CODEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Abbott:1992:PSU

[Abb92] P. C. Abbott. Problem solving using Mathematica. IFIP Transactions. A. Computer Science and Technology, A-2:33–43, 45, 1992. CODEN ITATEC. ISSN 0926-5473.

Ananthanarayan:2023:FMP

[ABBD23] B. Ananthanarayan, Sumit Banik, Souvik Bera, and Sudeepan Datta. FeynGKZ:

a Mathematica package for

solving Feynman integrals using GKZ hypergeometric systems. Computer Physics Communications, 287(??):Article 108699, June 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465523000449.

Abell:1994:STA

[ABD94] M. L. Abell, J. P. Braselton, and J. W. Davenport. (SUN-M38) topics in applied mathematics with Mathematica. In Lum [Lum94], pages 453–489. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Aberer:1990:NFF

[Abe90] K. Aberer. Normal forms in function fields. In Watanabe and Nagata [WN90], pages 1–7. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Abe:1994:IMPa

[Abe94a] Yutaka Abe. An Introduction to Mathematical Physics with Mathematica, Part I. Kodansha, ??, Japan, 1994. ISBN 4-06-153215-4. 196 pp. LCCN ???? In Japanese.

Abe:1994:IMPb

[Abe94b] Yutaka Abe. An Introduction to Mathematical Physics with Mathematica, Part II. Kodansha, ??, Japan, 1994. ISBN 4-06-153216-2. 174 pp. LCCN ???? In Japanese.

Amore:2014:ACS

[ABF14] Paolo Amore, John P. Boyd, and Francisco M. Fernández. Accurate calculation of the solutions to the Thomas–Fermi equations. Applied Mathematics and Computation, 232(??): 929–943, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0096300314001829.

Ananthanarayan:2024:OWR

[ABFP24] B. Ananthanarayan, Souvik Bera, S. Friot, and Tanay Pathak. Olsson.wl and ROC2.wl: Mathematica packages for transformations of multivariable hypergeometric functions and regions of convergence for their series representations in the two variables case. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465524000857.

Abell:1996:MCR

[ABGA96] Martha L. Abell, James P.

Braselton, John W. (John Walker) Gray, and Arnold O. Allen. Mathematica CD-ROM library. Academic Press, New York, NY, USA, 1996. ISBN ???? ???? pp. LCCN ???? One computer optical disc.

Avitzur:1995:HIP

[ABK95] Ron Avitzur, Olaf Bachmann, and Norbert Kajler. From honest to intelligent plotting. In Levelt [Lev95], pages 32–41. ISBN 0-89791-699-9. LCCN QA 76.95 I59

1995.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/

                                        p32-avitzur/. ACM order

number: 505950.

Abad:2015:AIN

[ABMBR15] A. Abad, R. Barrio, M. Marco-Buzunariz, and M. Rodríguez. Automatic implementation of the numerical Taylor series method: a Mathematica and Sage approach. Applied Mathematics and Computation, 268 (??):227–245, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300315008231.

Axler:1994:HFT

[ABR94] Sheldon Axler, Paul Bourdon, and Wade Ramey. Harmonic Function Theory, volume 137 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-97875-5. xii + 231 pp. LCCN QA405.A95 1992.

Abell:1999:SM

[ABR99] Martha L. Abell, James P. Braselton, and John A. Rafter. Statistics with Mathematica. Academic Press, New York, NY, USA, 1999. ISBN 0-12-041554-2 (book), 0-12-041555-0 (CD-ROM). xiv + 632 pp. LCCN QA276.4 .A24 1999. CD-ROM contains all Mathematica inputs from the text and also numerous procedures to extend Mathematica’s built-in, statistical capabilities.

Alex:1991:MFA

[AC91] M. Alex and J. Crain. Meanfield analysis on the Macintosh (magnetic ordering). Computers in Physics, 5(2):193–197, March/April 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Angeli:2013:AEM

[AC13] C. Angeli and R. Cimiraglia.

Automated evaluation of matrix elements between contracted wavefunctions: a Mathematica version of the FRODO program. Computer Physics Communications, 184(2):443–444, February 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465512003323.

Ayala:2015:AMP

[AC15] César Ayala and Gorazd Cvetic.

anQCD : a Mathematica package for calculations in general analytic QCD models. Computer Physics Communications, 190(??):182–199, May 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465515000120.

Ayala:2018:AMP

[AC18] César Ayala and Gorazd Cvetic.

anQCD: a Mathematica package for calculations in general analytic QCD models (new version announcement). Computer Physics Communications, 222 (??):413, January 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465517302928.

Andrew:1993:CPUa

[ACCM93a] Alfred D. Andrew, George L. Cain, Sheryl Crum, and Thomas Morley. Calculus projects using Mathematica. Technical report, School of Mathematics, Georgia Institute of Technology, Atlanta, GA, USA, 1993. 193 pp.

Andrew:1993:CPUb

[ACCM93b] Alfred D. Andrew, George L. Cain, Sheryl Crum, and Thomas Morley. Calculus Projects Using Mathematica. McGraw-Hill, New York, NY, USA, revised edition, 1993. ISBN 0-07-001867-7. 193 pp. LCCN ????

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC ’94: Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20–22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Adamchik:1992:LCF

[Ada92] V. Adamchik. Limits of continued fractions and nested radicals. Mathematica Journal, 2 (2):54–57, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Adamchik:1995:EIB

[Ada95] V. Adamchik. The evaluation of integrals of Bessel functions via G-function identities. Journal of Computational and Applied Mathematics, 64(3):283– 290, December 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Adami:1996:GNM

[Ada96] Chris Adami. Gaylord and Nishidate’s modelling nature with cellular automata using Mathematica. SIGSAM Bulletin, 30(3):24–25, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Angulo:1994:CDR

[ADG94] F. Angulo, A. Devia, and G. A.

Gonzales. Calculations of the dispersion relation for polynomials involving the dispersion function Z(ζ). In Sakanaka et al. [SBA94], pages 1–4 vol.2. 3 vol.

Adickes:1989:ITE

[Adi89] Martin D. Adickes. Integrating technology and education: using Mathematica as a tutorial to teach selected EDA techniques. Thesis (Engineering report, M.S.E.), Arizona State University, Tempe, AZ, USA, 1989. ix + 198 pp.

Alfano:1996:AAM

[AdSR96] G. Alfano, F. M. de Sciarra, and L. Rosati. Automatic analysis of multicell thin-walled sections. Computers and structures, 59(4):641–655, May 1996. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Andrew:1992:STG

[AF92] K. Andrew and C. G. Fleming. Space-time geometries characterized by solutions to the geodesic equations. Computers in Physics, 6(5):498–505, September/October 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

AguilarVilla:1997:EDP

[AF97a] Gloria Aguilar Villa and Ana Fernandez-Ferreiros Erviti. Ecuaciones diferenciales: practicas con Mathematica, volume 51 of Coleccion Textos Docentes. Prensas Universitarias de Zaragoza, Zaragoza, Spain, 1997. ISBN 84-7733-470-6. 305 pp. LCCN ????

AsensioSevilla:1997:CNP

[AF97b] Maria Isabel Asensio Sevilla and Luis Ferragut Canals. Calculo numerico: planteamiento y resolucion de problemas con “Mathematica”. Plaza Universitaria, Salamanca, Spain, 1997. ISBN 84-89109-12-5. v + 117 pp. LCCN ????

Agarwal:1990:IMI

[Aga90] Krishna K. Agarwal, editor. 1990 IEEE MTT-S International Microwave Symposium digest: May 8–10, 1990, Dallas Convention Center, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. LCCN TK 7876 I11 1990. Three volumes.

Aguirregabiria:1990:CRE

[Agu90] J. Aguirregabiria. Computing the Ricci and Einstein tensors. Mathematica Journal, 1(2):51– 54, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Apte:1995:ETI

[AGW95] P. Apte, M. C. Gower, and B. Ward. Experimental testing of integral truncation algorithms for the calculation of beam widths by proposed ISO standard methods. In Anonymous [Ano95a], pages 349–358. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Arnold:1990:PIC

[AH90] Stephen Arnold and Graham Hair, editors. Proceedings: International Computer Music Conference (1990): Glasgow, Scotland. Computer Music Association, San Francisco, CA, USA, 1990. LCCN ML1381 .I6 1990.

Allen:1991:SQM

[AH91] A. Allen and G. Hynes. Solving a queueing model with Mathematica. Mathematica Journal, 1(3):108–112, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Alur:1993:ASV

[AHH93] Rajeev Alur, Thomas A. Henzinger, and Pei-Hsin Ho. Automatic symbolic verification of embedded systems. Proceedings — Real-Time Systems Symposium, pages 2–11, 1993. CO-DEN PRSYEA. ISBN 0-8186-4480-X. IEEE catalog number 93CH3376-1. Aguirregabiria:1994:WCE

[AHR94] J. M. Aguirregabiria, A. Hernandez, and M. Rivas. Are we careful enough when using computer algebra? Computers in Physics, 8(1):56–61, January/February 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Albert:1995:PAF

[AJ95] Bernhard Albert and Anura P. Jayasumana. Performance analysis of FDDI LANs using numerical methods. In IEEE [IEE95d], pages 379–386. CO-DEN CLCPDN. ISBN 0-8186-7162-9. ISSN 0742-1303. LCCN TK5105.7 .C66 1995. IEEE Catalog No. 95TB100005.

Alexopoulos:1995:WSC

[AKLG95] C. Alexopoulos, K. Kang, W. R. Lilegdon, and D. Goldsman, editors. 1995 Winter Simulation Conference: proceedings, December 3–6, 1995. San Diego, CA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN WSCPDK. ISBN 0-7803-3018-8, 0-7803-3017-X. ISSN 0275-0708, 0743-1902. LCCN QA76.9.C65 W56 1995. IEEE Catalog No. 95CB35865.

Alberty:2006:BTA

[Alb06] Robert A. Alberty. Biochemical thermodynamics: applications of Mathematica, volume 48 of Methods of biochemical analysis. Wiley-Interscience, New York, NY, USA, 2006. ISBN 0-471-75798-5. xii + 464 pp. LCCN QA76.95 .A43 2006. Includes CD-ROM.

Alcaide:1997:ANM

[Alc97] Adrian Alcaide. P -adic numbers and Mathematica. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1997. 11 pp.

Allen:1992:ICP

[All92] Arnold O. Allen. Introduction to computer performance analysis with Mathematica. CMG Proceedings, pages 1189–1191, 1992. CODEN CMPREY.

Allen:1994:ICP

[All94] Arnold O. Allen. Introduction to Computer Performance Analysis with Mathematica. Computer science and scientific computing. Academic Press, New York, NY, USA, 1994. ISBN 0-12-051070-7. xx + 356 pp. LCCN QA76.9.E94 A43 1994. US$49.95.

Andrew:1994:LAP

[AM94] A. D. Andrew and T. D. Morley. Linear Algebra Projects Using Mathematica. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-001868-5. 128 pp. LCCN ????

Adickes:1991:STU

[AMR91] M. Adickes, W. Moor, and R. Rucker. Structuring tutorials using Mathematica. Mathematica Journal, 1(3):86–91, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1988:MS

[Ano88a] Anonymous. Mathematica is the solution. Computers in Physics, 2(5):97–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/10.
                                        1063/1.4822783.

Anonymous:1988:MV

[Ano88b] Anonymous. Mathematica, version 1.03. Macworld, ??(??): 174, 176, December 1988. CO-DEN MACWEA. ISSN 0741-8647. Requires Mac Plus, SE, or II; 2.5MB RAM, hard disk.

Anonymous:1989:MS

[Ano89] Anonymous. Mathematica (software). Mikrocomputer Zeitschrift, 11:42–59, November 1989.

Anonymous:1990:MMT

[Ano90a] Anonymous. Maple: $695 math toolbox is worthy competitor for Mathematica. PC Magazine, pages 48–49, February 27, 1990. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1990:MJE

[Ano90b] The Mathematica journal electronic supplement, 1990. Miller Freeman Publications, San Francisco, CA, USA. Computer disks.

Anonymous:1990:NOS

[Ano90c] Anonymous, editor. Northcon, October 9–11, 1990, Seattle, Washington, conference record, volume 9 of Northcon conference record. Western Periodicals Co.,, North Hollywood, CA, 1990. LCCN TK 7801 N67 1990.

Anonymous:1991:MPO

[Ano91a] Anonymous. Mathematica 2.0 is plagued with operational problems, high memory demands. InfoWorld, 13(??):81, September 1991. CODEN INWODU. ISSN 0199-6649.

Anonymous:1991:MWWa

[Ano91b] Anonymous. Mathematica for Windows 3.0: Wolfram Research’s great addition. BYTE Magazine, 16(2):127–135, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PAC

[Ano91c] Anonymous, editor. Proceedings 1991 Annual Conference, American Society for Engineering Education. Challenges of a Changing World. ASEE, Washington, DC, USA, 1991. Two volumes.

Anonymous:1992:MSI

[Ano92a] Anonymous. Math software improves ‘drudgery to insight’ ratio. T H E journal (Technological Horizons in Education), 20:

44–46, November 1992. CODEN THEJD4. ISSN 0192-592X.

Anonymous:1992:MW

[Ano92b] Anonymous. Mathematica for windows. InfoWorld, 14(9):85– 86, March 1992. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1992:PCI

[Ano92c] Anonymous, editor. Proceedings of the 1992 Conference on Information Sciences and Systems. Dept. of Electrical Engineering, Princeton University, Princeton, NJ, USA, 1992. LCCN Q 350 C66 1992. Two volumes.

Anonymous:1992:TNN

[Ano92d] Anonymous. Tandy 4860HD Notebook, Norton, Desktop for Windows, A-Train, NEC Silentwriter Model 95, Grand Prix Unlimited, Window Phone, dBASE IV 1.5, Solitaire’s Journey, Home Alone, Mathematica, and 22 more. Compute, 14(11): 114–??, December 1992. CO-DEN COMPER. ISSN 0194-357X.

Anonymous:1992:TST

[Ano92e] Anonymous. Tempra show, tempra pro. PC Week, 9(18):37, May 1992. ISSN 0740-1604.

Anonymous:1992:WYS

[Ano92f] Anonymous. What you see is what you solve: Mathematica and MathCAD for Windows. BYTE Magazine, 17 (5):263–268, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CS

[Ano93a] Anonymous. Computers and software. Laser focus world, 29 (5):107–??, May 1993. CODEN LFWOE8. ISSN 1043-8092.

Anonymous:1993:M

[Ano93b] Anonymous. Mathematica 2.2.

Macworld, 10(9):63–??, September 1993. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1993:MWR

[Ano93c] Anonymous. Mathematica for Windows receives a revamped interface. PC Week, 10(13):43– ??, April 1993. ISSN 0740-1604.

Anonymous:1993:SR

[Ano93d] Anonymous. Software reviews.

Mathematics and computer education, 27(2):168–??, Spring 1993. CODEN MCEDDA. ISSN 0730-8639.

Anonymous:1994:ISF

[Ano94a] Anonymous, editor. 1994 International Sherwood Fusion Theory Conference. University of Texas at Austin, Austin, TX, USA, 1994.

Anonymous:1994:AMA

[Ano94b] Anonymous, editor. 78th Annual meeting: April 1994, Jonesboro, AR, volume 48. Arkansas Academy of Science, ????, 1994. ISSN 0097-4374.

Anonymous:1994:BRBd

[Ano94c]  Anonymous.  Book review:

Mathematica als Werkzeug: Stephan Kaufmann (1992):

Basel: Birkhäuser Verlag, ISBN 3-7643-2832-0, pp. 396, DM 54. Computational Statistics & Data Analysis, 17(1): 104–105, January 1994. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0167947394900353.

Anonymous:1994:BR

[Ano94d]  Anonymous.  Book reviews.

Mathematics and computer education, 28(2):213–??, Spring 1994. CODEN MCEDDA. ISSN 0730-8639.

Anonymous:1994:MVW

[Ano94e] Anonymous. Mathematica version 2.2 for Windows. Australian Personal Computer, 15(9):32– ??, September 1994. ISSN 0725-4415.

Anonymous:1994:PTN

[Ano94f] Anonymous, editor. Proceeding of the Twelfth National Conference on Artificial Intelligence: July 31–August 4, 1994, Seattle, Washington. MIT Press, Cambridge, MA, USA, 1994. CO-DEN PNAIEE. ISBN 0-262-61102-3. LCCN Q334.N36 1994. Two volumes.

Anonymous:1994:PES

[Ano94g] Anonymous, editor. Proceedings / EUFIT ’94: Second European Congress on Intelligent Techniques and Soft Computing, Aachen, Germany, September 20–23, 1994, volume 1. Verlag der Augustinus-Buchhandel, Aachen, Germany, 1994. ISBN 3-86073-286-2. LCCN ???? DM237.00. Three volumes.

Anonymous:1994:SCA

[Ano94h] Anonymous. Symbolic circuit analysis using Mathematica. International Journal of Electrical Engineering Education, 31 (4):324–333, October 1994. CO-DEN IJEEAF. ISSN 0020-7209.

Anonymous:1994:TTE

[Ano94i] Anonymous, editor. Technology and Teacher Education Annual, 1994. Assoc. Adv. Comput. Educ, Charlottesville, VA, USA, 1994.

Anonymous:1994:TSP

[Ano94j] Anonymous, editor. Teoria si practica mecanismelor Theory and practice of mechanisms: 6 Simpozion international — June 1993, Bucharest, volume 6(1) of Simpozion International Teoria si Practica Mecanismelor 1994. Asociatia, Bucharest, Romania, 1994. ISBN ???? LCCN ????

Anonymous:1995:BCD

[Ano95a]  Anonymous, editor.  Beam

Control, Diagnostics, Standards, and Propagation, volume 2375 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Anonymous:1995:BRMg

[Ano95b]  Anonymous.  Book review:

Mathematica(R) for scientists and engineers: By Thomas B. Bahder. Addison-Wesley Publishing Company, reading, MA. (1995). 846 pages. $42.95. Computers and Mathematics with Applications, 29(11):108, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122195902597.

Anonymous:1995:CMT

[Ano95c]  Anonymous.  Computerized

mathematical tools in electrical engineering courses. American Society for Engineering Education Annual Conference. Proceedings, 2:2510–2512, 1995. CODEN ACOPDW. ISSN 0190-1052.

Anonymous:1995:ESA

[Ano95d] Anonymous, editor. Education: 4th Symposium; 75th Annual meeting — January 1995, Dallas, TX, Conference on Education — American Meteorological Association. American Mathematical Society, Providence, RI, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:MPN

[Ano95e] Anonymous. Mathematica 2.2.2 PowerPC-native version is surprisingly easy to use and a speed demon at graphing highlevel math operations. MacUser, 11(8):56–??, 1995. CO-DEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:NPR

[Ano95f] Anonymous. New products:

Red Hat Linux Developers Package; CE Editor for Linux; Mathematica for Linux; Directories & References Corrections; Linux Resources; Advertisers Index; Consultants Directory; Subscription Information. Linux Journal, 19:??, November 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic).  URL http://www.
linuxjournal.com/index.html;
 http://www.linuxjournal.
com/issue19/ad19.html;   http:
//www.linuxjournal.com/lj/
Extras/consultants.dir.html;
 http://www.linuxjournal.
com/lj/index.html;   http:
//www.linuxjournal.com/lj/
ljstaff.html;   http://www.
linuxjournal.com/lj/ljsubsorder.
html.

Anonymous:1995:OTA

[Ano95g] Anonymous. Optische Tauschung?: Animationen in Mathematica programmiert. Elektronik, 44(23):164–??, 1995. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:1995:SPS

[Ano95h] Anonymous, editor. SICE ’95:

proceedings of the 34th SICE Annual Conference, International Session papers, Hokkaido University, July 26–28, 1995, volume 1. SICE, ????, 1995. ISBN 0-7803-2781-0. LCCN TK7870.K37 1995. SICE catalog number 95 PR 0001-3. IEEE catalog number 95 TH 8107.

Anonymous:1995:SNC

[Ano95i] Anonymous. SSP, a new computer/math tool for EEs. American Society for Engineering Education Annual Conference. Proceedings, 2:2523–2526, 1995. CODEN ACOPDW. ISSN 0190-1052.

Anonymous:1995:TIS

[Ano95j] Anonymous, editor. Tenth International Symposium on Gas Flow and Chemical Lasers, volume 2502 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Anonymous:1996:IIC

[Ano96a] Anonymous, editor. 3rd IMACS International Conference on Computational Physics: Nonlinear Dynamic Phenomena in Physical, Chemical and Biological Systems, volume 40(3-4) of Mathematics and Computers in Simulation. Elsevier, Amsterdam, The Netherlands, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1996:ASS

[Ano96b] Anonymous. Automated synthesis of surface generation equations with application to gearing. Computer methods in applied mechanics and engineering, 135(1-2):63–83, August 1996. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Anonymous:1996:BRId

[Ano96c]  Anonymous.  Book review:

An introduction to programming with Mathematica(R) (second edition): By Richard Gaylord, Samuel Kamin and Paul Wellin. Springer-Verlag, New York. (1996). 452 pages. DM 72.00 (diskette included). Computers and Mathematics with Applications, 31(11):135, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122196873305.

Anonymous:1996:ESC

[Ano96d]  Anonymous, editor.  Euro-

pean Symposium on Computer Aided Process Engineering -6. ESCAPE-6, volume 20 (pt.B) of Computers & Chemical Engineering. Elsevier, Amsterdam, The Netherlands, 1996. CODEN CCENDW. ISSN 0098-1354.

Anonymous:1996:FPF

[Ano96e] Anonymous. Functional presentation of Fourier series convergence. Visual Computer, 12(7): 350–359, 1996. CODEN VI-COE5. ISSN 0178-2789 (print), 1432-2315 (electronic).

Anonymous:1996:HOT

[Ano96f] Anonymous. Higher order Timoshenko quotient in the stability and dynamic analysis of smoothly tapered beams. Journal of Sound and Vibration, 196 (3):253–262, September 1996. CODEN JSVIAG. ISSN 0022-460X.

Anonymous:1996:ICS

[Ano96g] Anonymous, editor. IEE Colloquium on Symbolic Computation for Control (Digest No.1996/ 078). IEE, London, UK, 1996.

Anonymous:1996:QMA

[Ano96h] Anonymous, editor. Quantitative Methods for Applied Sciences, volume 9(3) of Computational Economics. Kluwer Academic Publishers, Dordrecht, The Netherlands, August 1996. CODEN CNOMEL. ISSN 0927-7099.

Anonymous:1996:WPM

[Ano96i] Anonymous. Wavelet plug-in for Mathematica. SunExpert Magazine, 7(9):73, September 1996. ISSN 1053-9239.

Anonymous:1996:WMC

[Ano96j] Anonymous. Worldwide Mathematica conference. ACM SIGNUM Newsletter, 31(4):36– ??, ???? 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1997:BRAf

[Ano97a] Anonymous. Book review: Animating calculus: Mathematica(R) notebooks for the laboratory: By Ed Packel and Stan Wagon. Springer-Verlag, New York. (1997). 292 pages. DM 54.00, öS 394.20, sFr 48.00. Computers and Mathematics with Applications, 33 (7):140, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122197846111.

Anonymous:1997:BRAh

[Ano97b] Anonymous. Book review: Atlas for computing mathematical functions: an illustrated guide for practitioners, with programs in C and Mathematica: William J. Thompson. John Wiley & Sons, New York. (1997). 903 pages. $99.00 (CD included). Computers and Mathematics with Applications, 33(9):145, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122197901292.

Anonymous:1997:BRCe

[Ano97c]  Anonymous.  Book review:

Computational economics and finance: Modeling and analysis with mathematica(R): Edited by Hal R. Varian. Springer-Verlag, New York. (1996). 468 pages. DM 88.00, öS 642.40, sFr 77.50 (diskette included). Computers and Mathematics with Applications, 33(5):129, March 1997.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122197829367.

Anonymous:1997:BRMh

[Ano97d]  Anonymous.  Book review:

Mathematica 3.0: Standard add-on packages: Edited by Emily Martin. Wolfram Media/ Cambridge University Press, Champaign, IL/Cambridge, U.K. (1996). 516 pages. $59.95 (hardback); $29.95 (paperback). Computers and Mathematics with Applications, 33 (5):130, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122197829392.

Anonymous:1997:BRMb

[Ano97e]  Anonymous.  Book review:

Mathematica(R) in theoretical physics: Selected examples from classical mechanics to fractals: By Gerd Baumann. Springer-Verlag, Heidelberg, Germany. (1996). 348 pages. DM 88.00, sFr 77.50 (diskette included). Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122197900092.

Anonymous:1997:BRMg

[Ano97f] Anonymous. Book review: Modeling nature: Cellular automata simulations with mathematica(R): By Richard Gaylord and Kazume Nishidate. Springer-Verlag, New York. (1996). 260 pages. DM 64.00; öS 467.20; sFr 56.50 (diskette included). Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122197829422.

Anonymous:1997:NPK

[Ano97g] Anonymous. New products:

Khoros Pro 2.2; Kai C++ Version 3.2; Open Sound System v3.8; ARKEIA; Mathematica 3.0; PerfectBACKUP+; Empress RDBMS 6.10; NetTracker 3.0; ScriptEase: Integration SDK. Linux Journal, 42:??, October 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:BRAk

[Ano98a] Anonymous. Book review: Atlas for computing mathematical functions: an illustrated guide for practitioners, with programs in Fortran 90 and Mathematica: By William J. Thompson. John Wiley & Sons, Inc., New York. (1997). $145.00 (CD-ROM included). Computers and Mathematics with Applications, 36 (8):125, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122198911671.

Anonymous:1998:BRDd

[Ano98b] Anonymous. Book review: Differential equations with Mathematica^{T M}: By Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. John Wiley & Sons, Inc., New York (1998). 240 pages. $19.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122198907192.

Anonymous:1998:BRIm

[Ano98c]  Anonymous.  Book review:

Introduction to scientific programming: Computational problem solving using Mathematica and C: By Joseph L. Zachary. Springer-Verlag, New York. (1998). $49.95 (diskette included). Computers and Mathematics with Applications, 36 (8):122, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S089812219891135X.

Anonymous:1998:BRMm

[Ano98d]  Anonymous.  Book review:

Modern differential geometry of curves and surfaces with mathematica (second edition): By Alfred Gray. CRC Press, Boca Raton, FL. (1998). 1053 pages. $84.95. Computers and Mathematics with Applications, 36 (8):121, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122198911336.

Anonymous:1998:MKB

[Ano98e] Anonymous. Mathematica — kurz und bundig. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1998. ISBN 3-7643-6008-9. 220 pp. LCCN ????

Anonymous:1999:BREh

[Ano99a] Anonymous. Book review: Exploring abstract algebra with mathematica(R): By Allen C. Hibbard and Kenneth M. Levasseur. Springer-Verlag, New York. (1999). 467 pages. $45.00, DM 84.00, öS 614.00, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 38(11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122199911938.

Anonymous:1999:WMM

[Ano99b] Anonymous. Web mechanics:

MathScript, Internet connection to Mathematica. Computing in Science and Engineering, 1(1):92–??, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://dlib.computer.org/
                                        cs/books/cs1999/pdf/c1092.
                                        pdf.

Anonymous:2000:BRMg

[Ano00a]  Anonymous.  Book review:

Mathematica(R) in action: Second Edition. By Stan Wagon. Springer-Verlag, New York. (2000). 592 pages. $49.95, DM 98, öS 716, sFr 89.50, GBP 34 (CD-ROM included). Computers and Mathematics with Applications, 40(2–3): 420, July/August 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122100902256.

Anonymous:2000:BRB

[Ano00b] Anonymous. Book review: The beginner’s guide to Mathematica(R) version 4: By Jerry Glynn and Theodore Gray. Cambridge University Press, Cambridge, U.K. (2000). 434 pages. $74.95 (hardback); $29.95 (paperback). Computers and Mathematics with Applications, 40(2–3):418, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122100902013.

Anonymous:2002:MIC

[Ano02] Anonymous, editor. MathML International Conference: Hickory Ridge Conference Center, Chicago, IL, USA, June 28–30, 2000. ????, ????, 2002. ISBN ???? LCCN ????

Anonymous:2017:RDJ

[Ano17] Anonymous. The 2017 Richard D. Jenks Memorial Award. ACM Communications in Computer Algebra, 51(4):111, December 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anthony:1998:MAC

[Ant98] Neil Richard Anthony. A Mathematica 3.0.1 alternating current circuit analysis tutorial. Thesis (M.A.), Ball State University, Muncie, IN, USA, 1998. various pp.

Albert:1992:ABC

[APL92] J. Albert, C. Page, and R. LePage. Algorithms for Bayesian computing and Mathematica. In Page and LePage [PL92], pages 286–290. ISBN 0-387-97719-8. LCCN QA276.4.C592 1991.

Appleyard:1994:MMA

[App94]  D. Appleyard.  Mathemat-

ica materials to accompany a traditional calculus text. In Lum [Lum94], pages 749–761. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Aptaker:1991:ODE

[Apt91] P. S. Aptaker. Ordinary differential equation models for eddycurrents using Fourier techniques. In IEE [IEE91a], pages 26–29. ISBN 0-85296-529-X. LCCN QC760.54 .I57 1991.

Akyildiz:1997:TMC

[APU97] Y. Akyildiz, E. D. Popova, and C. P. Ullrich. Towards a more complete interval arithmetic in Mathematica. In Keranen et al. [KMH97], pages 26–36. ISBN 1-85312-502-9 (Southampton) (invalid checksum??), 952-5153-02-9 (Rovaniemi Polytechnic). ISSN 1239-7725. LCCN QA76.73.M29 I585 1997.

Adamchik:1993:GSM

[AR93] Victor Adamchik and Wolfram Research. Guide to standard Mathematica packages: version 2.2. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, third edition, 1993. ISBN ???? 459 pp. LCCN ????

Archetti:2020:DMP

[Arc20] Marco Archetti. DeFinetti: a Mathematica program to analyze the replicator dynamics of 3-strategy collective interactions. SoftwareX, 11(??): Article 100415, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S235271101930216X.

Aslaksen:1996:MVC

[Asl96] Helmer Aslaksen. Multiplevalued complex functions and computer algebra. SIGSAM Bulletin, 30(2):12–20, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Asmar:2005:PDE

[Asm05] Nakhlé H. Asmar. Partial differential equations with Fourier series and boundary value problems. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2005. ISBN 0-13-148096-0. various pp. LCCN QA577 .A85 2004.

Ablinger:2021:QMP

[AU21] Jakob Ablinger and Ali Ke-

mal Uncu. qFunctions —

a Mathematica package for qseries and partition theory applications. Journal of Symbolic Computation, 107(??):145–166, November/December 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0747717121000134.

Auciello:1996:FVL

[Auc96] N. M. Auciello. Free vibrations of a linearly tapered cantilever beam with constraining springs and tip mass. Journal of Sound and Vibration, 192(4):905–911, May 1996. CODEN JSVIAG. ISSN 0022-460X.

Adamchik:1997:SF

[AW97] Victor Adamchik and Stan Wagon. A simple formula for π. American Mathematical Monthly, 104(9):852–855, November 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). The authors employ Mathematica to extend earlier work of Bailey, Borwein, and Plouffe, [BBP97], done in 1995, but only just published, that discovered an amazing formula for π as is a power series in 16^{−k}, enabling any base-16 digit of π to be computed without knowledge of any prior digits. In this paper, Mathematica is used to find several simpler formulas having powers of 4^{−k}. They also note that it has been proven that their methods cannot be used to exhibit similar formulas in powers of 10^{−k}.

Azadegan:2013:MPC

[Aza13] B. Azadegan. A Mathematica package for calculation of planar channeling radiation spectra of relativistic electrons channeled in a diamond-structure single crystal (quantum approach). Computer Physics Communications, 184(3):1064–1069, March 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S001046551200389X.

Braden:1992:DCM

[B^{+}92] Bart Braden et al. Discovering Calculus with Mathematica. John Wiley, New York, NY, USA, 1992. ISBN 0-471-53969-4. 170 pp. LCCN QA303.5.D37 D57 1992.

Blachman:1996:CM

[B^{+}96] Nancy Blachman et al. CalcLabs with Mathematica. Brooks/Cole symbolic computation series. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1996. ISBN 0-534-34086-5. xvi + 245 pp. LCCN QA303.5.D37C343 1996.

Boersma:1996:CML

[BA96] J. Boersma and M. J. H.

Anthonissen. Calculations in Mathematica on low-frequency diffraction by a circular disk. Applied Computational Electromagnetics Society Journal, 11 (1):47–56, January 1996. CO-DEN ACEJEX. ISSN 1054-4887.

Baglivo:1995:CAS

[Bag95] Jenny Baglivo. Computer algebra systems: Maple and Mathematica. The American Statistician, 49(1):86–??, 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baglivo:2005:MLM

[Bag05] Jenny A. (Jenny Antoinette)

Baglivo. Mathematica laboratories for mathematical statistics: emphasizing simulation and computer intensive methods. ASA-SIAM series on statistics and applied probability. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2005. ISBN 0-89871-566-0 (paperback). ???? pp. LCCN QA276.4 .B34 2005.

Bahder:1992:ADR

[Bah92] T. Bahder. Analytic dispersion relations near the gamma-point in strained zincblende crystals. Physical Review B: Condensed Matter and Materials Physics, 45:1629, January 1992. CODEN PRBMDO. ISSN 1098-0121.

Bahder:1994:MSE

[Bah94] Thomas B. (Thomas Bazyli) Bahder. Mathematica for Scientists and Engineers. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54090-8. xv + 846 pp. LCCN QA76.95 .B35 1994.

Banks:2001:DEG

[Ban01] Bernard W. Banks. Differential equations with graphical and numerical methods. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-084376-8. xiii + 462 pp. LCCN QA371 .B26 2001.

Barnett:1990:FDD

[Bar90] M. Barnett. The formal derivation of distributed programs from systolic arrays. In IEEE [IEE90e], pages 363–366. ISBN 0-8186-2087-0. LCCN QA 76.58 I42 1990.

Barnett:1991:SSW

[Bar91] M. P. Barnett. Some simple ways to construct and to use formulas mechanically. SIGSAM Bulletin, 25(2):28–36, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Baran:1992:WYS

[Bar92] N. Baran. What you see is what you solve: Mathematica [2.0] and MathCAD [3.0] for Windows. (includes related article on Maple V) (software review). BYTE Magazine, 17:263– 265, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barnett:1993:IRF

[Bar93] M. P. Barnett. Implicit rule formation in symbolic computation. Computers and Mathematics with Applications, 26(10): 35–50, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Barnett:1998:MCM

[Bar98a] Michael P. Barnett. mathscape — Combining Mathematica and TEX. TUGboat, 19(2):147–156, June 1998. ISSN 0896-3207.

Barrow:1998:CMS

[Bar98b] David Barrow. CalcLabs with Mathematica: for Stewart’s Calculus, concepts and contexts, single variable. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1998. ISBN 0-534-34967-6. x + 228 pp. LCCN ????

Barnett:2010:ATD

[Bar10]  Michael Barnett.  Aligning

text in diagrams exported by Mathematica: a question about the PostScript infrastructure. TUGboat, 31(3):223–226, 2010. ISSN 0896-3207.

Bates:1993:SMR

[Bat93] Harry E. Bates. Selected Mathematica References for the sciences and mathematics. ????, ????, 1993. ISBN ???? 33 pp. LCCN ????

Baumann:1993:MTP

[Bau93] Gerd Baumann. Mathematica in der Theoretischen Physik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56210-9. 240 pp. LCCN ????

Baumann:1996:MTP

[Bau96] Gerd Baumann. Mathematica in Theoretical Physics: selected examples from classical mechanics to fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94424-9. xi + 384 pp. LCCN QC20.7.E4B3813 1996. US$59.00. Revised, translated, expanded and updated edition of [Kau92].

Baumann:1997:SAD

[Bau97] G. Baumann. Symmetry analysis of differential equations with Mathematica. Mathematical and computer modelling, 25(8):25– ??, ???? 1997. CODEN MC-MOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Baumann:1998:SAD

[Bau98] Gerd Baumann. Symmetry analysis of differential equations with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98552-2 (hardcover). ???? pp. LCCN QA371.B36 1998.

Baumann:1999:SAD

[Bau99] G. Baumann. Symmetry analysis of differential equations with Mathematica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98552-2. 570 pp. LCCN QA371.B36 1998.

Baumann:2000:SAD

[Bau00] Gerd Baumann. Symmetry analysis of differential equations with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 2000. ISBN 0-387-98552-2 (hardcover). xii + 521 pp. LCCN QA371 .B36 2000. Includes CD-ROM.

Baumann:2005:MTP

[Bau05] Gerd Baumann. Mathematica in theoretical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 0-387-01674-0. xvi + 544 pp. LCCN QC20.7.E4 B3813 2005.

Besancon:1995:PLS

[BB95] G. Besancon and G. Bornard.

Performances and limitations of symbolic computation in nonlinear analysis and control: an example using MATHEMATICA. In Guglielmi [Gug95], pages 617–622.

Banerjee:1993:SOA

[BBD93] Bappaditya Banerjee, Anil K.

Bajaj, and Patricia Davies. Second order averaging study of an autoparametric system. American Society of Mechanical Engineers, Design Engineering Division (Publication) DE, 54:127– 138, 1993. CODEN CHSPA4. ISBN 0-7918-1171-9.

Bailey:2014:ASL

[BBK14] David H. Bailey, Jonathan M. Borwein, and Alexander D. Kaiser. Automated simplification of large symbolic expressions. Journal of Symbolic Computation, 60(??):120–136, 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855x (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S074771711300117X.

Bailey:1997:RCV

[BBP97] David Bailey, Peter Borwein, and Simon Plouffe. On the rapid computation of various polylogarithmic constants. Mathematics of Computation, 66(218): 903–913, April 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.
org/mcom/1997-66-218. See

[AW97].

Blachman:1992:MQR

[BC92] Nancy Blachman and Robert Cambell. Mathematica Quick Reference, Version 2. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-62880-5. 304 pp. LCCN QA76.95 .B427 1992. US$18.25. See [BC94, BCxx].

Blachman:1994:MQR

[BC94] Nancy Blachman and Robert Cambell. Mathematica Quick Reference, Version 2. Toppan, Tokyo, Japan, 1994. ISBN 4-8101-8901-5. ???? pp. LCCN ???? Japanese translation of [BC92].

Blachman:19xx:MGV

[BCxx] Nancy Blachman and Robert Cambell. Mathematica griffbereit, Version 2. Friedrich Vieweg und Sohn, Braunschweig, Germany, 19xx. ISBN 3-528-06524-9. ???? pp. LCCN ???? German translation of Mathematica: Quick Reference, Version 2., [BC92].

Barnett:2007:MMP

[BC07] Michael P. Barnett and Joseph F.

Capitani. The MATHSCOUT Mathematica package to postprocess the output of other scientific programs. Computer Physics Communications, 177 (12):944–950, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465507003670.

Businger:1994:CMS

[BCDH94] W. Businger, P.-A. Chevalier, N. Droux, and W. Hett. Computing minimal surfaces on a transputer network. Mathematica Journal, 4(2):70–75, Spring 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Burbulla:1991:STC

[BD91] D. Burbulla and C. Dodson. Self-Tutor For Computer Calculus Using Mathematica 1.2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-803784-1. ?? pp. LCCN ????

Burbulla:1992:STC

[BD92] D. Burbulla and C. Dodson. Self-Tutor For Computer Calculus Using Mathematica 2.0. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-015280-3. 310 pp. LCCN ???? See [BDxx].

Berube:1993:CCG

[Bd93a] D. Berube and M. de Montigny. Computer calculation of graded contractions of Lie algebras and their representations. Computer Physics Communications, 76(3): 389–410, August 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Burbulla:1993:ILO

[BD93b] D. Burbulla and C. Dodson. An informal look at order and disorder. Mathematica in Education, 2(4):11–13, 1993. ISSN 1065-2965.

Burbulla:19xx:STC

[BDxx] D. Burbulla and C. Dodson. Self-Tutor for Computer Calculus Using Mathematica. Prentice-Hall and Toppan, Englewood Cliffs, NJ 07632, USA and Tokyo, Japan, 19xx. ISBN 4-8101-8558-3. ???? pp. LCCN ???? Japanese translation of [BD92].

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (ebook). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL http:/

/www.springer.com/us/book/
9783319641096.

Belsley:1989:MCA

[Bel89] D. A. Belsley. Mathematica: a critical appraisal (software review). Computer Science in Economics and Management, 2(2): 171–178, 1989. CODEN CSM-NEF. ISSN 0921-2736.

Belsley:1990:EMB

[Bel90] D. Belsley. Econometrics.m: a basic package for econometrics. Mathematica Journal, 1(1):95– 101, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Belknap:1991:CKS

[Bel91] S. Belknap. The Chicago Kinetic Simulator. Mathematica Journal, 1(4):68–86, Spring 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Belsley:1993:EMP

[Bel93]  David A. Belsley.  Econo-

metrics.m: a package for doing econometrics in Mathematica. In Varian [Var93], pages 300–343. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Bamberger:1995:IAS

[BEL^{+}95a] Roberto H. Bamberger, Brian L. Evans, Edward A. Lee, James H. McClellan, and Mark A. Yoder. Integrating analysis, simulation, and implementation tools in electronic courseware for teaching signal processing. In IEEE [IEE95b], pages 2873–2876 (vol. 5). CODEN IPRODJ. ISBN 0-7803-2431-5. ISSN 0736-7791. LCCN TK7882.S65 .I16 1995. IEEE catalog number 95CH35732.

Beltzer:1995:EAM

[Bel95b] Abraham I. Beltzer. Engineering analysis with Maple/ Mathematica. Academic Press, New York, NY, USA, 1995. ISBN 0-12-085570-4. xi + 282 pp. LCCN QA76.95 .B45 1995.

Benninga:1993:UMD

[Ben93] S. Benninga. Using Mathematica to do cost of capital calculations. Mathematica in Education, 2(3):11–14, 1993. ISSN 1065-2965.

Benker:1998:ICS

[Ben98a] Hans Benker. Ingenieurmathematik mit Computeralgebra-Systemen: AXIOM, DERIVE, MACSYMA, MAPLE, MATH-CAD, MATHEMATICA, MATH-LAP und MuPAD in der Anwendung. (German) [Engineering mathematics with computer algebra systems: AXIOM, DE-RIVE, MACSYMA, MAPLE, MATHCAD, MATHEMATICA, MATHLAP and MuPAD in their use]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. ISBN 3-528-05673-8. xiii + 439 pp. LCCN ????

Bennett:1998:BRP

[Ben98b] Charles A. Bennett. Book review: Patrick Tam, A Physicist’s Guide to Mathematica. Computers in Physics, 12(1): 41–??, January 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.168685.

Bernardin:1996:RSS

[Ber96] L. Bernardin. A review of symbolic solvers. SIGSAM Bulletin, 30(1):9–20, March 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bera:2024:MMP

[Ber24] Souvik Bera. MultiHypExp: a Mathematica package for expanding multivariate hypergeometric functions in terms of multiple polylogarithms. Computer Physics Communications, 297 (??):Article 109060, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465523004058.

Beskrovnyi:1998:AMA

[Bes98] V. N. Beskrovnyi. Applying Mathematica to the analytical solution of the nonlinear Heisenberg operator equations. Computer Physics Communications, 111(1–3):76–86, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465598000290.

Beaudoin:1994:IAC

[BESP94] Gaetan Beaudoin, Philippe Etchecopar, and Celine Saint-Pierre. Initiation au calcul differentiel et integral I avec Mathematica. Presses pedagogiques de l’Est, Rimouski, PQ, Canada, version experimentale. edition, 1994. ISBN 2-921676-02-8. various pp. LCCN ????

Betts:1990:MPM

[Bet90] K. Betts. Math packages multiply. Mechanical Engineering, 112(8):32–38, August 1990. CO-DEN MEENAH. ISSN 0025-6501.

Beneke:1993:CSP

[BFS93] T. W. Beneke, S. Fingerling, and W. W. Schwippert. Computer software packages ‘Maple’, ‘Mathcad’ and ‘Mathematica’ compared. Elektronik, 42(13): 50, 63–68, June 1993. CODEN EKRKAR. ISSN 0013-5658.

Badger:2014:MLI

[BFZ14] Simon Badger, Hjalte Frellesvig, and Yang Zhang. Multi-loop integrand reduction with computational algebraic geometry. Journal of Physics: Conference Series, 523(1):012061, 2014. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596 (elec-

tronic). URL http://stacks.
iop.org/1742-6596/523/i=1/
a=012061.

Benamara:1992:ADL

[BG92] A. Benamara and J. P. Guerin.

Algebraic design of linear multivariable systems. use of symbolic computation (Mathematica). In Strejc [Str92a], pages 396–399. ISBN 0-08-042057-5. LCCN TJ212.2 .S97 1992.

Bebbington:2001:KMP

[BG01] David Bebbington and Manfred Göbel. KLEIN: a Mathematica package for radar polarimetry based on spinor and tensor algebra. Journal of Symbolic Computation, 31(6):745–751, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http://www.

idealibrary.com/links/doi/
10.1006/jsco.2000.0460;
http://www.idealibrary.com/
links/doi/10.1006/jsco.2000.
0460/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0460/ref.

Breslin:1996:MMB

[BGH96] S. G. Breslin, M. J. Grimble, and C. H. Houpist. MIMOQCAD: a Mathematica based multivariable control system CAD package. In Anonymous [Ano96g], pages 3/1–5.

Blum:1993:IOC

[BGS93] W. Blum, V. Ganzha, and W. Strampp. An introduction to ODE’s by CAS. In G. Jacob, N. E. Oussous, and S. Steinberg, editors, Proceedings of the 1993 International IMACS Symposium on Symbolic Computation, pages 110–119. IMACS, Laboratoire d’Informatique Fondamentale de Lille, France, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1993.

Barker:1996:CSA

[BGZ96] H. A. Barker, P. W. Grant, and M. Zhuang. Control system analysis with Mathematica. In Anonymous [Ano96g], pages 2/1–5.

Barvik:1993:ASP

[BH93] I. Barvik and P. Herman. Application of the software package Mathematica in generalized master equation method. In De Groot and Nadrchal [DN93], pages 248–249. ISBN 981-02-1245-3. LCCN QC19.2.I53 1992.

Budny:1995:PFE

[BH95] D. Budny and B. Herrick, editors. Proceedings. Frontiers in Education 1995 25th Annual Conference. Engineering Education for the 21st Century. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-3022-6. LCCN ???? Two volumes. IEEE Catalog No. 95CH35867.

Brown:2007:RAM

[BH07] J. Randall Brown and Milton E.

Harvey. Rational arithmetic Mathematica functions to evaluate the one-sided one-sample K-S cumulative sample distribution. Journal of Statistical Software, 19(6):1–32, March 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v19/i06.

Bauer:2008:CFG

[BH08a] Friedrich L. Bauer and Christoph Haenel. Carl Friedrich Gauß, das 17-Eck und MATHEMAT-ICA. (German) [Carl Friedrich Gauß, the 17-gon and Mathematica]. Informatik Spektrum, 31(5):??, ???? 2008. CO-DEN INSKDW. URL http://

link.springer.com/article/
10.1007/s00287-008-0268-7.

Brown:2008:APM

[BH08b] J. Randall Brown and Milton E.

Harvey. Arbitrary precision Mathematica functions to evaluate the one-sided one sample K-S cumulative sampling distribution. Journal of Statistical Software, 26(3):1–55, June 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v26/i03.

Brown:2008:RAM

[BH08c] J. Randall Brown and Milton E. Harvey. Rational arithmetic Mathematica functions to evaluate the two-sided one sample K-S cumulative sampling distribution. Journal of Statistical Software, 26(2):1–40, June 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v26/i02.

Bhatti:1996:OCD

[Bha96] M. Asghar Bhatti. Optimum cost design of partially composite steel beams using LRFD. Engineering Journal, 33(1):18–29, First Quarte 1996. CODEN EJASAR. ISSN 0013-8029.

Bhatti:2000:POM

[Bha00] M. Asghar Bhatti. Practical optimization methods: with Mathematica applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98631-6. xiii + 715 pp. LCCN QA402.5.B49 2000.

Bhatti:2005:FFE

[Bha05] M. Asghar Bhatti. Fundamental Finite Element Analysis and Applications: with Mathematica and MATLAB Computations: Fundamental Concepts. John Wiley and Sons, Inc., Chichester, UK, 2005. ISBN 0-471-64808-6. xx + 700 pp. LCCN TA646 .B56 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;

http://www.loc.gov/catdir/
toc/ecip0414/2004002270.html.

Bhatti:2006:ATF

[Bha06] M. Asghar Bhatti. Advanced topics in finite element analysis of structures: with Mathematica and MATLAB computations. John Wiley and Sons, Inc., Chichester, UK, 2006. ISBN 0-471-64807-8 (cloth). xvi + 590 pp. LCCN TA647 .B494 2006. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip058/2005005179.html.

Brown:1991:BRM

[BHH91] D. Brown, B. A. Hassan, and E. Hjelmfelt. A basic review of Mathematica. International journal of mathematical education in science and technology, 22(2):207–??, March 1991. CO-DEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Birkl:2000:CTR

[BHJ^{+}00] Gerhard Birkl, Stephan Hartmann, Manfred Jacobi, Hans Christoph Wolf, Erhard Keppler, Armin Bunde, Jens Eisert, and Max Rauner. Crichton: Timeline. Eine Reise in die Mitte der Zeit/Dyson: The Sun, The Genome, and The Internet. Tools of Scientific Revolutions/ Pais: Raffiniert ist der Herrgott Albert Einstein. Eine wissenschaftliche Biografie/Kuhn/ Försterling: Principles of Physical Chemistry/Wiser: Energy Resources. Occurrence, Production, Conversion, Use/Moss de Oliveira, de Oliveira und Stauffer: Evolution, Money, War, and Computers/Canty: Konfliktlösungen mit Mathematica. Zweipersonenspiele/Adobe Acrobat 4.0/Buchtipp. Physikalische Blätter, 56(10):69–73, October 2000. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/
doi/10.1002/phbl.20000561016/
abstract.

Bisoi:1995:NRM

[Bis95] A. K. Bisoi. Newton Raphson method, scaling at fractal boundaries and MATHE-MATICA. Mathematical and computer modelling, 21(10):91– 102, 1995. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Bittl:1993:MES

[Bit93] R. Bittl. Matrix elements of spin operators in exchange coupled tetrameric metal clusters. Chemical Physics Letters, 215 (4):279–284, December 3, 1993. CODEN CHPLBC. ISSN 0009-2614.

Boyer:1995:RSI

[BK95] F. Boyer and W. Khalil. Recursive solution of inverse and forward dynamics of flexible manipulators. In Isidori et al. [IBM^{+}95], pages 2696–2701 vol.3. 4 vol.

Boyer:1996:SFM

[BK96] F. Boyer and W. Khalil. Simulation of flexible manipulators using Newton-Euler inverse dynamic model. In IEEE [IEE96], pages 1947–1952 vol.3. ISBN 0-7803-2988-0. LCCN TJ210.3 .I44 1996. Four volumes. IEEE Catalog No. 96CH35857.

Bakulev:2013:FMP

[BK13] Alexander P. Bakulev and Vyacheslav L. Khandramai. FAPT: a Mathematica package for calculations in QCD Fractional Analytic Perturbation Theory. Computer Physics Communications, 184(1):183–193, January 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465512002809.

Bytev:2015:HHF

[BK15] Vladimir V. Bytev and Bernd A.

Kniehl. HYPERDIRE HYPERgeometric functions DIfferential REduction: Mathematicabased packages for the differential reduction of generalized hypergeometric functions: Horntype hypergeometric functions of two variables. Computer Physics Communications, 189 (??):128–154, ???? 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465514004056.

Bytev:2016:HHF

[BK16] Vladimir V. Bytev and Bernd A. Kniehl. HYPERDIRE — HY-PERgeometric functions DIfferential REduction: Mathematicabased packages for the differential reduction of generalized hypergeometric functions: Lauricella function F_{c} of three variables. Computer Physics Communications, 206(??):78– 83, September 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465516301059.

Bytev:2013:HHF

[BKK13] Vladimir V. Bytev, Mikhail Yu. Kalmykov, and Bernd A. Kniehl. HYPERDIRE, HYPERgeometric functions DIfferential REduction: Mathematicabased packages for differential reduction of generalized hypergeometric functions pF_{p−1}, F_{1}, F_{2}, F_{3}, F_{4}. Computer Physics Communications, 184 (10):2332–2342, October 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465513001690.

Bytev:2014:HFD

[BKM14] Vladimir V. Bytev, Mikhail Yu. Kalmykov, and Sven-Olaf Moch. HYPERgeometric functions DIfferential REduction (HYPER-DIRE): MATHEMATICA based packages for differential reduction of generalized hypergeometric functions: F_{D} and F_{S} Horntype hypergeometric functions of three variables. Computer Physics Communications, 185 (11):3041–3058, November 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465514002574.

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D. Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin, 25(3):28–37, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bau:1993:TTM

[BL93] H. H. Bau and K. Lebiedzinski. Teaching thermodynamics with Mathematica. CoED, 3(4): 47–55, October 1993. CODEN CWLJDP. ISSN 0736-8607.

Blachman:1990:WM

[Bla90] Nancy Blachman. Working with Mathematica. Visual Symbols, Inc., Berkeley, CA, USA, 1990. ISBN ???? 173 + 11 + 3 pp. LCCN ????

Blachman:1992:MPAa

[Bla92a] Nancy Blachman. Mathematica, a Practical Approach. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-563826-7. xv + 365 pp. LCCN QA76.95.B53 1992. See [Bla92b, Blaxx, UB93].

Blachman:1992:MPAb

[Bla92b] Nancy Blachman. Mathematica: a Practical Approach. Prentice-Hall and Toppan, Englewood Cliffs, NJ 07632, USA and Tokyo, Japan, 1992. ISBN 4-8101-8538-9. 413 pp. LCCN ???? Japanese translation of [Bla92a].

Blachman:19xx:M

[Blaxx] Nancy Blachman. Mathematica. Academic Service, Den Haag, The Netherlands, 19xx. ISBN 90-395-0016-9. 362 pp. LCCN ???? Dutch translation of Mathematica: A Practical Approach, [Bla92a].

Barrio:1997:CVP

[BLM97] Roberto Barrio, Carlos Lopez, and Eduardo Martinez. Calculo vectorial: practicas con Mathematica y Maple, volume 54 of Textos Docentes. Prensas Universitarias de Zaragoza, Zaragoza, Spain, 1997. ISBN 84-7733-488-9. viii + 229 pp. LCCN ????

Binous:1992:MM

[BM92] H. Binous and B. J. McCoy. Moments with Mathematica. CEE. Chemical engineering education, 26(1):12–17, Winter 1992. CO-DEN CHEDAY. ISSN 0009-2479.

Baldwin:1994:FTM

[BM94] J. F. Baldwin and T. P. Martin. Fuzzifying a target motion analysis model using Fril and Mathematica. IEEE International Conference on Fuzzy Systems, 2:1171–1175, 1994. IEEE catalog number 94CH3430-6.

Baras:1990:FEA

[BMM90] John S. Baras, David C. MacEnany, and Robert L. Munach. Fast error-free algorithms for polynomial matrix computations. In IEEE [IEE90d], pages 941–946. LCCN TJ 217 I11c 1990. IEEE catalog number 90CH2917-3.

Bar-Natan:1991:RDS

[BN91] D. Bar-Natan. Random-dot stereograms. Mathematica Journal, 1(3):69–75, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Bobeszko:1993:LMN

[Bob93] A. Bobeszko. The layer model of the nucleus. Mathematica in Education, 2(2):3–5, Winter 1993. ISSN 1065-2965.

Boss:1996:TRM

[Bos96] E. Boss. Tsunami research on the Macintosh. SciTech Journal, 6(1):14–16, January 1996. CO-DEN SITJE8. ISSN 1072-0995.

Boyland:1991:GSM

[Boy91] Philip Boyland. Guide to standard Mathematica packages. Technical report, Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1991. 306 pp.

Barnett:1994:SCE

[BP94] Michael P. Barnett and Kevin R. Perry. Symbolic computation for electronic publishing. TUGboat, 15(3):285–292, September 1994. ISSN 0896-3207.

Bombardi:1995:NMD

[BP95] T. Bombardi and L. Piancastelli. A numerical method for determining the tritangential circumference to three regular primitive planes. Pixel. Computer Graphics, CAD/CAM, Image Processing, 16(12):19–26, 1995. CODEN PIXEDD. ISSN 0392-8217.

Bellomo:1999:MDS

[BPR99] Nicola Bellomo, Luigi Preziosi, and Antonio Romano. Mechanics and dynamical systems with Mathematica. Modeling and simulation in science, engineering, and technology Modeling and simulation in science, engineering and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4007-X. ???? pp. LCCN QA805.B44 1999.

Bellomo:2000:MDS

[BPR00] N. Bellomo, Luigi Preziosi, and Antonio Romano. Mechanics and dynamical systems with Mathematica. Modeling and simulation in science, engineering, and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 0-8176-4007-X, 3-7643-4007-X. xiii + 417 pp. LCCN QA805 .B44 2000.

Brown:1990:CMC

[BPU90a] D. Brown, H. Porta, and J. Uhl. Calculus & Mathematica: Courseware for the nineties. Mathematica Journal, 1(1):43– 50, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Brown:1990:CMPa

[BPU90b] Donald P. Brown, Horacio Porta, and J. Jerry Uhl. Calculus & Mathematica: Part 1, Basics, Tutorials, and Literacy Sheets. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-57270-2. xxi + 426 pp. LCCN QA303 .B8848 1990. 6 computer disks.

Brown:1990:CMPb

[BPU90c] Donald P. Brown, Horacio Porta, and J. Jerry Uhl. Calculus & Mathematica: Part 1, Give it a Try. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-57271-0. iv + 286 pp. LCCN QA303 .B8849 1990.

Brown:1991:CMP

[BPU91] D. Brown, H. Porta, and J. Uhl.

Calculus & Mathematica: Part 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN ???? ?? pp. LCCN ????

Bischof:1998:ACR

[BQO98] C. H. Bischof and G. Quintana-Ortí. Algorithm 782: Codes for rank-revealing QR factorizations of dense matrices. ACM Transactions on Mathematical Software, 24(2):254–257, June 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toms/1998-24-2/p254-
bischof/.

Bramley:1996:TNR

[Bra96] Randall Bramley. Technical news & reviews: Maple V Release 4; Mathematica 3.0 from Wolfram; HotGAMS from NIST; Informant scans Web and e-mails you; more on the web. . .. IEEE Computational Science & Engineering, 3(3):80, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses a Java-based front end for the Guide to Available Mathematical Software, with a pointer to

URL http://math.nist.gov/
HotGAMS/.

Brown:1991:CMb

[Bro91] Donald P. Brown. Calculus and Mathematica. Addison-Wesley, Reading, MA, USA, next preliminary edition. edition, 1991. ISBN 0-201-56811-X. ???? pp. LCCN ???? 2 computer disks.

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. IS-SAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6–8, 1993, Kiev, Ukraine. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Brockway:1994:DME

[Bro94] Jack W. Brockway. Determining the matrix elements of the linearized cross-collision operators of gaseous mixtures using Mathematica. Thesis (M.S.), University of North Carolina at Greensboro, Greensboro, NC, USA, 1994. iv + 38 pp.

Broughan:2009:AEE

[Bro09] Kevin A. Broughan. An algorithm for the explicit evaluation of GL(nR) Kloosterman sums. ACM Communications in Computer Algebra, 43(1):1– 10, March 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Beneke:1994:DTF

[BS94] T. Beneke and W. Schwippert. Double-track into the future: MathCAD will gain new users with Standard and Plus versions. Elektronik, 43(15): 107–110, July 1994. CODEN EKRKAR. ISSN 0013-5658.

Borwein:2011:SVGa

[BS11a] Jonathan M. Borwein and Armin Straub. Special values of generalized log-sine integrals. ArXiv e-prints, March 2011. URL http:

//adsabs.harvard.edu/abs/
2011arXiv1103.4298B.

Borwein:2011:SVGb

[BS11b] Jonathan M. Borwein and Armin Straub. Special values of generalized log-sine integrals. In Schost and Emiris [SE11], pages 43–50. ISBN 1-4503-0675-

6. LCCN QA76.95 .I59 2011.

Borwein:2012:IMM

[BS12] Jonathan M. Borwein and Matthew P. Skerritt. An introduction to modern mathematical computing: with Mathematica(R). Springer undergraduate texts in mathematics and technology. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-4252-4 (print), 1-4614-4253-2 (ebook). ISSN 1867-5506 (print), 1867-5514 (electronic). LCCN QA76.95 .B67 2012. URL http:

//public.eblib.com/choice/
publicfullrecord.aspx?p=1030664.

Bitterfeld:1994:CPC

[BSB94] G. Bitterfeld, J. Steiner, and J. Browne. Construction of parametric curves using exterior calculus and Mathematica symbolic computation programme. In Anonymous [Ano94j], pages 043–050. ISBN ???? LCCN ????

Broschat:1993:ISU

[BSHS93] S. Broschat, J. Schneider, F. Hastings, and M. Steeds. Interactive software for undergraduate electromagnetics. IEEE transactions on education, 36(1): 123–126, February 1993. CO-DEN IEEDAB. ISSN 0018-9359.

Benninga:1993:INO

[BSS93] S. Benninga, R. Steinmetz, and J. Stroughhair. Implementing numerical option pricing models. Mathematica Journal, 3(4):66– 73, Fall 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Buchberger:1992:GBM

[Buc92] B. Buchberger. Gröbner bases in Mathematica: enthusiasm and frustration. IFIP Transactions. A. Computer Science and Technology, A-2:119–129, 132, 1992. CODEN ITATEC. ISSN 0926-5473.

Buchberger:1993:MSD

[Buc93] B. Buchberger. Mathematica:

a system for doing mathematics by computer? Lecture Notes in Computer Science, ??(722): 1, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1991:CMa

[BUP91] Donald P. Brown, J. J. (J. Jerry)

Uhl, and Horacio Porta. Calculus and Mathematica. Addison-Wesley, Reading, MA, USA, preliminary edition, 1991. ISBN 0-201-57273-7 (disk 1), 0-201-57274-5 (disk 2), 0-201-57275-3 (disk 3), 0-201-57276-1 (disk 4) 0-201-57277-X (disk 5), 0-201-57278-8 (disk 6), 0-201-57271-0 (Give it a try manual), 0-201-57270-2 (Basics, tutorials manual). LCCN QA303. 6 computer disks.

Burkhardt:1993:ESM

[Bur93] W. Burkhardt. Erste Schritte mit Mathematica. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56650-3. 120 pp. LCCN ????

Burkhardt:1994:ESM

[Bur94a]  Werner Burkhardt.  Erste

Schritte mit Mathematica (English: First Steps in Mathematica). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-19875-X (Berlin), 0-387-19875-X (New York). viii + 100 pp. LCCN QA76.95 .B8713 1994.

Burkhardt:1994:FSM

[Bur94b] Werner Burkhardt. First Steps in Mathematica. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-19875-X. viii + 100 pp. LCCN QA76.95 .B8713 1994. US$19.95.

Blachman:1999:MPA

[BW99] Nancy Blachman and Colin P. Williams. Mathematica: a practical approach. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-259201-0. xiv + 631 pp. LCCN QA76.95 .B53 1999.

Ben-Yaakov:1994:AMT

[BYA94] Sam Ben-Yaakov and Daniel Adar. Average models as tools for studying the dynamics of switch mode DC-DC converters. PESC Record — IEEE Annual Power Electronics Specialists Conference, 2:1369–1376, 1994. CODEN PRICDT. ISSN 0275-9306. IEEE catalog number 94CH3418-1.

Bykhovsky:2019:MCN

[Byk19]  D. Bykhovsky.  Mathemat-

ica code for numerical generation of random process with given distribution and exponential autocorrelation function. SoftwareX, 8(??):18–20, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S235271101730033X.

Bzowski:2021:TMP

[Bzo21] Adam Bzowski. TripleK: a Mathematica package for evaluating triple-K integrals and conformal correlation functions. Computer Physics Communications, 258(??):Article 107538, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465520302563.

Calistru:1995:SAT

[Cal95a] C. Calistru. A symbolic approach to tuning PID controllers via MATHEMATICA software. Systems analysis, modelling, simulation, 18(??): 341–344, 1995. CODEN SAM-SEC. ISSN 0232-9298.

Calistru:1995:SOA

[Cal95b] Catalin-Nicolae Calistru. Symbolic optimization approach for tuning of PID controllers. IEEE Conference on Control Applications — Proceedings, pages 174– 175, 1995. CODEN ICOAE8. ISSN 1085-1992.

Calvis:2004:MDE

[Cal04] David Calvis. Mathematica for differential equations: projects, insights, syntax and animations. Pearson Prentice Hall, Upper Saddle River, NJ, USA, 2004. ISBN 0-13-143976-6. xx + 160 pp.

Christian:1994:BRJ

[CAMM94] Wolfgang Christian, G. Andrew Antonelli, A. John Mallinckrodt, and Susan McKay. Book review: James M. Feagin, Quantum Methods with Mathematica(R). Computers in Physics, 8(4):427–??, July 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

                                        URL https://aip.scitation.
                                        org/doi/10.1063/1.4823318.

Canty:2000:KMH

[Can00]  Morton J. Canty.  Kon-

fliktlösungen mit Mathematica: Zweipersonenspiele. (German) [Resolving conflicts with Mathematica: two-person games]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-642-57107-7, 3-540-65827-0. xiii + 247 pp. LCCN HB144; QA269-272.

URL http://zbmath.org/?q=
an:0995.91001.

Canty:2003:RCM

[Can03] Morton John Canty. Resolving conflicts with Mathematica: algorithms for two-person games. Academic Press, New York, NY, USA, 2003. ISBN 0-12-158855-6. 327 pp. LCCN ???? Revised and updated edition of the popular German textbook. Translation from the German language edition: “Konfliktlösungen mit Mathematica”. Includes CD-ROM.

Cap:2003:MMP

[Cap03] Ferdinand Cap. Mathematical methods in physics and engineering with Mathematica. Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 1-58488-402-9. 340 pp. LCCN QC20.6 .C36 2003. Also issued online.

Carroll:1992:RCS

[Car92] John Carroll. The role of computer software in numerical analysis teaching. ACM SIGNUM Newsletter, 27(2): 2–31, April 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carducci:1994:GSL

[Car94a] O. Carducci. Graphical solutions to linear programming problems. Mathematica in Education, 3(2):20–22, Spring 1994. ISSN 1065-2965.

Carter:1994:LPM

[Car94b] Michael Carter. Linear programming with Mathematica. Discussion paper 9406, Dept. of Economics, University of Canterbury, Christchurch, New Zealand, November 1994. 61 pp.

Carriere:1996:PRW

[Car96] L. R. Oudin A. Carrière, editor. Proceedings of the 5th RHINE workshop on computer algebra. Institut Franco-Allemand de Recherches de Saint-Louis, 68300 Saint-Louis, France, April 1996. ISBN ???? LCCN ????

Carter:2004:M

[Car04] J. D. Carter. Mathematica 5.0. SIAM Review, 46(3):564–568, ???? 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:2008:FRM

[Car08] John D. Carter. Featured review: Mathematica 6. SIAM Review, 50(1):149–152, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Caudill:1990:NNA

[Cau90] M. Caudill. Neural networks application Mathematica v.1.2. AI Expert, 5(5):61–65, May 1990. ISSN 0888-3785.

Cavallini:1993:CUH

[Cav93] Fabio Cavallini. Computing the unit hydrograph via linear programming. Computers and Geosciences, 19(9):1285–1294, October 1993. CODEN CGEODT. ISSN 0098-3004 (print), 1873-7803 (electronic).

Cook:1995:BSU

[CB95] Peyton Cook and Lyle D.

Broemeling. Bayesian statistics using Mathematica. The American Statistician, 49(1):70– ??, 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:1992:DAL

[CD92] J. Cohen and D. DeBaun. Discrete approximations of linear functionals. Mathematica Journal, 2(2):62–65, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Cabral:2002:CCM

[CD02] Hildeberto Cabral and Florin Diacu, editors. Classical and celestial mechanics: the Recife lectures. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05022-

8. xviii + 385 pp. LCCN QB362.M3.C52 2002. URL

http://www.loc.gov/catdir/
description/prin022/2002072263.
html.

Cook:1992:CSS

[CDS^{+}92] D. Cook, R. Dubish, G. Sowell, P. Tam, and D. Donnelly. A comparison of several symbolmanipulating programs: Part I. Computers in Physics, 6(4):411– 420, July/August 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Caballero:1996:OSM

[CF96] Carlos A. Caballero and Rui Fragassi Souza. Observation of solitons with MATHEMAT-ICA. IEEE transactions on education, 39(1):46–49, February 1996. CODEN IEEDAB. ISSN 0018-9359.

Cordero:19xx:GDC

[CFGxx] Luis A. Cordero, Marisa Fernandez, and Alfred Gray. Geometria differencial de curvas y superfices. Addison-Wesley IberoAmericana, ????, 19xx. ISBN 0-201-65364-8. ???? pp. LCCN ???? Spanish translation of Modern Differential Geometry of Curves and Surfaces.

Clark:2017:MMN

[CGO17] D. B. Clark, E. Godat, and F. I. Olness. ManeParse: a Mathematica reader for parton distribution functions. Computer Physics Communications, 216 (??):126–137, July 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465517300802.

Chowdhury:1994:SMS

[CH94] K. L. Chowdhury and R. C. Huntsinger. Simulation of minimal surfaces by isovector methods. Transactions of the Society for Computer Simulation, 11(3): 159–177, October 1994. CODEN TSCSEV. ISSN 0740-6797.

Caferra:1995:GGF

[CH95] R. Caferra and M. Herment.

A generic graphic framework for combining inference tools and editing proofs and formulae. Journal of Symbolic Computation, 19(1/2/3):217–244 (or 217–243??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chatterjee:1991:DSF

[Cha91] T. K. Chatterjee. A dynamical study of the frequency of merging galaxies. III. merges involving disk galaxies in a single crossing time or orbital period. Astrophysics and space science, 184(2):273–280, October 1991. CODEN APSSBE. ISSN 0004-640X.

Chaitin:2001:ER

[Cha01] Gregory J. Chaitin. Exploring randomness. Discrete mathematics and theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 1-85233-417-7. x + 164 pp. LCCN QA267 .C485 2001.

Cheng:1991:SFE

[Che91] K. J. Cheng. Symbolic finite element analysis using computer algebra: heat transfer in rectangular duct flow. Computers and Mathematics with Applications, 22(12):15–22, 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Cheng:1994:FCS

[Che94]  Ching-Hsue Cheng.  Fuzzy

consecutive-k-out-of-n:F system reliability. Microelectronics and Reliability, 34(12):1909–1922, December 1994. CODEN MCR-LAS. ISSN 0026-2714 (print), 1872-941X (electronic).

Chevallereau:1996:FTN

[Che96] C. Chevallereau. Feasible trajectories for a non redundant robot at a singularity. Proceedings — IEEE International Conference on Robotics and Automation, 2: 1871–1876, 1996. CODEN PI-IAET. ISSN 1050-4729. IEEE catalog number 96CB35857.

Cheung:1998:GSM

[Che98] Chi Keung Cheung. Getting started with Mathematica. John Wiley and Sons, Inc., Chichester, UK, 1998. ISBN 0-471-24050-8. x + 183 pp. LCCN QA76.95 .G47 1998.

Cheung:2005:GSM

[Che05] Chi Keung Cheung, editor. Getting started with Mathematica. John Wiley and Sons, Inc., Chichester, UK, second edition, 2005. ISBN 0-471-47815-6 (paperback). 231 pp. LCCN QA76.95 .G47 2005.

Chen:2025:BRD

[Che25] Hao Chen. Book review: Differential Equations: Solving Ordinary and Partial Differential Equations with Mathematica. SIAM Review, 67(3): 654–655, ???? 2025. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://epubs.siam.
org/doi/10.1137/24M170096X.

Chin:1992:AMD

[Chi92] J. Chin. Applying Mathematica in decision support systems. Mathematica in Education, 2(1): 15–18, Fall 1992. ISSN 1065-2965.

Coombs:1995:DEM

[CHL^{+}95] Kevin R. Coombs, Brian R.

Hunt, Ronald L. Lipsman, John E. Osborn, and Garrett J. Stuck. Differential Equations with Mathematica. John Wiley and Sons, Inc., Chichester, UK, 1995. ISBN 0-471-10874-X. v + 218 pp. LCCN QA371.5.D37 D54 1995.

Christensen:1993:PSP

[Chr93] A. C. Christensen. Prospects symbolic processing for modelling in control system design. In Proc. on Information Control Problems in Manufacturing Technology Symposium (IFAC), volume VI, pages 193–196. ????, ????, July 1993.

Christensen:1994:PFJ

[Chr94] A. C. Christensen. Symbolics in control design: Prospects and research issues. In Mattsson et al. [MGC94], page ???? ISBN 0-7803-1800-5 (softbound), 0-7803-1801-3 (microfiche). LCCN TJ 212.2 I3256 1994.

Chung:2014:PSC

[Chu14] Youngjoo Chung. Symbolic Computing package for Mathematica for versatile manipulation of mathematical expressions. In Hong and Yap [HY14], pages 21–25. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (ebook). LCCN QA76.9.M35.

Castillo:1993:M

[CIG^{+}93] Enrique Castillo, Andrés Iglesias, José Manuel Gutiérrez, Elena Alvarez, and Angel Cobo. Mathematica. Editorial Paraninfo, ????, 1993. ISBN 84-283-2017-9. 534 pp. LCCN ????

Castellvi:1994:TST

[CJL94] P. Castellvi, X. Jaen, and E. Llanta. TTC: symbolic tensor and exterior calculus. Computers in Physics, 8(3):360– 367, 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Conceicao:2013:CSC

[CKP13] Ana C. Conceição, Viktor G. Kravchenko, and José C. Pereira. Computing some classes of Cauchy type singular integrals with Mathematica software. Advances in Computational Mathematics, 39(2):273–288, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-012-9279-7.

Cole:1990:VEU

[CKS90] Rodney Cole, David Krull, and Monica Sweitzer. Visual electromagnetics using Mathematica and HyperCard. In IEEE [IEE90a], pages 1671–1674. CO-DEN IAPSBG. ISSN 0272-4693. LCCN TK 7871.6 N56 1990. IEEE catalog number 90CH2776-3.

Chetyrkin:2000:RMP

[CKS00] K. G. Chetyrkin, J. H. Kühn, and M. Steinhauser. Run-Dec: a Mathematica package for running and decoupling of the strong coupling and quark masses. Computer Physics Communications, 133 (1):43–65, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465500001557.

Cox:1992:IVA

[CLO92] David A. Cox, John B. Little, and Donal O’Shea. Ideals, varieties, and algorithms: An introduction to computational algebraic geometry and commutative algebra. Undergraduate Texts in Mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97847-X. xii + 513 pp. LCCN QA564 .C688 1992.

Cox:2007:IVA

[CLO07] David A. Cox, John B. Little, and Donal O’Shea. Ideals, varieties, and algorithms: an introduction to computational algebraic geometry and commutative algebra. Undergraduate texts in mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2007. ISBN 0-387-35650-9, 0-387-35651-7 (ebook). xv + 551 pp. LCCN QA564 .C688 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0814/2006930875-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0814/
2006930875-t.html;   http:
//www.springerlink.com/openurl.
asp?genre=book&isbn=978-0-
387-35650-1.

Coombes:1998:MCM

[CLR98] Kevin Robert Coombes, Ronald L. Lipsman, and J. (Jonathan) Rosenberg. Multivariable calculus and Mathematica: with applications to geometry and physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98360-0 (paperback). xiii + 283 pp. LCCN QA303.5.C65 C66 1998.

Cyrol:2017:FMT

[CMS17] Anton K. Cyrol, Mario Mitter, and Nils Strodthoff. Form-Tracer: A Mathematica tracing package using FORM. Computer Physics Communications, 219(??):346–352, October 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0010465517301662.

Campbell:2012:NSI

[CMT12] David K. Campbell, Sebastian M. Marotta, and Thomas A. Tanury. Nonlinear science: an interactive Mathematica notebook. Cambridge University Press, Cambridge, UK, 2012. ISBN 0-521-13882-5. LCCN ???? One DVD.

Choi:1992:ASC

[CN92] D. K. Choi and S. Nomura. Application of symbolic computation to two-dimensional elasticity. Computers and structures, 43(4):645–649, May 1992. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Constantinescu:1994:ADA

[CN94] F. Constantinescu and M. Nitescu.

Analysis and design of analog circuits using a symbol manipulation technique. Revue Roumaine des Sciences Techniques, Serie Electrotechnique et Energetique, 39(1):73–77, January–March 1994. CODEN RTEEAE. ISSN 0035-4066.

Coffee:1989:MV

[Cof89] Peter Coffee. Mathematica, version 1.2. PC Week, ??(??):52–53, 56–58, October 30, 1989. ISSN 0740-1604. Requires Mac; 4MB; RAM.

Coffee:1991:MLS

[Cof91] P. Coffee. Mathematica 2.0 for DOS lacks speed, stability. PC Week, 8:23–24, August 12, 1991. ISSN 0740-1604.

Coffee:1992:MPT

[Cof92a] P. Coffee. Math packages take distinctive approaches: Mathcad 3.0 is more compact, stable; Mathematica 2.0 has edge in 3-D graphics. PC Week, 9:81–82, February 17, 1992. ISSN 0740-1604.

Coffee:1992:MWG

[Cof92b] P. Coffee. Mathematica for Windows greater than DOS version. PC Week, 9:33–34, January 27, 1992. ISSN 0740-1604.

Cohen:1992:PLS

[Coh92] J. Cohen. Packages for logic and set theory. Mathematica Journal, 2(1):91–93, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Cohen:1994:AEN

[Coh94a] J. Cohen. Arclength — an example of the new calculus. Mathematica in Education, 3(2):23–28, Spring 1994. ISSN 1065-2965.

Cohen:1994:MCM

[Coh94b]  R. Cohen.  MathLink con-

nects Mathematica and Excel. MacWEEK, 8(5):6, January 31, 1994. CODEN MWEEEI. ISSN 0892-8118.

Comello:1996:PST

[Com96] Vic Comello. Physics software that’s suitable for engineers, too. Research & Development, 38 (13):36–38, 40, December 1996.

CODEN REDEEA. ISSN 0746-9179. Reviews several software packages, including some from Mathematica and Maple.

Coombes:1998:DEM

[Coo98a] Kevin Robert Coombes. Differential equations with Mathematica: revised for Mathematica 3.0. John Wiley and Sons, Inc., Chichester, UK, second edition, 1998. ISBN 0-471-17696-6. vi + 240 pp. LCCN QA371.5.D37 D542 1998.

Coombes:1998:MP

[Coo98b] Kevin Robert Coombes. The Mathematica primer. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63130-0 (hardcover), 0-521-63715-5 (paperback). xvii + 214 pp. LCCN QA76.95.M3884 1998.

Coxeter:1993:RPP

[Cox93] H. S. M. (Harold Scott Macdonald) Coxeter. The Real Projective Plane. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97890-9 (Mac disk), 0-387-97889-5 (PC disk). xiii + 222 pp. LCCN QA471.C68 1993.

Christensen:1990:MSP

[CP90] S. Christensen and L. Parker.

MathTensor: a system for performing tensor analysis by computer. Mathematica Journal, 1 (1):51–61, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Corder:1993:CSP

[CP93a] P. R. Corder and R. Persh. Castigliano and symbolic programming in finite element analysis. American Society of Mechanical Engineers, Petroleum Division (Publication) PD, 52:173– 180, 1993. CODEN ASMPEX. ISBN 0-7918-0948-X.

Corder:1993:TDR

[CP93b] P. R. Corder and R. Persh. Transient dynamic response of a whip antenna to transverse shock loading. American Society of Mechanical Engineers, Petroleum Division (Publication) PD, 52:181–183, 1993. CO-DEN ASMPEX. ISBN 0-7918-0948-X.

Crooke:1991:GCM

[CR91] Philip Crooke and John Ratcliffe. A Guidebook to Calculus and Mathematica. Wadsworth new directions in mathematics series. Wadsworth Pub. Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-15483-2. 256 pp. LCCN QA303.5.D37C76 1991.

Crandall:1990:NCP

[Cra90] R. E. Crandall. The NeXT computer as physics machine. Computers in Physics, 4(2):132– 141, March/April 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Crandall:1991:SSS

[Cra91a]  R. E. Crandall.  Symbolic

software supports new adventures in physics. Computers in Physics, 5(6):576–579, November/December 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Crandall:1991:MSa

[Cra91b] Richard E. Crandall. Mathematica for the Sciences. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51001-4. xiii + 300 pp. LCCN Q172 .C73 1991. See [Cra91c].

Crandall:1991:MSb

[Cra91c] Richard E. Crandall. Mathematica for the Sciences. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 1991. ISBN 4-8101-8036-0. 315 pp. LCCN ???? Japanese translation of [Cra91b].

Crandall:1994:PSC

[Cra94] Richard E. Crandall. Projects in Scientific Computation. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-97808-9. xxiv + 470 pp. LCCN Q183.9.P733 1994. US$49.95. Includes computer disk.

Crandall:1996:TAS

[Cra96] Richard E. Crandall. Topics in Advanced Scientific Computation. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1996. ISBN 0-387-94473-7. ix + 340 pp. LCCN QA9.58.C73 1996. US$39.95. Includes programs in C and Mathematica.

Crocker:1991:QSI

[Cro91] Deborah Ann Crocker. A qualitative study of interactions, concept development and problem solving in a calculus class immersed in the computer algebra system Mathematica. Thesis (Ph.D.), College of Education, Ohio State University, Columbus, OH, USA, 1991. ix + 274 pp. Advisor: Alan Osborne.

Crocker:1993:DCD

[Cro93] D. A. Crocker. Development of the concept of derivative in a calculus class using the computer algebra system Mathematica. In Lum [Lum93], pages 251– 255. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Cropper:1998:MCP

[Cro98] William H. Cropper. Mathematica computer programs for physical chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98337-6 (softcover). xi + 246 pp. LCCN QD455.3.E4 C76 1998.

CaballeroP:1996:OSM

[CS96a] C. A. Caballero P. and R. F. Souza. Observation of solitons with MATHEMATICA. IEEE transactions on education, 39(1): 46–49, February 1996. CODEN IEEDAB. ISSN 0018-9359.

Cheng:1996:GWF

[CS96b] A. H-D Cheng and P. Sidauruk. A ground-water flow Mathematica tool package. Ground water, 34(1):41–??, 1996. CODEN GR-WAAP. ISSN 0017-467X.

Cheng:1994:AIL

[CSA94] A. H. Cheng, P. Sidauruk, and Y. Abousleiman. Approximate inversion of the Laplace transform. Mathematica Journal, 4(2):76–82, Spring 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Carney:1994:DNP

[CT94] Sean N. Carney and John R.

Thompson. Dispersive and nonlinear propagation of dualfrequency pulses in optical fiber. American Journal of Physics, 62 (8):744–749, August 1994. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Clendenning:1997:BRG

[CTMC97] Lester M. Clendenning, Patrick T.

Tam, Susan R. McKay, and Wolfgang Christian. Book review: Gerd Baumann, Mathematica in Theoretical Physics: Selected Examples from Classical Mechanics to Fractals. Computers in Physics, 11(6):595– ??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4822608.

Culioli:1991:IM

[Cul91] Jean-Christophe Culioli. Introduction à Mathematica. Ellipses: Edition Marketing, Paris, France, 1991. ISBN 2-7298-9144-7. 266 pp. LCCN ???? In French.

Culioli:1993:IM

[Cul93] Jean-Christophe Culioli. Introduction à Mathematica. Ellipses: Edition Marketing, Paris, France, deuxième edition, 1993. ISBN 2-7298-9351-2. 237 pp. LCCN ???? In French.

Currie:1995:MLE

[Cur95] I. D. Currie. Maximum likelihood estimation and Mathematica. Applied Statistics, 44 (3):379–??, 1995. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cabrera:1992:IM

[CW92] J. Cabrera and A. Wilks. An interface from S to Mathematica. Mathematica Journal, 2(1): 66–74, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Chonacky:2005:R

[CW05a] N. Chonacky and D. Winch. 3Ms: a response. Computing in Science and Engineering, 7(5):7–9, September/October 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        5992/32219/01501733.pdf?isnumber=
                                        32219&prod=JNL&arnumber=1501733&
                                        arSt=+7&ared=+9&arAuthor=+
                                        Chonacky%2C+N.%3B++Winch%
                                        2C+D.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=32219&arnumber=1501733&
                                        count=14&index=1.

Chonacky:2005:IPM

[CW05b] N. Chonacky and D. Winch.

3Ms for instruction, Part 2: Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(4):14–23, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://ieeexplore.ieee.
org/iel5/5992/31456/01463131.
pdf?isnumber=31456&prod=JNL&
arnumber=1463131&arSt=+14&
ared=+23&arAuthor=Chonacky%
2C+N.%3B+Winch%2C+D.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=31456&
arnumber=1463131&count=14&
index=1.

Chonacky:2005:IRM

[CW05c] Norman Chonacky and David Winch. 3Ms for instruction: Reviews of Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7 (3):7–13, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://csdl.computer.
org/comp/mags/cs/2005/03/
c3007.pdf;   http://csdl.
computer.org/comp/mags/cs/
2005/03/c3007abs.htm.

Chonacky:2005:MMM

[CW05d] Norman Chonacky and David Winch. Maple, Mathematica, and Matlab: The 3M’s without the tape. Computing in Science and Engineering, 7(1):8–16, January/February 2005. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/cs/2005/01/c1008.pdf;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1008.
                                        htm.

Chonacky:2005:RMM

[CW05e] Norman Chonacky and David Winch. Reviews of Maple, Mathematica, and Matlab: Coming soon to a publication near you. Computing in Science and Engineering, 7(2):9–10, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/cs/2005/02/c2009.pdf;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2009.
                                        htm.

Clarke:1992:ATP

[CZ92] Edmund Clarke and Xudong Zhao. Analytica — a theorem prover in Mathematica. In Kapur [Kap92], pages 761– 765. ISBN 0-387-55602-8 (New York), 3-540-55602-8 (Heidelberg). LCCN QA76.9.A96 I57 1992.

Chu:2009:ISH

[CZ09] Wenchang Chu and Deyin Zheng. Infinite series with harmonic numbers and central binomial coefficients. International Journal of Number Theory (IJNT), 5(3):429– 448, May 2009. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.
worldscientific.com/doi/10.
1142/S1793042109002171.

deAlwis:1993:MPM

[dA93] T. de Alwis. Mathematica and the power method. International journal of mathematical education in science and technology, 24(6):813–824, November 1993. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Darrough:1994:BRE

[Dar94] Masako Darrough. Book review:

Economic and financial modeling with Mathematica, by Hal R. Varian. Journal of economic literature, 32(4):1868–1870, December 1994. CODEN JECLB3. ISSN 0022-0515 (print), 1547-1101 (electronic).

Dasgupta:1990:BMM

[Das90] G. Dasgupta. Boundary modulation with MATHEMATICA. Software for Engineering Workstations, 6(1):2–10, January 1990. CODEN SEWOEA. ISSN 0266-9463, 0952-8768.

Dauphine:2017:GMM

[Dau17] André.In Dauphiné. Geographical Models with Mathematica. Elsevier, Amsterdam, The Netherlands, 2017. ISBN 1-78548-225-4, 0-08-102230-1 (ebook). xvii + 296 pp. LCCN G70.23 .D374 2017. Diaz-Contreras:1996:GFA

[DCN96] R. E. Diaz-Contreras and S. Nomura. Green’s function applied to solution of Mindlin plates. Computers and structures, 60 (1):41–48, July 1996. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

DeAlwis:1994:MCM

[De 94] T. De Alwis. Mathematica as a conjecture making and a multimedia tool. In Ottmann and Tomek [OT94], pages 642–643. ISBN 1-880094-10-X. LCCN ????

DeJong:1999:MCB

[De 99] Marvin L. De Jong. Mathematica for calculus-based physics. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-60339-X. 257 pp. LCCN QC20.D33 1999.

Deppisch:2017:EMP

[Dep17] Thomas Deppisch. E6Tensors: a Mathematica package for E_{6} tensors. Computer Physics Communications, 213(??):130– 135, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0010465516302818.

Devlin:1992:AMC

[Dev92a] Keith Devlin. An article by M. C. Nucci precedes two reviews of the new version 2.0 of the Mathematica system. Notices of the American Mathematical Society, 39(5):427–437, May 1992. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1992:CM

[Dev92b] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 39(8):825–843, October 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMa

[Dev94a] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(5):455–??, May 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMc

[Dev94b] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(7):772–??, September 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunbar:1991:CTM

[DF91] S. Dunbar and D. Fowler. Cinematic thinking and Mathematica notebooks. Mathematica in Education, 1(2):3–8, Fall 1991. ISSN 1065-2965.

Dingle:1994:BCC

[DF94] Adam Dingle and Richard J.

Fateman. Branch cuts in computer algebra. In ACM [ACM94], pages 250–257. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/

proceedings/issac/190347/
p250-dingle/.

DeRosa:1996:ORM

[DF96] M. A. De Rosa and C. Franciosi. The optimized Rayleigh method and Mathematica in vibrations and buckling problems. Journal of Sound and Vibration, 191(5): 795–808, April 1996. CODEN JSVIAG. ISSN 0022-460X.

Dominguez:1994:MFA

[DFPA94] J. A. Dominguez, A. Fernandez, F. J. Plaza, and M. A. Asensio. Mathematica: fundamentos y aplicaciones de la informatica en matematicas. Plaza Universitaria Ediciones, ????, 1994. ISBN 84-89109-04-4. ???? pp. LCCN ????

Dominguez:1995:ALP

[DFPA95] J. A. Dominguez, A. Fernandez, F. J. Plaza, and M. A. Asensio. Algebra Lineal: Planteamiento y resolucion de problemas con Mathematica. Plaza Universitaria Ediciones, ????, 1995. ISBN 84-89109-06-0. ???? pp. LCCN ???? In Spanish.

Dudas:1992:AQM

[DHE92] M. Dudas, H. Hsieh, and W. Ermler. Application of quantum mechanical perturbation theory to molecular vibrationalrotational analysis. Mathematica Journal, 2(2):66–69, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dill:1991:CPT

[Dil91a] D. Dill. Creating packages with TEX/Mathematica. Mathematica Journal, 1(4):16–17, Spring 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dill:1991:ITM

[Dil91b]  D. Dill.  Interactive TEX/

Mathematica documents. Mathematica Journal, 1(3):104–107, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dill:1991:IMM

[Dil91c] Dan Dill. Implementing matrix mechanics in Mathematica: Determination of Clebsch–Gordan coefficients by matrix diagonalization. Computers in Physics, 5(5):616–625, 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Dikhaut:1991:PFN

[DK91] J. Dikhaut and T. Kaplan.

A program for finding Nash equilibria. Mathematica Journal, 1(4):87–93, Spring 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Durfee:1993:CCP

[DK93] A. Durfee and I. Kronenfeld, N., Munson, H., Roy, J., Westby. Counting critical points of real polynomials in two variables. American Mathematical Monthly, 100(3):255–271, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

DeGroot:1993:PCP

[DN93] R. A. De Groot and J. Nadrchal, editors. Physics computing ’92: proceedings of the 4th international conference, Prague, Czechoslovakia, August 24–28, 1992. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1245-3. LCCN QC19.2.I53 1992.

Doering:1995:SVC

[Doe95] E. R. Doering. Scientific visualization in the circuits curriculum: enhancing student insight. In Budny and Herrick [BH95], pages 2c6.13–16 vol.1. ISBN 0-7803-3022-6. LCCN ???? Two volumes. IEEE Catalog No. 95CH35867.

DominguezPerez:1998:MFA

[Dom98] Jose Angel Dominguez Perez. Mathematica: fundamentos y aplicaciones de la informatica en matematicas. Plaza Universitaria, Salamanca, Spain, 2a. edition, 1998. ISBN 84-89109-04-4. 157 pp. LCCN ????

Don:2001:SOT

[Don01] Eugene Don. Schaum’s outline of theory and problems of Mathematica. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-135719-X. viii + 342 pp. LCCN QA76.95 .D66 2001.

Don:2009:M

[Don09]  Eugene Don.  Mathematica.

Schaum’s outline series. Mc-Graw-Hill, New York, NY, USA, second edition, 2009. ISBN 0-07-160828-1. xii + 359 pp. LCCN QA76.95 .D66 2009.

Doty:1990:MAP

[Dot90] L. T. Doty. Mathematica with applications to plans. In IEEE [IEE90c], pages 143–145. CO-DEN RPLSDW. LCCN TL695 .I44 1990. IEEE catalog number 90CH2811-8.

Davis:1994:SCM

[DPU94a] B. Davis, H. Porta, and J. Uhl.

(SUN-S115) calculus and Mathematica: Addressing fundamental questions about technology. In Lum [Lum94], pages 305– 314. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Davis:1994:CMb

[DPU94b] Bill Davis, Horacio Porta, and J. J. (J. Jerry) Uhl. Calculus and Mathematica. Addison-Wesley, Reading, MA, USA, Windows version 1.0 edition, 1994. ISBN 0-201-92776-4. LCCN ???? 1 computer disk.

Davis:1994:CMA

[DPU94c] Bill Davis, Horacio Porta, and J. J. (J. Jerry) Uhl. Calculus and Mathematica: approximations: measuring nearness. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-58468-9. ix + 261 pp. LCCN QA 221 D3228 1994.

Davis:1994:CMc

[DPU94d] Bill Davis, Horacio Porta, and J. Jerry Uhl. Calculus & Mathematica. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-58461-1 (Windows version), 0-201-58153-1 (Macintosh/NeXT version). ???? pp. LCCN ????

Davis:1994:CMD

[DPU94e] Bill Davis, Horacio Porta, and J. Jerry Uhl. Calculus & Mathematica: Derivatives: Measuring Growth. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-58466-2. x + 388 pp. LCCN QA306 .D38 1994.

Davis:1994:CMV

[DPU94f] Bill Davis, Horacio Porta, and J. Jerry Uhl. Calculus & Mathematica: Vector Calculus: Measuring in Two and Three Dimensions. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-58469-7. xi + 541 pp. LCCN QA433.D379 1994.

Davis:1994:CMI

[DPU94g] Bill Davis, Horacio Porta, and Jerry Uhl. Calculus & Mathematica: Integrals: Measuring Accumulated Growth. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-58467-0. ix + 342 pp. LCCN QA311 .D38 1994.

Duran:2014:MTM

[DPV14] Antonio J. Durán, Mario Pérez, and Juan L. Varona. The misfortunes of a trio of mathematicians using computer algebra systems. Can we trust in them? Notices of the American Mathematical Society, 61(10):1249–1252, November

2014. URL http://www.ams.
org/notices/201410/rnoti-
p1249.pdf.

Dick:1997:ML

[DRS97] Samuel Dick, Alfred Riddle, and Douglas Stein. Mathematica in the Laboratory. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58137-0 (hardcover), 0-521-49906-2 (paperback). xviii + 320 pp. LCCN TS158.6.D53 1997. US$74.95 (hardcover), US$29.95 (paperback).

duToit:1993:ESA

[dT93] C. F. du Toit. Evaluation of some algorithms and programs for the computation of integerorder Bessel functions of the first and second kind with complex arguments. IEEE Antennas and Propagation Magazine, 35 (3):19–25, June 1993. CODEN IAPMEZ. ISSN 1045-9243.

Davis:2000:LAL

[DT00] H. Ted (Howard Ted) Davis and Kendall T. Thomson. Linear algebra and linear operators in engineering with applications in Mathematica, volume 3 of Process systems engineering. Academic Press, New York, NY, USA, 2000. ISBN 0-12-206349-X. xi + 547 pp. LCCN QA184.D38 2000.

Dubish:1990:TKN

[Dub90] R. Dubish. The tool kit: a notebook subclass. Mathematica Journal, 1(2):55–64, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dubin:2003:NAM

[Dub03] Daniel Dubin. Numerical and analytical methods for scientists and engineers using Mathematica. Wiley-Interscience, New York, NY, USA, 2003. ISBN 0-471-26610-8 (cloth). xvi + 633 pp. LCCN QC20.6 .D83 2003. Includes CD-ROM.

Durand:1998:PMR

[Dur98]  Robert Durand.  Problèmes

de mathématiques résolus avec Maple et Mathematica. Ellipses, Paris, France, 1998. ISBN 2-7298-4821-5. x + 180 pp. LCCN ????

Dvorak:1990:MPM

[Dvo90] P. Dvorak. Math package makes equations more visual. Machine Design, 62:119, March 22, 1990. CODEN MADEAP. ISSN 0024-9114.

Devlin:1992:CMC

[DW92] K. Devlin and S. X. Wei. Computers and Mathematica column. Notices of the American Mathematical Society, 39(5): 428–437, May/June 1992. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Denker:1998:ISR

[DWY98] Manfred Denker, W. A. Woyczynski, and Bernard Ycart. Introductory statistics and random phenomena: uncertainty, complexity, and chaotic behavior in engineering and science. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1998. ISBN 0-8176-4031-2. xxiv + 509 pp. LCCN QA276.12 .D45 1998. URL

http://www.birkhauser.com/
book/isbn/08176-4031-2. In-

cludes bibliographical references and index.

Evans:1994:RMM

[EBM94] B. Evans, R. Bamberger, and J. McClellan. Rules for multidimensional multirate structures. IEEE Trans. Signal Processing, 42(4):762–771, April 1994. CO-DEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Edelen:1991:ECV

[Ede91] D. Edelen. Exterior calculus for vectors and forms. Mathematica Journal, 1(3):92–95, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Edelman:1992:CPC

[Ede92] A. Edelman. The complete pivoting conjecture for Gaussian eliminations is false. Mathematica Journal, 2(2):58–61, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Ediger:1990:GNM

[Edi90] M. Ediger. A Gauss–Newton method for nonlinear regression. Mathematica Journal, 1(2):42– 44, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Eilbeck:1996:SAC

[EE96] J. C. Eilbeck and V. Z. Enol‘skii. Some applications of computer algebra to problems in theoretical physics. In Anonymous [Ano96a], pages 443–452. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Evans:1994:ISS

[EGM94] B. Evans, T. Gardos, and J. Mc-Clellan. Imposing structure on Smith form decompositions of rational resampling matrices. IEEE Trans. Signal Processing, 42(4):970–973, April 1994. CO-DEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Eames:1994:CAL

[EJ94] I. W. Eames and A. R. Johnson, editors. Computer aided learning in engineering: Conference — September 1994, Sheffield. University of Sheffield, Department of Mechanical and Process Engineering, Sheffield, UK, 1994. ISBN 1-899402-00-4. LCCN ????

Ekstrand:2011:LMP

[Eks11] Joel Ekstrand. Lambda: a Mathematica package for operator product expansions in vertex algebras. Computer Physics Communications, 182(2):409– 418, February 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465510003760.

Evans:1993:LSS

[EKWM93] Brian L. Evans, Lina J. Karam, Kevin A. West, and James H. McClellan. Learning signals and systems with Mathematica. IEEE transactions on education, 36(1):72–78, February 1993. CODEN IEEDAB. ISSN 0018-9359.

Ellis:1990:TIM

[EL90] Wade Ellis, Jr. and Ed Lodi. A Tutorial Introduction to Mathematica. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1990. ISBN 0-534-15588-X. xiii + 92 pp. LCCN QA76.95.E45 1990.

Ellis:1991:TIM

[EL91] Wade Ellis and Ed Lodi. A Tutorial Introduction to Mathematica. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-15588-X. xiii + 92 pp. LCCN QA76.95 .E45 1991.

Evans:1992:SAS

[EM92] B. Evans and J. McClellan.

Symbolic analysis of signals and systems. In Oppenheim and Nawab [ON92], pages 88–141. ISBN 0-13-880444-3. LCCN TK5102.5 .S959 1992. Chapter 3.

El-Misiery:1993:CWD

[EM93] A. E. M. El-Misiery. Calculating Wadati–Deguchi–Akutsu N = 3 knot polynomials. International Journal of Theoretical Physics, 32(5):713–725, May 1993. CO-DEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Enns:2001:NPM

[EM01] Richard H. Enns and George McGuire. Nonlinear physics with Mathematica for scientists and engineers. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2001. ISBN 0-8176-4223-4, 3-7643-4223-4. xiii + 691 pp. LCCN QC20.7.N6 E57 2001.

Evans:1992:SAL

[EMB92] B. Evans, J. McClellan, and R. Bamberger. A symbolic algebra for linear multidimensional multirate systems. In Anonymous [Ano92c], pages 387–393. LCCN Q 350 C66 1992. Two volumes.

El-Misiery:1996:ACJ

[EMEH96] A. E. M. El-Misiery and E.- S. M. El-Horbaty. An algorithm for calculating Jones polynomials. Applied Mathematics and Computation, 74(2-3): 249–259, February 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1990:STA

[EMM90a] B. L. Evans, J. H. McClellan, and W. B. McClure. Symbolic transforms with applications to signal processing. Mathematica Journal, 1(2):70–80, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Evans:1990:SZU

[EMM90b] Brian L. Evans, James H. Mc-Clellan, and Wallace B. Mc-Clure. Symbolic z-transforms using DSP knowledge bases. In IEEE [IEE90b], pages 1775– 1778. CODEN IPRODJ. ISSN 0736-7791. LCCN TK 7882 S65 I16 1990. IEEE catalog number 90CH2847-2.

Evans:1993:ISP

[EMT93] Brian L. Evans, James H. Mc-Clellan, and H. Joel Trussell. Investigating signal processing theory with Mathematica. In IEEE [IEE93], pages 12–15. CO-DEN IPRODJ. ISBN 0-7803-0948-0 (microfiche), 0-7803-0947-2 (casebound), 0-7803-0946-4 (softbound), 0-7803-0949-9 (CD ROM). ISSN 0736-7791. LCCN TK 7882 S65 I16 1993. IEEE catalog number 93CH3252-4.

Evans:1991:MET

[EMW91] Brian L. Evans, James H. Mc-Clellan, and Kevin A. West. Mathematica as an educational tool for signal processing. In IEEE [IEE91c], pages 1162– 1166. CODEN CPISDM. ISBN 0-7803-0033-5 (softbound), 0-7803-0034-3 (casebound), 0-7803-0035-1 (microfiche). ISSN 0734-7502. LCCN TK 7801 I117 1991. IEEE catalog number 91CH2998-3.

Emert:1992:CM

[EN92] John Emert and Roger Nelson.

Calculus & Mathematica. Saunders College Pub., Fort Worth, TX, USA, 1992. ISBN 0-03-076154-9. 64 pp. LCCN ????

Erkal:1996:TTS

[EN96] A. Erkal and D. T. Numbere. Tracking thermal saturation fronts by a high level PC programming language. Proceedings — Petroleum Computer Conference, pages 83–89, 1996. CODEN PPCMEG.

Edwards:1996:CPD

[EP96a] C. H. Edwards, Jr. and David E. Penney. Computing Projects for Differential Equations: Computing and Modeling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-504465-0. 304 pp. URL http:

                                        //www.prenhall.com/books/
                                        esm_0135044650.html.  Lab

supplement to [EP96b].

Edwards:1996:DEB

[EP96b] C. H. Edwards, Jr. and David E. Penney. Differential Equations and Boundary Value Problems: Computing and Modeling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-382094-7. 718 pp. LCCN QA371.E28 1996. URL http:

                                        //www.prenhall.com/books/
                                        esm_0133820947.html. See also

lab supplement [EP96a].

Edwards:2003:DEC

[EP03] C. H. (Charles Henry) Edwards and David E. Penney. Differential equations: computing and modeling. Pearson Education, Upper Saddle River, NJ, USA, third edition, 2003. ISBN 0-13-067337-4. various pp. LCCN QA371 .E29 2003.

Edwards:2004:EDE

[EP04] C. H. (Charles Henry) Edwards and David E. Penney. Elementary differential equations. Pearson Education, Upper Saddle River, NJ, USA, fifth edition, 2004. ISBN 0-13-145773-X. xi + 613 + 6 pp. LCCN QA371 .E33 2004b.

Erlebacher:1990:SMS

[Erl90] G. Erlebacher. Symbolic manipulation and supersonic transition. In Noor et al. [NEH90], pages 285–299. CODEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Erocal:2010:SPU

[ES10] Burçin Eröcal and William Stein. The Sage Project: Unifying free mathematical software to create a viable alternative to Magma, Maple, Mathematica and MATLAB. Lecture Notes in Computer Science, 6327:12–27, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-15582-6_
4.pdf.

Evans:1993:KED

[Eva93a] B. Evans. A Knowledge-Based Environment for the Design and Analysis of Multidimensional Multirate Signal Processing Algorithms. Ph.d. thesis, Georgia Institute of Technology, Atlanta, GA, USA, June 1993. 167 pp.

Evans:1993:XXA

[Eva93b] K. Evans, Jr. Xorbit — an X-Windows accelerator simulation. In Ryne [Ryn93], pages 450–458. ISBN 1-56396-222-5. LCCN QC787.L5 C66 1993.

Falco:1993:MPA

[Fal93]  Giulio Falco.  Mathematica:

Principi e applicazioni. Addison-Wesley, Reading, MA, USA, 1993. ISBN 88-7192-037-6. 119 pp. LCCN ???? In Italian.

Fateman:1990:ATD

[Fat90a] R. J. Fateman. Advances and trends in the design and construction of algebraic manipulation systems. In Watanabe and Nagata [WN90], pages 60–67. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Fateman:1990:LLM

[Fat90b] R. J. Fateman. A Lisp-language Mathematica-to-Lisp translator. SIGSAM Bulletin, 24(2):19–21, April 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1990:SCA

[Fat90c] R. J. Fateman. On the systematic construction of algebraic manipulation systems. In Noor et al. [NEH90], pages 3–13. CO-DEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Fateman:1992:RM

[Fat92]  R. J. Fateman.  A review

of Mathematica. Journal of Symbolic Computation, 13(5): 545–579, May 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fateman:1996:SMS

[Fat96] Richard J. Fateman. Symbolic mathematics system evaluators (extended abstract). In Lakshman Y. N. [Lak96], pages 86–94. ISBN 0-89791-796-0. LCCN QA 76.95 I59

1996.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/236869/
p86-fateman/.

Fateman:2015:PAS

[Fat15] Richard Fateman. Partitioning of algebraic subexpressions in computer algebra systems: an alternative to matching with an application to symbolic integration. ACM Communications in Computer Algebra, 49(2):38– 47, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Feagin:1994:QMM

[Fea94] James M. Feagin. Quantum Mechanics with Mathematica. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-97973-5 (new York), 3-540-97973-5 (Berlin). xix + 482 pp.

LCCN QC174.17.M35 F4 1994. US$54.95. With a foreword by Eugen Merzbacher. With 1 Macintosh/IBM-PC floppy disk (3.5 inch).

Feng:2012:AGM

[Fen12] Feng Feng. $apart : a generalized Mathematica Apart function. Computer Physics Communications, 183(10):2158– 2164, October 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465512001300.

Feng:2016:AGM

[Fen16] Feng Feng. APart 2: a generalized Mathematica Apart function. Computer Physics Communications, 198(??):260– 261, January 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465515003501.

Ferrari:2007:SMP

[Fer07]  A. F. Ferrari.  SusyMath:

a Mathematica package for quantum superfield calculations. Computer Physics Communications, 176(5):334–346, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465506003973.

Fritzson:1993:VHI

[FEV93] P. Fritzson, V. Engelson, and L. Viklund. Variant handling, inheritance and composition in the ObjectMath computer algebra environment. In Miola [Mio93], page ?? ISBN 3-540-57235-X (Berlin), 0-387-57235-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1993. “Sponsored by the Research Institute for Symbolic Computation (University J. Kepler, Linz, Austria) and by the Dipartimento di informatica e sistemistica (University ‘La Sapienza’, Roma, Italy)”– Foreword.

Faught:1993:CM

[FF93] Leslie Faught and Ginger Freeze.

Calendar Mathematica. Arithmetic Teacher, 41(3):154–??, November 1993. CODEN AR-ITBF. ISSN 0004-136X.

Fritzson:1994:OOM

[FFVH94] D. Fritzson, P. Fritzson, L. Viklund, and J. Herber. Objectoriented mathematical modelling — applied to machine elements. Computers and structures, 51(3):241–253, May 1994. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Fateman:1996:SLB

[FH96] R. J. Fateman and M. Hayden.

Speeding up Lisp-based symbolic mathematics. SIGSAM Bulletin, 30(1):25–30, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Feger:2015:LMA

[FK15] Robert Feger and Thomas W. Kephart. LieART — a Mathematica application for Lie algebras and representation theory. Computer Physics Communications, 192(??):166–195, July 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465515000119.

Fournier:1995:VMS

[FKM95] Robert Fournier, Norbert Kajler, and Bernard Mourrain. Visualization of mathematical surfaces: the IZIC server approach. Journal of Symbolic Computation, 19(1/2/3):159– 174 (or 159–173??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Feger:2020:LMA

[FKS20] Robert Feger, Thomas W. Kephart, and Robert J. Saskowski. LieART 2.0 — a Mathematica application for Lie algebras and representation theory. Computer Physics Communications, 257 (??):Article 107490, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465520302290.

Finch:1992:EMM

[FL92a] J. Finch and M. Lehmann. Exploring Mathematics with Mathematica. Addison-Wesley, Reading, MA, USA, 1992. ISBN ???? 341 pp. LCCN ???? See [FL94].

Finch:1992:ECM

[FL92b] James K. Finch and Millianne Lehmann. Exploring Calculus with Mathematica. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55572-7. x + 341 pp. LCCN QA303.5.D37F56 1992.

Finch:1994:EMM

[FL94] J. Finch and M. Lehmann. Exploring Mathematics with Mathematica. Addison-Wesley, Reading, MA, USA, 1994. ISBN 4-8101-8045-X. ???? pp. LCCN ???? Japanese translation of [FL92a].

Fukuzawa:1993:CAD

[FM93] Mitsunori Fukuzawa and Masami Masubuchi. Computer-aided design system for linearization and control of nonlinear systems using ‘Mathematica’. Nippon Kikai Gakkai Ronbunshu, C Hen/Transactions of the Japan Society of Mechanical Engineers, Part C, 59(563):178–185, July 1993. CODEN NKCHDB. ISSN 0387-5024.

Feagin:2002:QMM

[FM02] James M. Feagin and Eugen Merzbacher. Quantum methods with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 2002. ISBN 0-387-95365-5 (paperback). xix + 482 pp. LCCN QC174.17.M35 F4 2002. Springer study edition.

Favata:2016:ABM

[FMRP16] Antonino Favata, Andrea Micheletti, Seunghwa Ryu, and Nicola M. Pugno. An analytical benchmark and a Mathematica program for MD codes: Testing LAMMPS on the 2nd generation Brenner potential. Computer Physics Communications, 207(??):426–431, October 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465516301679.

Foley:2002:ICE

[Fol02] Henry C. Foley. Introduction to chemical engineering analysis using Mathematica. Academic Press, New York, NY, USA, 2002. ISBN 0-12-261912-9. xxvii + 509 pp. LCCN QD75.22 .F65 2002. Accompanying CD-ROM contains a nearly identical color version of the book.

Fonseca:2021:GMP

[Fon21] Renato M. Fonseca. GroupMath: a Mathematica package for group theory calculations. Computer Physics Communications, 267(??):Article 108085, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465521001971.

Foster:1991:MU

[Fos91a] K. Foster. Mathematica update. IEEE Spectrum, 28 (??):73, November 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1991:PMO

[Fos91b] K. Foster. Prepackaged math.

(one of seven articles). IEEE Spectrum, 28(11):44–47, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1991:MUM

[Fos91c] Kenneth R. Foster. Mathematics^{T M} 2.0: An updated math program for MS-DOS platforms. Computers in Physics, 5(6):643–644, November/December 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4823034.

Foster:1991:PM

[Fos91d] Kenneth R. Foster. Prepackaged math. IEEE Spectrum, 28(11): 44–50, November 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1992:ESM

[Fos92] K. Foster. Engineering softwaremath and graphics. IEEE Spectrum, 29(11):72, 74, 76, 78, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1993:RSM

[Fos93] K. Foster. Report on symbolic mathematics software tools for engineering analysis and design. IEEE Spectrum, 30(11):42–46, 54, 57–79, November 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:MI

[Fos99] Kenneth R. Foster. Math on the Internet. IEEE Spectrum, 36(4):36–40, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:CMS

[Fos01]  Kenneth R. Foster.  Con-

trolling Mathematica [software]. IEEE Spectrum, 38(5):94–95, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2011:MMT

[Fos11] Kenneth R. Foster. Mathematica 8 and Maple 15 [tools toys]. IEEE Spectrum, 48(12):23, December 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fowler:1996:SPM

[Fow96] D. Fowler. Symbolic power in mathematics education. In Robin et al. [RPWW96], pages 226–228.

Fatic:1990:DVS

[FP90] V. Fatic and Y. Park. Direct variational solutions of the Hamiltonian state model of linear RLC circuits. part I. overdamped case. In Vogt and Mickle [VM90], pages 1855– 1863. ISBN 1-55617-271-0. LCCN TA 343 M62 1990. Five volumes.

Fitzgerald:1995:TEQ

[FP95] S. Fitzgerald and J. Place.

Teaching elementary queueing theory with a computer algebra system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):350–354, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freeman:1992:NNM

[Fre92] J. Freeman. Neural networks in Mathematica. AI Expert, 7:26– 35, November 1992. ISSN 0888-3785.

Freeman:1994:AIF

[Fre94a] J. Freeman. Artificial intelligence: Fuzzy systems for control applications. Mathematica Journal, 4(1):64–69, Winter 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Freeman:1994:ENN

[Fre94b] James A. Freeman. Exploring Neural Networks with Mathematica. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-56629-X. x + 341 pp. LCCN QA76.87 .F724 1994.

Friedman:1994:IME

[Fri94] Ben Friedman. Introduction to Mathematica in engineering. In IEEE [IEE94b], pages 180–?? CODEN WCREDI. ISBN 0-7803-9993-5, 0-7803-9992-7. ISSN 1044-6036, 0083-8837. LCCN TK 7801 W47 1994.

Falloon:2017:QMP

[FRW17] Peter E. Falloon, Jeremy Rodriguez, and Jingbo B. Wang. QSWalk: a Mathematica package for quantum stochastic walks on arbitrary graphs. Computer Physics Communications, 217(??):162–170, August 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465517301029.

Fernandez-FerreirosErviti:1995:ALP

[FSC95a] Ana Fernandez-Ferreiros Erviti, Ma. Luisa Sein-Echaluce Lacleta, and Vicente Camarena. Algebra lineal: practicas con Mathematica. Number 43 in Coleccion Textos Docentes. Prensa Universitarias de Zaragoza, Zaragoza, Spain, 1995. ISBN 84-7733-452-8. 178 pp. LCCN ????

Fernandez-FerreirosErviti:1995:CPC

[FSC95b] Ana Fernandez-Ferreiros Erviti, Ma. Luisa Sein-Echaluce Lacleta, and Vicente Camarena. Calculo: practicas con Mathematica. Number 42 in Coleccion Textos Docentes. Prensa Universitarias de Zaragoza, Zaragoza, Spain, 1995. ISBN 84-7733-451-X. 216 pp. LCCN ????

Fine:1991:SPS

[FW91] J. Fine and L. Waite. Spatial problem solving in an integrated first-year curriculum. Mathematica Journal, 1(4):62– 67, Spring 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Gallant:1991:SSC

[Gal91] J. Gallant. Software smooths complex computations; over- $400 math software packages. (technology update) (buyers guide). EDN, 36:75–77, November 7, 1991. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Galbraith:1998:PRM

[Gal98] Patrick Galbraith. Product review: Mathematica version 3.0 for Linux. Linux Journal, 56:75– 77, December 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Gloor:1994:IMVb

[GAM94a] Oliver Gloor, Beatrice Amrhein, and Roman Maeder. Illustrierte Mathematik: Visualisierung von mathematischen Gegenständen. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 3-7643-5100-4. ???? pp. LCCN ????

Gloor:1994:IMVa

[GAM94b] Oliver Gloor, Beatrice Amrhein, and Roman E. Maeder. Illustrated mathematics: Visualizations for classroom use. IFIP Transactions. A. Computer Science and Technology, A-48:51– 65, 1994. CODEN ITATEC. ISSN 0926-5473.

Gloor:1995:IMC

[GAM95] Oliver Gloor, Beatrice Amrhein, and Roman Maeder. Illustrated Mathematics (CD-ROM). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-14222-3. 64 pp. LCCN ???? US$69.95.

Gass:1998:MSE

[Gas98] Richard Gass. Mathematica for scientists and engineers: using Mathematica to do science. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-227612-7 (paperback). xii + 498 pp. LCCN QA76.95.G37 1998.

Gray:2006:MDG

[GAS06] Alfred Gray, Elsa Abbena, and Simon Salamon. Modern differential geometry of curves and surfaces with Mathematica. Studies in advanced mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2006. ISBN 1-58488-448-7. 984 pp. LCCN QA641 .G72 2006.

Gaylord:1992:CCS

[Gay92a] R. Gaylord. Catastrophes in complex systems. Mathematica in Education, 2(1):19–22, Fall 1992. ISSN 1065-2965.

Gaylord:1992:DLA

[Gay92b] R. Gaylord. Diffusion-limited aggregation. Mathematica in Education, 1(3):6–10, Spring 1992. ISSN 1065-2965.

Gaylord:1992:SP

[Gay92c] R. Gaylord. Spreading phenomena. Mathematica in Education, 1(4):17–19, Summer 1992. ISSN 1065-2965.

Gaylord:1998:SSM

[GD98] Richard J. Gaylord and Louis J.

D’Andria. Simulating society: a Mathematica toolkit for modeling socioeconomic behavior. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98532-8 (paperback). xxiii + 218 pp. LCCN HM291 .G375 1998.

Gorodkov:1996:ASD

[GDD^{+}96] A. Gorodkov, N. B. Dobrova, J.-Ph. Dubernard, G. I. Kiknadze, I. A. Gatchetchiladze, V. G. Oleinikov, N. B. Kuzmina, J. L. Barat, and Ch. Baquey. Anatomical structures determining blood flow in the heart left ventricle. Journal of Materials Science: Materials in Medicine, 7(3):153–160, March 1996. CODEN JSMMEL. ISSN 0957-4530.

Gear:1991:RFM

[Gea91] J. A. (John Anthony) Gear.

Root finding with Mathematica. Research report 20, Royal Melbourne Institute of Technology, Dept. of Mathematics, Melbourne, Victoria, Australia, 1991. ISBN 0-86444-231-9. 8 pp.

Green:1994:ECM

[GEJHH94] Edward Green, Benny Evans, Jerry Johnson, and Deborah Hughes-Hallett. Exploring calculus with Mathematica: to accompany Calculus, Deborah Hughes-Hallett, Andrew M. Gleason et al. John Wiley and Sons, Inc., Chichester, UK, 1994. ISBN 0-471-09718-7. ix + 322 pp. LCCN ????

Gray:1991:EMM

[GG91] Theodore W. Gray and Jerry Glynn. Exploring Mathematics with Mathematica: Dialogs Concerning Computers and Mathematics. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52809-6 (hardcover), 0-201-52818-5 (paperback). 535 + 16 pp. LCCN QA76.95 .G73 1991. US$34.95. Includes CD-ROM. See translation [GG94].

Gray:1992:BGM

[GG92] Theodore W. Gray and Jerry Glynn. The Beginner’s Guide to Mathematica, Version 2. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-58221-X. ix + 225 pp. LCCN QA76.95 .G72 1992. US$18.95. See [GG93, GGxx].

Gray:1993:GIM

[GG93] Theodore W. Gray and Jerry Glynn. Guide d’initiation à Mathematica 2. Addison-Wes-ley, France, Paris, France, 1993.

ISBN 2-87908-026-6. vii + 262 pp. LCCN ???? French translation of The Beginner’s Guide to Mathematica Version 2, [GG92].

Gray:1994:EMM

[GG94] Theodore W. Gray and Jerry Glynn. Exploring Mathematics with Mathematica. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 1994. ISBN 4-8101-8045-X. 580 pp. LCCN ???? Japanese translation of [GG91].

Glynn:1997:BGM

[GG97] Jerry Glynn and Theodore W.

Gray. The beginner’s guide to Mathematica version 3. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-62202-6 (hardcover), 0-521-62734-6 (paperback). viii + 347 pp. LCCN QA76.95 .G55 1997.

Gray:19xx:BGM

[GGxx] Theodore W. Gray and Jerry Glynn. The Beginner’s Guide to Mathematica Version 2. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 19xx. ISBN 4-8101-8056-5. ???? pp. LCCN ???? Japanese translation of [GG92].

Glynn:2000:BGM

[GG00] Jerry Glynn and Theodore W.

Gray. The beginner’s guide to Mathematica version 4. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77153-6 (hardcover), 0-521-77769-0 (paperback). viii + 434 pp. LCCN QA76.95 .G552 2000.

URL http://store.wolfram.
com/view/ISBN0521771536.

Guenther:1996:HAC

[GGK96] R. B. Guenther, J. A. Gottsch, and D. B. Kramer. The Herglotz algorithm for constructing canonical transformations. SIAM Review, 38(2):287–293, June 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/p287.htm.

Gantes:1994:USM

[GGV94] C. Gantes, A. Giakoumakis, and P. Vousvounis. Using symbolic manipulation for the geometric design of deployable domes shaped as platonic or Archimedean polyhedra. International Conference on Computational Structures Technology — Proceedings, pages 171–185, 1994.

Gidh:1993:TTA

[GH93] K. Gidh and M. E. Hanyak, Jr. A thermodynamics teaching aid for undergraduate engineers. International Journal of Engineering Education, 9(2):162–185, 1993. CODEN IEEDEF. ISSN 0742-0269.

Getz:2004:GMF

[GH04] Chonat Getz and Janet Helmstedt. Graphics with Mathematica: fractals, Julia sets, patterns and natural forms. Elsevier, Amsterdam, The Netherlands, 2004. ISBN 0-444-51760-X (hardcover), 0-444-51774-X (CD-ROM). vii + 322 pp. LCCN QA76.95 .G48 2004. Includes CD-ROM.

Gray:2009:SMP

[GHIL09] James Gray, Yang-Hui He, Anton Ilderton, and André Lukas. STRINGVACUA. A Mathematica package for studying vacuum configurations in string phenomenology. Computer Physics Communications, 180 (1):107–119, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465508002841.

Gutierrez:1998:MPA

[GI98] José Manuel Gutiérrez and Andrés Iglesias. Mathematica package for analysis and control of chaos in nonlinear systems. Computers in Physics, 12 (6):608–??, November 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.168743.

Gilliot:1995:OTI

[Gil95] J. M. Gilliot. Optimization of trajectories for industrial robots using Mathematica. In Johannsen [Joh95], pages 305–310. ISBN 0-08-042361-2. LCCN T55.45 .I5 1995.

Gray:1992:UMF

[GK92] A. Gray and R. Knill. Using Mathematica to find closed form expressions for approximations to the square root of x. Mathematica in Education, 1(4):12–13, Summer 1992. ISSN 1065-2965.

Goktas:2011:MMS

[GK11] Ünal Göktaş and Devendra Kapadia. Methods in Mathematica for solving ordinary differential equations. Mathematical and Computational Applications, 16(4):784–796, December 2011. CODEN ???? ISSN 2297-

8747. URL https://www.mdpi.
com/2297-8747/16/4/784.

Gluza:2007:AMP

[GKR07] J. Gluza, K. Kajda, and T. Riemann. AMBRE — a Mathematica package for the construction of Mellin–Barnes representations for Feynman integrals. Computer Physics Communications, 177(11):879– 893, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465507003505.

Gaylord:1993:IPM

[GKW93] Richard J. Gaylord, Samuel N. Kamin, and Paul R. Wellin. Introduction to Programming with Mathematica. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-94048-0 (Berlin), 0-387-94048-0 (New York). xix + 302 pp. LCCN QA76.73.M29 G39 1993. US$39.95. See [GKW94].

Gaylord:1994:IPM

[GKW94] Richard J. Gaylord, Samuel N.

Kamin, and Paul R. Wellin. Introduction to Programming with Mathematica. Kindai Kagaku, ??, Japan, 1994. ISBN 4-7649-0228-1. 304 pp. LCCN ???? Japanese translation of [GKW93].

Gaylord:1996:IPM

[GKW96] Richard J. Gaylord, Samuel N.

Kamin, and Paul R. Wellin. An introduction to programming with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, second edition, 1996. ISBN 0-387-94434-6. xix + 452 pp. LCCN QA76.73.M29G39 1996.

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN QA155.7.E4 C64954 2003. URL http:/

/www.springer.com/sgw/cda/
frontpage/0,11855,1-102-22-
1477871-0,00.html. Includes

CD-ROM.

Gladd:1990:UMA

[Gla90] N. T. Gladd. Using MATHE-MATICA to assist plasma stability analysis. In IEEE International Conference on Plasma Science, pages 112–?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. IEEE catalog number 90CH2857-1.

Glidden:1994:MME

[Gli94] P. L. Glidden. Mathematicabased mathematics for elementary teachers: an experimental course. In Anonymous [Ano94i], pages 529–533.

Glidden:1995:MBM

[Gli95] P. L. Glidden. Mathematicabased mathematics for elementary teachers: the role of student attitudes and beliefs in an experimental course. Journal of Technology Teacher Education, 3(2-3):189–204, 1995. CODEN JT-EDFU.

Gomez-Lobo:2012:SMP

[GLMG12] Alfonso García-Parrado Gómez-Lobo and José M. Martín-García. Spinors : a Mathematica package for doing spinor calculus in General Relativity. Computer Physics Communications, 183 (10):2214–2225, October 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465512001634.

Gilbert:1993:UMT

[GM93] R. C. Gilbert and J. Mathews. Using Mathematica to teach Bessel functions. International journal of mathematical education in science and technology, 24(1):45–54, January 1993. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Garrett:1995:UME

[GMM95] C. A. Garrett and R. Malek-Madani. Using Mathematica to enhance learning of oceanographic processes: Breaking of waves and Burgers’ equation. In Anonymous [Ano95d], pages 50– 53. ISBN ???? LCCN ????

Gray:1997:IOD

[GMP97] Alfred Gray, Michael Mezzino, and Mark A. Pinsky. Introduction to ordinary differential equations with Mathematica: an integrated multimedia approach. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1997. ISBN 0-387-94481-8 (hardcover). xxii + 890 pp. LCCN QA 371.5 D37 G74 1997.

Gomes:2009:BBA

[GMP09] Abel J. P. Gomes, José F. M.

Morgado, and Edgar S. Pereira. A BSP-based algorithm for dimensionally nonhomogeneous planar implicit curves with topological guarantees. ACM Transactions on Graphics, 28(2):17:1– 17:24, April 2009. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gaylord:1996:MNC

[GN96] Richard J. Gaylord and Kazume Nishidate. Modeling nature with cellular automata using Mathematica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94620-9. ???? pp. LCCN QA267.5.C45G39 1996.

Gray:1993:MDG

[Gra93] Alfred Gray. Modern differential geometry of curves and surfaces. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-7872-9. xviii + 664 pp. LCCN QA641.G73 1993. See [Gra94a].

Gray:1994:DKT

[Gra94a]  Alfred Gray.  Differential-

geometrie: Klassische Theorie in moderner Darstellung. Spektrum Akademischer Verlag, ????, 1994. ISBN 3-86025-141-4. 618 pp. LCCN ???? German translation of [Gra93].

Gray:1994:MM

[Gra94b]  John W. Gray.  Mastering

Mathematica. Academic Press, New York, NY, USA, 1994. ISBN 0-12-296040-8. xx + 644 pp. LCCN QA76.95 .G68 1994. US$44.95. Includes computer disk.

Gray:1994:MMP

[Gra94c] John W. (John Walker) Gray. Mastering Mathematica: Programming Methods and Applications. Academic Press, New York, NY, USA, 1994. ISBN 0-12-296040-8. xx + 644 pp. LCCN QA76.95 .G68 199.

Grabe:1996:APS

[Grä96] Hans-Gert Gräbe. About the polynomial system solve facility of Axiom, Macsyma, Maple, Mathematica, MuPAD, and Reduce. Technical report 96-11, Institut für Informatik, Universität Leipzig, Leipzig, Germany, 1996. 28 pp.

Gray:1997:MDG

[Gra97a]  Alfred Gray.  Modern Dif-

ferential Geometry of Curves and Surfaces with Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 1997. ISBN 0-8493-7164-3. 1056 pp. LCCN QA641.G72 1997. US$79.95. URL http://

www.crcpress.com/;   http:
//www.crcpress.com/PRODS/
7164.HTM.

Gray:1997:MMP

[Gra97b] John W. Gray. Mastering Mathematica: Programming Methods and Applications. Academic Press, New York, NY, USA, second edition, 1997. ISBN 0-12-296105-6. xx + 629 pp. LCCN QA76.95.G68 1998. US$44.95. Includes computer disk.

Gray:1998:MDG

[Gra98a] Alfred Gray. Modern differential geometry of curves and surfaces with Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 1998. ISBN 0-8493-7164-3. xxiv + 1053 pp. LCCN QA641.G72 1998.

Gray:1998:MMP

[Gra98b] John W. (John Walker) Gray. Mastering Mathematica: programming methods and applications. Academic Press, New York, NY, USA, second edition, 1998. ISBN 0-12-296105-6 (paperback). xx + 629 pp. LCCN QA76.95 .G68 1998.

Graf:2004:ALT

[Gra04] Urs Graf. Applied Laplace transforms and z-transforms for scientists and engineers: a computational approach using a Mathematica package. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2004. ISBN 3-7643-2427-9, 0-8176-2427-9. x + 500 pp. LCCN QA432 .G74 2004. Includes CD-ROM.

Gregory:2005:BLD

[Gre05] P. C. (Philip Christopher) Gregory. Bayesian logical data analysis for the physical sciences: a comparative approach with Mathematica support. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-84150-X. xvii + 468 pp. LCCN QA279.5 .G74 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/cam041/2004045930.
html;   http://www.loc.gov/
catdir/toc/cam041/2004045930.
html.

Grodzicky:2001:CPW

[Gro01] Roman Grodzicky. A computer program written in Mathematica for calculating H_{2} quasicrystals and their diffraction patterns. Computer Physics Communications, 136(3):236– 249, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465500002484.

Guntert:1993:PCM

[GSOW93] Peter Guntert, Niklaus Schaefer, Gottfried Otting, and Kurt Wuthrich. POMA: a complete Mathematica implementation of the NMR product-operator formalism. Journal of magnetic resonance. Series A, 101(1):103– 105, January 1993. CODEN JM-RAE2. ISSN 1064-1858.

Gronlund:1990:IGM

[GSR90] S. D. Gronlund, Ching-Fan Sheu, and R. Ratcliff. Implementation of global memory models with software that does symbolic computation. Behavior research methods, instruments, and computers, 22(2):228–235, April 1990. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Gruber:1994:PPC

[GT94] Ralf Gruber and Marco Tomassini, editors. PC ’94: physics computing ’94: proceedings of the 6th Joint EPS-APS International Conference on Physics Computing, Palazzo dei Congressi, Lugano, Switzerland, 22– 26 Aug. 1994. European Physical Society, Petit-Lancy, France (??), 1994. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Guenter:2007:ESD

[Gue07] Brian Guenter. Efficient symbolic differentiation for graphics applications. ACM Transactions on Graphics, 26(3):108:1– 108:??, July 2007. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guglielmi:1995:ICS

[Gug95] M. Guglielmi, editor. IFAC Conference on System Structure and Control. Univ. Nantes, Nantes, France, 1995.

Guiasu:2006:USC

[Gui06] Silviu Guiasu. Using symbolic computing in building probabilistic models for atoms. International Journal of Quantum Chemistry, 106(1):4–26, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1991:UMM

[Gup91] S. K. Gupta. Using Mathematica in materials science. In Anonymous [Ano91c], pages 1927–1929 vol.2. Two volumes.

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-827-X. LCCN ????

Ganzha:1996:NSP

[GV96] V. G. (Victor Grigorevich)

Ganzha and E. V. (Evgenii Vasilevich) Vorozhtsov. Numerical solutions for partial differential equations: problem solving using Mathematica. Symbolic and numeric computation series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-7379-4. ???? pp. LCCN QA377.G2345 1996.

Garnerin:1991:VTR

[GVV91] P. Garnerin, A. Valleron, and J. Vidal. A visualization tool for representing epidemic spread. Mathematica Journal, 1(4):59– 61, Spring 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Gaylord:1995:CSMa

[GW95] Richard J. Gaylord and Paul R.

Wellin. Computer Simulations with Mathematica: Explorations in the Physical, Biological, and Social Sciences. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94274-2, 3-540-94274-2. xix + 297 pp. LCCN QA76.9.C65 G39 1994. US$54.95. Includes CD ROM.

Gaylord:1995:CSMb

[GWWJS95] Richard J. Gaylord, Paul R.

Wellin, Tom Wickham-Jones, and Alec J. Schramm. Computer simulations with Mathematica: Explorations in complex physical and biological systems and Mathematica graphics: Techniques and applications. Physics Today, 48(11):77–??, November 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/link.aip.org/link/phtoad/
v48/i11/p88/s1.

Hahn:2008:MIF

[Hah08]  T. Hahn.  A Mathemat-

ica interface for FormCalcgenerated code. Computer Physics Communications, 178 (3):217–221, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465507004201.

Hamza:1994:PTI

[Ham94] M. H. Hamza, editor. Proceedings of the Twelfth IASTED International Conference Applied Informatics. IASTED, Anaheim, CA, USA, 1994. ISBN 0-88986-190-0. LCCN ????

Haneberg:2004:CGM

[Han04] William C. Haneberg. Computational geosciences with Mathematica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40245-4. 381 pp. LCCN QE33.2.M3 H36 2004. Includes CD-ROM. System requirements for accompanying disc: Windows 98/ME/NT 4.0/2000/XP, Macintosh OS X 10.2 or 10.3 (or later), Macintosh 8.1 or later and a variety of Linux and Unix platforms Accompanying disc contains notebooks of all text and graphics plus an appendix on color graphics and specialized functions.

Harris:1994:IRR

[Har94] Jason F. Harris. Inheritance of rewrite rule structures applied to symbolic computation. In ACM [ACM94], pages 318–323. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/190347/
p318-harris/.

Harris:2000:ANM

[Har00] Jason Harris. Advanced notations in Mathematica. In Traverso [Tra00], pages 153– 160. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/
pubs/contents/proceedings/
issac/345542/. ACM order

number 505000.

Harris:2002:IMM

[Har02] Jason Harris. Implementing MathML in Mathematica. In Anonymous [Ano02], page ?? ISBN ???? LCCN ????

Hartle:2003:GIE

[Har03] J. B. (James B.) Hartle. Gravity: an introduction to Einstein’s general relativity. Addison-Wes-ley, Reading, MA, USA, 2003.

ISBN 0-8053-8662-9. xxii + 582 pp. LCCN QC173.6 .H38

2003. URL http://wps.aw.
com/aw_hartle_gravity_1/0,
6533,512494-,00.html.

Harris:2014:MMU

[Har14] Frank E. Harris. Mathematics for physical science and engineering: symbolic computing applications in Maple and Mathematica. Academic Press, New York, NY, USA, 2014. ISBN 0-12-801000-2 (hardcover), 0-12-801049-5 (e-book). xiv + 768 pp. LCCN QA37.3 .H36 2014. URL http://www.

sciencedirect.com/science/
book/9780128010006.

Harker:2015:NSI

[Har15]  A. H. Harker.  Nonlin-

ear Science: An Interactive Mathematica^{T M} Notebook, by David K. Campbell, Sebastian M. Marotta and Thomas A. Tanury, Scope: guide. Level: undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 56(2):236– 237, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2019:BRG

[Har19] A. H. Harker. Book review: Group theory in solid state physics and photonics: problem solving with Mathematica. Contemporary Physics, 60(1):108– 109, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hastings:2001:IPM

[Has01] Kevin J. Hastings. Introduction to probability with Mathematica. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-109-7. 380 pp. LCCN QA273.19.E4 H37 2001.

Hassani:2003:MMU

[Has03] Sadri Hassani. Mathematical methods using Mathematica: for students of physics and related fields. Undergraduate texts in contemporary physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95523-2 (softcover). xv + 235 pp. LCCN QC20 .H393 2003 UCB. Includes CD-ROM.

Hastings:2006:IMO

[Has06] Kevin J. Hastings. Introduction to the mathematics of operations research with Mathematica. Monographs and textbooks in pure and applied mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2006. ISBN 1-57444-612-6. xix + 567 pp. LCCN T57.6 .H385 2006.

Hastings:2010:IPM

[Has10] Kevin J. Hastings. Introduction to probability with Mathematica. Textbooks in mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2010. ISBN 1-4200-7938-7 (hardcover). 451 pp. LCCN QA273.19.E4 H37 2010.

Hayes:1992:PPR

[Hay92a] A. Hayes. Performance of programs revisited. Mathematica Journal, 2(2):50–53, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Hayes:1992:SCD

[Hay92b] A. Hayes. Sums of cubes of digits — driven to abstraction. Mathematica in Education, 1(4):3–11, Summer 1992. ISSN 1065-2965.

Hayes:1993:MSD

[Hay93] Allan Hayes. Mathematica: a system for doing mathematics by computer. Mathematical Gazette, 77(478):52–70, March 1993. CODEN MAGAAS. ISSN 0025-5572.

Hayes:1994:LG

[Hay94a] A. Hayes. A little geometry. Mathematica in Education, 3(1): 12–19, 1994. ISSN 1065-2965.

Hayes:1994:MEE

[Hay94b] A. Hayes. MATHEMATICA in engineering education. In Eames and Johnson [EJ94], pages 109– 124. ISBN 1-899402-00-4. LCCN ????

Hazrat:2010:MPC

[Haz10] Roozbeh Hazrat. Mathematica: a problem-centered approach, volume 53 of Springer undergraduate mathematics series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84996-250-2 (softcover). xiv + 188 pp. LCCN ????

Haager:1995:ADP

[HBN95] G. Haager, G. Baumann, and T. F. Nonnenmacher. An algorithm to determine potential systems in Mathematica. Journal of Symbolic Computation, 20 (2):179–196, August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Huang:1997:MME

[HC97] Cliff J. Huang and Philip Crooke. Mathematics and Mathematica for economists. Blackwell, Malden, MA, 1997. ISBN 1-57718-034-8. xiv + 674 pp. LCCN HB143.5 .H8 1997. Accompanying computer disk contains a package of functions called MathEcon.

Hunt:1991:INS

[HD91] L. R. Hunt and R. D. De-Groat. Identifying nonlinear systems from experimental data. In IEEE [IEE91b], pages 3517– 3520. ISBN 0-7803-0003-3 (softbound), 0-7803-0004-1 (casebound), 0-7803-0005-X (microfiche). LCCN TK 7882 S65 I16 1991. Five volumes.

Hench:1994:SMA

[Hen94] D. L. Hench. Software for mathematical analysis. WESCON Conference Record, pages 258– 263, 1994. CODEN WCREDI. ISSN 1044-6036, 0083-8837.

Hensley:1996:PTA

[Hen96] D. Hensley. A polynomial time algorithm for the Hausdorff dimension of continued fraction Cantor sets. Journal of Number Theory, 58(1):9–45, May 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Herman:1988:MR

[Her88] Eugene A. Herman. Mathematica — a review. Notices of the American Mathematical Society, 35:1334–1344, November 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hata:1994:TAS

[HFO94] Toshiaki Hata, Masashi Furumaki, and Kazuhumi Ohenoki. Transient analysis of stress wave penetration in a plate subjected to a stress pulse (application of symbolic manipulation system to ray theory). Nippon Kikai Gakkai Ronbunshu, A Hen/Transactions of the Japan Society of Mechanical Engineers, Part A, 60(576):1800– 1806, August 1994. CODEN NKGADA. ISSN 0387-5008.

Hoft:1998:CM

[HH98] Hartmut F. W. Hoft and Margret Hoft. Computing with Mathematica. Academic Press, New York, NY, USA, 1998. ISBN 0-12-351660-9 (paperback). xiii + 290 pp. LCCN QA 76.95 H64 1998.

Hoft:2003:CM

[HH03] Hartmut F. W. Hoft and Margret Hoft. Computing with Mathematica. Academic Press, New York, NY, USA, second edition, 2003. ISBN 0-12-351666-8. xviii + 313 pp. LCCN QA76.95 .H64 2003. Includes CD-ROM.

Husa:2006:KMP

[HHL06] Sascha Husa, Ian Hinder, and Christiane Lechner. Kranc: a Mathematica package to generate numerical codes for tensorial evolution equations. Computer Physics Communications, 174 (12):983–1004, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465506001020.

Hoschele:2014:MMP

[HHP^{+}14] Maik Höschele, Jens Hoff, Alexey Pak, Matthias Steinhauser, and Takahiro Ueda. MT: a Mathematica package to compute convolutions. Computer Physics Communications, 185(2):528–539, February 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465513003354.

Henzinger:1995:HNG

[HHWT95] T. A. Henzinger, Pei-Hsin Ho, and H. Wong-Toi. HYTECH: the next generation. In IEEE [IEE95a], pages 56–65. ISBN 0-8186-7337-0. LCCN QA76.54 .R43 1995.

Hidding:2021:DMP

[Hid21] Martijn Hidding. DiffExp, a Mathematica package for computing Feynman integrals in terms of one-dimensional series expansions. Computer Physics Communications, 269 (??):Article 108125, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S001046552100237X.

Hilbe:2006:M

[Hil06] Joseph M. Hilbe. Mathematica 5.2. The American Statistician, 60(2):176–186, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Henrich:1993:MA

[HJ93] Elkedagmar Henrich and Hans-Dieter Janetzko. Das Mathematica Arbeitsbuch. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-06528-1. 259 pp. LCCN ????

He:1993:SCV

[HKM93] Xiaoyi He, David N. Ku, and James E. Moore, Jr. Simple calculation of the velocity profiles for pulsatile flow in a blood vessel using Mathematica. Annals of biomedical engineering, 21(1): 45–49, January–February 1993. CODEN ABMECF. ISSN 0090-6964. See erratum: Ann Biomed Eng 1993 Sep–Oct, 21(5):557– 558.

Hibbard:1999:EAA

[HL99] Allen C. Hibbard and Kenneth Levasseur. Exploring abstract algebra with Mathematica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98619-7 (softcover). xiii + 467 pp. LCCN QA162 .H52 1999.

Haxton:2008:SMS

[HL08] Wick Haxton and Cecilia Lunardini. SevenOperators, a Mathematica script for harmonic oscillator nuclear matrix elements arising in semileptonic electroweak interactions. Computer Physics Communications, 179 (5):345–358, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465508001112.

Hunt:2009:DEM

[HLO^{+}09] Brian R. Hunt, Ronald L. Lipsman, John E. Osborn, Donald A. Outing, and Jonathan Rosenberg, editors. Differential equations with Mathematica: updated for Mathematica 6. John Wiley and Sons, Inc., Chichester, UK, third edition, 2009. ISBN 0-471-77316-6. viii + 271 pp. LCCN QA371.5.D37; QA371.5.D37 D54 2009.

Howell:1992:ITP

[HM92] R. Howell and J. H. Mathews. Investigation of tangent polynomials with a computer algebra system. The American Mathematical Association of Two Year Colleges Review, 14 (1):20–27, Fall 1992.

Huber:2006:HMP

[HM06] T. Huber and D. Maître. Hypexp, a Mathematica package for expanding hypergeometric functions around integervalued parameters. Computer Physics Communications, 175 (2):122–144, July 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465506001135.

Hastings:2015:HSW

[HMM15] Cliff Hastings, Kelvin Mischo, and Michael Morrison. Handson start to Wolfram Mathematica: and programming with the Wolfram language. Wolfram Media, Champaign, IL, USA, 2015. ISBN 1-57955-077-0 (paperback). ix + 469 pp. LCCN QA76.95 .H377

                                        2015.  URL http://www.
                                        handsonstart.com/.

Hastings:2016:HSW

[HMM16] Cliff Hastings, Kelvin Mischo, and Michael Morrison. Handson start to Wolfram Mathematica: and programming with the Wolfram language. Wolfram Media, Champaign, IL, USA, second edition, 2016. ISBN 1-57955-012-6 (paperback), 1-57955-017-7 (Kindle). ix + 484 pp. LCCN QA76.95 .H377 2016. URL http://www.

                                        wolfram.com/books/profile.
                                        cgi?id=9522.

Hoft:1991:IND

[Hof91] H. Hoft. Implementation of a non-deterministic loop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):24–28, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hoft:1992:LCU

[Höf92] Margret H. Höft. Laboratories for Calculus I Using Mathematica. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54345-1. 153 pp. LCCN ????

Hogben:2007:HLA

[Hog07] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete Mathematics and its Applications (Boca Raton). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). xxx + 1370 pp. LCCN QA184.2 .H36 2007.

URL http://www.crcnetbase.
com/isbn/9781420010572;
http://www.crcnetbase.com/
isbn/9781584885108;   http:
//www.loc.gov/catdir/enhancements/
fy0647/2006045491-d.html.

Associate editors: Richard Brualdi, Anne Greenbaum and Roy Mathias.

Holmgren:1994:FCD

[Hol94] Richard A. Holmgren. A First Course in Discrete Dynamical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94208-4. xi + 214 pp. LCCN QA614.8.H65 1994.

Hollis:1998:CMS

[Hol98] Selwyn L. Hollis. CalcLabs with Mathematica: for Stewart’s Multivariable calculus, concepts and contexts. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1998. ISBN 0-534-35741-5. ix + 262 pp. LCCN QA303.5.D37H65 1998.

Hollis:2003:CMS

[Hol03a] Selwyn L. Hollis. CalcLabs with Mathematica: for Stewart’s Multivariable Calculus, fifth edition. Thomson Brooks/Cole, Pacific Grove, CA, USA, 2003. ISBN 0-534-39362-4. vii + 267 pp. LCCN QA303.5.D37 H65 2003.

Hollis:2003:MCD

[Hol03b] Selwyn L. Hollis. A Mathematica companion for differential equations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-046329-9. 271 pp. LCCN QA371.5.D37 H65 2003.

Hong:1994:IM

[Hon94] Wien Hong. Introduction to Mathematica. Unalis Corporation, ????, 1994. ISBN 957-22-1697-X. 560 pp. LCCN ???? In Chinese.

Hoste:2009:MD

[Hos09] Jim Hoste. Mathematica demystified. Demystified series.

McGraw-Hill, New York, NY, USA, 2009. ISBN 0-07-159144-3. xv + 386 pp. LCCN QA76.95 .H67 2009. URL http://www.

loc.gov/catdir/toc/fy0903/
2008037478.html.

Hehl:1996:RSC

[HPR96] F. W. Hehl, R. A. Puntigam, and H. Ruder, editors. Relativity and scientific computing. Computer algebra, numerics, visualization. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60361-1. xix + 389 pp. LCCN QC173.55 .R445 1996.

Herman:2005:VLA

[HPS05] Eugene A. Herman, Michael D.

Pepe, and Eric P. Schulz. Visual linear algebra. John Wiley and Sons, Inc., Chichester, UK, 2005. ISBN 0-471-68299-3 , 0-471-70628-0 (CD-ROM). xix + 550 pp. LCCN QA185.C65 H47 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/fy053/2004059088.html.

Hebisch:2011:ERM

[HR11] Waldemar Hebisch and Martin Rubey. Extended rate, more GFUN. Journal of Symbolic Computation, 46(8): 889–903, August 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0747717111000162.

Houstis:1992:ESS

[HRV92] E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors. Expert systems for scientific computing: proceedings of the Second IMACS International Conference on Expert Systems for Numerical Computing, Purdue University, USA, 24–26 April, 1990. Elsevier Science Publishers, Amsterdam, The Netherlands, 1992. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Hill:1990:MEN

[HS90] F. S. Hill, Jr. and Lawrence M. Seiford. Mathematica for engineers: numeric, symbolic and graphical computation, 1990. 3 videocassettes.

Headrick:2007:NCG

[HSH07] Todd C. Headrick, Yanyan Sheng, and Flaviu-Adrian Hodis. Numerical computing and graphics for the power method transformation using Mathematica. Journal of Statistical Software, 19(3):1–17, April 2007. CO-DEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v19/i03.

Helton:1994:CSE

[HSW94] J. W. Helton, M. Stankus, and J. Wavrik. Computer simplification of engineering systems formulas. In IEEE, editor, Proceedings of the 33rd IEEE Conference on Decision and Control, 14–16 December 1994, volume 2, pages 1893–1898. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1968-0. LCCN TJ217 .I11c 1994. IEEE catalog number 94CH3460-3.

Helton:1998:CSF

[HSW98] J. W. Helton, M. Stankus, and J. J. Wavrik. Computer simplification of formulas in linear systems theory. IEEE Transactions on Automatic Control, 43 (3):302–314, March 1998. CO-DEN IETAA9. ISSN 0018-9286.

Hakula:2013:MIH

[HT13] Harri Hakula and Tomi Tuominen. Mathematica implementation of the high order finite element method applied to eigenproblems. Computing, 95(1s):277–301, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer.com/article/
10.1007/s00607-012-0262-4.

Hughes:1993:UVM

[Hug93] W. L. Hughes. Use of the vector magnetic potential concept and very high precision computation software to study the electromagnetic fields of accelerated charged particles. IEEE Transactions on Magnetics, 29 (2):1918–1922, March 1993. CO-DEN IEMGAQ. ISSN 0018-9464.

Hounslow:1992:MPP

[HW92] M. J. Hounslow and E. J. W.

Wynn. Modelling particulate processes: full solutions and short cuts. reprint of May 1992 conference paper. Computers & Chemical Engineering, 16S:411– 420, 1992. CODEN CCENDW. ISSN 0098-1354.

Harper:1991:GCA

[HWH91] David Harper, Chris Wooff, and David Hodgkinson. A Guide to Computer Algebra Systems. John Wiley and Sons, Inc., Chichester, UK, 1991. ISBN 0-471-92910-7. xii + 148 pp. LCCN QA155.7.E4 H37 1991. US$31.85.

Hsieh:1992:HSH

[HY92] A. Hsieh and E. Yehudai. HIP: symbolic high-energy physics calculations. Computers in Physics, 6(3):253–261, May/ June 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Hong:2014:MSI

[HY14] Hoon Hong and Chee Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5–9, 2014, Proceedings, volume 8592. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (ebook). LCCN QA76.9.M35.

Ioakimidis:1993:CED

[IA93] N. I. Ioakimidis and E. G. Anastasselou. Constructing elementary databases and using mechanics-related functions and object types in fracture mechanics with MATHEMATICA. Computers and structures, 47 (2):233–238, April 1993. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1995:SPE

[IA95] N. I. Ioakimidis and G. T. Anastasselos. Solution of plane elasticity problems with Mathematica. Computers and structures, 55(2):229–236, April 1995. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Isidori:1995:PTE

[IBM^{+}95] A. Isidori, S. Bittanti, E. Mosca, A. De Luca, M. D. Di Benedetto, and G. Oriolo, editors. Proceedings of the Third European Control Conference. ECC 95. Eur. Union Control Assoc, Rome, Italy, 1995. 4 vol.

IEEE:1990:ISD

[IEE90a] IEEE, editor. 1990 International symposium digest: antennas and propagation: merging technologies for the 90’s: Dallas, TX, May 7–11, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN IAPSBG. ISSN 0272-4693. LCCN TK 7871.6 N56 1990. Four volumes. IEEE catalog no. 90CH2776-3.

IEEE:1990:IIC

[IEE90b] IEEE, editor. ICASSP 90 / 1990 International Conference on Acoustics, Speech and Signal Processing, April 3–6, 1990, Albuquerque Convention Center, Albuquerque, New Mexico, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN IPRODJ. ISSN 0736-7791. LCCN TK 7882 S65 I16 1990. Five volumes.

IEEE:1990:IPP

[IEE90c] IEEE, editor. IEEE Plans ’90, Position Location and Navigation Symposium record: the 1990’s — a decade of excellence in the navigation sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN RPLSDW. LCCN TL695 .I44 1990. IEEE catalog number 90CH2811-8.

IEEE:1990:PIC

[IEE90d] IEEE, editor. Proceedings of the 29th IEEE Conference on Decision and Control, December 5–7, 1990, Hilton Hawaiian Village, Honolulu, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. LCCN TJ 217 I11c 1990. Six volumes.

IEEE:1990:PSI

[IEE90e] IEEE, editor. Proceedings of the Second IEEE Symposium on Parallel and Distributed Processing, 1990. Sponsored by the IEEE Computer Society, IEEE-Dallas Section, in cooperation with Southern Methodist University, University of Texas at Dallas, Dallas, Texas, December 9– 13, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 1990. ISBN 0-8186-2087-0. LCCN QA 76.58 I42 1990.

IEE:1991:ICC

[IEE91a] IEE, editor. International Conference on Computation in Electromagnetics, 25–27 November 1991, volume 350 of IEE Conference. IEE, London, UK, 1991. ISBN 0-85296-529-X. LCCN QC760.54 .I57 1991.

IEEE:1991:IIC

[IEE91b]  IEEE, editor.  ICASSP 91

/ 1991 International Conference on Acoustics, Speech, and Signal Processing, May 14–17, 1991, the Sheraton Centre Hotel & Towers, Toronto, Ontario, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-7803-0003-3 (softbound), 0-7803-0004-1 (casebound), 0-7803-0005-X (microfiche). LCCN TK 7882 S65 I16 1991. Five volumes.

IEEE:1991:IPS

[IEE91c] IEEE, editor. IEEE proceedings of the Southeastcon ’91, Apr. 7-10, 1991 in Williamsburg, VA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN CPISDM. ISBN 0-7803-0033-5 (softbound), 0-7803-0034-3 (casebound), 0-7803-0035-1 (microfiche). ISSN 0734-7502. LCCN TK 7801 I117 1991. Two volumes.

IEEE:1991:PSC

[IEE91d] IEEE, editor. Proceedings: the Seventh Conference on Artificial Intelligence Applications, Miami Beach, Florida, February 24–28, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2135-4 (paperback), 0-8186-6135-6 (microfiche), 0-8186-9135-2 (case). LCCN Q 334 C66 1991. Two volumes.

IEEE:1992:IIC

[IEE92a] IEEE, editor. 1992 IEEE International Conference on Systems, Man and Cybernetics. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0720-8. LCCN TA168.I19 1992. Two volumes. IEEE Cat. No.92CH3176-5.

IEEE:1992:III

[IEE92b] IEEE, editor. ICASSP-92 / 1992 IEEE International Conference on Acoustics, Speech, and Signal Processing, March 23–26, 1992, the San Francisco Marriott, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0533-7 (case), 0-7803-0532-9 (pbk). LCCN TK 7882 S65 I16 1992. Five volumes.

IEEE:1993:III

[IEE93] IEEE, editor. ICASSP 93:

1993 IEEE International Conference on Acoustics, Speech, and Signal Processing, April 27– 30, 1993, Minneapolis Convention Center, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN IPRODJ. ISBN 0-7803-0948-0 (microfiche), 0-7803-0947-2 (casebound), 0-7803-0946-4 (softbound), 0-7803-0949-9 (CD ROM). ISSN 0736-7791. LCCN TK 7882 S65 I16 1993. Five volumes.

IEE:1994:ICC

[IEE94a] IEE, editor. International Conference on Control ’94, 21– 24 March 1994, University of Warwick, Coventry, UK, volume 389 of Conference publication. IEE, London, UK, March 1994. ISBN 0-85296-612-1 (set), 0-85296-610-5 (v. 1), 0-85296-611-3 (v. 2). ISSN 0537-9989. LCCN TJ212.2 .I546 1994 v.1-2 (1994). Two volumes.

IEEE:1994:WIM

[IEE94b]  IEEE, editor.  Wescon/94:

Idea/microelectronics: conference record, Anaheim Convention Center, Anaheim, California, September 27–29, 1994, number 38 in Wescon Conference Record. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN WCREDI. ISBN 0-7803-9993-5, 0-7803-9992-7. ISSN 1044-6036, 0083-8837. LCCN TK 7801 W47 1994.

IEEE:1995:IRT

[IEE95a] IEEE, editor. 16th IEEE Real-Time Systems Symposium: proceedings: December 5–7, 1995, Pisa, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7337-0. LCCN QA76.54 .R43 1995.

IEEE:1995:IPI

[IEE95b] IEEE, editor. ICASSP ’95: Proceedings of the 1995 IEEE International Conference on Acoustics, Speech, and Signal Processing, May 9–12, 1995, Detroit, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN IPRODJ. ISBN 0-7803-2431-5. ISSN 0736-7791. LCCN TK7882.S65 .I16 1995. Five volumes. IEEE Catalog No. 95CH35732.

IEEE:1995:ICR

[IEE95c] IEEE, editor. IEEE conference record–abstracts / 1995 IEEE International Conference on Plasma Science, 5–8 June 1995, Madison, Wisconsin, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2669-5. LCCN TA 2020 I612c 1995. IEEE Catalog No. 95CH35796.

IEEE:1995:PCL

[IEE95d] IEEE, editor. Proceedings: 20th Conference on Local Computer Networks, October 16–19, 1995, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN CLCPDN. ISBN 0-8186-7162-9. ISSN 0742-1303. LCCN TK5105.7 .C66 1995. IEEE Catalog No. 95TB100005.

IEEE:1995:PIC

[IEE95e] IEEE, editor. Proceedings of the 34th IEEE Conference on Decision and Control, December 13–15, 1995, New Orleans Hilton Riverside, New Orleans, Louisiana, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN PCDCDZ. ISBN 0-7803-2685-7. ISSN 0191-2216. LCCN TJ217 .I17 1995. Four volumes. IEEE Catalog No. 95CH35803.

IEEE:1996:PII

[IEE96]  IEEE, editor.  Proceedings:

IEEE International Conference on Robotics and Automation, April 22–28, 1996, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-2988-0. LCCN TJ210.3 .I44 1996. Four volumes. IEEE Catalog No. 96CH35857.

Izquierdo:2018:EMC

[IIS18] Luis R. Izquierdo, Segismundo S. Izquierdo, and William H. Sandholm. EvoDyn-3s: a Mathematica computable document to analyze evolutionary dynamics in 3-strategy games. SoftwareX, 7(??):222–225, January/ June 2018. CODEN ???? ISSN 2352-7110. URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S2352711018300864.

Itoh:1995:IIS

[IMAS95] T. Itoh, K. Mineshita, M. Asakura, and M. Seiki. Interactive instructional system for robotics base on Mathematica. In Anonymous [Ano95h], pages 105–106. ISBN 0-7803-2781-0. LCCN TK7870.K37 1995. SICE catalog number 95 PR 0001-3. IEEE catalog number 95 TH 8107.

Ioakimidis:1992:AMDa

[Ioa92a] N. I. Ioakimidis. Application of MATHEMATICA to the direct solution of torsion problems by the energy method. Computers and structures, 43(4):803– 807, May 1992. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1992:MAS

[Ioa92b] N. I. Ioakimidis. Minimax approximation to stress intensity factors with Mathematica. Computers and structures, 43 (1):181–183, April 1992. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1992:SNI

[Ioa92c] N. I. Ioakimidis. Semi-numerical iterative series solution of linear algebraic equations with MATHEMATICA’. Communications in Applied Numerical Methods, 8(7):421–430, July 1992. CODEN CANMER. ISSN 0748-8025.

Ioakimidis:1992:AMDb

[Ioa92d] N. L. Ioakimidis. Application of Mathematica to the direct semi-numerical solution of finite element problems. Computers and structures, 45(5-6):833–839, December 1992. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1992:AMI

[Ioa92e] Nikolaos I. Ioakimidis. Application of MATHEMATICA to the iterative SAN solution of singular integral equations appearing in crack problems. Advances in Engineering Software, 14(2): 151–156, 1992. CODEN AE-SODT. ISSN 0141-1195, 0965-9978.

Ioakimidis:1992:DSI

[Ioa92f] Nikolaos I. Ioakimidis. Derivation of the singular integral equations for curvilinear cracks with computer algebra software with computer algebra software. Engineering fracture mechanics, 43(4):671–676, November 1992. CODEN EFMEAH. ISSN 0013-

Ioakimidis:1993:EAM

[Ioa93] N. I. Ioakimidis. Elementary applications of MATHEMAT-ICA to the solution of elasticity problems by the finite element method [CMA 274]. Computer methods in applied mechanics and engineering, 102(1): 29–40, January 1993. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Iyoda:1995:FAC

[IOA^{+}95] M. Iyoda, H. Okaya, M. Akiyama, M. Taniwaki, and S. Sato. FFT analysis of CO laser beam mode based on PC and workstation. In Anonymous [Ano95j], pages 157–160. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Ishak:2016:ATC

[Ish16] B. Ishak. Astrophysics Through Computation: With Mathematica^{(R)} Support, by Brian Koberlein and David Meisel, Scope: textbook. Level: advanced undergraduates, graduate students. Contemporary Physics, 57(1): 129, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Issaevitch:1996:WSC

[Iss96] Tom Issaevitch. What should computer scientists teach to physical scientists and engineers? part 2. response to Wilson: forget multiple tools; use Mathematica. IEEE Computational Science & Engineering, 3(3):59–61, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3059abs.htm.

Iverson:1992:SSA

[Ive92] W. Iverson. The sound of silence; adding data-driven sound to scientific visualizations can help researchers better understand what they’re seeing. Computer Graphics World, 15(1):54, January 1992. CODEN CG-WODH. ISSN 0271-4159.

Jacob:2001:IEC

[Jac01] Christian Jacob. Illustrating evolutionary computation with Mathematica. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2001. ISBN 1-55860-637-8. 578 pp. LCCN QA76.618 .J3313 2001.

Jacobse:2019:MMP

[Jac19]  Peter H. Jacobse.  Mathe-

maTB: a Mathematica package for tight-binding calculations. Computer Physics Communications, 244(??):392– 408, November 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0010465519301821.

Jaeku:1994:IM

[Jae94] Ryu Jaeku. Introduction to Mathematica. Crown Publishing, ??, Korea, 1994. ISBN 89-406-6217-2. 722 pp. LCCN ???? In Korean.

Janiak:1994:WDM

[Jan94] Wlodzimerz Janiak. Wstep do Mathematica. ????, ????, 1994. ISBN 83-7101-192-X. ???? pp. LCCN ???? In Polish.

Julia-Diaz:2006:QMQ

[JDBT06] Bruno Juliá-Díaz, Joseph M. Burdis, and Frank Tabakin. QDENSITY — a Mathematica quantum computer simulation. Computer Physics Communications, 174(11):914–934, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465506000543.

Julia-Diaz:2009:QMQ

[JDBT09] Bruno Juliá-Díaz, Joseph M. Burdis, and Frank Tabakin. QDENSITY — a Mathematica quantum computer simulation. Computer Physics Communications, 180(3):474, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465508003457.

Jeffrey:1993:IOE

[Jef93]  D. J. Jeffrey.  Integration

to obtain expressions valid on domains of maximum extent. In Bronstein [Bro93], pages 34–41. ISBN 0-89791-604-2. LCCN QA 76.95 I59

                                        1993.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/164081/

p34-jeffrey/. ACM order

number: 505930.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, ISSAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Jones:1996:SAM

[JKP96] J. Jones, N. P. Karampetakis, and A. C. Pugh. Some applications of Maple in linear systems analysis. In Anonymous [Ano96g], pages 7/1–5.

Jamin:1993:TVM

[JL93] Matthias Jamin and Markus E.

Lautenbacher. TRACER version 1.1. A MATHEMATICA package for γ-algebra in arbitrary dimensions. Computer Physics Communications, 74(2): 265–288, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Jeanneau:1999:EMR

[JLP99] Xavier Jeanneau, Daniel Lignon, and Jean-Louis Poss. Exercices de mathématiques résolus à l’aide de Maple et Mathematica. Ellipses-Marketing, Paris, France, 1999. ISBN 2-7298-9909-X. viii + 230 pp. LCCN ????

Johnson:1993:SRP

[JM93] Scott E. Johnson and Ross R. Moore. Surface reconstruction from parallel serial sections using the program Mathematica: example and source code. Computers and Geosciences, 19(7): 1023–1032, August 1993. CO-DEN CGEODT. ISSN 0098-3004 (print), 1873-7803 (electronic).

Joyner:2008:SCT

[JM08] David Joyner and Robert Miller. SAGE and coding theory (abstract only). ACM Communications in Computer Algebra, 42 (1–2):74–78, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:1994:LAM

[Joh94]  Eugene W. Johnson.  Lin-

ear Algebra with Mathematica. Brooks/Cole symbolic computation series. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1994. ISBN 0-534-13068-2. xi + 174 pp. LCCN QA185.D37J65 1995.

Johannsen:1995:ISE

[Joh95] G. Johannsen, editor. Integrated systems engineering: a postprint volume from the IFAC Conference, Baden-Baden, Germany, 27–29 September 1994. Pergamon Press, New York, NY, USA, 1995. ISBN 0-08-042361-2. LCCN T55.45 .I5 1995.

Jones:1992:LFO

[Jon92] Herbert W. Jones. Löwdin αfunction, overlap integral, and computer algebra. International Journal of Quantum Chemistry, 41(5):749–754, March 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1993:BVT

[Jon93] H. W. Jones. Benchmark values for two-center Coulomb integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 45(1):21–30, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jacob:1993:PSI

[JOS93] G. Jacob, N. E. Oussous, and S. Steinberg, editors. Proceedings SC 93. International IMACS Symposium on Symbolic Computation. New Trends and Developments. LIFL Univ. Lille, Lille, France, 1993.

Janhunen:1993:MMB

[JS93a] P. Janhunen and D. Stein.

MathHDF: MathLink-based distributed visualization between Mathematica and HDF files. Computers in Physics, 7(3):290– 294, May/June 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Janhunen:1993:VMM

[JS93b] Pekka Janhunen and Douglas Stein. VISUALIZATION MathHDF: MathLink-based distributed visualization between Mathematica and HDF files. Computers in Physics, 7(3): 290–295, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Janjunen:1993:MMB

[JS93c] P. Janjunen and D. Stein. MathHDF: MathLink-based distributed visualization between Mathematica and HDF files. Computers in Physics, 7(3):290– 294, May/June 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Jurs:1992:M

[Jur92] Peter C. Jurs. Mathematica. Journal of Chemical Information and Computer Sciences, 32 (4):388–390, July 1992. CODEN JCISD8. ISSN 0095-2338.

Javanainen:1993:SLC

[JY93] J. Javanainen and S. Yoo. Semiclassical laser-cooling theory for a trapped multistate ion. Physical Review A (Atomic, Molecular, and Optical Physics), 48 (5):3776–3785, November 1993. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Kehayopulu:1993:ALA

[KA93] Niovi Kehayopulu and Philip Argyris. An algorithm for Light’s associativity test using Mathematica. J. Comput. Inform., 3(1):87–98, 1993. ISSN 1180-3886.

Khan:1999:CSS

[KA99] Rao F. H. Khan and N. Ahmad. A comparison of symbolic solution of radioactive decay chains using Mathematica. SIGSAM Bulletin, 33 (3):20, September 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaoud:1996:VAC

[KAG96] M. Kaoud and J. Ari-Gur. Vibration of annular circular plate with intermediate ring support. Collection of Technical Papers — AIAA/ASME/ASCE/AHS Structures, Structural Dynamics & Materials Conference, 3:1515– 1521, 1996. CODEN CPSCDO. ISSN 0273-4508.

Kant:1993:SMM

[Kan93] E. Kant. Synthesis of mathematical modeling software. IEEE Software, 10(3):30–41, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kapur:1992:ADC

[Kap92] Deepak Kapur, editor. Automated Deduction, CADE-11: 11th International Conference on Automated Deduction, Saratoga Springs, NY, USA, June 15–18, 1992: Proceedings, number 607 in Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June 15–18, 1992. ISBN 0-387-55602-8 (New York), 3-540-55602-8 (Heidelberg). LCCN QA76.9.A96 I57 1992.

Kaufmann:1992:MWE

[Kau92]  Stephan Kaufmann.  Math-

ematica als Werkzeug: Eine Einführung mit Anwendungsbeispielen. (German) [Mathematica as a tool: an introduction with application examples]. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1992. ISBN 3-7643-2832-0. 396 pp. LCCN ???? See [Bau96].

Kaufmann:1994:MWE

[Kau94] Stephan Kaufmann. Mathematica als Werkzeug (English: Mathematica as a tool: an introduction with practical examples). Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 3-7643-5031-8 (Basel), 0-8176-5031-8 (Boston). ix + iv + 429 pp. LCCN QA76.95 .K3813 1994. URL http:/

/www.birkhauser.com/Books/
kaufmann/summary.htm.

Kaufmann:1999:CCM

[Kau99] Stephan Kaufmann. A crash course in Mathematica. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-6127-1, 0-8176-6127-1. 200 pp. LCCN QA76.95.K3913 1999.

Kaw:1993:USM

[Kaw93] A. K. Kaw. Use of a symbolic manipulator in mechanics of composites. Computers in education journal, 3(1):61–65, January/March 1993. CODEN CE-JOE7. ISSN 1069-3769.

Korayem:1993:PBS

[KB93] M. H. Korayem and A. Basu.

PC-based symbolic modeling and dynamic simulation for flexible manipulators. American Society of Mechanical Engineers, Dynamic Systems and Control Division (Publication) DSC, 49: 163–168, 1993. CODEN AS-MDEV. ISBN 0-7918-1019-4.

Kleene:1994:SBP

[KC94] S. J. Kleene and H. C. Cejtin.

Solving buffering problems with Mathematica software. Analytical biochemistry, 222(2):310– 314, November 1, 1994. CODEN ANBCA2. ISSN 0003-2697.

Kant:1992:KPG

[KDM^{+}92] E. Kant, F. Daube, W. Macgregor, J. Wald, E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Knowledge-based program generation for mathematical modeling. In Houstis et al. [HRV92], pages 371–392. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Keiper:1993:IAM

[Kei93] Jerry B. Keiper. Interval arithmetic in Mathematica. Interval Computations, 3:76–87, 1993. ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Keiper:1996:ZRD

[Kei96] J. B. Keiper. On the zeros of the Ramanujan τ-Dirichlet series in the critical strip. Mathematics of Computation, 65(216):1613– 1619, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL http://www.ams.org/
                                        jourcgi/jour-pbprocess?fn=
                                        110&arg1=S0025-5718-96-00734-
                                        X&u=/mcom/1996-65-216/.

Kendall:1993:IDS

[Ken93] Wilfrid S. Kendall. Itovsn3: doing stochastic calculus with Mathematica. In Varian [Var93], pages 214–238. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Kendall:2001:SIC

[Ken01] Wilfrid S. Kendall. Symbolic Itô calculus in AXIOM: an ongoing story. Statistics and Computing, 11(1):25–35, 2001. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Kerley:1996:AVS

[Ker96] L. M. Kerley. Animation of the vibrating string. Mathematics and computer education, 30(2): 158–163, Spring 1996. CODEN MCEDDA. ISSN 0730-8639.

Khera:1990:IAE

[KG90] V. Khera and H. Greenside.

An interface for accessing external numerical libraries. Mathematica Journal, 1(2):84–88, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Kirkwood:1994:CPM

[Kir94] James R. Kirkwood. Calculus projects for Mathematica. Wm. C. Brown Publishers, Dubuque, IA, USA, 1994. ISBN 0-697-16736-4. vi + 192 pp. LCCN ????

Kim:1992:SMS

[KL92] Changsik Kim and Noam Lior.

Simplified model for spectral radiative properties in pulverized coal combustors. American Society of Mechanical Engineers, Heat Transfer Division, (Publication) HTD, 212:113– 122, 1992. CODEN ASMHD8. ISBN 0-7918-1052-6. ISSN 0272-5673.

Knopfmacher:2002:MER

[KL02] A. Knopfmacher and D. S. Lubinsky. Mathematica evidence that Ramanujan kills Baker– Gammel–Wills. Applied Mathematics and Computation, 128(2– 3):289–302, May 25, 2002. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:1996:SME

[KLO96] Changsik Kim, N. Lior, and K. Okuyama. Simple mathematical expressions for spectral extinction and scattering properties of small size-parameter particles, including examples for soot and TiO_{2}. Journal of Quantitative Spectroscopy and Radiative Transfer, 55(3):391–411, March 1996. CODEN JQSRAE. ISSN 0022-4073.

Keranen:1995:MVP

[KM95a] V. (Veikko) Keranen and P. (Peter) Mitic. Mathematics with vision: proceedings of the First International Mathematica Symposium. Computational Mechanics Publications, Southampton, UK, 1995. ISBN 1-85312-386-2 (Southampton), 1-56252-310-4 (Boston). 404 pp. LCCN QA76.73.M29 I58 1995.

Kutylowski:1995:BSC

[KM95b] R. Kutylowski and K. Myslecki. Buckling study of cylindrical shell using Mathematica computing system. Z. Angew. Math. Mech, 75(suppl.2):S539– 540, 1995. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Kulenovic:2002:DDS

[KM02] M. R. S. (Mustafa R. S.) Kulenovic and Orlando Merino. Discrete dynamical systems and difference equations with Mathematica. Chapman and Hall/CRC, Boca Raton, FL, USA, 2002. ISBN 1-58488-287-5. xv + 344 pp. LCCN QA431 .K838 2002.

Koberlein:2013:ATC

[KM13] Brian Koberlein and David D.

Meisel. Astrophysics through computation: with Mathematica support. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-01074-8. ix + 372 pp. LCCN QB462.2 .K63 2013.

URL http://www.cambridge.
org/gb/academic/subjects/
astronomy/astrophysics/astrophysics-
through-computation-mathematica-
support.

Keranen:1997:IMP

[KMH97] V. Keranen, P. Mitic, and A. Hietamaki, editors. Innovation in mathematics: proceedings of the Second International Mathematica Symposium, volume 2 of Rovaniemen ammattikorkeakoulun julkaisuja. Computational Mechanics Publications, Boston, MA; Southampton, UK, 1997. ISBN 1-85312-502-9 (Southampton) (invalid checksum??), 952-5153-02-9 (Rovaniemi Polytechnic). ISSN 1239-7725. LCCN QA76.73.M29 I585 1997.

Kijewski:1990:VHP

[KMRC90] M. Kijewski, S. Mueller, F. Rybicki, and P. Carvalho. Very high-precision inversion of illconditioned matrices: a medical imaging application. Mathematica Journal, 1(1):69–74, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Kerckhove:1993:CLM

[KN93] M. G. Kerckhove and V. C.

Nall. Calculus Laboratories With Mathematica. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-034220-2. ???? pp. LCCN ???? This book requires Mathematica version 2.0.

Kreyszig:1994:MCM

[KN94] Erwin Kreyszig and E. J. (Edward J.) Norminton. Mathematica Computer Manual to accompany Advanced Engineering Mathematics, Seventh Edition, by Erwin Kreyszig. John Wiley and Sons, Inc., Chichester, UK, 1994. ISBN 0-471-11719-6. xii + 389 + 76 + 16 pp. LCCN QA401 .K71 1995.

Kreyszig:2002:MCG

[KN02] Erwin Kreyszig and E. J. Norminton. Mathematica computer guide: a self-contained introduction for Erwin Kreyszig, Advanced engineering mathematics, eighth edition. John Wiley and Sons, Inc., Chichester, UK, 2002. ISBN 0-471-38669-3 (paperback). various pp. LCCN QA401 .K75 2002.

Knight:1995:DSM

[Kni95] Kelly J. Knight. Development of the stiffness matrices for a line, beam, quadratic displacement triangle, and linear displacement quadrilateral finite element using Mathematica. Thesis (M.S.), Department of Mechanical Engineering, University of Utah, Salt Lake City, UT, USA, 1995. vi + 228 + 1 pp.

Kang:1993:MPI

[KNJ93] S. G. Kang, G. W. Nam, and G. C. H. Jun. Mathematica: a practical introductory guide for science or engineering majors. Sungandang, ??, Korea, 1993. ISBN 89-315-4164-3. 840 pp. LCCN ???? In Korean.

Knuth:1992:CP

[Knu92] D. Knuth. Convolution polynomials. Mathematica Journal, 2 (4):67–78, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Kobayashi:1994:LMA

[Kob94a] Michimasa Kobayashi. Learning Math Again. Sanseido, ??, Japan, 1994. ISBN 4-385-35629-7. 257 pp. LCCN ???? In Japanese.

Kobayashi:1994:MIS

[Kob94b] Michimasa Kobayashi. Mathematica: An Introduction to Statistics and Probability. Toppan, Tokyo, Japan, 1994. ISBN 4-8101-8932-5. 232 pp. LCCN ???? In Japanese.

Koepf:1993:EAC

[Koe93] Wolfram Koepf. Examples for the algorithmic calculation of formal Puisieux, Laurent and power series. SIGSAM Bulletin, 27(1):20–32, January 1993. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Koepf:1994:PFP

[Koe94] W. Koepf. A package on formal power series. Mathematica Journal, 4(2):62–69, Spring 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Kofler:1993:MEL

[Kof93] Michael Kofler. Mathematica: Einführung und Leitfaden für den Praktiker. Addison-Wesley, Reading, MA, USA, 1993. ISBN 3-89319-485-1. 472 pp. LCCN ????

Korsan:1993:DAE

[Kor93] Robert J. Korsan. Decision analytica: an example of Bayesian inference and decision theory using Mathematica. In Varian [Var93], pages 407–458. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Korsan:1994:DUA

[Kor94] R. Korsan. Decisions, uncertainty, and all that: Nothing ventured, nothing gained: Modeling venture capital decisions. Mathematica Journal, 4(1):74– 80, Winter 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Korsan:1995:DSS

[Kor95] Robert J. Korsan. Decision Support Systems in Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94183-5. 400 pp. LCCN ?? Includes DOS diskette.

Korelc:1997:AGF

[Kor97] Jože Korelc. Automatic generation of finite-element code by simultaneous optimization of expressions. Theoretical Computer Science, 187(1–2):231– 248, November 15, 1997. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1997&volume=
187&issue=1-2&aid=2603.

Kormanyos:2011:APC

[Kor11] Christopher Kormanyos. Algorithm 910: a portable C++ multiple-precision system for special-function calculations. ACM Transactions on Mathematical Software, 37(4):45:1– 45:27, February 2011. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Koutschan:2009:AAH

[Kou09] Christoph Koutschan. Advanced applications of the holonomic systems approach. ACM Communications in Computer Algebra, 43(4):119, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2013:HFM

[Kou13] Christoph Koutschan. Holonomic functions in Mathematica. ACM Communications in Computer Algebra, 47(3–4):179– 182, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2016:MPP

[Kou16] Christoph Koutschan. Motion polynomials and planar linkages. ACM Communications in Computer Algebra, 50(3):109– 112, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kripfjanz:1994:AME

[KP94] Jochen Kripfjanz and Holger Perlt. Arbeiten mit Mathematica: Eine Einführung mit Beispielen. Carl Hanser, München, Germany, 1994. ISBN 3-446-17649-7. 184 pp. LCCN ????

Kythe:1997:PDE

[KPS97] Prem K. Kythe, Pratap Puri, and Michael R. Schaferkotter. Partial differential equations and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1997. ISBN 0-8493-7853-2. xx + 378 pp. LCCN QA374.K97 1997.

Kythe:2003:PDE

[KPS03] Prem K. Kythe, Puri, Pratap, and Michael R. Schaferkotter. Partial differential equations and boundary value problems with Mathematica. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2003. ISBN 1-58488-314-6. 418 pp. LCCN QA374 .K97 2003 UCB.

Kinzel:1998:PCP

[KR98] Wolfgang Kinzel and Georg Reents. Physics by computer: programming physical problems using Mathematica and C. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-62743-X. viii + 289 pp. LCCN QC32 .K626 1998. Revised translation of: Physik per Computer.

Krattenthaler:1993:HHM

[Kra93] C. Krattenthaler. HYP and HYPQ: Mathematica packages for the manipulation of binomial sums and hypergeometric series respectively q-binomial sums and basic hypergeometric series. Technical report, Séminaire Lotharingien de Combinatoire (Gerolfingen, 1993), 1993. 61–76 pp. Prépubl. Inst. Rech. Math. Av., 1993/34, Univ. Louis Pasteur, Strasbourg.

Krattenthaler:1995:HHM

[Kra95] C. Krattenthaler. HYP and HYPQ: Mathematica packages for the manipulation of binomial sums and hypergeometric series, respectively q-binomial sums and basic hypergeometric series. Journal of Symbolic Computation, 20(5–6):737–744, November–December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Kroll:1989:RMS

[Kro89]  Lawrence S. Kroll.  Re-

views: Mathematica — A System for Doing Mathematics by Computer, by Wolfram Research. American Mathematical Monthly, 96(9):855–861, November 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kroeker:1998:NTS

[Kro98] Kirk L. Kroeker. New tools: Software development: Rational Software’s development tools for Ada, Artlandia’s Mathematica graphics tool, ISE’s Eiffel-Base library available for free, Annasoft System’s Windows CE development tool; hardware development: White Mountain’s DSP tools, Analogy’s mixedsignal IC simulator, Dynamic Soft Analysis’ thermal analysis tool; new development: Framework Technologies’ project management system, CAI’s message broker, Gordian’s network redirector software; component technology: Black & White’s usage control software for multi-ORB CORBA implementations, IONA Technologies’ middleware development tools, GreenTree Technologies’ ActiveX controls, KL Group’s JavaBeans with automatic data binding. Computer, 31(10):110– 114, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/

co/books/co1998/pdf/rx110.
pdf.

Kuyacak:1993:CTM

[KS93] S. Kuyacak and A. E. Stuchbery.

Critical test of multi-j supersymmetries from magnetic moment measurements. Physical Review C (Nuclear Physics), 48(1):13– 16, July 1993. CODEN PRV-CAN. ISSN 0556-2813, 1089-4918.

Knapp:1997:DEC

[KS97] Robert Knapp and Mark Sofroniou. Difference equations and chaos in Mathematica: Symbolic and numerical mathematics at work. Dr. Dobb’s Journal of Software Tools, 22(11):84, 85, 88–90, 95, 98, 99, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Kitamoto:1992:CPC

[KSA92] T. Kitamoto, M. Saeki, and K. Ando. CAD package for control system on Mathematica. In IEEE [IEE92a], pages 448– 451. ISBN 0-7803-0720-8. LCCN TA168.I19 1992. Two volumes. IEEE Cat. No.92CH3176-5.

Knoll:1995:DCM

[KSJE95] Cecilia Knoll, Michael Shaw, Jerry Johnson, and Benny Evans. Discovering Calculus with Mathematica. John Wiley and Sons, Inc., Chichester, UK, second edition, 1995. ISBN 0-471-00976-8. viii + 340 pp. LCCN ????

Klima:2007:AAA

[KSS07] Richard E. Klima, Neil P. Sigmon, and Ernest L. Stitzinger. Applications of abstract algebra with Maple and MAT-LAB. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2007. ISBN 1-58488-610-2, 1-4200-1119-7. xii + 505 pp. LCCN QA162 .K65 2007. Includes CD-ROM.

Kollar:1994:GFS

[KU94] J. Kollar and B. Ujfalussy. General formulas for the Slater– Koster tables. J. Comput. Phys., 110(1):187–189, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Kullberg:1991:DCS

[Kul91] C. Kullberg. Desktop calculation strategies for reactor analysis using Mathematica. Transactions of the American Nuclear Society, 64:742–743, 1991.

Kuska:1997:MCM

[Kus97]  Jens-Peer Kuska.  Mathe-

matica und C in der modernen Theoretischen Physik: mit Schwerpunkt Quantenmechanik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61489-3. xii + 359 pp. LCCN ????

Kiselev:1999:FFM

[KVF99] S. P. (Sergei Petrovich) Kiselev, E. V. (Evgenii Vasilevich) Vorozhtsov, and Vasily M. Fomin. Foundations of fluid mechanics with applications: problem solving using Mathematica. Modeling and simulation in science, engineering and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 1999. ISBN 0-8176-3995-0, 3-7643-3995-0. xiv + 575 pp. LCCN QA901 .K578 1999; QA901.K58 1999.

Kajiwara:1992:RZH

[KW92] Joji Kajiwara and Hideharu Watanabe. Real zeros of holomorphic functions with parameters. I. complex numerical integration using Mathematica. Mem. Fac. Sci. Kyushu Univ. Ser. A, 46(1):1–30, 1992. CO-DEN MFKAAF. ISSN 0373-6385.

Kutzler:1992:MEP

[KWW92] B. (Bernhard) Kutzler, Bernhard Wall, and Franz Winkler. Mathematische Expertensysteme: Praktisches Arbeiten mit den Computer Algebra-System MACSYMA, Mathematica und DERIVE. Number 430 in Kontakt und Studium. Expert-Verlag, Ehningen bei Boeblingen, Germany, 1992. ISBN 3-8169-0908-6. 119 pp. LCCN ????

Korayem:1994:ASM

[KYB94] M. H. Korayem, Y. Yao, and A. Basu. Application of symbolic manipulation to inverse dynamics and kinematics of elastic robots. International Journal of Advanced Manufacturing Technology, 9(5):343–350, 1994. CODEN IJATEA.

Korotyeyev:2010:ESC

[KZK10] Igor Korotyeyev, V. I. (Valerii Iakovlevich) Zhuikov, and Radoslaw Kasperek. Electrotechnical systems: calculation and analysis with Mathematica and PSpice. CRC Press/Taylor and Francis, Boca Raton, FL, USA, 2010. ISBN 1-4200-8709-6 (hardcover). xi + 256 pp. LCCN TK7870 .K6526 2010.

LaBone:1994:MEV

[LaB94] T. R. LaBone. Mathematica, enhanced version 2.2 for Windows. Health physics: official journal of the Health Physics Society, 67 (4):419–??, 1994. CODEN HLT-PAO. ISSN 0017-9078.

LakshmanYN:1996:IPI

[Lak96] Lakshman Y. N., editor. ISSAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24–26, 1996, Zurich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Lebiedzinski:1991:TUT

[LB91] Karen Lebiedzinski and Halm H. Bau. Teaching undergraduate thermodynamics with Mathematica. American Society of Mechanical Engineers. Advanced Energy Systems Division, 24:65– 72, 1991. CODEN AMEAE8. ISBN 0-7918-0867-X. ISSN 1071-6947.

Laudanski:1996:UCA

[LBSA96] T. Laudanski, B. Bulkszas, J. Szamatowicz, and M. Akerlund. Uterine contractilitydigital analysis of intra-uterine pressure recording. In Richards and de Glanville [RdG96], pages 94–99. ISBN 0-948198-24-9. LCCN ????

Levine:1990:MV

[LC90] Minna Levine and Linda Custer.

Mathematica, version 1.2. MacUser, pages 148–154, 156, 158, 160, 162, 164, November 1990. CODEN MCUSEY. ISSN 0884-0997. Requires IBM PC or; compatibles; 4MB RAM, DOS 3.0 or later. Discusses Mathematica, Theorist, Maple, and Milo.

Lim:1991:CLP

[LEF91] P. Lim, M. L. Epstein, and E. H. Freeman. A constraint logic programming language for combinatorial optimization and linear programming. In IEEE [IEE91d], pages 299–306. ISBN 0-8186-2135-4 (paperback), 0-8186-6135-6 (microfiche), 0-8186-9135-2 (case). LCCN Q 334 C66 1991. Two volumes.

Levasseur:1992:HNC

[Lev92a] K. Levasseur. A human name collision. Mathematica in Education, 2(1):7–8, Fall 1992. ISSN 1065-2965.

Levy:1992:LLM

[Lev92b] S. Levy. LaTEX labels in Mathematica graphics. Mathematica Journal, 2(4):65–66, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Lev:1994:BR

[Lev94] Benjamin Lev. Book reviews. Interfaces, 24(4):129–??, July 1994. CODEN INFAC4. ISSN 0092-2102 (print), 1526-551X (electronic).

Levelt:1995:IPI

[Lev95] A. H. M. (Antonius Henricus Maria) Levelt, editor. IS-SAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montreal, Canada. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Lewis:1991:MSJ

[Lew91] P. H. Lewis. Math software’s not just numbers; programs look at concepts. San Jose Mercury News, page 5, January 27, 1991. ISSN 0747-2099.

Lewin:1994:MDA

[Lew94] David I. Lewin. Mathematica’s developer asks, “can algorithms replace equations?”. Computers in Physics, 8(3): 238–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4823291.

Lewis:2000:OVU

[Lew00] Alan L. Lewis. Option valuation under stochastic volatility: with Mathematica code. Finance Press, Newport Beach, CA, 2000. ISBN 0-9676372-0-1. vii + 350 pp. LCCN HG6024.A3 L49 2000.

Lichtblau:2011:SDI

[Lic11] Daniel Lichtblau. Symbolic definite (and indefinite) integration: methods and open issues. ACM Communications in Computer Algebra, 45(1):1– 16, March 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Linderholm:1991:MDW

[Lin91a] O. Linderholm. Mathematica does windows. BYTE Magazine, 16:128, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1991:MW

[Lin91b] Owen Linderholm. Mathematica for Windows 3.0. BYTE Magazine, ??(??):128, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lipovetsky:2012:BRBd

[Lip12] Stan Lipovetsky. Book review:

Bayesian Logical Data Analysis for the Physical Sciences:

A Comparative Approach With Mathematica^{T M} Support by Phil Gregory. Technometrics, 54(4): 443–444, November 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.
jstor.org/stable/41714935.

Leu:1991:SRK

[LJW91] M. C. Leu, Z. Ji, and Y. S. Wang. Studying robot kinematics and dynamics with the aid of MATHEMATICA. The International journal of mechanical engineering education, 19(3): 213–228, July 1991. CODEN IMEEB3. ISSN 0306-4190.

Lassez:1990:PQL

[LM90a] J. L. Lassez and A. Miola. Parametric queries, linear constraints and variable elimination. In Miola [Mio90], pages 164–173. ISBN 3-540-52531-9 (Berlin), 0-387-52531-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1990.

Lopez:1990:ECS

[LM90b] R. Lopez and J. H. Mathews. Exploring conic sections with a computer algebra system. Collegiate Microcomputer, 8(3):215– 219, August 1990. CODEN CMICDL. ISSN 0731-4213.

Lopez:1990:UCA

[LM90c] Robert Lopez and John Mathews. Using a computer algebra system to solve for maxima and minima. College Mathematics Journal, 21(5):410–414, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/07468342.1990.
11973346.

Lassig:1991:HSS

[LM91] M. Lassig and G. Mussardo.

Hilbert space and structure constants of descendant fields in two-dimensional conformal theories. Computer Physics Communications, 66(1):71–88, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Lopez:1993:SPP

[LM93] R. Lopez and J. H. Mathews.

The shape of a projectile’s path: Explorations with a computer algebra system. The American Mathematical Association of Two Year Colleges Review, 14 (2):38–47, Spring 1993.

Liu:2023:AMP

[LM23] Xiao Liu and Yan-Qing Ma.

AMFlow: a Mathematica package for Feynman integrals computation via auxiliary mass flow. Computer Physics Communications, 283(??):Article 108565, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465522002843.

Loris:2008:LMP

[Lor08] Ignace Loris. L1Packv2: a Mathematica package for minimizing an ℓ_{1}-penalized functional. Computer Physics Communications, 179(12):895–902, December 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S001046550800252X.

Landau:2005:FCS

[LP05] Rubin H. Landau and M. J. Páez. A first course in scientific computing: symbolic, graphical, and numeric problem solving using Maple, Java, Mathematica, and Fortran. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12183-4 (hardcover). xxiv + 481 pp. LCCN Q183.9 .L36 2005.

Ladroue:2013:DPC

[LP13] Christophe Ladroue and Anastasia Papavaviliou. A distributed procedure for computing stochastic expansions with Mathematica. Journal of Statistical Software, 53(11):??, May 2013. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v53/i11.

Larsen:2018:MMPa

[LR18a] Kasper J. Larsen and Robbert Rietkerk. MultivariateResidues: a Mathematica package for computing multivariate residues. Computer Physics Communications, 222(??):250–262, January 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465517302886.

Larsen:2018:MMPb

[LR18b] Kasper J. Larsen and Robbert Rietkerk. MultivariateResidues: a Mathematica package for computing multivariate residues. Computer Physics Communications, 233 (??):269–270, December 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465518302212.

Lustig:1995:TFM

[LRW95] S. R. Lustig, S. Rastogi, and N. Wagner. Telescoping fast multipole methods using Chebyshev economization. J. Comput. Phys., 122(2):317–322, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Lefton:1995:CME

[LS95] Lew E. Lefton and Enid M.

Steinbart. Calculus and Mathematica: An end-user’s point of view. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), 5(1):80–96, March 1995. ISSN 1051-1970.

Lu:1996:SCD

[LS96] X. Y. Lu and S. K. Spurgeon.

Symbolic computation for dynamic sliding mode controller design. In Anonymous [Ano96g], pages 4/1–5.

Lea:1994:PSA

[LT94] R. Mike Lea and Stuart Tewksbury, editors. 1994 proceedings / Sixth Annual IEEE International Conference on Wafer Scale Integration, San Francisco, California, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1850-1 (casebound), 0-7803-1849-8 (softbound), 0-7803-1851-X (microfiche). LCCN TK 7874 I578 1994.

Lutovac:2001:FDS

[LTE01] Miroslav D. Lutovac, Dejan V. Tosic, and Brian L. Evans. Filter design for signal processing using MATLAB and Mathematica. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-201-36130-2. xxvi + 756 pp. LCCN TK7872.F5 L88 2001.

Lorig:1995:ERP

[LU95] Tyler S. Lorig and Thomas P. Urbach. Event-related potential analysis using Mathematica. Behavior research methods, instruments, and computers, 27 (3):358–366, August 1995. CO-DEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Lucic:1995:DAS

[Luc95] Vladan Lucic. Dill: an algorithm and a symbolic software package for doing classical supersymmetry calculations. Computer Physics Communications, 92(1):

90–110, November 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Lum:1993:PFA

[Lum93] Lewis Lum, editor. Proceedings of the Fourth Annual International Conference on Technology in Collegiate Mathematics, Portland State University, Department of Mathematics, Portland, Oregon, November 15–17, 1991. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Lum:1994:PFA

[Lum94] Lewis Lum, editor. Proceedings of the Fifth Annual International Conference on Technology in Collegiate Mathematics: William Rainey Harper College and Northern Illinois University, Rosemont, Illinois, November 12–15, 1992. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Ligh:1996:LFI

[LW96] S. Ligh and R. Wills. On linear functions. III. Mathematics and computer education, 30 (1):9–18, Winter 1996. CODEN MCEDDA. ISSN 0730-8639.

Lewis:1999:CPO

[LW99] Robert H. Lewis and Michael Wester. Comparison of polynomialoriented computer algebra systems. SIGSAM Bulletin, 33 (4):5–13, December 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lewis:2022:BDE

[LW22] Robert Y. Lewis and Minchao Wu. A bi-directional extensible interface between Lean and Mathematica. Journal of Automated Reasoning, 66(2):215–238, May 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10817-021-09611-1.

Lynch:2007:DSA

[Lyn07] Stephen Lynch. Dynamical systems with applications using Mathematica. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 0-8176-4482-2 (paperback), 0-8176-4586-1 (ebook). xv + 484 pp. LCCN QA614.8 .L963 2007.

Logan:1996:TPT

[LZ96] J. D. Logan and V. Zlotnik. Time-periodic transport in heterogeneous porous media. Applied Mathematics and Computation, 75(2-3):119–138, March 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:1995:SSM

[Ma95] Jiefu Ma. Spin-up of a stratified magnetofluid as a model of planetary interiors. Geophysical and Astrophysical Fluid Dynamics, 81(3-4):159–191, 1995.

CODEN GAFDD3. ISSN 0309-1929.

Miyaji:2000:MNP

[MA00] Chikara Miyaji and Paul Abbott. MathLink: network programming with Mathematica. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-64172-1 (hardcover), 0-521-64598-0 (paperback). xix + 243 pp. LCCN QA76.625 .M59 2001. US$85.00 (hardcover), US$29.95 (paperback).

Miyaji:2001:MNP

[MA01] Chikara Miyaji and Paul Abbott. MathLink: network programming with Mathematica. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-64172-1, 0-521-64598-0. xix + 243 pp. LCCN QA76.625 .M59 2001 Accompanying CD-ROM shelved in Reserves.

MacDonald:1992:PUR

[Mac92] William M. MacDonald. Postuse review. Mathematica: a practical approach by Nancy Blachman. American Journal of Physics, 60(12):1157–??, December 1992. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Macedo:2008:AGM

[MAC08] Raquel S. Macedo, Marcelo F.

Alfradique, and Marcelo Castier. Automatic generation of Matlab functions using Mathematica and Thermath. Computing in Science and Engineering, 10(4):

41–49, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Maddox:1991:GYS

[Mad91] John Maddox. Growing up with your software. (plans of Steven Wolfram, designer of “Mathematica,” computationalist’s software package). Nature, 351(6326):437, June 6, 1991. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maeder:1989:PM

[Mae89] Roman E. Maeder. Programming in Mathematica. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-51002-2. xiv + 267 pp. LCCN QA76.95 .M34 1989.

Maeder:1990:MC

[Mae90a] R. E. Maeder. Mathematics on the computer. Output, 19(11): 29–32, October 31, 1990. What journal is this? Several have the same name.

Maeder:1990:PM

[Mae90b] Roman Maeder. Programming in Mathematica. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51002-2. xiv + 267 pp. LCCN QA76.95 .M34 1989. Prepared with TEX.

Maeder:1991:PM

[Mae91] Roman E. Maeder. Programming in Mathematica. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-54877-1 (hardcover), 0-201-54878-X (paperback). xiv + 279 pp. LCCN Q183.9 .M34 1991. See [Maexx].

Maeder:1992:DMP

[Mae92] Roman E. Maeder. The design of the Mathematica programming language: a single paradigm provides surprising diversity. (tutorial). Dr. Dobb’s Journal of Software Tools, 17(4):86–88, 90, 92, 94, 97, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Maeder:1993:WCM

[Mae93a] R. E. Maeder. Writing courseware with Mathematica. In Jacob et al. [JOS93], pages 237– 239.

Maeder:1993:IMN

[Mae93b] Roman Maeder. Informatik für Mathematiker und Naturwissenschaftler: Eine Enführung mit Mathematica (Computer Science for Mathematicians and Scientists: an Introduction with Mathematica). Addison-Wesley, Reading, MA, USA, 1993. ISBN 3-89319-519-X. 408 pp. LCCN ????

Maeder:1994:MPL

[Mae94a] R. Maeder. The Mathematica programmer: Logic programming I: The interpreter. Mathematica Journal, 4(1):53– 63, Winter 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Maeder:1994:SI

[Mae94b] R. Maeder. The stellated icosahedra. Mathematica in Education, 3(1):5–11, 1994. ISSN 1065-2965.

Maeder:1994:MP

[Mae94c] Roman E. Maeder. The Mathematica Programmer. Academic Press, New York, NY, USA, 1994. ISBN 0-12-464990-4. xv + 199 + 8 pp. LCCN QA76.73.M29 M33 1994. US$44.95.

Maeder:1996:MPI

[Mae96a] Roman Maeder. The Mathematica programmer II. Academic Press, New York, NY, USA, 1996. ISBN 0-12-464992-0. xvii + 296 pp. LCCN QA76.73.M29M34 1996. US$44.95.

Maeder:1996:PM

[Mae96b] Roman Maeder. Programming in Mathematica. Addison-Wes-ley, Reading, MA, USA, third edition, 1996. ISBN 0-201-85449-X. ???? pp. LCCN Q183.9.M34 1996.

Maeder:1997:PM

[Mae97] Roman Maeder. Programming in Mathematica. Addison-Wes-ley, Reading, MA, USA, third edition, 1997. ISBN 0-201-85449-X. xvi + 366 pp. LCCN Q183.9 .M34 1997.

Maeder:19xx:PM

[Maexx] Roman E. Maeder. Programming in Mathematica. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 19xx. ISBN 4-8101-8043-3. ???? pp. LCCN ???? Japanese translation of [Mae91].

Maeder:2000:CSM

[Mae00] Roman Maeder. Computer Science with Mathematica: Theory and Practice for Science, Mathematics, and Engineering. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-63172-6 (hardcover), 0-521-66395-4 (paperback). xx + 389 pp. LCCN QA76.95 .M34 2000. US$64.95 (hardcover), US$27.95 (paperback).

Maitre:2006:HMI

[Maî06] D. Maître. HPL, a Mathematica implementation of the harmonic polylogarithms. Computer Physics Communications, 174(3):222–240, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465505005771.

Majumdar:1996:AEI

[Maj96] A. Majumdar. Analytical expressions for isotropic mixing in three- and four-spin topologies in /sup 13/C systems. Journal of magnetic resonance. Series A, 121(2):121–126, August 1996. CODEN JMRAE2. ISSN 1064-1858.

Makhlouf:1997:AAC

[Mak97] Abdenacer Makhlouf. Algèbres associatives et calcul formel.

(French) [Associative algebras and computer algebra]. Theoretical Computer Science, 187(1– 2):123–145, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        187&issue=1-2&aid=2597.

Martin:1992:CMD

[MAL92] R. A. Martin, A. J. Ahumada, Jr., and J. O. Larimer. Color matrix display simulation based upon luminance and chromatic contrast sensitivity of early vision. In Rogowitz [Rog92], pages 336–342. ISBN 0-8194-0820-4. LCCN TS510.S63 v.1666.

Mangano:2010:MC

[Man10] Sal Mangano. Mathematica Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-52099-9. 782 (est.) pp. LCCN ????

Maruszewski:1994:APL

[Mar94] Richard F. Maruszewski. Approximating the parameters for the logistic model. Mathematics and computer education, 28(1): 16–19, Winter 1994. CODEN MCEDDA. ISSN 0730-8639.

Marasco:2000:LPM

[Mar00]  A. Marasco.  Lindstedt–

Poincaré method and Mathematica applied to the motion of a solid with a fixed point. Computers and Mathematics with Applications, 40(2–3):333– 343, July/August 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122100001644.

Mathews:1989:CAS

[Mat89a] J. H. Mathews. Computer algebra systems approach to teaching Taylor polynomials. The American Mathematical Association of Two Year Colleges Review, 11(1, (part 2)):61–66, Fall 1989.

Mathews:1989:CSA

[Mat89b] J. H. Mathews. Computer symbolic algebra solutions of second order differential equations. Collegiate Microcomputer, 7(4): 311–317, November 1989. CO-DEN CMICDL. ISSN 0731-4213.

Mathews:1989:UCA

[Mat89c] J. H. Mathews. Using a computer algebra system to teach second order differential equations. CoED, 9(4):7–10, October/December 1989. CODEN CWLJDP. ISSN 0736-8607.

Wolfram:1990:MJ

[MAT90a] Stephen Wolfram, editor. The Mathematica Journal, 1990.

ISSN 1047-5974 (print), 1097-1610 (electronic). Addison-Wes-ley, Reading, MA, USA.

Mathews:1990:PFT

[Mat90b] J. H. Mathews. Partial fractions, theory and practice with Mathematica. Mathematics and computer education, 24(3):257–266, Fall 1990. CODEN MCEDDA. ISSN 0730-8639.

Mateosian:1991:MHS

[Mat91a] R. Mateosian. Mathematica help stack. IEEE Micro, 11 (4):43, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wellin:1991:MER

[Mat91b] Paul R. Wellin, editor. Mathematica in Education and Research, 1991. ISSN 1065-2965. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA. Published quarterly. Also available electronically in PostScript form.

Mathews:1990:SDE

[Mat91c] J. H. Mathews. Solving differential equations with computer algebra systems software. Journal of Computers in Mathematics and Science Teaching, 10(2): 79–86, Winter 1990–1991. CO-DEN JCMTDV. ISSN 0731-9258.

Mathews:1991:FLS

[Mat91d] J. H. Mathews. Finding least squares lines with Mathematica. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), I(1): 103–111, 1991. ISSN 1051-1970.

Mathews:1991:LRM

[Mat91e] J. H. Mathews. L’Hôpital’s rule with Mathematica. The American Mathematical Association of Two Year Colleges Review, 13 (1):40–46, 1991.

Mathews:1991:MNT

[Mat91f] J. H. Mathews. Mathematica, a new tool for studying optimization. International journal of mathematical education in science and technology, 22 (4):569–576, July 1991. CO-DEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Mathews:1991:UCS

[Mat91g] J. H. Mathews. Using computer symbolic algebra in applied engineering mathematics to teach Laplace transforms. CoED, 1 (1):71–76, January/March 1991. CODEN CWLJDP. ISSN 0736-8607.

Matthews:1991:MDO

[Mat91h] K. Matthews. Mathematica during office hours. Mathematica in Education, 1(2):16–20, Fall 1991. ISSN 1065-2965.

Mathews:1992:BPG

[Mat92a] J. H. Mathews. Bounded population growth: a curve fitting lesson. Mathematics and computer education, 26(2):169– 176, Spring 1992. CODEN MCEDDA. ISSN 0730-8639.

Mathews:1992:SSMb

[Mat92b] J. H. Mathews. Sequences, series and Mathematica. School science and mathematics, 92(6): 318–324, October 1992. CODEN SSMAAC. ISSN 0036-6803.

Mathews:1992:SSMa

[Mat92c] John H. Mathews. Sequences, series, and Mathematica. School science and mathematics, 92(6): 249–??, October 1992. CODEN SSMAAC. ISSN 0036-6803.

Mathews:1993:SGS

[Mat93] John H. Mathews. Summing geometric series recursively. Mathematics and computer education, 27(2):125–??, Spring 1993. CO-DEN MCEDDA. ISSN 0730-8639.

Anonymous:1994:MER

[Mat94] Anonymous, editor. Mathematica in education, 1994. CO-DEN ???? ISSN 1065-2965. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA.

Hunt:1995:MW

[Mat95a] Stephen M. Hunt, editor. Mathematica World, 1995. ISSN 1062-7030. URL http:

                                        //www.vut.edu.au/~steveh/
                                        MW/;   http://www.wolfram.
                                        com/mathematica/books/bookmathworld.
                                        html;   mailto:mathematica@
                                        matilda.vut.edu.au. Parkville

Victoria 3052, Australia.

Wolfram:1995:M

[Mat95b] Wolfram Research, Inc. Math-User, 1995. 100 Trade Center Drive, Champaign, IL 61820-7237, USA.

Mathematica:1998:MJ

[Mat98] The Mathematica journal, 1998.

ISSN 1097-1610. URL http:

//www.idealibrary.com. Aca-

demic Press, New York, NY, USA. Online journal, accessible to qualified subscribers via the publisher’s IDEAL Project.

Anonymous:2005:MER

[Mat05] Anonymous, editor. Mathematica in education and research, 2005. CODEN ???? ISSN 1096-3324. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA.

May:1994:UMD

[May94a] Susan Kathy May. Using Mathematica to discover calculus. Thesis (M.A.), University of Texas at Austin, Austin, TX, USA, 1994. iii + 125 pp.

Mayes:1994:VEP

[May94b] E. Mayes. Visualizing electrostatic phenomena using Mathematica. In Anonymous [Ano94b], pages 116–122. ISSN 0097-4374.

Martin:1994:UFM

[MB94] T. P. Martin and J. F. Baldwin.

Using FRIL and Mathematica to give advice on engineering models. In Anonymous [Ano94g], pages 101–105. ISBN 3-86073-286-2. LCCN ???? DM237.00. Three volumes.

Mertig:1991:FCC

[MBD91] R. Mertig, M. Bohm, and A. Denner. Feyn-Calc-computer-algebraic calculation of Feynman amplitudes. Computer Physics Communications, 64 (3):345–359, June 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Meisel:1993:ELP

[MC93] L. V. Meisel and P. J. Cote. Exponents in lifetime and power spectral density forms in selforganized critical systems — a Mathematica application. Computers in Physics, 7(6):710– 713, November 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Mikhailov:1996:ORD

[MC96] M. D. Mikhailov and R. M. Cotta. Ordering rules for double and triple eigenseries in the solution of multidimensional heat and fluid flow problems. International Communications in Heat and Mass Transfer, 23 (2):299–303, March–April 1996. CODEN IHMTDL. ISSN 0735-1933.

McAndrew:2008:TCO

[McA08] Alasdair McAndrew. Teaching cryptography with open-source software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):325–329, March 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McCullough:2000:AMS

[McC00] B. D. McCullough. The accuracy of Mathematica 4 as a statistical package. Computational Statistics, 15(2):279–299, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/PL00022713.

McCullough:2002:BRM

[McC02] B. D. McCullough. Book review: Mathematical Statistics With Mathematica, by Colin Rose and Murray Smith. New York: Springer-Verlag, 2002. ISBN 0-387-95234-9. xiii + 481 pp. (+2 CDs). $79.97. Journal of the American Statistical Association, 97(460):1202– 1203, December 2002. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://fidelio.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/01621459/v97n460/s25/
p1202.

McClain:2009:STM

[McC09] William McClain. Symmetry theory in molecular physics with Mathematica: a new kind of tutorial book. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-73469-4 (hardcover). xv + 689 pp. LCCN ????

Monagan:2013:IPI

[MCG13] Michael Monagan, Gene Cooperman, and Mark Giesbrecht, editors. ISSAC 2013: Proceedings of the 2013 International Symposium on Symbolic and Algebraic Computation, June 26– 29, 2013, Boston, MA, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

McPherson:1992:MBF

[McP92] A. McPherson. Modal behavior of the four-layer planar waveguide. Mathematica Journal, 2 (1):75–77, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Maroni:2013:CCO

[MdR13] Pascal Maroni and Zélia da Rocha. Connection coefficients for orthogonal polynomials: symbolic computations, verifications and demonstrations in the Mathematica language. Numerical Algorithms, 63(3):507–520, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s11075-012-9634-2.

Mezzino:1992:ITE

[Mez92] M. Mezzino. Including TEX expressions in a notebook. Mathematica Journal, 2(4):61–64, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Monjoie:1994:MMT

[MG94] F. S. Monjoie and H. P. Garnir. Mathematica as a multimedia teaching aid. In Gruber and Tomassini [GT94], pages 559– 562. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Mattsson:1994:IIJ

[MGC94] Sven Erik Mattsson, J. O. Gray, and François E. Cellier, editors. IEEE/IFAC Joint Symposium on Computer-Aided Control System Design: proceedings, March 7–9, 1994, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1800-5 (softbound), 0-7803-1801-3 (microfiche). LCCN TJ 212.2 I3256 1994.

Montero:1996:SEA

[MGS96] G. A. Montero, T. D. Giorgio, and K. B. Schnelle, Jr. Scaleup and economic analysis for the design of supercritical fluid extraction equipment for remediation of soil. Environmental Progress, 15(2):112–121, Summer 1996. CODEN ENVPDI. ISSN 0278-4491.

Michaels:1992:AF

[Mic92a] J. Michaels. Arbing on the fly.

Wall Street and Technology, 9: 85–86, April 1992. CODEN WSTEE5. ISSN 1060-989X.

Michaels:1992:TF

[Mic92b] J. Michaels. Trader of the future. Wall Street Computer Review, 9:32–36, January 1992. ISSN 0738-4343.

Mielke:1995:DSV

[Mie95] Craig John Mielke. The design and simulation of a vector quantizer using the Mathematica software package. Thesis (M.S.), University of Texas at Arlington, Arlington, TX, USA, 1995. viii + 59 pp.

Mikhailov:1994:FDM

[Mik94a] M. D. Mikhailov. Finite difference method by using Mathematica. International journal of heat and mass transfer, 37 (Suppl 1):375–379, March 1994. CODEN IJHMAK. ISSN 0017-9310.

Mikhailov:1994:FEM

[Mik94b] M. D. Mikhailov. Finite element method by using Mathematica. NATO ASI series. Series F, Computer and system sciences, 132(??):129–??, 1994. CODEN NASFEG. ISSN 0258-1248.

Miller:1990:CAF

[Mil90] R. Miller. Computer-aided financial analysis: An implementation of the Black–Scholes model. Mathematica Journal, 1(1):75–79, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Miola:1990:DIS

[Mio90] A. Miola, editor. Design and implementation of symbolic computation systems: International Symposium DISCO ’90, Capri, Italy, April 10–12, 1990: proceedings, volume 429 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-52531-9 (Berlin), 0-387-52531-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1990.

Miola:1993:DIS

[Mio93]  A. Miola, editor.  Design

and implementation of symbolic computation systems: International Symposium, DISCO ’93, Gmunden, Austria, September 15–17, 1993: proceedings, volume 722 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57235-X (Berlin), 0-387-57235-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88 I576 1993. “Sponsored by the Research Institute for Symbolic Computation (University J. Kepler, Linz, Austria) and by the Dipartimento di informatica e sistemistica (University ‘La Sapienza’, Roma, Italy)”– Foreword.

Miszczak:2012:GUT

[Mis12] Jaroslaw Adam Miszczak. Generating and using truly random quantum states in Mathematica. Computer Physics Communications, 183(1):118–124, January 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465511002748.

Miszczak:2013:EOQ

[Mis13] Jaroslaw Adam Miszczak. Employing online quantum random number generators for generating truly random quantum states in Mathematica. Computer Physics Communications, 184(1):257–258, January 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465512002780.

Mitsui:1989:TIN

[Mit89] T. Mitsui. Towards an integration of numerical and algebraic computations on computer. Journal of the Institute of Electronics, Information and Communication Engineers. (JAPAN), 72(10):1066– 1073, October 1989.

Miyaoka:1994:MG

[Miy94] Etsuo Miyaoka. Mathematica Guidebook. Brain, ??, Japan, 1994. ISBN 4-89242-144-8. 546 pp. LCCN ???? In Japanese.

Mulquiney:2003:MMM

[MK03] Peter J. Mulquiney and Philip W. Kuchel. Modelling metabolism with Mathematica: detailed examples including erythrocyte metabolism. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2003. ISBN 0-8493-1468-2. 309 pp. LCCN QH634.5 .M85 2003. Includes CD-ROM.

Mitchell:1993:BLS

[MM93] C. E. Mitchell and A. B.

McLean. Bravais lattices, surface nets, and Buckminsterfullerenes. Mathematica Journal, 3(3):65–68, 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Malek-Madani:1997:AEM

[MM97] Reza Malek-Madani. Advanced engineering mathematics with Mathematica and Matlab. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-30822-3. 6 + 3 + 79–705 pp. LCCN ????

Malaina:1998:MAC

[MM98a] Jose Luis Malaina and Ana Isabel Martin. Matematicas avanzadas con Mathematica. Universidad del Pais Vasco. Servicio editorial, Bilbão, Spain, 1998. ISBN 84-7585-967-4. 345 pp. LCCN ????

Malek-Madani:1998:AEM

[MM98b] Reza Malek-Madani. Advanced engineering mathematics with Mathematica and Matlab. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-59881-7 (vol. 1), 0-201-32549-7 (vol. 2). various pp. LCCN TA345.M34 1997. Two volumes.

Maitre:2008:SMI

[MM08] D. Maître and P. Mastrolia.

S@M, a Mathematica implementation of the spinor-helicity formalism. Computer Physics Communications, 179(7):501– 534, October 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465508001793.

Markushevich:1991:DDR

[MN91] V. Markushevich and E. Nyland. Dispersion dependence for rayleigh-like vibrations in a lowdensity channel. J. of Geophysical Research, 96(B12):20325– 20330, November 10, 1991. CO-DEN JGREA2. ISSN 0148-0227.

Majamemi:1994:MWV

[MO94] Antti Majamemi and Tapani Ojanperä. Mathematica for Windows Version 2.2. Kymdata, ??, Finland, 1994. ISBN 951-559-137-6. 92 pp. LCCN ???? In Finnish.

Moller:2007:OLC

[Möl07] Karl Dieter Möller. Optics: learning by computing with examples using MathCAD, Matlab, Mathematica, and Maple. Undergraduate texts in contemporary physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 0-387-69492-7, 0-387-26168-0. xvi + 453 pp. LCCN QC381 .M66 2007eb.

Monasse:1993:SGG

[Mon93] D. Monasse. Symbolic geometry, golden spirals, and Descartesgeometry. Mathematica in Education, 2(3):3–6, 1993. ISSN 1065-2965.

Moriya:1993:LMMa

[Mor93a] Ryoji Moriya. Learning Mathematics with Mathematica: Introduction. Kaibundo, ??, Japan, 1993. ISBN 4-303-72780-6. 180 pp. LCCN ???? In Japanese.

Moriya:1993:LMMb

[Mor93b] Ryoji Moriya. Learning Mathematics with Mathematica: Linear Algebra. Kaibundo, ??, Japan, 1993. ISBN 4-303-72800-4. 167 pp. LCCN ???? In Japanese.

Moriya:1994:DIC

[Mor94] Ryoji Moriya. Differential and Integral Calculus with Mathematica. Kaibundo, ??, Japan, 1994. ISBN 4-303-72790-3. ???? pp. LCCN ???? In Japanese.

Morales:1995:ECM

[Mor95]  Francisco Morales.  Explor-

ing computational models with Mathematica. Thesis (M.S.), State University of New York Institute of Technology at Utica/ Rome, Utica, NY, USA, May 1995. ii + 139 pp.

MorenoFlores:1998:PCP

[Mor98a] Joaquin Moreno Flores. Problemas de calculo paso a paso con Mathematica. Coleccion Libro docente. Universidad Politecnica de Valencia, Valencia, Spain, 1998. ISBN 84-7721-691-6. 261 pp. LCCN ????

Morgan:1998:BRB

[Mor98b] Geoffrey Morgan. Book review: Computational Economics and Finance: Modelling and Analysis with Mathematica, by H. R. Varian. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (2):409–410, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/2988700.

Monagan:2010:SPM

[MP10] Michael Monagan and Roman Pearce. Sparse polynomial multiplication and division in Maple 14. ACM Communications in Computer Algebra, 44(4):205– 209, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mone:1994:UMF

[MR94] Robert Mone and Michael I. Ratliff. Using Mathematica to find minimal length confidence intervals. The Pentagon, 53(2): 25–??, Spring 1994. ISSN 0031-4870.

Marasco:2001:SCM

[MR01] Addolorata Marasco and Antonio Romano. Scientific computing with Mathematica: mathematical problems for ordinary differential equations. Modeling and simulation in science, engineering, and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2001. ISBN 0-8176-4205-6, 3-7643-4205-6. xiv + 270 pp. LCCN QA371.5.D37 M37 2001.

Mertig:1998:TMP

[MS98] R. Mertig and R. Scharf.

TARCER — a Mathematica program for the reduction of two-loop propagator integrals. Computer Physics Communications, 111(1–3):265–273, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465598000423.

McGrail:2008:TCM

[MS08] Robert W. McGrail and Mary Sharac. Tricoloring as a corrective measure (abstract only). ACM Communications in Computer Algebra, 42(1–2):83–85, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mihalic:1994:NST

[MT94a] Z. Mihalic and N. Trinajstica.

On the number of spanning trees in fullerenes. Fullerene Science and Technology, 2(1):89– 95, 1994. CODEN FTECEG. ISSN 1064-122X.

Munro:1994:SRRa

[MT94b] N. Munro and P. Tsapekis.

Some recent results using symbolic algebra. In IEE [IEE94a], pages 7–12. ISBN 0-85296-612-1 (set), 0-85296-610-5 (v. 1), 0-85296-611-3 (v. 2). ISSN 0537-9989. LCCN TJ212.2 .I546 1994 v.1-2 (1994). Two volumes.

Munro:1994:SRRb

[MT94c] N. Munro and P. Tsapekis. Some recent results using symbolic algebra. In Mattsson et al. [MGC94], pages 109–116. ISBN 0-7803-1800-5 (softbound), 0-7803-1801-3 (microfiche). LCCN TJ 212.2 I3256 1994.

Meredith:1996:BRH

[MT96] Dawn C. Meredith and Bill Titus. Book reviews: Harvey Gould and Jan Tobochnik, An Introduction to Computer Simulation Methods: Applications to Physical Systems, 2nd Edition. Richard L. Gaylord and Paul R. Wellin, Mathematica: Explorations in Complex Physical and Biological Systems, Springer-Verlag, New York, 1995; ISBN 0-387-94274-2; 297 pp. plus CD-ROM, cloth, $54.95. Computers in Physics, 10(4):349–??, July 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.4822415.

McMahon:2006:BGM

[MT06] David McMahon and Dan Topa. A beginners guide to Mathematica. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-467-3. ix + 725 pp. LCCN QA76.95 .M44 2005. Includes CD-ROM.

Miner:1994:MDA

[MTWW94] W. H. Miner, Jr., H. Y. W. Tsui, J. C. Wiley, and A. J. Wootton. Mathematica as a data analysis tool. In Anonymous [Ano94a], page 1C54.

Munro:1996:SAT

[Mun96]  N. Munro.  Symbolic alge-

bra tools for control teaching. In Anonymous [Ano96g], pages 1/1–7.

Murrell:1992:ARR

[Mur92] H. Murrell. Animation of rotating rigid bodies. Mathematica Journal, 2(1):62–65, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Murrell:1993:MGS

[Mur93] H. Murrell. A mathematical golf swing. Mathematica Journal, 3 (4):62–65, Fall 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Mende:1995:MTA

[MW95] R. Mende and O. Wanierke.

‘Mathematica’: a tool for animation in symbolic mathematics. Elektronik, 44(23):164, 166– 171, November 1995. CODEN EKRKAR. ISSN 0013-5658.

Marchak:1992:RSV

[MZ92] F. M. Marchak and D. D. Zulager. A review of scientific visualization software for the Macintosh computer. Behavior research methods, instruments, and computers, 24(2):328–335, May 1992. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Marquard:2014:SMI

[MZ14] Peter Marquard and Nikolai Zerf. SLAM, a Mathematica interface for SUSY spectrum generators. Computer Physics Communications, 185 (3):1153–1171, March 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465513004177.

Noor:1990:ASC

[NA90] Ahmed K. Noor and Carl M. Anderson. Application of symbolic computation to geometrically nonlinear analysis of curved beams. In Noor et al. [NEH90], pages 115–148. CO-DEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Noor:1992:HAT

[NA92] A. K. Noor and C. M. Andersen. Hybrid analytical technique for the nonlinear analysis of curved beams. Computers and structures, 43(5):823–830, June 1992. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Nagasaka:2007:SPM

[Nag07] Kosaku Nagasaka. SNAP package for Mathematica. ACM Communications in Computer Algebra, 41(3):105–106, September 2007. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

Naiman:2022:AVS

[Nai22] Aaron E. Naiman. Automated vector space proofs using Mathematica. ACM Communications in Computer Algebra, 56 (1):1–13, March 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL

https://dl.acm.org/doi/10.
1145/3572865.3572866.

Nam:1994:ENM

[Nam94] Tran Quoc Nam. Extended Newton’s method for finding the roots of an arbitrary system of equations and its applications. In Hamza [Ham94], pages 344– 347. ISBN 0-88986-190-0. LCCN ????

Nazarov:2012:AMM

[Naz12] Anton Nazarov. Affine.m — Mathematica package for computations in representation theory of finite-dimensional and affine Lie algebras. Computer Physics Communications, 183 (11):2480–2493, November 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465512002263.

Noor:1990:SCT

[NEH90] Ahmed K. Noor, Isaac Elishakoff, and Greg Hulbert, editors. Symbolic computations and their impact on mechanics: presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Dallas, Texas, November 25–30, 1990, volume 205 of PVP. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1990. CODEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Neunhauserer:2011:FEP

[Neu11] Jörg Neunhäuserer. A family of exceptional parameters for non-uniform self-similar measures. Electronic Communications in Probability, 16:19:192– 19:199, 2011. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/
view/1618.

Neshyba:1993:ILB

[NG93] S. P. Neshyba and R. R. Gamache. Improved linebroadening coefficients for asymmetric rotor molecules with application to ozone lines broadened by nitrogen. Journal of Quantitative Spectroscopy and Radiative Transfer, 50(5):443– 453, November 1993. CODEN JQSRAE. ISSN 0022-4073.

Nguyen:1993:SSP

[Ngu93] T. V. Nguyen. A super software package for linear systems. Computers in education journal, 3 (1):52–57, January/March 1993. CODEN CEJOE7. ISSN 1069-

Niefield:2003:IFS

[Nie03] Susan B. Niefield. Implementing finite structures in Mathematica via a skeletal topos of finite sets. Journal of Symbolic Computation, 35(2):137–151, February 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nishida:1994:QRA

[Nis94] Toyoaki Nishida. Qualitative reasoning for automated exploration for chaos. In Anonymous [Ano94f], pages 1211–1216 (vol. 2). CODEN PNAIEE. ISBN 0-262-61102-3. LCCN Q334.N36 1994. Two volumes.

Noguchi:1991:TSG

[Nog91] A. Noguchi. The two-sector general equilibrium model. Mathematica Journal, 1(3):96–103, Winter 1991. ISSN 1047-5974 (print), 1097-1610 (electronic).

Noll:1997:MIP

[Nol97] D. Justus Noll. Mathematica interaktiv: Programmierung — Grafik — Sound. Hanser, München, Germany, 1997. ISBN 3-446-18649-2. 450 pp. LCCN ????

Nordemann:1994:IAM

[Nor94] Daniel Nordemann. Introdução ao Mathematica for Windows. Transtec Editorial, ????, 1994. ISBN 85-85417-06-4. ???? pp. LCCN ????

Novotny:2017:PMG

[Nov17] Filip Novotný. ParticleRecognition, a Mathematica GUI interface for analysis of complex shaped nanoparticles in micrographs. Computer Physics Communications, 214(??):98–104, May 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/S0010465516303137.

Noyes:1995:TMN

[Noy95]  James L. Noyes.  Teaching

a modern numerical analysis course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1): 355–359, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nemes:1995:RMP

[NP95] István Nemes and Marko Petkovšek. RComp: a mathematica package for computing with recursive sequences. Journal of Symbolic Computation, 20(5–6):745–754 (or 745–753??), November–December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Nayfeh:1996:ANA

[NR96] Jamal F. Nayfeh and Nicholas J. Rivieccio. Analysis of the nonlinearity associated with the free vibration of an orthotropic shell. Proceedings of the International Conference on Engineering, Construction, and Operations in Space, 2:1115–1121, 1996.

Nethery:1994:RMP

[NS94] John F. Nethery and Mark W.

Spong. Robotica: a Mathematica package for robot analysis. IEEE Robotics & Automation Magazine, 1(1):13–20, March 1994. CODEN IRAMEB. ISSN 1070-9932.

Ozay:2024:NOR

[ÖAÖ24] S. Özay, S. Akdemir, and E. Öztekin. New orthogonality relationships of the Gaunt coefficients. Computer Physics Communications, 298 (??):Article 109118, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465524000419.

Ozay:2025:CCS

[ÖAÖ25] S. Özay, S. Akdemir, and E. Öztekin. The coupling coefficients with six parameters and the generalized hypergeometric functions. Computer Physics Communications, 315 (??):??, October 2025. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465525001584.

Ohtani:1994:CSN

[OFM94] T. Ohtani, M. Fukuzawa, and M. Masubuchi. A CAD system for nonlinear control system design using Mathematica. In Mattsson et al. [MGC94], pages 197–204. ISBN 0-7803-1800-5. LCCN TJ212.2.I3248 1994.

Ohmiya:1994:SAP

[Ohm94] Mayumi Ohmiya. Rational Darboux transformations—utilizing the computer algebra system Mathematica. (Japanese). In State of the art and perspectives in studies on nonlinear integrable systems (Japanese) (Kyoto, 1993), volume 868, pages 169–178. Sūrikaisekikenkyūsho Kōkyūroku, ????, 1994.

Olds:1992:MTW

[Old92] M. Olds. Mathematica tries the Windows environment: Wolfram’s sophisticated math package is promising but still suffers from old bugs. InfoWorld, 14: 85–86, March 2, 1992. CODEN INWODU. ISSN 0199-6649.

Olness:1998:BRW

[Oln98] F. Olness. Book review: W. Kinzel, G. Reents, Physics by Computer: Programming of Physical Problems Using Mathematica and C (1998) Springer,New York 3-540-62743-X. Computer Physics Communications, 112(2–3):270–272, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/S0010465598000484.

Ohtani:1992:CSN

[OM92] T. Ohtani and M. Masubuchi.

A CAD system for nonlinear dynamic compensators using “Mathematica”. In Strejc [Str92a], pages 364–367. ISBN 0-08-042057-5. LCCN TJ212.2 .S97 1992.

Oppenheim:1992:SKB

[ON92] Alan V. Oppenheim and S. Hamid Nawab, editors. Symbolic and Knowledge-Based Signal Processing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, April 1992. ISBN 0-13-880444-3. xviii + 340 pp. LCCN TK5102.5 .S959 1992.

Ono:1992:MDC

[Ono92] Hiroyuki Ono. Mathematica:

DSP and Control. Toppan, Tokyo, Japan, 1992. ISBN 4-8101-8549-4. 330 pp. LCCN ???? In Japanese.

ORourke:1992:MAS

[O’R92] Joseph O’Rourke. Mathematica in action (Stan Wagon). SIAM Review, 34(3):515–516, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OShea:1996:GMP

[O’S96] D. C. O’Shea. Generation of mask patterns for diffractive optical elements using Mathematica. Computers in Physics, 10 (4):391–399, July–August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Ostovic:1994:CAA

[Ost94] Vlado Ostovic. Computer-aided Analysis of Electric Machines: a Mathematica Approach. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-068859-2. xiv + 520 pp. LCCN TK2211 .O78 1994. US$39.98.

Ottmann:1994:EMH

[OT94] T. Ottmann and I. Tomek, editors. Educational multimedia and hypermedia, 1994: ED-MEDIA 94; 2nd World conference on educational multimedia and hypermedia — June 1994, Vancouver, Canada. AACE, Charlottesville, 1994. ISBN 1-880094-10-X. LCCN ????

Ozbasaran:2017:SVS

[Ozb17] Hakan Ozbasaran. solveTruss v1.0: Static, global buckling and frequency analysis of 2D and 3D trusses with Mathematica. SoftwareX, 6(??):128–134, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S235271101730016X.

Pao:1999:EAI

[Pao99] Yen-Ching Pao. Engineering Analysis: Interactive Methods and Programs with FORTRAN, QuickBASIC, MATLAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2016-X. 360 pp. LCCN TA345.P36 1999. US$79.95.

Pao:2001:EAI

[Pao01] Y. C. Pao. Engineering analysis: interactive methods and programs with FORTRAN, Quick-BASIC, MATLAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001.

Park:1993:CST

[Par93] Kyungmee Park. A comparative study of the traditional calculus course vs. the calculus and Mathematica course. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1993. x + 207 pp.

Pathak:1993:ADA

[Pat93] Dhiraj K. Pathak. Agent design for automatic use of a software system: a case study with a Soar agent for Mathematica. Thesis (Ph.D.), Carnegie Mellon University, The Robotics Institute, Pittsburgh, PA, USA, May 1993. x + 159 pp. Also issued as technical report CMU-RI-TR-93-30.

Patel:2015:PXM

[Pat15] Hiren H. Patel. Package-X: a Mathematica package for the analytic calculation of one-loop integrals. Computer Physics Communications, 197(??):276– 290, December 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465515003033.

Patel:2017:PXM

[Pat17] Hiren H. Patel. Package-X 2.0: a Mathematica package for the analytic calculation of one-loop integrals. Computer Physics Communications, 218 (??):66–70, September 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465517301297.

Polking:2006:DE

[PBA06] John C. Polking, Albert Boggess, and David Arnold. Differential Equations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2006. ISBN 0-13-143738-0. ???? pp. LCCN QA371 .P565 2006.

Powell:1993:LDS

[PC93] J. Powell and R. Crandall. Lattice dynamics for a simplified DNA model. Mathematica in Education, 2(3):15–20, 1993. ISSN 1065-2965.

Parker:1994:MTS

[PC94] Leonard Parker and Steven M. Christensen. Math Tensor: a System for Doing Tensor Analysis by Computer. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-56990-6. xv + 379 pp. LCCN QA433 .P37 1994. US$49.50.

Prieto:2000:MOA

[PDP00] Hélène Prieto, Stéphane Dalmas, and Yves Papegay. Mathematica as an OpenMath application. SIGSAM Bulletin, 34(2):22–26, June 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Perez:1995:CSN

[Pér95] César Pérez. Cálculo simbólico y numérico con Mathematica. RA-MA Editorial, ????, 1995. ISBN 84-7897-181-5. 803 pp. LCCN ???? In Spanish.

PerezSainzdeRozas:2003:UMB

[Per03] Gloria Perez Sainz de Rozas.

Using Mathematica to build non-parametric statistical tables. Journal of Statistical Software, 8(4):1–26, ???? 2003. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v08/i04/
tablasjss2.pdf;   http://
www.jstatsoft.org/v08/i4.

Peterson:1989:DMT

[Pet89] N. C. Peterson. DERIVE and Mathematica: two programs that are high achievements in computer mathematics. Science Software, 5(4):329–338, 1989. CODEN SCISE6. ISSN 0893-9101.

Petkovsek:2008:SML

[Pet08]  Marko Petkovsek.  Solving

multivariate linear recurrences in wedges (abstract only). ACM Communications in Computer Algebra, 42(1–2):18–19, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Place:1995:UCA

[PF95] I. Place and S. Fitzgerald. Using a computer algebra system (Maple) to teach elementary queueing theory. Computer Applications in Engineering Education, 3(1):65–73, 1995. CODEN CAPEED. ISSN 1061-3773.

Polyakov:1994:SNC

[PGB94] V. Polyakov, R. Ghanadan, and G. L. Blankenship. Symbolic numerical computational tools for nonlinear and adaptive control. In Mattsson et al. [MGC94], pages 117–122. ISBN 0-7803-1800-5 (softbound), 0-7803-1801-3 (microfiche). LCCN TJ 212.2 I3256 1994.

Pilipovic:1995:NSB

[PH95] Sanja Pilipović and Djordje Herceg. On numerical solving of boundary value problems in Mathematica. In IX Conference on Applied Mathematics (Budva, 1994), pages 111–126. Univ. Novi Sad, Novi Sad, Russia, 1995.

Phillips:1990:UMI

[Phi90] M. Phillips. Using Mathematica as an interface to an interactive graphics program. Mathematica Journal, 1(2):45–50, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Phillips:1991:IMV

[Phi91] R. L. Phillips. An interpersonal multimedia visualization system. IEEE Computer Graphics and Applications, 11(3):20– 27, May 1991. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Pidgeon:1996:ATB

[Pid96a] C. (Charles) Pidgeon, editor.

Advanced tutorials for biomedical scientists: animations, simulations, and calculations using Mathematica. VCH Publishers, New York, NY, USA, 1996. ISBN 1-56081-950-2. ???? pp. LCCN R859.7.C65A396 1996.

Pidgeon:1996:TBS

[Pid96b] C. (Charles) Pidgeon, editor.

Tutorials for biomedical scientists: animations, simulations, and calculations using Mathematica. VCH Publishers, New York, NY, USA, 1996. ISBN 1-56081-928-6. ???? pp. LCCN R859.7.C65T88 1996.

Piessens:1996:BRM

[Pie96] R. Piessens. Book review: Mathematica graphics techniques and applications: T. Wickham-Jones, Wolfram Research Inc., Champaign, IL, USA Springer, Berlin 1994, 721 pp., ISBN 3-540-94047-2, Hardcover, £35, DM 80. Journal of Computational and Applied Mathematics, 66(1–2):N6, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0377042796804754.

Popova:2010:CFE

[PK10] Evgenija D. Popova and Walter Krämer. Communicating functional expressions from Mathematica to C-XSC. Lecture Notes in Computer Science, 6327:354–365, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-15582-6_
56.pdf.

Page:1992:CSS

[PL92] Connie Page and Raoul LePage, editors. Computing science and statistics: statistics of many parameters: curves, images, spatial models: Proceedings of the 22nd Symposium on the Interface: Computing Science and Statistics, held May 16–19, 1990, at the Kellogg Center on the campus of Michigan State University. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97719-8. LCCN QA276.4.C592 1991.

Pichler:1992:CAS

[PMD92] Franz Pichler and Roberto Moreno-Diaz, editors. Computer aided systems theory, EURO-CAST ’91: a selection of papers from the Second International Workshop on Computer Aided Systems Theory, Krems, Austria, April 15–19, 1991: proceedings, volume 585 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55354-1 (New York), 3-540-55354-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1991.

Pichler:1996:CAS

[PMDA96] F. Pichler, R. Moreno-Diaz, and R. Albrecht, editors. Computer Aided Systems Theory — EU-ROCAST ‘95. A Selection of Papers from the Fifth International Workshop on Computer Aided Systems Theory. Proceedings: Innsbruck, Austria, May 22–25, 1995. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60748-X. LCCN TA345 .I62 1995.

Poitevineau:1998:MPP

[Poi98] Yvon Poitevineau. Mathematica 3: par la pratique. Eyrolles, Paris, France, 1998. ISBN 2-212-08979-1. ix + 358 pp. LCCN ????

Preti:2018:WMP

[Pre18] M. Preti. WiLE: A Mathematica package for weak coupling expansion of Wilson loops in ABJ(M) theory. Computer Physics Communications, 227 (??):126–147, June 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465517304162.

Pria:1993:MPI

[Pri93] A. Pria. ‘Mathematica’, a powerful informatics environment for the solution of engineering problems. Revista española de electronica, June 1993. CODEN RVEEBT. ISSN 0482-6396.

Provasi:1996:UAS

[Pro96] C. Provasi. Using algebraic software to compute the moments of order statistics. In Anonymous [Ano96h], pages 199–213. CODEN CNOMEL. ISSN 0927-7099.

Pai:1993:SCR

[PS93] Dinesh K. Pai and Tony H. S. Ser. Simultaneous computation of robot kinematics and differential kinematics with automatic differentiation. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0823-9. 775–780 pp. LCCN TJ210.3.I447 1993. IEEE catalog number 93CH3213-6.

Paule:1995:MVZ

[PS95] P. Paule and M. Schorn. A Mathematica version of Zeilberger’s algorithm for proving binomial coefficient identities. Journal of Symbolic Computation, 20(5-6):673–698, November–December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pemmaraju:2003:CDM

[PS03] Sriram V. Pemmaraju and Steven S. Skiena. Computational discrete mathematics: combinatorics and graph theory with Mathematica. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-80686-0. xiii + 480 pp. LCCN QA164 .P45 2003.

Petkovic:2006:IAL

[PS06] Marko D. Petković and Predrag S. Stanimirović. Interpolation algorithm of Leverrier– Faddeev type for polynomial matrices. Numerical Algorithms, 42(3–4):345–361, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=1017-
1398&volume=42&issue=3&spage=
345.

Preyer:1995:UME

[PSMM95] J. A. Preyer, D. R. Smith, and R. Malek-Madani. Using Mathematica to enhance learning of atmospheric processes: Entrainment into cumulus clouds. In Anonymous [Ano95d], pages 41– 43. ISBN ???? LCCN ????

Pittenger:1993:GAS

[PSW93] K. Pittenger, D. Schlimingmann, and D. Whitecotton. Graphical and audio simulations using Mathematica. Mathematica in Education, 2(4):3–6, 1993. ISSN 1065-2965.

Pawlowski:2023:QDM

[PSW23] Jan M. Pawlowski, Coralie S.

Schneider, and Nicolas Wink.

QMeS-Derivation: Mathematica

package for the symbolic derivation of functional equations. Computer Physics Communications, 287(??):Article 108711, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/S0010465523000565.

Picariello:2004:STM

[PTL04] Marco Picariello and Emilio Torrente-Lujan. Slavnov-Taylor1.0: a Mathematica package for computation in BRST formalism. Computer Physics Communications, 156(2):171– 179, January 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465503004624.

Packel:1994:RPA

[PW94a] E. Packel and S. Wagon. Rearrangement patterns for the alternating harmonic series. Mathematica in Education, 3(2):5–10, Spring 1994. ISSN 1065-2965.

Packel:1994:ACM

[PW94b] Edward W. Packel and Stan Wagon. Animating Calculus: Mathematica Notebooks for the Laboratory. W. H.Freeman and Company, New York, NY, USA, 1994. ISBN 0-7167-2428-6. xii + 289 pp. LCCN QA303.5.D37P33 1994.

Packel:1997:ACM

[PW97] Edward W. Packel and S. Wagon. Animating calculus: Mathematica notebooks for the laboratory. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, revised edition, 1997. ISBN 0-387-94748-5 (soft cover). xiv + 292 + 2 pp. LCCN QA 303.5 C65 P33 1997.

Pointer:1990:MDS

[PWW90] S. Pointer, J. Wawrzynek, and D. Wessel. A multimedia digital signal processing tutoring system. In Arnold and Hair [AH90], pages 35–38. LCCN ML1381 .I6 1990.

Postel:1996:ROS

[PZ96] F. Postel and P. Zimmermann.

A review of the ODE solvers of Maple, Mathematica, Macsyma and MuPAD. In Carrière [Car96], page ???? ISBN ???? LCCN ???? updated version includes Axiom, Derive and Reduce.

Rowe:1995:DWM

[RA95] Alistair C. H. Rowe and Paul C.

Abbott. Daubechies wavelets and Mathematica. Computers in Physics, 9(6):635–648, November–December 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Raade:1994:SSG

[Raa94] L. Raade. (SAT-S94) stochastic geometry and Mathematica. In Lum [Lum94], pages 276– 280. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Rade:1992:PLM

[Rad92] L. Rade. A probability lesson with Mathematica. Mathematica in Education, 1(4):14–16, Summer 1992. ISSN 1065-2965.

Rade:1993:PM

[Rad93a] L. Rade. Probability and Mathematica. In Lum [Lum93], pages 142–146. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Rade:1993:RPC

[Rad93b] L. Rade. Random points in circles and Mathematica. International journal of mathematical education in science and technology, 24(3):329–336, May 1993. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Raade:1993:RPC

[Råd93c] Lennart Råde. Random points in circles and Mathematica. International journal of mathematical education in science and technology, 24:329–335, 1993. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Rade:1993:RSE

[Rad93d] Lennart Rade. Reliability survival equivalence. Microelectronics and Reliability, 33(6):881– 894, May 1993. CODEN MCR-LAS. ISSN 0026-2714 (print), 1872-941X (electronic).

Rade:1995:MUA

[Rad95] L. Rade. Marbles in urns, applied probability and computer algebra. European Journal of Engineering Education, 20(3):347–352, 1995. CODEN EJEED8. ISSN 0304-3797.

RamirezGonzalez:1997:MCM

[Ram97] Victoriano Ramirez Gonzalez.

Matematicas con Mathematica: introduccion a Mathematica y primeras aplicaciones. Proyecto Sur, Armilla, Spain, 2. rev. edition, 1997. ISBN 84-8254-107-2. x + 286 pp. LCCN ????

Raouf:1993:QAN

[Rao93]  R. A. Raouf.  A qualita-

tive analysis of the nonlinear dynamic characteristics of curved orthotropic panels. Composites engineering, 3(12):1101– 1110, 1993. CODEN CMENEZ. ISSN 0961-9526.

Rao:1998:BRB

[Rao98] T. Subba Rao. Book review:

Computational Economics and Finance-Modelling and Analysis with Mathematica, by H. R. Varian. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (3):558–559, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/2988658.

Rashid:1991:CCS

[Ras91] Richard F. Rashid, editor. CMU Computer Science: a 25th Anniversary Commemorative. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52899-1. xxxii + 519 pp. LCCN QA75.5 .C548 1990.

Ravi:2007:BRBc

[Rav07] Sreenivasan Ravi. Book review: Bayesian Logical Data Analysis for the Physical Sciences: A Comparative Approach with Mathematica^{(}R) Support, by P. C. Gregory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1182–1183, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic).  URL http://www.
jstor.org/stable/4623237.

Rogers:1993:FST

[RB93] Steven Hill Rogers and Adedeji B. Badiru. A fuzzy set theoretic framework for knowledgebased simulation. Computers & industrial engineering, 25(1–4): 119–122, September 1993. CO-DEN CINDDL. ISSN 0360-8352 (print), 1879-0550 (electronic). Proceedings of the 15th Annual Conference on Computers and Industrial Engineering.

Rehm:1995:MCG

[RB95] R. G. Rehm and H. R. Baum. Modeling of combustion of a gaseous sphere using Mathematica. Lecture notes in physics, ??(449):209–220, 1995. CO-DEN LNPHA4. ISSN 0075-8450 (print), 1616-6361 (electronic).

Ramsden:2003:CBS

[RCM03] Philip Ramsden, Janet Carne, and P. (Peter) Mitic, editors. Challenging the boundaries of symbolic computation: proceedings of the 5th International Mathematica Symposium. Imperial College Press, London, UK, 2003. ISBN 1-86094-363-2. xiii + 420 pp. LCCN QA76.73.M29 I58 2003. Includes CD-ROM.

Riddle:1994:AEE

[RD94] Alfred Riddle and James Samuel Dick. Applied Electronic Engineering with Mathematica. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-53477-0 (package) 0-201-82689-5 (book) 0-201-94252-6 (disk). xv + 375 pp. LCCN TK7835 .R525 1994.

Richards:1996:CPH

[RdG96] B. Richards and H. de Glanville, editors. Current Perspectives in Healthcare Computing Conference. BJHC, Weybridge, UK, 1996. ISBN 0-948198-24-9. LCCN ????

Reed:1998:BRB

[Ree98] R. J. Reed. Book review: The Beginner’s Guide to Mathematica Version 3, by J. Glynn; T. W. Gray. Journal of the Royal Statistical Society. Series D (The Statistician), 47 (4):715, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/2988384.

Reed:2000:BRBb

[Ree00] R. J. Reed. Book review:

The Student’s Introduction to MATHEMATICA: A Handbook for Precalculus, Calculus and Linear Algebra, by B. F. Torrence; E. A. Torrence. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):635, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2681056.

Reiter:1993:LAL

[Rei93] C. Reiter. Linear algebra and laboratories. Mathematica in Education, 2(2):7–10, Winter 1993. ISSN 1065-2965.

Reilly:1995:ESM

[Rei95] C. Reilly. Exploring specifications with Mathematica. Lecture Notes in Computer Science, 967:408–422, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

StaffReview:1993:MW

[Rev93] Staff Review. Mathematica for Windows. Social science computer review, 11(2):254–255, Summer 1993. CODEN SS-CREH. ISSN 0894-4393.

Rossi:1995:SAR

[RF95] M. Rossi and A. P. Flitney. Symbolic algebra and renormalization of gauge theories. Computer Physics Communications, 90(2-3):189–200, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Rutz:1996:AGG

[RF96] O. E. Rutz and P. M. Ferreira. Algebraic geometry and group theory in geometric constraint satisfaction for computer-aided design and assembly planning. IIE Transactions, 28(4):281– 294, April 1996. CODEN IIETDM. ISSN 0740-817X (print), 1573-9724 (electronic).

Rial:1991:GTM

[Ria91]  J. Rial.  Geomathematics:

Teaching math and geoscience with Mathematica. Mathematica in Education, 1(2):9–15, Fall 1991. ISSN 1065-2965.

Rich:1991:CS

[Ric91] J. Rich. Computing software.

Automotive Engineering, 99:69, April 1991. ISSN 0097-711X.

Riddle:1990:BSC

[Rid90a] A. Riddle. Benefits of symbolic computations in mechanoelectrical network analysis. In Noor et al. [NEH90], pages 327– 337. CODEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Riddle:1990:CPS

[Rid90b] A. Riddle. A CAD program for the symbolic and numerical analysis of microwave electronic circuit. In Agarwal [Aga90], pages 355–358. LCCN TK 7876 I11 1990. Three volumes.

Riddle:1990:NCA

[Rid90c] A. Riddle. A nodal circuit analysis program. Mathematica Journal, 1(1):62–68, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Rudolph-Lilith:2019:CMT

[RL19] Michelle Rudolph-Lilith. ChessY: a Mathematica toolbox for the generation, visualization and analysis of positional chess graphs. SoftwareX, 9(??):39– 43, January/June 2019. CO-DEN ???? ISSN 2352-

7110.  URL http://www.
sciencedirect.com/science/
article/pii/S2352711018301687.

Romano:2006:CMU

[RLM06] Antonio Romano, Renato Lancellotta, and Addolorata Marasco. Continuum mechanics using Mathematica: fundamentals, applications, and scientific computing. Modeling and simulation in science, engineering, and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 0-8176-3240-9. xii + 388 pp. LCCN QA808.2 .R66 2004. Includes CD-ROM.

Roanes-Lozano:2004:GAS

[RLRML04] Eugenio Roanes-Lozano, Eugenio Roanes-Macías, and Luis M. Laita. The geometry of algebraic systems and their exact solving using Gröbner bases. Computing in Science and Engineering, 6(2):76–79, March/ April 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2004/02/c2076abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/c2076.
htm;   http://csdl.computer.

org/dl/mags/cs/2004/02/c2076.
pdf.

Rodriguez-Millan:1996:BAE

[RM96] J. Rodriguez-Millan. Basins of attraction estimation through symbolic graphical computing techniques. In Pichler et al. [PMDA96], pages 137–150. ISBN 3-540-60748-X. LCCN TA345 .I62 1995. Rodriguez-Millan:1996:SCA

[RMC96] J. Rodriguez-Millan and J. Cardillo.

Symbolic computing aided design of nonlinear PID controllers. In Pichler et al. [PMDA96], pages 253–270. ISBN 3-540-60748-X. LCCN TA345 .I62 1995.

Roach:1996:HFR

[Roa96] Kelly Roach. Hypergeometric function representations. In Lakshman Y. N. [Lak96], pages 301–308. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/236869/
p301-roach/.

Robertson:1995:EMM

[Rob95] John S. Robertson. Engineering mathematics with Mathematica. International series in pure and applied mathematics. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-053171-4. xiv + 288 pp. LCCN TA345.R55 1995.

Rogowitz:1992:HVV

[Rog92] Bernice Ellen Rogowitz, editor.

Human vision, visual processing, and digital display III: Papers from a meeting held 10–13 February 1992, San Jose, California, volume 1666 of Proceedings of SPIE—the International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, February 1992. ISBN 0-8194-0820-4. LCCN TS510.S63 v.1666.

Rohrl:1999:BRB

[Röh99] Armin Röhrl. Book review: The Mathematica Primer, by K. R. Coombes; B. R. Hunt; R. L. Lipsman; J. E. Osborn; G. K. Stuck. Journal of the Royal Statistical Society. Series D (The Statistician), 48 (3):452–453, ???? 1999. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/2681014.

Roman:1994:IWM

[Rom94] T. A. Roman. The inflating wormhole: a MATHEMAT-ICA animation. Computers in Physics, 8(4):480–??, 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Romano:2009:GOT

[Rom09] Antonio Romano. Geometric optics: theory and design of astronomical optical systems using Mathematica, volume 4960 of Modeling and simulation in science, engineering and technology. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2009. ISBN 0-8176-4871-2 (hardcover). xi + 224 pp. LCCN QC383 .R66 2009.

Ross:1995:DEI

[Ros95]  Clay C. Ross.  Differential

Equations: An Introduction with Mathematica. Textbooks in mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94301-3 (New York), 3-540-94301-3 (Berlin). xiii + 503 pp. LCCN QA371 .R575 1995.

Ross:2004:DEI

[Ros04]  Clay C. Ross.  Differential

equations: an introduction with Mathematica. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 0-387-21284-1. xiii + 431 pp. LCCN QA371 .R595 2004.

Rosiek:2016:MMP

[Ros16]  Janusz Rosiek.  MassToMI

— a Mathematica package for an automatic Mass Insertion expansion. Computer Physics Communications, 201 (??):144–158, April 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S001046551500449X.

Raouf:1992:NLF

[RP92] R. A. Raouf and A. N. Palazotto. Non-linear free vibrations of symmetrically laminated, slightly compressible cylindrical shell panels. Composite Structures, 20(4):249–257, 1992. CODEN COMSE2. ISSN 0263-8223.

Robin:1996:TTE

[RPWW96] B. Robin, J. D. Price, J. Willis, and D. A. Willis, editors. Technology and Teacher Education Annual, 1996 Proceedings of SITE 96 — Seventh International Conference of the Society for Information Technology and Teacher Education (SITE). Assoc. Advancement of Comput. Educ, Charlottesville, VA, USA, 1996 xvi+ 1996.

Rinaldo:1992:MEF

[RRIR^{+}92] A. Rinaldo, I. Rodriguez-Iturbe, R. Rigon, R. Bras, E. Ijjasz-Vasquez, and A. Marani. Minimum energy and fractal structure of drainage networks. Water Resources Research, 28(9): 2183–2196, September 1992.

Redish:1990:CCP

[RRM90] Edward F. Redish, John S. Risley, and Nancy Margolis, editors. The Conference on Computers in Physics Instruction: proceedings. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-16306-3. LCCN QC30 .C636 1988.

Rose:2002:MSM

[RS02] Colin Rose and Murray D. Smith. Mathematical statistics with Mathematica. Springer texts in statistics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95234-9. xiii + 481 pp. LCCN QA276.4 .R67 2002. Includes CD-ROM.

Raz:2026:EAU

[RSL^{+}26] Tomer Raz, Michael Shalyt, Elyasheev Leibtag, Rotem Kalisch, Shachar Weinbaum, Yaron Hadad, and Ido Kaminer. From Euler to AI: Unifying formulas for mathematical constants. arXiv.org, ?? (??):1–82, February 24, 2026.

URL https://arxiv.org/abs/
2502.17533;   https://arxiv.
org/pdf/2502.17533.

Rueda:2007:BRB

[Rue07] Raul Rueda. Book review:

Bayesian Logical Data Analysis for the Physical Sciences: A Comparative Approach with Mathematica Support by P. C. Gregory. Hardcover: 486 pages. Cambridge University Press. ISBN 0-521-84150-X, $75.00. Bulletin of Mathematical Biology, 69(3):1119–1120, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/
10.1007/s11538-006-9144-2;
 http://link.springer.com/
content/pdf/10.1007/s11538-
006-9144-2.pdf.

Rundel:1989:MNS

[Run89] R. D. Rundel. Mathematica: a new standard in equation processing. Science Software, 5 (4):339–342, 1989. CODEN SCISE6. ISSN 0893-9101.

Ruskeepaa:1999:MNG

[Rus99] Heikki Ruskeepaa. Mathematica navigator: graphics and methods of applied mathematics. Academic Press, New York, NY, USA, 1999. ISBN 0-12-603640-3 (paperback manual), 0-12-603641-1 (CD-ROM). xx + 848 pp. LCCN QA76.95 .R87 1999. The CD-ROM contains the entire book, including all animations and data sets.

Ruskeepaa:2004:MNM

[Rus04] Heikki Ruskeepää. Mathematica navigator: mathematics, statistics, and graphics. Elsevier Academic Press, Amsterdam, The Netherlands, second edition, 2004. ISBN 0-12-603642-X , 0-12-603643-8 (CD-ROM). xx + 844 pp. LCCN QA76.95 .R87 2004.

                                        URL http://www.loc.gov/
                                        catdir/description/els041/
                                        2003058330.html;   http://
                                        www.loc.gov/catdir/toc/els051/
                                        2003058330.html.

Ruskeepaa:2009:MNM

[Rus09] Heikki Ruskeepää. Mathematica navigator: mathematics, statistics, and graphics. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2009. ISBN 0-12-374164-5 (paperback). xxii + 1111 pp. LCCN QA76.95 .R87 2009.

Ryne:1993:CAP

[Ryn93] Robert Ryne, editor. Computational accelerator physics: Feb. 22–26, 1993 in Los Alamos, New Mexico, volume 297 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 1-56396-222-5. LCCN QC787.L5 C66 1993.

Safer:1991:MLH

[Saf91] A. Safer. Math labs [in high school]. Electronic Learning, 10:54, March 1991. CODEN ELEADA. ISSN 0278-3258.

Saha:1996:NET

[Sah96] R. K. Saha. Note on the evaluation of the total integral for continuous systems. IEEE Transactions on Aerospace and Electronic Systems, 32(2):840–842, April 1996. CODEN IEARAX. ISSN 0018-9251.

Sakakibara:1995:BGW

[Sak95] Susumu Sakakibara. A Beginner’s Guide to Wavelets. Tokyo Denki Daigaku, Tokyo, Japan, 1995. ISBN 4-501-52270-4. 225 pp. LCCN ???? In Japanese.

Salvy:1989:EAA

[Sal89] B. Salvy. Examples of automatic asymptotic expansions. Technical Report 114, Inst. Nat. Recherche Inf. Autom., Le Chesnay, France, December 1989. 18 pp.

Salvy:1991:EAA

[Sal91] B. Salvy. Examples of automatic asymptotic expansions. SIGSAM Bulletin, 25(2):4–17, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Salgado:1994:GRC

[Sal94] M. Salgado. 3+1 general relativity by computer. Computer Physics Communications, 79(2):309–328, April 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Samson:1994:ORT

[Sam94]  J. R. Samson, Jr.  Opti-

mizing reliability in a twolevel distributed architecture for wafer scale integration. In Lea and Tewksbury [LT94], pages 292–314. ISBN 0-7803-1850-1 (casebound), 0-7803-1849-8 (softbound), 0-7803-1851-X (microfiche). LCCN TK 7874 I578 1994.

Sandbakken:1995:API

[San95] Jan Sandbakken. Analyzing prime and irreducible algebraic integers with Mathematica (Norwegian). Normat, 43(1):1– 17, 1995. ISSN 0029-1412, 0801-3500.

Sanns:1999:MWI

[San99] Werner Sanns. Mathematik für Wirtschaftswissenschaftler und Ingenieure mit Mathematica. R. Oldenbourg, München, Germany, 1999. ISBN 3-486-25074-4. ca. 330 pp. LCCN ????

Sarid:1997:ESP

[Sar97] Dror Sarid. Exploring scanning probe microscopy with Mathematica. John Wiley and Sons, Inc., Chichester, UK, 1997. ISBN 0-471-16818-1 (cloth). xx + 262 pp. LCCN QH212.S33 S27 1997.

Savory:1995:UMA

[Sav95] Paul A. Savory. Using Mathematica to aid simulation analysis. In Alexopoulos et al. [AKLG95], pages 1324–1328. CODEN WSCPDK. ISBN 0-7803-3018-8, 0-7803-3017-X. ISSN 0275-0708, 0743-1902. LCCN QA76.9.C65 W56 1995. IEEE Catalog No. 95CB35865.

Sakakibara:1992:MPA

[SB92] S. Sakakibara and N. Blachman. Mathematica: a Practical Approach (Japanese translation). Toppan and Prentice-Hall, Tokyo, Japan and Englewood Cliffs, NJ 07632, USA, 1992. ISBN 4-8101-8538-9. ???? pp. LCCN ????

Smith:1995:MGG

[SB95] Cameron Smith and Nancy Blachman. The Mathematica Graphics Guidebook. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-53280-8, 0-201-82656-9 (book), 0-201-82655-0 (disk). xxi + 339 pp. LCCN T385 .S622 1995. US$39.75.

Sakanaka:1994:ICP

[SBA94] P. H. Sakanaka, E. Del Bosco, and M. V. Alves, editors. 1994 International Conference on Plasma Physics. Joint Conference of the 10th Kiev International Conference on Plasma Theory, 10th International Congress on Waves and Instabilities in Plasmas, Combined with 6th Latin American Workshop on Plasma Physics. Proceedings. INPE, Sao Jose des Campos, Brazil, 1994. 3 vol.

Shacham:1996:AFC

[SBP96] M. Shacham, N. Brauner, and M. Pozin. Application of feedback control principles for solving differential-algebraic systems of equations in process control education. In Anonymous [Ano96d], pages S1329– 1334. CODEN CCENDW. ISSN 0098-1354.

Squire:2014:VAV

[SBQ14] J. Squire, J. Burby, and H. Qin. VEST: Abstract vector calculus simplification in Mathematica. Computer Physics Communications, 185(1):128–135, January 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465513002944.

Stosic:2009:UMM

[SC09] Borko D. Stosić and Gauss M. Cordeiro. Using Maple and Mathematica to derive bias corrections for two parameter distributions. Journal of Statistical Computation and Simulation, 79(6):751–767, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarid:2010:MIS

[SC10] Dror Sarid and William Albert Challener. Modern introduction to surface plasmons: theory, Mathematica modeling, and applications. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76717-2 (hardcover). xiv + 371 pp. LCCN QC176.8.P55 S37

2010. URL http://assets.
cambridge.org/97805217/67170/
cover/9780521767170.jpg.

Scales:1994:GIP

[Sca94] J. Scales. Geopsychical inverse problems. Mathematica in Education, 3(2):11–19, Spring 1994. ISSN 1065-2965.

Schulman:1990:UUP

[Sch90] Andrew Schulman. DOS unbounded: uses of protected mode. BYTE Magazine, 15(11): 250–256, 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schwartz:1992:CDH

[Sch92] K. D. Schwartz. Computerdriven humvee is DOT’s own Terminator. (Department of Transportation uses computerdriven jeep with vision processing systems to study traffic patterns). Government computer news, 11:55, November 1992. ISSN 0738-4300.

Schaper:1994:GM

[Sch94a] Ralph Schaper. Grafik mit Mathematica. Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-612-9. ???? pp. LCCN ????

Schneider:1994:RRS

[Sch94b] Dean L. Schneider. Review of the Robotica software package for robotic manipulators. IEEE Robotics & Automation Magazine, 1(1):21–22, March 1994. CODEN IRAMEB. ISSN 1070-9932.

Schweitzer:1995:QAG

[Sch95] Gunther Schweitzer. Quickly available graphics of static, electromagnetic field distributions given by conformal maps using Mathematica. Thesis, Technische Universität Graz, Graz, Austria, January 1995. iii + 134 pp.

Schnitz:1998:LPA

[Sch98a] Sarah Schnitz. Linear programming: an application in Mathematica. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1998. i + 5 pp.

Schwalbe:1998:RCE

[Sch98b]  Daniel Schwalbe.  Reviews:

Computational Economics and Finance: Modeling and Analysis with Mathematica, by Hal R. Varian. American Mathematical Monthly, 105(5):477–481, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_
may98_toc.html.

Schmitt:2002:ASM

[Sch02] Hermann Schmitt. An application server for Mathematica. In Anonymous [Ano02], page ?? ISBN ???? LCCN ????

Scott:1991:EM

[Sco91]  D. S. Scott.  Exploration

with Mathematica. In Rashid [Ras91], pages 505–519. ISBN 0-201-52899-1. LCCN QA75.5 .C548 1990.

Sparks:1993:ICF

[SDB93a] A. G. Sparks, J. W. Davenport, and J. P. Braselton. Implementing calculus as formal laboratory courses using Mathematica. In Lum [Lum93], pages 408– 412. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Sparks:1993:CLU

[SDB93b] Arthur G. Sparks, John W. Davenport, and James P. Braselton. Calculus Labs using Mathematica. HarperCollins College Publishers, New York, NY, USA, 1993. ISBN 0-06-501196-1. 250 (est.) pp. LCCN QA303.5.D37 S63 1993.

Schost:2011:IPI

[SE11] Éric Schost and Ioannis Z.

Emiris, editors. ISSAC 2011: Proceedings of the 2011 International Symposium on Symbolic and Algebraic Computation, June 7–11, 2011, San Jose, CA, USA. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Seiter:1990:MV

[Sei90] Charles Seiter. Mathematica, version 1.2. Macworld, ??(??): 158–165, January 1990. CODEN MACWEA. ISSN 0741-8647.

Seiter:1992:MAF

[Sei92] C. Seiter. Mathematics analysis: five math programs that help scientists and engineers solve complex computations. Macworld, 9:172–175, June 1992. CODEN MACWEA. ISSN 0741-8647.

Seiter:1993:M

[Sei93] C. Seiter. Mathematica 2.2. Macworld, 10(9):63–65, September 1993. CODEN MACWEA. ISSN 0741-8647.

Senn:1995:CRP

[Sen95] P. Senn. The computation of RKR potential energy curves of diatomic molecules using Mathematica. Computers and Chemistry, 19(4):437–??, 1995. CO-DEN COCHDK. ISSN 0097-8485.

Sfeir:1993:MNA

[Sfe93]  A. A. Sfeir.  Mathematica:

Une Nouvelle Approche du Calcul Scientifique. Angkor Editeur and RITME Informatique, Paris, France, 1993. ISBN 2-87892-005-8. 232 pp. LCCN ???? In French.

Sfeir:1998:MLM

[Sfe98] A. A. Sfeir. Mathematica 3: langage et méthodes. Global Design: Ritme Informatique, Paris, France, 1998. ISBN 2-911502-05-1. iv + 289 pp. LCCN ????

Strampp:1995:DM

[SG95] W. Strampp and V. Ganzha.

Differentialgleichungen mit Mathematica. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1995. ISBN 3-528-06618-0. ???? pp. LCCN ????

Strong:1995:UME

[SGMM95] B. M. Strong, C. R. Gunderson, and R. Malek-Madani. Using Mathematica to enhance learning of oceanographic processes: Wind-driven circulation. In Anonymous [Ano95d], pages 44– 46. ISBN ???? LCCN ????

Skidmore:1998:GTD

[SH98] Alexandra Skidmore and Margie Hale. A guided tour of differential equations using computer technology. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-592767-6. ca. 100 pp.

Steeb:2010:QMU

[SH10] Willi-Hans Steeb and Yorick Hardy. Quantum mechanics using computer algebra: including sample programs in C++, SymbolicC++, Maxima, Maple, and Mathematica. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2010. ISBN 981-4307-16-5 (hardcover). x + 234 pp. LCCN QC174.17.D37 S74 2010.

Shampine:1996:BRD

[Sha96] Lawrence F. Shampine. Book reviews: Differential equations with Mathematica, by Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. Mathematics of Computation, 65(215):??, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
jourcgi/jour-pbprocess?fn=
110&arg1=S0025-5718-96-00705-
3&u=/mcom/1996-65-215/.

Shaw:1998:MFD

[Sha98] William T. Shaw. Modelling financial derivatives with Mathematica: mathematical models and benchmark algorithms. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-59233-X. xi + 537 pp. LCCN HG6024.A3 S533 1998. Includes CD-ROM.

Shaw:2006:CAM

[Sha06] William T. Shaw. Complex analysis with Mathematica. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-83626-3 (hardcover). xxv + 571 pp. LCCN QA76.95 .S523 2006 UCB. Includes CD-ROM.

Sating:1993:DMQ

[SHH93] Richard R. Sating, C. H. Houpis, and I. M. Horowitz. Development of a MIMO QFT CAD package (version 2). In American Control Conference (Jun 2–4 1993: San Francisco, CA, USA), pages 3081–3083 (of xxxviii + 3201). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0861-1, 0-7803-0860-3 (paperback). LCCN TJ212 .A53 1993. IEEE catalog number 93CH3225-0.

Shirer:1997:SLL

[Shi97] Donald L. Shirer. Symbolic I/O leads list of new Mathematica 3.0.1 features. Computers in Physics, 11(6):624– ??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4822613.

Shirer:2002:MSW

[Shi02]  Donald L. Shirer.  Mathe-

matica spins Web wizardry. Computing in Science and Engineering, 4(2):9–10, March/ April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/computer.org/cise/cs2001/
c2009abs.htm;   http://dlib.
computer.org/cs/books/cs2002/
pdf/c2009.pdf.

Shingareva:2009:MMP

[Shi09] Inna K. Shingareva. Maple and Mathematica: a problem solving approach for mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 3-211-99431-9 (softcover). xviii + 483 pp. LCCN ????

Shriver:1991:NPO

[Shr91] J. W. Shriver. NMR productoperator calculations in Mathematica. Journal of Magnetic Resonance, 94(3):612–616, October 1991. CODEN JOMRA4. ISSN 0022-2364.

Shute:1993:AAB

[Shu93] Malcolm J. Shute. Abotec: an automatic back-of-the-envelope calculator. ACM SIGPLAN Notices, 28(8):90–98, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shuchat:1994:JMP

[Shu94] Alan Shuchat. The Joy of Mathematica: a Point-and-click Way to Use and Learn Mathematica. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-59145-6 (package), 0-201-56723-7 (book), 0-201-56724-5 (disk). xviii + 318 pp. LCCN QA76.95 .S527 1994.

Simon:1989:WMW

[Sim89] B. Simon. Wolfram’s Mathematica: wonderful, revolutionary and bug-ridden. (evaluation). PC Magazine, pages 33– 35, March 28, 1989. CODEN PCMGEP. ISSN 0888-8507.

Simon:1990:NWH

[Sim90] Barry Simon. The new world of higher math. Mathematica, version 1.2. PC Magazine, pages 323–329, May 29, 1990. CODEN PCMGEP. ISSN 0888-8507.

Simon:1991:MAS

[Sim91] B. Simon. The mathematical alternative. (software review) (technical graphing capabilities of mathematical software packages). PC Magazine, pages 154– 155, March 26, 1991. CODEN PCMGEP. ISSN 0888-8507.

Simon:1992:JMA

[Sim92a] B. Simon. It’s not just for mainframes anymore: symbolic math software. PC Magazine, 11: 405–420, August 1992. CODEN PCMGEP. ISSN 0888-8507.

Simon:1992:MWR

[Sim92b] B. Simon. Mathematica for Windows: reshaping math software with graphics and sound. PC Magazine, 11:37–38, March 31, 1992. CODEN PCMGEP. ISSN 0888-8507.

Simon:1992:WRI

[Sim92c] B. Simon. Wolfram Research Inc.: Mathematica for DOS, Mathematica for windows. PC Magazine, 11:425–429, August 1992. CODEN PCMGEP. ISSN 0888-8507.

Simon:1992:SMS

[Sim92d] Barry Simon. Symbolic math software: It’s not just for mainframes anymore. PC Magazine, 11(14):405–??, August 1992. CODEN PCMGEP. ISSN 0888-8507.

Simon:1993:EST

[Sim93] B. Simon. The essential scientific toolkit. Windows Sources, 1(8):307–318, September 1993.

Sit:1997:MNC

[Sit97] W. Y. Sit. Mathematica notebooks for a conventional differential equations course. Journal of Symbolic Computation, 23 (5–6):589–623, May–June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Skeel:1993:ENC

[SK93] Robert D. Skeel and Jerry B. Keiper. Elementary Numerical Computing with Mathematica. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-057820-6. xiv + 434 pp. LCCN QA297 .S545 1993. See [SK95].

Skeel:1995:ENC

[SK95] Robert D. Skeel and Jerry B. Keiper. Elementary Numerical Computing with Mathematica. Kyoritsu Suppan, ??, Japan, 1995. ISBN 4-320-01488-X. 410 pp. LCCN ???? Japanese translation of [SK93].

Skiena:1990:IDM

[Ski90] Steven S. Skiena. Implementing Discrete Mathematics: Combinatorics and Graph Theory with Mathematica, with programs by Steven Skiena and Anil Bhansali. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-50943-1. x + 334 pp. LCCN QA164 .S56 1990. See [Skixx].

Skiena:19xx:IDM

[Skixx] Steven S. Skiena. Implementing Discrete Mathematics: Combinatorics and Graph Theory with Mathematica. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 19xx. ISBN 4-8101-8050-6. ???? pp. LCCN ???? Japanese translation of [Ski90].

Skottun:1994:APP

[Sko94] B. C. Skottun. On amplitude and phase in printed characters. Mathematica Journal, 4(2):83– 86, Spring 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Skrinar:1993:UPM

[Skr93] Matjaz Skrinar. Use of program MATHEMATICA in the finite element method for thin plate analysis. Strojniski Vestnik, 39 (1-2):35–42, January–February 1993. CODEN STJVAX. ISSN 0039-2480.

Seliger:1995:RNO

[SKS95] R. Seliger and B. Koeppen-Seliger. Robust nonlinear observer-based fault detection for an U-tube steam generator. Proceedings of the American Control Conference, 2:1134– 1135, 1995. CODEN PRACEO. ISSN 0743-1619.

Slaney:1990:ISP

[Sla90] Malcolm Slaney. Interactive signal processing documents. IEEE ASSP magazine: a publication of the IEEE Acoustics, Speech, and Signal Processing Society, 7(2):8–20, April 1990. CO-DEN IAMAEI. ISSN 0740-7467 (print), 1558-1284 (electronic).

Shingareva:2007:MMP

[SLC07] Inna Shingareva and Carlos Lizárraga-Celaya. Maple and Mathematica: a problem solving approach for mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-211-73265-9, 3-211-73264-0. xiii + 263 pp. LCCN QA155.7.E4 S55 2007eb.

Sloan:1993:TRC

[Slo93]  J. Sloan.  Two rectangles

are constructible with tangrams: An enumeration proof using Mathematica. Mathematica in Education, 2(4):3–6, 1993. ISSN 1065-2965.

Su:1995:ACT

[SLR^{+}95] Ernesto Su, Antonio Lain, Shankar Ramaswamy, Daniel J. Palermo, Eugene W. Iv Hodges, and Prithviraj Banerjee. Advanced compilation techniques in the PARADIGM compiler for distributed-memory multicomputers. Proceedings of the International Conference on Supercomputing, pages 424–433, 1995.

Snowman:1998:BRJ

[SMC98] Daniel P. Snowman, Susan R.

Mackay, and Wolfgang Christian. Book review: Jerry Glynn and Theodore Gray, The Beginner’s Guide to Mathematica Version 3. Computers in Physics, 12(5):451–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.168737.

Smith:1991:MW

[Smi91] Cathy Smith. Mathematica for Windows 3.0. PC World, ?? (??):114, March 1991. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

Smith:1993:NBL

[Smi93]  C. Smith.  Notebooks into

books via LaTEX. Mathematica Journal, 3(3):69–73, 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Shtabovenko:2020:FNF

[SMO20] Vladyslav Shtabovenko, Rolf Mertig, and Frederik Orellana. FeynCalc 9.3: New features and improvements. Computer Physics Communications, 256 (??):Article 107478, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S001046552030223X.

Snell:1994:MFM

[Sne94] Lionel Snell. Maths functionality for maths-phobics. Electronics world + wireless world, pages 395–397, May 1994. CODEN EWWWE6. ISSN 0959-8332.

Shiakolas:1992:CFS

[SNLR92] P. S. Shiakolas, R. V. Nambiar, K. L. Lawrence, and W. A. Rogers. Closed-form stiffness matrices for the linear strain and quadratic strain tetrahedron finite elements. Computers and structures, 45(2):237–242, November 1992. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Sobelman:1992:CAF

[Sob92] G. Sobelman. Computer algebra and fast algorithms. In IEEE [IEE92b], pages 89–92. ISBN 0-7803-0533-7 (case), 0-7803-0532-9 (pbk). LCCN TK 7882 S65 I16 1992. Five volumes.

Sobelman:1995:SRS

[Sob95] Gerald E. Sobelman. Simulation of Reed–Solomon encoder/ decoder systems using computer algebra. Transactions of the Society for Computer Simulation, 12(2):179–190, June 1995. CO-DEN TSCSEV. ISSN 0740-6797.

Sofroniou:1993:PWS

[Sof93a]  M. Sofroniou.  An efficient

symbolic-numeric environment using Mathematica. In Proceedings of the Workshop on Symbolic and Numerical Computation, page ???? ????, ????, 1993.

Sofroniou:1993:EBF

[Sof93b] M. Sofroniou. Extending the built-in format rules. Mathematica Journal, 3(3):74–??, 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Sofroniou:1996:OSL

[Sof96] M. Sofroniou. Order stars and linear stability theory. Journal of Symbolic Computation, 21(1): 101–131, January 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Soiffer:1995:MTM

[Soi95] Neil Soiffer. Mathematical typesetting in Mathematica. In Levelt [Lev95], pages 140–149. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Solomon:1992:RMA

[Sol92]  Bruce Solomon.  Reviews:

Mathematica in Action, by Stan Wagon; Exploring Mathematics with Mathematica, by Theodore W. Gray and Jerry Glynn. American Mathematical Monthly, 99(6):581–589, June/ July 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Soleng:1996:MPC

[Sol96] H. H. Soleng. The Mathematica packages CARTAN and Math-Tensor for tensor analysis. In Hehl et al. [HPR96], pages 210– 230. ISBN 3-540-60361-1. LCCN QC173.55 .R445 1996.

Squire:1992:MPA

[Squ92] William Squire. Mathematica: A Practical Approach. Nancy Blachman. Applied mechanics reviews, 45(11):B149, November 1992. CODEN AMREAD. ISSN 0003-6900.

Scholl:1990:TTW

[SR90] J. Scholl and D. Rogovin. Transient two-wave mixing in Kerr media. Mathematica Journal, 1 (2):81–83, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Steyn-Ross:1991:TMD

[SR91] Alistair Steyn-Ross. A tour of Mathematica: Doing mathematics by computer. The New Zealand mathematics magazine, 27(3):37–46, May 1991. CODEN NZMAB7. ISSN 0549-0510.

Sparrow:1994:IDF

[SRR94] Victor W. Sparrow, Daniel A. Russell, and Judith L. Rochat. Implementation of discrete fuzzy structure models in Mathematica. International Journal for Numerical Methods in Engineering, 37(17):3005–3014, September 1994. CODEN IJNMBH. ISSN 0029-5981. Symposium on Computational Structural Acoustics (Washington, DC, 1993).

Steele:1993:MD

[SS93] J. Michael Steele and Robert A. Stine. Mathematica and diffusions. In Varian [Var93], pages 192–213. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Shuchat:1994:FJM

[SS94] A. Shuchat and F. Schultz.

(FRI-S14) the joy of Mathematica: a point-and-click interface for Mathematica. In Lum [Lum94], pages 338–341. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Santarius:1995:PPI

[SS95a] J. F. Santarius and K. H. Simmons. Performance of Polywell inertial-electrostatic confinement for applications. In IEEE [IEE95c], page 258. ISBN 0-7803-2669-5. LCCN TA 2020 I612c 1995. IEEE Catalog No. 95CH35796.

Shuchat:1995:JM

[SS95b] Alan Shuchat and Fred Shulz.

The Joy of Mathematica. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 1995. ISBN 4-8101-8081-6. 188 pp. LCCN ???? In Japanese.

Sofroniou:2005:PNC

[SS05] Mark Sofroniou and Giulia Spaletta. Precise numerical computation. Journal of Logic and Algebraic Programming, 64 (1):113–134, July 2005. ISSN 1567-8326 (print), 1873-5940 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S1567832604000785.

Stys:2014:LNN

[SS14] Tadeusz Styś and Krystyna Styś. Lecture Notes in Numerical Analysis with Mathematica. Bentham Science Publishers, Inc., Sharjah, United Arab Emirates, 2014. ISBN 1-60805-942-1 (e-book), 1-60805-943-X. 243 pp. LCCN QA298 .S797

                                        2014. URL http://ebooks.
                                        benthamsciencepublisher.org/
                                        book/9781608059423/.

Sullivan:1990:AHE

[ST90] J. Sullivan and M. Thomas. Animating the heat equation: a case study in optimization. Mathematica Journal, 1(1):80– 84, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Shaw:1994:AMG

[ST94] William T. Shaw and Jason Tigg. Applied Mathematica: Getting Started, Getting it Done. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54217-X. xiv + 432 + 8 pp. LCCN QA76.95 .S52 1994. US$28.95.

Stavroulakis:1999:PDE

[ST99] Ioannis P. Stavroulakis and Stepan A. Tersian. Partial Differential Equations: An Introduction With Mathematica and Maple. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3891-6. x + 297 pp. LCCN QA377 .S79 1999.

Stavroulakis:2004:PDE

[ST04] Ioannis P. Stavroulakis and Stepan A. Tersian. Partial differential equations an introduction with Mathematica and MAPLE. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2004. ISBN 981-256-244-3 (electronic book). xii + 306 pp. LCCN QA377 .S79 2004.

Strzebonski:2011:URR

[ST11] Adam Strzebonski and Elias Tsigaridas. Univariate real root isolation in an extension field. In Schost and Emiris [SE11], pages 321–328. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Strzebonski:2012:URR

[ST12] Adam Strzeboński and Elias P.

Tsigaridas. Univariate real root isolation in multiple extension fields. In van der Hoeven and van Hoeij [vdHvH12], pages 343–350. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Stapleton:1990:VM

[Sta90] L. Stapleton. Visualizing math.

Computer Graphics World, 13 (6):59–60, 63–64, June 1990. CODEN CGWODH. ISSN 0271-

Stancevic:2003:M

[Sta03] Dragan Stancevic. Mathematica 4.2. Linux Journal, 107:??, March 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Staub:2019:XHA

[Sta19] Florian Staub. xSLHA: an Les Houches Accord reader for Python and Mathematica. Computer Physics Communications, 241(??):132–138, August 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465519300918.

Stelzer:1993:MSL

[Ste93] Ernst H. K. Stelzer. Mathematica: Ein systematisches Lehrbuch mit Anwendungsbeispielen. Addison-Wesley, Reading, MA, USA, 1993. ISBN 3-89319-495-9. 408 pp. LCCN ????

Steeb:1994:QMU

[Ste94] Willi-Hans Steeb. Quantum mechanics using computer algebra: includes sample programs for REDUCE, MAPLE, MATHE-MATICA and C++. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1770-6. viii + 189 pp. LCCN QC174.17.D37 S74 1994.

Stein:2008:CWC

[Ste08] William A. Stein. Can we create a viable free open source alternative to Magma, Maple, Mathematica and Matlab? In Jeffrey [Jef08], pages 5–6. ISBN 1-59593-904-0. LCCN ????

Stein:2013:SCV

[Ste13] William Stein. Sage: creating a viable free open source alternative to Magma, Maple, Mathematica, and MATLAB. In Felipe Cucker, Teresa Krick, Allan Pinkus, and Agnes Szanto, editors, Foundations of computational mathematics, Budapest 2011, volume 403 of London Mathematical Society lecture note series, pages 230–238. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-60407-9 (paperback), 1-139-09540-4 (e-book), 1-139-61690-0. LCCN QA297 .F635 2011.

Stine:1993:TSM

[Sti93] Robert A. Stine. Time series models and Mathematica. In Varian [Var93], pages 368–406. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Stine:1995:DAU

[Sti95] Robert A. Stine. Data analysis using Mathematica. Sociological methods and research, 23(3):352–??, February 1995. ISSN 0049-1241 (print), 1552-8294 (electronic).

Stinespring:2002:MML

[Sti02] John Robert Stinespring. Mathematica for microeconomics: learning by example. Academic Press, New York, NY, USA, 2002. ISBN 0-12-670961-0, 0-12-670962-9 (CD-ROM). xii + 222 pp. LCCN QA76.95 .S75 2002. Includes CD-ROM.

Stojanovic:2003:CFM

[Sto03]  Srdjan Stojanovic.  Compu-

tational financial mathematics using Mathematica: optimal trading in stocks and options. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2003. ISBN 0-8176-4197-1 (Boston : hardcover), 3-7643-4197-1 (Basel : hardcover). xi + 481 pp. LCCN HG106 .S76 2003. Includes CD-ROM.

Strejc:1992:SSC

[Str92a] V. Strejc, editor. System structure and control: preprints of the IFAC workshop, Prague, Czechoslovakia, 3–5 September 1992. Pergamon Press, New York, NY, USA, 1992. ISBN 0-08-042057-5. LCCN TJ212.2 .S97 1992.

Stroyan:1992:CUM

[Str92b] K. D. Stroyan. Calculus Using Mathematica. Academic Press, New York, NY, USA, preliminary edition, 1992. ISBN 0-12-672970-0. v + 491 + ii + 97 + ii + 107 + 12 pp. LCCN QA303.5.D37 S77 1992. Includes 3 computer disks.

Strnadl:1993:SCP

[Str93a] C. F. Strnadl. Symbolic computation of the pair-distribution function for hard-sphere systems in the whole R-range. Computer Physics Communications, 75(1-2):47–54, April 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Stroyan:1993:CUMa

[Str93b]  K. D. Stroyan.  Calculus

using Mathematica. Academic Press, New York, NY, USA, 1993. ISBN 0-12-672971-9 (text and disk), 0-12-672976-X (Macintosh), 0-12-672972-7 (IBM/DOS), 0-12-672973-5 (NeXT), 0-12-672977-8 (IBM/DOS version), 0-12-672978-6 (NeXT version). xxv + 532 pp. LCCN QA303.5.D37 S77 1993.

Stroyan:1993:CUMb

[Str93c]  K. D. Stroyan.  Calculus

using Mathematica: Scientific Projects and Mathematical Background. Academic Press, New York, NY, USA, 1993. ISBN 0-12-672975-1. xi + 353 pp. LCCN QA303.5.D37 S78 1993.

Stroyan:1994:SCU

[Str94] K. D. Stroyan. (SAT-M31)

calculus using Mathematica at large universities. In Lum [Lum94], pages 356–359. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Strampp:1997:HMM

[Str97]  Walter Strampp.  Höhere

Mathematik mit Mathematica. Bd. 2: Analysis. Vieweg-Lehrbuch Computeralgebra. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997. ISBN 3-528-06789-6. viii + 328 pp. LCCN ????

Stroyan:1998:CLC

[Str98] K. D. Stroyan. Calculus: the language of change. Academic Press, New York, NY, USA, second edition, 1998. ISBN 0-12-673030-X. xix + 555 pp. LCCN QA303.5.C65S77 1998.

Stroyan:1999:PCL

[Str99] K. D. Stroyan. Projects for calculus: the language of change. Academic Press, New York, NY, USA, second edition, 1999. ISBN 0-12-673031-8. xii + 326 pp. LCCN QA303.5.C65S77 1998.

Stuart:1995:EAR

[Stu95a] R. G. Stuart. Erratum: Algebraic reduction of Feynman diagrams to scalar integrals: a Mathematica implementation of LERG-I (Comput. Phys. Commun. 85 (1995) 267–277). Computer Physics Communications, 88(2):347–??, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). See [Stu95b].

Stuart:1995:ARF

[Stu95b] Robin G. Stuart. Algebraic reduction of Feynman diagrams to scalar integrals: a Mathematica implementation of LERG-I. Computer Physics Communications, 85(2):267–277, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). See erratum [Stu95a].

Stuart:1995:EBR

[Stu95c] Robin G. Stuart. Erratum: Algebraic reduction of Feynman diagrams to scalar integrals: a Mathematica implementation of LERG-I [Comput. Phys. Commun. 85 (1995) 267–277]. Computer Physics Communications, 88(2–3):347, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559500063L.

See [Stu95b].

Studt:1996:MAE

[Stu96] Tim Studt. Mathematica 3.0 adds equation editor for presentation graphics. Research & Development, 38(8):33, July 1996. CODEN REDEEA. ISSN 0746-9179.

Sulejmanpasic:2018:APT

[SÜ18] Tin Sulejmanpasic and Mithat Ünsal. Aspects of perturbation theory in quantum mechanics: The BenderWu Mathematica(R) package. Computer Physics Communications, 228 (??):273–289, July 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465518300420.

Sutner:1992:FSM

[Sut92] K. Sutner. Finite state machines and syntactic semigroups. Mathematica Journal, 2(1):78– 87, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Swaine:1993:SWM

[SV93] M. Swaine and R. Valdes. Stephen Wolfram: Multiparadigm man. (interview). Dr. Dobb’s Journal of Software Tools, 17(4):86–88, 90, 92, 94, 97, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Schoenefeld:1993:IDM

[SW93] Dale A. Schoenefeld and Roger L. Wainwright. Integration of discrete mathematics topics into the secondary mathematics curriculum using Mathematica — a summer institute for high school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):78–82, March 1993. CO-DEN SIGSD3. ISBN 0-89791-565-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schwalbe:1997:VVD

[SW97] Dan Schwalbe and S. Wagon. VisualDSolve: visualizing differential equations with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1997. ISBN 0-387-94721-3. xiv + 271 pp. LCCN QA 371.5 D37 S39 1997.

Swaine:1991:DBU

[Swa91] M. Swaine. The desktop blackboard. (use of Macintoshes in education). MacUser, 7:41–43, October 1991. CODEN MCUSEY. ISSN 0884-0997.

Swaine:1993:PPa

[Swa93a] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(1): 109–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPb

[Swa93b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(2):105– ??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Swenson:1991:MAF

[Swe91] C. Swenson. Making animation and flipbooks. Mathematica in Education, 1(1):2–7, Fall 1991. ISSN 1065-2965.

Swenson:1992:EAF

[Swe92] C. Swenson. Enhanced animations and flipbooks. Mathematica in Education, 2(1):9–14, Fall 1992. ISSN 1065-2965.

Spada:1990:CAA

[SYS^{+}90] G. Spada, D. Yuen, R. Sabadini, P. Morin, and P. Gasperini. A computer-aided, algebraic approach to the post-glacial rebound problem. Mathematica Journal, 1(2):65–69, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Sandler:1996:UMI

[SZ96] M. Sandler and E. Zeheb. On using Mathematica to implement the zero set technique. In Anonymous [Ano96g], pages 8/1–5.

Szabo:2000:LAI

[Sza00] Fred Szabo. Linear algebra: an introduction using Mathematica. Academic Press, New York, NY, USA, 2000. ISBN 0-12-680137-1. various pp. LCCN QA184 .S94 2000.

Szabo:2015:LAS

[Sza15] Fred Szabo. The Linear Algebra Survival Guide: Illustrated with Mathematica. Academic Press, New York, NY, USA, 2015. ISBN 0-12-409520-8 (paperback)), 0-12-409531-3 (ebook). xii + 423 pp. LCCN QA184.2 .S94 2015.

Tabakin:2016:QQM

[Tab16]  Frank Tabakin.  QDEN-

SITY/QCWAVE: a Mathematica quantum computer simulation update. Computer Physics Communications, 201 (??):171–172, April 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465515004531.

Tam:1991:PMIa

[Tam91a] P. Tam. Physics and Mathematica I. Computers in Physics, 5(3):342–348, May/June 1991.

CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Tam:1991:PMIb

[Tam91b] P. Tam. Physics and Mathematica II. Computers in Physics, 5 (4):438–442, July/August 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Tam:1991:PM

[Tam91c] Patrick Tam. Physics and Mathematica^{T M}. Computers in Physics, 5(3):342–348, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.4822994.

Tam:1991:PMU

[Tam91d] Patrick Tam. Physics and Mathematica^{T M}. Computers in Physics, 5(4):438–442, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.4823005.

Tamaanis:1995:MMW

[Tam95a] Doug Tamaanis. Mathematica meets Warp: It’s no surprise that the new OS/2 version of Mathematica runs faster than the Windows version on IBM’s 32-bit OS. BYTE Magazine, 20(5):137–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamasanis:1995:MMW

[Tam95b] Doug Tamasanis. Mathematica meets Warp. BYTE Magazine, 20(5), May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tam:1997:PGM

[Tam97] Patrick Tam. A Physicist’s Guide to Mathematica. Academic Press, New York, NY, USA, 1997. ISBN 0-12-683190-4. xxi + 506 pp. LCCN QC20.7.E4T36 1997. US$59.00.

Tam:2008:PGM

[Tam08] Patrick Tam. A physicist’s guide to Mathematica. Elsevier Academic Press, Amsterdam, The Netherlands, second edition, 2008. ISBN 0-12-683192-0 (paperback). xix + 728 pp. LCCN QC20.7.E4

                                        T36 2008. URL http://www.
                                        loc.gov/catdir/toc/fy0906/
                                        2008044787.html.

Taylor:1991:MC

[Tay91] Cyrus Taylor. Mathematica in the classroom. Computers in Physics, 5(1):16–21, 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Taylor:1992:BRB

[Tay92] Cyrus Taylor. Book reviews: Mathematica for the Sciences, by Richard E. Crandall and Mathematica in Action, by Stan Wagon. Physics Today, 45 (1):71–??, January 1992. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic).  URL http://www.
physicstoday.org/resource/
1/phtoad/v45/i1/p71_s1.

Taylor:1999:BRC

[Tay99] Peter R. Taylor. Book review: Y. C. Pao, Engineering analysis: Interactive methods and programs with FOR-TRAN, QuickBasic, MATLAB, and Mathematica (1999) CRC Press, Bristol 0-8493-2016-X. Computer Physics Communications, 120(2–3):271–272, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465599002374.

Tsamasphyros:1995:SIS

[TD95] George Tsamasphyros and George Dimou. Stress intensities in a strip reinforced by stiffeners at the edges. Engineering fracture mechanics, 51(6): 897–914, August 1995. CODEN EFMEAH. ISSN 0013-7944. terHaarRomeny:2003:FEV

[ter03] Bart M. ter Haar Romeny.

Front-end vision and multi-scale image analysis: multi-scale computer vision theory and applications, written in Mathematica, volume 27 of Computational imaging and vision. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2003. ISBN 1-4020-1503-8 (hardcover), 1-4020-1507-0 (paperback). xviii + 464 pp. LCCN R856 .H33 2003. Includes CD-ROM.

Teng:1993:MSB

[TF93] Jianfu Teng and J. K. Fidler. Multi-solutions in Butterworth LC filters. International Journal of Electronics Theoretical & Experimental, 75(6):1201– 1207, December 1993. CODEN IJELA2. ISSN 0020-7217.

Tavouktsoglou:1994:PRS

[TF94] A. N. Tavouktsoglou and B. Freed. Parametric representations of surfaces over arbitrary domains. Mathematica in Education, 3(1):20–23, 1994. ISSN 1065-2965.

Teng:1994:SCA

[TFS94] J. Teng, J. K. Fidler, and Y. Sun. Symbolic circuit analysis using Mathematica. International Journal of Electrical Engineering Education, 31(4):324– 333, October 1994. CODEN IJEEAF. ISSN 0020-7209.

Thacher:1989:HPV

[Tha89] H. C. Thacher, Jr. The Hewlett– Packard 28S vs Mathematica (mathematics computing). Access, 8(1):8, 10–13, January/ February 1989. There are about a dozen journals with this name. Is this Access (Research Triangle Park, NC), ISSN 0733-8074, which does have a volume 8 in 1989?

Thielemans:1991:MPC

[Thi91] K. Thielemans. A Mathematica package for computing operator product expansions. International Journal of Modern Physics C [Physics and Computers], 2(3):787–798, 1991. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Thorkildsen:1990:IEA

[Tho90] Gunnar Thorkildsen. An integral equation approach to threebeam diffraction using Laue and Bragg diffraction examined with Takagi’s equations. Arbeidspapirer fra Hogskolesenteret i Rogaland = Working papers from Rogaland Center 111, Hogskolesenteret i Rogaland, Stavanger, Norway, 1990. 99 + 8 pp.

Thomas:1991:ISD

[Tho91] M. Thomas. Illustrating solutions to 2-D partial differential equations. Mathematica in Education, 1(1):11–14, Fall 1991. ISSN 1065-2965.

Thompson:1992:VPW

[Tho92]  W. Thompson.  Visualizing

plane wave partial wave expansions. Mathematica in Education, 2(1):3–6, Fall 1992. ISSN 1065-2965.

Thompson:1994:MMC

[Tho94] W. J. Thompson. MATH-EMATICA and MAPLE: Has Champaign met its Waterloo? Computers in Physics, 8(3):269– 278, 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Thompson:1997:ACMb

[Tho97a] William J. (William Jackson)

Thompson. Atlas for computing mathematical functions: an illustrated guide for practitioners with programs in Fortran 90 and Mathematica. John Wiley and Sons, Inc., Chichester, UK, 1997. ISBN 0-471-18171-4 (cloth). xiv + 888 pp. LCCN QA331.T386 1997. Includes CD-ROM.

Thompson:1997:ACMa

[Tho97b] William J. (William Jackson) Thompson. Atlas for computing mathematical functions: an illustrated guidebook for practitioners: with programs in C and Mathematica. John Wiley and Sons, Inc., Chichester, UK, 1997. ISBN 0-471-00260-7 (cloth). xiv + 903 pp. LCCN QA331.T385 1997.

Tadic:1996:EWF

[TIM96] M. Tadic, Z. Ikonic, and V. Milanovic. Electron-wave-function calculation in the continuous part of the spectrum: the case of slowly varying potential asymptotics. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 53(6): 6523–6529, June 1996. CO-DEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Titus:1994:QMM

[Tit94] W. J. Titus. Quantum Mechanics with Mathematica, by James M. Feagin. American Journal of Physics, 62(10):957–??, October 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//ajp.aapt.org/resource/1/
ajpias/v62/i10/p957_s1.

Tabakin:2011:QMQ

[TJD11] Frank Tabakin and Bruno Juliá-Díaz. QCWAVE — a Mathematica quantum computer simulation update. Computer Physics Communications, 182 (8):1693–1707, August 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465511001287.

Tagliaboschi:2014:PMI

[TM14] Aaron M. Tagliaboschi and Jeremy B. Maddox. POTLIB2Math: a Mathematica interface to the potential energy surface library for chemical systems. Computer Physics Communications, 185 (1):294–301, January 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465513002725.

Titus:1996:BRR

[TMC96] Bill Titus, Susan R. McKay, and Wolfgang Christian. Book review: Richard J. Gaylord and Paul R. Wellin, Computer Simulations with Mathematica: Explorations in Complex Physical and Biological Systems. Computers in Physics, 10(4):349–??, July 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

                                        URL https://aip.scitation.
                                        org/doi/10.1063/1.4822455.

Todd:1992:NMF

[Tod92] D. Todd. New Mathematica: faster, leaner, linkable and QuickTime-compatible: Math-Link kit allows ties to other apps. Macworld, 6:6, June 15, 1992. CODEN MACWEA. ISSN 0741-8647.

Torsello:2020:BMP

[Tor20] Francesco Torsello. bimEX: a Mathematica package for exact computations in 3 + 1 bimetric relativity. Computer Physics Communications, 247 (??):Article 106948, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465519303030.

Tosiek:2010:FPM

[Tos10] Jaromir Tosiek. The Fedosov ⋆- product in Mathematica. Computer Physics Communications, 181(3):704, March 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465509003439.

Tsantilis:1996:QCL

[TPH96] E. Tsantilis, R. A. Puntigam, and F. W. Hehl. A quadratic curvature Lagrangian of Pawlowski and Raczka: a finger exercise with MathTensor. In Hehl et al. [HPR96], pages 231–240. ISBN 3-540-60361-1. LCCN QC173.55 .R445 1996.

Traverso:2000:IAU

[Tra00] Carlo Traverso, editor. IS-SAC 2000: 7–9 August 2000, University of St. Andrews, Scotland: proceedings of the 2000 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/
pubs/contents/proceedings/
issac/345542/. ACM order

number 505000.

Treat:1990:UMS

[Tre90] H. Treat. Using Mathematica in support of LabVIEW: power in the laboratory. In Anonymous [Ano90c], pages 353–358. LCCN TK 7801 N67 1990.

Treat:1991:ULM

[Tre91] Treat. Using LabVIEW and Mathematica in combination: Data collection to graphical presentation in a single setting. CoED, II(2):19–23, April 1991. CODEN CWLJDP. ISSN 0736-8607.

Trott:1994:MDI

[Tro94] Michael Trott. Mathematica:

a Detailed Introduction. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-94282-3. ???? pp. LCCN QA76.95 .T77 1994.

Trott:1995:TLM

[Tro95a]  M. Trott.  Three lectures

on Mathematica. In Vandoni [Van95], pages 211–236. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

Trott:1995:MGC

[Tro95b] Michael Trott. The Mathematica Guidebook: Concepts, Examples and Applications. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94282-3. 896 pp. LCCN QA76.95.T77 1994. US$59.95. Includes CD ROM.

Trott:1997:MSI

[Tro97] M. Trott. Mathematica solutions to the ISSAC system challenge 1997. SIGSAM Bulletin, 31(4):2–35, December 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trott:1998:GMC

[Tro98] Michael Trott. The guide to Mathematica: CD-ROM and applications book. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94282-3 (hardcover). ???? pp. LCCN QA76.95 .T767 1998.

Trott:2000:MGM

[Tro00] Michael Trott. The Mathematica guidebook: mathematics in Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 2000. ISBN 0-387-95011-7. ???? pp. LCCN QA76.95 .T773 2000.

Trott:2004:MGG

[Tro04a] Michael Trott. The Mathematica guidebook for graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-95010-9. xxxv + 1340 pp. LCCN T385 .T76 2004. Includes DVD-ROM.

Trott:2004:MGP

[Tro04b] Michael Trott. The Mathematica guidebook for programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-94282-3. xxxvii + 1028 pp. LCCN QA76.73.M29 T76 2004. Includes DVD-ROM.

Trott:2005:MGN

[Tro05] Michael Trott. The Mathematica guidebook for numerics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-95011-7. 1000 (est.) pp. LCCN QA76.95 .T773 2000.

Trott:2006:MGN

[Tro06a] Michael Trott. The Mathematica Guidebook for Numerics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-95011-7. xxxv + 1208 pp. LCCN QA76.95 .T773 2006. Includes DVD-ROM.

Trott:2006:MGS

[Tro06b] Michael Trott. The Mathematica Guidebook for Symbolics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-95020-6. xxxviii + 1453 pp. LCCN QA76.73.M29 T765 2006. Includes CD-ROM.

Turetken:2000:CSC

[TS00] B. Turetken and S. Eren San. Comparison of symbolic computation techniques for problems in electromagnetics. In IEEE, editor, International Conference on Mathematical Methods in Electromagnetic Theory, 2000. MMET 2000, 12–15 September 2000, volume 1, pages 172– 174. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-6347-7. LCCN QC760 .I65 2000. IEEE catalog number 00EX413.

Tien:1993:SAU

[TSC93] Win-Min Tien, N. Sri Namachchivaya, and V. T. Coppola. Stochastic averaging using elliptic functions to study nonlinear stochastic systems. Nonlinear Dynamics, 4(4):373–387, August 1993. CODEN NODYES. ISSN 0924-090X (print), 1573-269X (electronic). URL http://

link.springer.com/article/
10.1007/BF00120672.

Torrence:1999:SIM

[TT99] Bruce F. (Bruce Follett) Torrence and Eve A. (Eve Alexandra) Torrence. The student’s introduction to Mathematica: a handbook for precalculus, calculus, and linear algebra. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-59445-6 (hardcover), 0-521-59461-8 (paperback). xvii + 280 pp. LCCN QA76.95.T67 1999.

Torrence:2009:SIM

[TT09] Bruce F. (Bruce Follett) Torrence and Eve A. (Eve Alexandra) Torrence. The student’s introduction to Mathematica: a handbook for precalculus, calculus, and linear algebra. Cambridge University Press, Cambridge, UK, second edition, 2009. ISBN 0-521-71789-2 (paperback). xi + 471 pp. LCCN QA76.95 .T67 2009.

Trappe:2006:ICC

[TW06] Wade Trappe and Lawrence C.

Washington. Introduction to Cryptography: with Coding Theory. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2006. ISBN 0-13-186239-1, 0-13-198199-4 (paperback). xiv + 577 pp. LCCN QA268 .T73 2006.

Tzavelas:2009:MLP

[Tza09] George Tzavelas. Maximum likelihood parameter estimation in the three-parameter gamma distribution with the use of Mathematica. Journal of Statistical Computation and Simulation, 79(12):1457–1466, 2009.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Udina:1993:MEP

[UB93] F. Udina and N. Blachman. Mathematica: Un Enfoque Práctico. Editorial Ariel and Prentice-Hall, Barcelona, Spain andEnglewood Cliffs, NJ 07632, USA, 1993. ISBN 84-344-0478-8. 369 pp. LCCN ???? Spanish translation of Mathematica: A Practical Approach, [Bla92a].

Uhl:1995:ICD

[Uhl95] Jerry Uhl. Interactive calculus at a distance. T H E journal (Technological Horizons in Education), 22(11):67–??, June 1995. CODEN THEJD4. ISSN 0192-592X.

Ufuktepe:2011:UCT

[UK11] Ünal Ufuktepe and Sinan Kapçak. Unification of calculus on Time Scales with Mathematica. Applied Mathematics and Computation, 218(3):1102– 1106, October 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300311003857.

Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Ueno:1994:BHS

[UOT94] Yoshiaki Ueno, Hisao Oikawa, and Takashi Tokita. Basic High School Mathematics with Mathematica. Brain, ??, Japan, 1994. ISBN 4-89242-145-6. 244 pp. LCCN ???? In Japanese.

Vantaggiato:1993:MFF

[Van93] A. Vantaggiato. Modeling finite fields with Mathematica — applications to the computation of exponential sums and to the solution of equations over finite fields. Lecture Notes in Computer Science, ??(722):364– 368, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandoni:1995:CSC

[Van95] C. E. Vandoni, editor. 1995 CERN School of Computing: Arles, France, 20 August – 2 September 1995: proceedings, number 5 in CERN European Organization for Nuclear Research — Reports. CERN, Geneva, Switzerland, 1995. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

VariableSymbols:1990:MHS

[Var90] Variable Symbols, Berkeley, CA, USA. Mathematica help stack, version 1.2 rel 1. edition, 1990. 2 computer disks.

Vardi:1991:CRM

[Var91a]  Ilan Vardi.  Computational

Recreations in Mathematica. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52989-0. xvii + 286 pp. LCCN QA76.95 .V36 1990. See [Varxx].

Vardi:1991:SEP

[Var91b] Ilan Vardi. A sampler of elegant programs. Addison-Wesley, Reading, MA, USA, 1991. ISBN ???? ???? pp. LCCN ???? Advanced tutorial notes, Mathematica conference.

Varney:1991:WBM

[Var91c] S. Varney. Wolfram boosts Mathematica performance. Digital Review, 8:9–11, January 21, 1991. CODEN DIRVE5. ISSN 0739-4314.

Varian:1992:EFM

[Var92] Hal R. Varian, editor. Economic and Financial Modeling with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1992. ISBN 0-387-97882-8. xxii + 458 pp. LCCN HB143 .E36 1992.

Varian:1993:EFM

[Var93] Hal R. Varian, editor. Economic and Financial Modeling with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1993. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (Berlin). xxii + 458 pp. LCCN HB143 .E36 1993. US$49.95. Primarily Mathematica Notebooks run through nb2tex conversion software. Includes MS-DOS diskette.

Vardi:1994:SCC

[Var94] Ilan Vardi. Symbolic Computation: Computers in Pure and Applied Mathematics. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-94147-9. 352 pp. LCCN ??

Vardi:1995:ISC

[Var95] Ilan Vardi. Introduction to Symbolic Computation. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94147-9. 320 pp. LCCN ???? US$49.95. Covers C, Fortran, Maple, and Mathematica.

Varian:1996:CEF

[Var96] Hal R. Varian, editor. Computational economics and finance: modeling and analysis with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1996. ISBN 0-387-94518-0. xiv + 468 pp. LCCN HB143 .C65 1996. URL

http://www.loc.gov/catdir/
enhancements/fy0815/96200216-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/
96200216-t.html.

Varga:1997:CMF

[Var97] Kálmán Varga. A combined Mathematica–Fortran program package for analytical calculation of the matrix elements of the microscopic cluster model. Computer Physics Communications, 104(1–3):259–274, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465597000313.

Vardi:19xx:CRM

[Varxx]  Ilan Vardi.  Computational

Recreations in Mathematica. Toppan, Tokyo, Japan, 19xx. ISBN 4-8101-8038-7. ???? pp. LCCN ???? Japanese translation of [Var91a].

vanderGenugten:1994:DLS

[vdG94] B. B. van der Genugten. Density of the least squares estimator in the multivariate linear model with arbitrarily normal variables. Computational Statistics, 9(1):25–39, 1994. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

vanderHoeven:2012:IPI

[vdHvH12] Joris van der Hoeven and Mark van Hoeij, editors. ISSAC 2012: Proceedings of the 2012 International Symposium on Symbolic and Algebraic Computation, July 22–25, 2012, Grenoble, France. ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Verdel:1997:LM

[Ver97] Thierry Verdel. L’essentiel de Mathematica 3. Global Design Marketing et Edition, Paris, France, 1997. ISBN 2-911502-03-5. v + 289 pp. LCCN ????

Vinod:2003:RMV

[Vin03] H. D. Vinod. Review of math-Statica (v.1): an add-on to Mathematica. Journal of Applied Econometrics, 18(4):485– 491, July–August 2003. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vogt:1990:PTF

[VM90] William G. Vogt and Marlin H.

Mickle, editors. Proceedings of the twenty-first annual Pittsburgh conference, held May 3– 4, 1990, University of Pittsburgh, volume 21 of Modeling and simulation. Instrument Society of America, Research Triangle Park, NC, USA, 1990. ISBN 1-55617-271-0. LCCN TA 343 M62 1990. Five volumes.

Valdes:1994:DMG

[VMP94] V. S. Valdes, W. R. McKinney, and C. Palmer. The differential method for grating efficiencies implemented in Mathematica. Nuclear instruments and methods in physics research. Section A, Accelerators, spectrometers, detectors and associated equipment, 347(1):216–219, 1994. CODEN NIMAER. ISSN 0168-9002, 0167-5087.

vonSeggern:1993:CSC

[von93] David H. (David Henry) von Seggern. CRC Standard Curves and Surfaces: a Mathematica Notebook. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-0761-9. 6 pp. LCCN QA643. Includes diskette.

Vorobev:1997:SAN

[Vor97] E. M. Vorob’ev. Symmetry analysis of nonlinear differential equations with the “Mathematica” programm SYMMAN. Mathematical and computer modelling, 25(8):141–??, ???? 1997. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Vossler:1999:EAG

[Vos99] Donald L. Vossler. Exploring analytic geometry with Mathematica. Academic Press, New York, NY, USA, 1999. ISBN 0-12-728255-6. xviii + 865 pp. LCCN QA551 .V67 1999.

Vossler:2000:EAG

[Vos00] Donald L. Vossler. Exploring analytic geometry with Mathematica. Academic Press, New York, NY, USA, 2000. ISBN 0-12-728255-6, 0-12-728256-4 (CD-ROM). xviii + 865 pp. LCCN QA551.5 .V67 2000.

vanAlmsick:1989:MS

[vS89] M. van Almsick and K. Schulten. Mathematica (software). Mikrocomputer Zeitschrift, November 1989. CODEN MDMZDL. ISSN 0720-4442.

Vvedensky:1993:PDE

[Vve93] Dimitri Dimitrievich Vvedensky. Partial Differential Equations with Mathematica. Physics Series. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-54409-1. xi + 465 pp. LCCN QA377 .V84 1993. See [Vve94].

Vvedensky:1994:PDE

[Vve94] Dimitri Dimitrievich Vvedensky.

Partial Differential Equations with Mathematica. Addison-Wesley and Toppan, Reading, MA, USA and Tokyo, Japan, 1994. ISBN 4-8101-8069-7. ???? pp. LCCN ???? Japanese translation of [Vve93].

Welfert:1996:ANM

[WA96] Bruno D. Welfert and Ricardo Aguilar. Applied numerical methods and graphical visualization. Computer Applications in Engineering Education, 4(2):127–143, 1996. CODEN CAPEED. ISSN 1061-3773.

Wagon:1991:MA

[Wag91] Stan Wagon. Mathematica in Action. W. H.Freeman and Company, New York, NY, USA, 1991. ISBN 0-7167-2229-1, 0-7167-2202-X. xiv + 419 pp. LCCN QA76.95 .W34 1991. US$14.95, US$24.95. See [Wagxxb, Wagxxa].

Wagon:1992:DDI

[Wag92] S. Wagon. A deceptive definite integral. Mathematica in Education, 1(3):3–5, Spring 1992. ISSN 1065-2965.

Wagon:1993:HIB

[Wag93] S. Wagon. A hyperbolic interpretation of the Banach-Tarski paradox. Mathematica Journal, 3(4):58–62, Fall 1993. ISSN 1047-5974 (print), 1097-1610 (electronic).

Wagon:1994:PVN

[Wag94] Stan Wagon. The Power of visualization: notes from a Mathematica course, Leadville, Colorado. Front Range Press, Copper Mountain, CO, USA, July 1994. ISBN 0-9631678-3-9. 117 pp. LCCN ????

Wagon:1999:MA

[Wag99] S. Wagon. Mathematica in action. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, second edition, 1999. ISBN 0-387-98252-3. xvi + 592 pp. LCCN QA76.95 .W34 1999.

Wagon:19xx:MAb

[Wagxxa] Stan Wagon. Mathematica in Action. Brain, ??, Japan, 19xx. ISBN 4-89242-143-X. ???? pp. LCCN ???? Japanese translation of Mathematica in Action [Wag91].

Wagon:19xx:MAa

[Wagxxb] Stan Wagon. Mathematica in Aktion. Spektrum Akademischer Verlag, ????, 19xx. ISBN 3-86025-041-8. ???? pp. LCCN ???? German translation of Mathematica in Action [Wag91].

Wagon:2010:MA

[Wag10] Stan Wagon. Mathematica in action. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-75366-4. xi + 578 pp. LCCN ????

Wang:1996:TED

[Wan96] J. B. Wang. Time evolution and decay of an excited atom in a weak electric field. Computers in Physics, 10(4):400– 406, July–August 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Wass:2005:AAC

[Was05] John A. Wass. Adroit analyst:

Calling all Mathematica enthusiasts! an introductory programming text for the masses. Scientific Computing, 23(1):16, 54, December 2005. ISSN 1524-2560. Review of [WGK05].

Watson:1990:ELP

[Wat90a] M. Watson. An elementary logic package. Mathematica Journal, 1(1):85–91, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Watt:1990:PSC

[Wat90b] S. M. Watt. Potential of symbolic computation in engineering education. the opportunity and the challenge. In Noor et al. [NEH90], pages 339–360. CO-DEN AMPPD5. ISBN 0-7918-0598-0. ISSN 0277-027X. LCCN TA350 .S88 1990.

Watson:1994:ICU

[Wat94] A. Watson. Image compression using the discrete cosine transform. Mathematica Journal, 4(1):81–88, Winter 1994. ISSN 1047-5974 (print), 1097-1610 (electronic).

Wang:1996:VAO

[WAW96] J. B. Wang, P. C. Abbott, and J. F. Williams. Visualizing atomic orbitals. Computers in Physics, 10(1):69–82, January–February 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Wayner:1989:SMM

[Way89] Peter Wayner. Symbolic math on the Mac (Mathematica 1.0). BYTE Magazine, 14(1):239– 242, 144, January 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1990:ZSM

[Way90] P. Wayner. The zen of symbolic math (mathematics package). BYTE Magazine, 15(6): 193–194, 196, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wolfram:1994:MSB

[WB94] Stephen Wolfram and George Beck. Mathematica: The Student Book. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-55479-8. xix + 501 pp. LCCN QA76.95 .W65 1994. US$28.95.

Williams:1998:EQC

[WC98] Colin P. Williams and Scott H. Clearwater. Explorations in quantum computing. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-94768-X. xx + 307 pp. LCCN TK7888.3 .W46 1998. Includes CD-ROM.

Wolfram:1997:MV

[WD97] Wolfram Research and Thierry Dubois. Mathematica (Version 3). International Thomson Publishing, London, New York, Boston, Madrid, …, 3eme edition, 1997. ISBN 2-84180-137-3. 1499 pp. LCCN ????

Webster:1992:RMS

[Web92] Michael Webster. Review of “Mathematica: A System for Doing Mathematics by Computer, Second Edition, by Stephen Wolfram”. SIAM Review, 34(3):519–522, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weerawarana:1996:BNR

[Wee96] Sanjiva Weerawarana. Book news & reviews: Mathematica as a Tool by Stephan Kaufmann; Mathematica for Scientists and Engineers by Thomas B. Bahder. IEEE Computational Science & Engineering, 3(2):93– 95, Summer 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Weis:1992:VMI

[Wei92] Stephen Weis. Visualization of modal irradiance patterns in an optical pattern. IEEE transactions on education, 35(2): 109–111, May 1992. CODEN IEEDAB. ISSN 0018-9359.

Wellin:1993:SCP

[Wel93] P. Wellin. Simple closed paths. Mathematica in Education, 2(3): 7–10, 1993. ISSN 1065-2965.

Wellin:2013:PMI

[Wel13] Paul R. Wellin. Programming with Mathematica: an introduction. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-00946-4 (hardcover), 0-511-97294-6 (e-book). xviii + 711 pp. LCCN QA76.73.M29 W454

2013. URL http://reference-
tree.com/book/programming-
with-mathematica-an-introduction.

Wester:1992:RMS

[Wes92] Michael Wester. Review of “Mathematica: A System for Doing Mathematics by Computer, Second Edition, by Stephen Wolfram”. SIAM Review, 34(3):519–522, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

West:1993:FTP

[Wes93] Todd H. West. FeynmanParameter and trace — programs for expressing Feynman amplitudes as integrals over Feynman parameters. Computer Physics Communications, 77(2): 286–298, October 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559390011Z.

Wessels:1995:SRM

[Wes95] P. Wessels. Software review:

Maple V. SciTech Journal, 5(10):15, 17–19, November– December 1995. CODEN SITJE8. ISSN 1072-0995.

Wessels:1996:EH

[Wes96a] P. S. Wessels. Exploring HiQ 2.2. SciTech Journal, 6(4):15– 17, May–June 1996. CODEN SITJE8. ISSN 1072-0995.

Wessels:1996:SRT

[Wes96b] P. S. Wessels. Software review:

Theorist. SciTech Journal, 6(3): 22–24, March 1996. CODEN SITJE8. ISSN 1072-0995.

Wester:1999:CAS

[Wes99] Michael J. Wester, editor. Computer algebra systems: a practical guide. John Wiley and Sons, Inc., Chichester, UK, 1999. ISBN 0-471-98353-5. xvi + 436 pp. LCCN QA155.7.E4 W48 1999. URL

http://www.loc.gov/catdir/
description/wiley033/99021149.
html;   http://www.loc.gov/
catdir/toc/onix03/99021149.
html.

Walker:1992:TMA

[WG92] James S. Walker and J. Gathright. A transfer-matrix approach to one-dimensional quantum mechanics using Mathematica. Computers in Physics, 6 (4):393–399, July 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Walker:1994:EOD

[WG94] J. S. Walker and J. Gathright. Exploring one-dimensional quantum mechanics with transfer matrices. American Journal of Physics, 62(5):408–422, May 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Wang:2004:MPT

[WGDZ04] Zhongcheng Wang, Yonghua Ge, Yongming Dai, and Deyin Zhao. A Mathematica program for the two-step twelfth-order method with multi-derivative for the numerical solution of a one-dimensional Schrödinger equation. Computer Physics Communications, 160(1):23– 45, June 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465504001092.

Wellin:2005:IPM

[WGK05] Paul R. Wellin, Richard J. Gaylord, and Samuel N. Kamin. An introduction to programming with Mathematica. Cambridge University Press, Cambridge, UK, third edition, 2005. ISBN 0-521-84678-1. xx + 550 pp. LCCN QA76.73.M29 W45 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/cam051/2004056143.
html;   http://www.loc.gov/

catdir/toc/cam051/2004056143.
html.

Wicks:1996:LAI

[Wic96] John R. Wicks. Linear algebra: an interactive laboratory approach with Mathematica. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-82642-9. ???? pp. LCCN QA185.C65W53 1996.

Wiebusch:2015:HMP

[Wie15] Martin Wiebusch. HEPMath 1.4: a Mathematica package for semi-automatic computations in high energy physics. Computer Physics Communications, 195(??):172–190, October 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465515001642.

Wiggins:1992:FIA

[Wig92] R. R. Wiggins. Figure it all out with Mathematica 2.0. Computer Shopper, 12:350–352, January 1992. ISSN 0886-0556.

Wilf:2006:G

[Wil06] Herbert S. Wilf. Generatingfunctionology. A. K. Peters, Ltd., Wellesley, MA, USA, third edition, 2006. ISBN 1-56881-279-5 (hardcover). x + 245 pp. LCCN QA353.G44 W55 2006.

Wimmer:2012:AEN

[Wim12] M. Wimmer. Algorithm 923:

Efficient numerical computation of the Pfaffian for dense and banded skew-symmetric matrices. ACM Transactions on Mathematical Software, 38(4): 30:1–30:17, August 2012. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Winkel:1991:LTS

[Win91] B. Winkel. Let THEM [the students] wow you [with their own notebooks]! Mathematica in Education, 1(1):8–11, Fall 1991. ISSN 1065-2965.

Withoff:1990:PSD

[Wit90] D. Withoff. Processing survey data. Mathematica Journal, 1(1):92–94, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Withoff:1991:MWM

[Wit91a] David Withoff. Mathematica warning messages. Technical report, Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1991. 151 pp.

Withoff:1991:UPM

[Wit91b] David Withoff. Upgrading packages to Mathematica version 2.0. Technical report, Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1991. 7 pp.

Wickham-Jones:1994:MGT

[WJ94] Tom Wickham-Jones. Mathematica Graphics: Techniques and Applications. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-94047-2 (New York), 3-540-94047-2 (Berlin). xiii + 721 pp. LCCN T385 .W543 1994. US$49.95.

Wickham-Jones:2002:WMH

[WJ02]  Tom Wickham-Jones.  Web

Mathematica: How to deliver computational and visualization services from a Web server. In Anonymous [Ano02], page ?? ISBN ???? LCCN ????

Wojas:2017:FSD

[WK17] Wlodzimierz Wojas and Jan Krupa. Familiarizing students with definition of Lebesgue integral: Examples of calculation directly from its definition using Mathematica. Mathematics in Computer Science, 11 (3–4):363–381, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11786-017-0321-5.
pdf.

Wojas:2020:FSD

[WKB20] Włodzimierz Wojas, Jan Krupa, and Jarosław Bojarski. Familiarizing students with definition of Lebesgue outer measure using Mathematica: Some examples of calculation directly from its definition. Mathematics in Computer Science, 14 (2):253–270, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.

                                        springer.com/content/pdf/
                                        10.1007/s11786-019-00435-
                                        2.pdf.

Wang:2011:MMP

[WL11] Lai Wang and Frank X. Lee. MathQCDSR: a Mathematica package for QCD sum rules calculations. Computer Physics Communications, 182(8):1721– 1731, August 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465511001482.

West:1993:SC

[WM93] K. West and J. McClellan. Symbolic convolution. IEEE transactions on education, 36(4):386– 393, November 1993. CODEN IEEDAB. ISSN 0018-9359.

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wes-ley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Wolfram:1988:MSDa

[Wol88a] Stephen Wolfram. Mathematica: a System for Doing Mathematics by Computer. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-19334-5 (hardcover), 0-201-19330-2 (paperback), 0-201-51502-4, 0-201-51507-5 (paperback). xviii + 749 pp. LCCN QA76.95 .W65 1988. US$24.95 (paperback), US$39.95 (hardcover). This book was prepared with TEX, LaTEX, and Post-Script.

Wolfram:1988:MBMa

[Wol88b] Stephen Wolfram. Mathematica for 386-based MS-DOS systems. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 386 version 1.1. edition, 1988. 1 pp.

Wolfram:1988:MMc

[Wol88c] Stephen Wolfram. Mathematica for the Macintosh. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, enhanced Macintosh version 1.2. edition, 1988. 8 computer disks for Macintosh experts + release notes + technical report + reference guide updates.

Wolfram:1988:MMa

[Wol88d] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica for the Macintosh, version 1.1. edition, 1988. 4 computer disks.

Wolfram:1988:MMb

[Wol88e] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica for the Macintosh, standard Macintosh version 1.1. edition, 1988. 5 computer disks for Macintosh experts.

Wolfram:1989:MM

[Wol89] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica for the Macintosh, standard Macintosh version 1.2. edition, 1989. 749 pp. 8 computer disks for Macintosh experts + release notes + technical report + reference guide updates + book.

Wolfram:1991:MSDa

[Wol91a] Stephen Wolfram. Mathematica: a System for Doing Mathematics by Computer. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-51507-5, 0-201-51502-4. xxii + 961 pp. LCCN QA76.95 .W65 1991. See [Wolxxa, Wolxxb].

Wolfram:1991:MSDc

[Wol91b] Stephen Wolfram. Mathematica: a system for doing mathematics by computer. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, enhanced version 2.0 for the Macintosh. edition, 1991. 15 computer disks technical reports.

Wolfram:1991:MSDd

[Wol91c] Stephen Wolfram. Mathematica: a system for doing mathematics by computer. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, student version 2.0. edition, 1991. 56 pp. 6 computer disks + 1 Release notes (7).

Wolfram:1991:MC

[Wol91d] Wolfram Research, Inc., editor. 1991 Mathematica Conference. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1991. ISBN 1-880083-00-0 (v.

1) 1-880083-01-9 (v. 2) 1-880083-02-7 (v. 3). LCCN ????

Wolfram:1991:MSDb

[Wol91e] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica: a system for doing mathematics by computer, standard version 2.0 for the Macintosh. edition, 1991. 15 computer disks and technical reports.

Wolfram:1991:MSDe

[Wol91f] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica a system for doing mathematics by computer, Windows version edition, 1991. 56 pp. 8 computer disks + 1 Release notes (7)+ 2 technical manuals.

Wolfram:1992:MRG

[Wol92] Stephen Wolfram. Mathematica Reference Guide. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-51502-4, 0-201-51507-5 (paperback), 0-201-19330-2, 0-201-51012-X. 305 pp. LCCN QA76.95 .W66 1992. US$20.50.

Wolfram:1993:MSDa

[Wol93a] Stephen Wolfram. Mathematica: a system for doing mathematics by computer. Addison-Wes-ley, Reading, MA, USA, second, reprinted with corrections July, 1993. edition, 1993. ISBN 0-201-51502-4, 0-201-51507-5. xxii + 961 pp. LCCN QA76.95 .W66 1992.

Wolfram:1993:GSM

[Wol93b] Wolfram Research, Inc. Guide to Standard Mathematica Packages, Version 2.2. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1993. ISBN 1-880083-09-4. 459 pp. LCCN ????

Wolfram:1993:MSDb

[Wol93c] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica a system for doing Mathematics by computer, version 2.2 edition, 1993. 7 computer disks and notes.

Wolfram:1993:MSDd

[Wol93d] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica a system for doing mathematics by computer, second edition, 1993. 118 pp. Computer disks + 2 technical manuals.

Wolfram:1993:MSDe

[Wol93e] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA.

Mathematica: a system for doing mathematics by computer, version 2.2: User’s guide for the Macintosh, 4th. edition, 1993. l + 273 pp.

Wolfram:1993:MSDc

[Wol93f] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica: a system for doing mathematics by computer: version 2.2: user’s guide for the Unix shell, fourth edition, 1993. 68 pp.

Wolfram:1993:MSDf

[Wol93g] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica: a system for doing mathematics by computer: version 2.2: User’s guide for the X front end, 1993. xxx + 186 pp.

Wolfram:1993:MRG

[Wol93h] Wolfram Research, Inc. Mathlink Reference Guide, Version 2.2. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, second edition, 1993. ISBN 1-880083-08-6. 80 pp. LCCN ????

Wolfram:1993:RNM

[Wol93i] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Release notes for Mathematica version 2.1 for Sun workstations, January 29, 1993. 8 pp.

Wolfram:1994:M

[Wol94a] Wolfram Research. Mathsource, 1994. CD ROM containing Mathematica-related programs, interactive applications, courseware, Mathlink programs, and supplements from many books and journals on Mathematica. Contact Wolfram Research at tel: +1 800 441-MATH, or e-

mail: ms-cdrom@wri.com.

Wolfram:1994:MSD

[Wol94b] Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA. Mathematica a system for doing mathematics by computer for Microsoft Windows, version 2.2.3. edition, 1994. 6 computer disks Windows (138) + 1 release notes (7) + 1 Mathlink reference guide (80) + new features (32) + guide to standard Mathematica packages (459).

Wolfram:1995:MSI

[Wol95a] Stephen Wolfram. Mathematica: Le système informatique pour les mathématiques. Addison-Wesley, France, Paris, France, deuxième édition edition, 1995. ISBN 2-87908-035-5. xxiv + 1018 pp. LCCN ????

Wolfram:1995:GMP

[Wol95b] Wolfram Research, Inc. Guide to Mathematica Packages, Version 2.3. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-55579-4. 520 pp. LCCN ????

Wolfram:1996:MBa

[Wol96a] Stephen Wolfram. The Mathematica book. Wolfram Media and Cambridge University Press, Champaign, IL, USA and Cambridge, UK, third edition, 1996. ISBN 0-9650532-0-2 (hardcover), 0-9650532-1-0 (paperback), 0-521-58889-8 (hardcover), 0-521-58888-X (paperback). xxiv + 1403 pp. LCCN QA76.95.W65 1996. US$59.95.

Wolfram:1996:MBb

[Wol96b] Stephen Wolfram. The Mathematica book. Cambridge University Press, Cambridge, UK, third edition, 1996. ISBN 0-521-58889-8 (hardcover), 0-521-58888-X (paperback). xxiv + 1403 pp. LCCN QA76.95.W65 1996.

Wolfram:1996:MSA

[Wol96c] Wolfram Research, Inc. Mathematica 3.0 Standard Add-on Packages. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-9650532-3-7 (hardcover), ISBN 0-9650532-2-9 (paperback), 0-521-58586-4 (hardcover), 0-521-58585-6 (paperback). v + 516 pp. LCCN QA76.95.M385 1996. US$59.95 (hardcover), US$29.95 (paperback).

Wolfram:1997:SS

[Wol97] Wolfram Research. Signals and systems, 1997. 2 computer disks.

Wolfram:1999:MB

[Wol99a] Stephen Wolfram. The Mathematica book. Cambridge University Press and Wolfram Research, Inc., Cambridge, UK and 100 Trade Center Drive, Champaign, IL 61820-7237, USA, fourth edition, 1999. ISBN 0-521-64314-7 (Cambridge: hardbound), 1-57955-004-5 (Wolfram: hardbound). xxvi + 1470 + 20 pp. LCCN QA76.95 .W65 1999.

Wolfram:1999:MSA

[Wol99b] Wolfram Research. Mathematica 4.0 standard add-on packages: the official guide to over a thousand additional functions for use with Mathematica 4. Wolfram Media, Champaign, IL, USA, 1999. ISBN 1-57955-006-1 (hardcover) 1-57955-007-X (paperback). ???? pp. LCCN QA76.95.M386 1999.

Wolfram:19xx:MSD

[Wolxxa] Stephen Wolfram. Mathematica: a System for Doing Mathematics by Computer. Addison-Wesley, Reading, MA, USA, second edition, 19xx. ISBN 4-7952-9614-6. ???? pp. LCCN ???? Japanese translation of [Wol91a].

Wolfram:19xx:MSM

[Wolxxb] Stephen Wolfram. Mathematica: Ein System für Mathematik auf dem Computer. Addison-Wes-ley, Reading, MA, USA, 19xx. ISBN 3-89319-371-5. ???? pp. LCCN ???? German translation of Mathematica: A System for Doing Mathematics by Computer [Second Edition] [Wol91a].

Wolfram:2002:NKS

[Wol02]  Stephen Wolfram.  A new

kind of science. Wolfram Media, Champaign, IL, USA, 2002. ISBN 1-57955-008-8. 1192 (est.) pp. URL http://www.

wolframscience.com/.

Wolfram:2003:MB

[Wol03] Stephen Wolfram. The Mathematica book. Wolfram Media, Champaign, IL, USA, fifth edition, 2003. ISBN 1-57955-022-3. xxiv + 1464 pp. LCCN QA76.95 .W65 2003.

Wolfram:2015:EIW

[Wol15] Stephen Wolfram. An Elementary Introduction to the Wolfram Language. Wolfram Media, Inc., Champaign, IL, USA, 2015. ISBN 1-944183-00-0 (paperback). xv + 324 pp. LCCN QA76.73.W65 W65 2015. URL

http://wolfr.am/eiwl.

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers:

personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (e-book), 1-57955-011-8. 250 (est.) pp. LCCN Q141 .W678562 2016. URL

http://www.wolfram-media.
com/products/idea-makers.
html.

Wolfram:1990:MCM

[WRR90] S. Wolfram, E. F. Redish, and J. S. Risley. Mathematics by computer (Mathematica). In Redish et al. [RRM90], pages 338–346. ISBN 0-201-16306-3. LCCN QC30 .C636 1988.

Wang:1994:TDG

[WSW94] J. B. Wang, A. T. Stelbovics, and J. F. Williams. Theoretical description of gamma1 and gamma2 coincidences in electron-atom collisions. Zeitschrift für Physik. D, Atoms, molecules, and clusters, 30(2-3):119–127, 1994. CODEN ZDACE2. ISSN 0178-7683.

Wolfram:1988:MSDb

[WW88a] Stephen Wolfram and Wolfram Research, Inc. Mathematica: a system for doing mathematics by computer. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, version 2.0, standard edition, 1988. xxii + 951 pp. 14 computer disks + 1 warning messages report + 1 standard packages guide + 1 upgrading report + 1 new features report + 1 release notes + 1 text.

Wolfram:1988:MBMb

[WW88b] Stephen Wolfram and Wolfram Research, Inc. Mathematica for 386-based MS-DOS systems. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 386 version 1.2. edition, 1988. 4 computer disks + 4 booklets + 1 version 1.2 sheet + 1 technical support card.

Wolfram:1988:MMIa

[WW88c] Stephen Wolfram and Wolfram Research, Inc. Mathematica Macintosh II version. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, version 1.1. edition, 1988. xviii + 749 pp. 5 computer disks + 1 summary (v + 16) + 1 setting-up guide (v + 16) + text.

Wolfram:1988:MMIb

[WW88d] Stephen Wolfram and Wolfram Research, Inc. Mathematica Macintosh II version. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, version 2.0, enhanced edition, 1988. xxii + 951 pp. 7 computer disks + warning messages report (151) + upgrading report (7) + release notes (8) + text.

Wolfram:1988:MSM

[WW88e] Stephen Wolfram and Wolfram Research, Inc. Mathematica standard Macintosh version. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, version 1.1. edition, 1988. xviii + 749 pp.

Wolfram:1989:MBM

[WW89a] Stephen Wolfram and Wolfram Research, Inc. Mathematica for 386-based MS-DOS systems: user’s guide. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1989. v + 62 pp.

Wolfram:1989:MMU

[WW89b] Stephen Wolfram and Wolfram Research, Inc. Mathematica for the Macintosh: user manual. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, third edition, 1989. ISBN ???? xi + 151 pp. LCCN ???? Third printing, May 1990.

Wang:1993:SHC

[WW93a] J. B. Wang and J. F. Williams. Study of the helium 3^{1}D → 2^{1}P → ^{1}S cascade with algebraic computing. Computer Physics Communications, 75 (3):275–282, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/001046559390044D.

Wang:1993:SHS

[WW93b] J. B. Wang and J. F. Williams. A study of the helium 3/sup 1/D to 2/sup 1/P to 1/sup 1/S cascade with algebraic computing. Computer Physics Communications, 75(3):275–282, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Wolfram:1993:STN

[WW93c] Stephen Wolfram and Wolfram Research, Inc. Selected Tutorial Notes. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1993. ISBN 1-880083-07-8. 405 pp. LCCN ????

Walsh:1995:MSD

[WW95a] Jerome Walsh and Timothy Williams. Mathematica: a system for doing mathematics by computer, version 2.2: user’s guide for Microsoft Windows. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, third edition, 1995. xxxviii + 139 pp.

Wolfram:1995:MBM

[WW95b] Stephen Wolfram and Wolfram Research, Inc. The Mathematica Book: For Mathematica Version 3. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-82290-3 (softcover), 0-201-82291-1 (hardcover). 1150 pp. LCCN ????

Wu:2006:MPA

[WW06] Dongmei Wu and Zhongcheng Wang. A Mathematica program for the approximate analytical solution to a nonlinear undamped Duffing equation by a new approximate approach. Computer Physics Communications, 174(6):447–463, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465505005138.

Wolfram:1989:IM

[WWH89] Stephen Wolfram, Gary Welz, and Livingston Hinckley. Introducing Mathematica. American Mathematical Society and Science Television Company, Providence, RI, USA and P. O. Box 2498, Times Square Station, New York, NY 10108, USA, 1989. ISBN 1-878310-01-1. LCCN ???? 60-minute video cassette.

Wolff:1993:VNP

[WY93] Robert S. Wolff and Larry Yaeger. Visualization of Natural Phenomena. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1993. ISBN 0-387-97809-7 (New York), 3-540-97809-7 (Berlin). xxvi + 374 pp. LCCN T385.W645 1993.

Xavier:2022:MCM

[Xav22] George D. Xavier. M-CHIPR: a Mathematica program for constructing multi-state coupled adiabatic potential energy functions in triatomic molecule using many body partitioning approach. Computer Physics Communications, 278 (??):Article 108419, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465522001382.

Xemard:1998:MAM

[Xem98] Jean-Pierre Xemard. Mathématiques avec Mathematica. Bibliothèque des sciences. Diderot Multimedia, Paris, France, 1998. ISBN 2-84352-106-8. 505 pp. LCCN ????

Xiaoyi:1993:SCV

[XKM93] He Xiaoyi, David N. Ku, and James E. Moore. Simple calculation of the velocity profiles for pulsatile flow in a blood vessel using Mathematica. Annals of biomedical engineering, 21(1): 45–50, January 1993. CODEN ABMECF. ISSN 0090-6964.

Xiao:2016:MPC

[XMLC16] Shuyuan Xiao, Xueli Mu, Tingting Liu, and Hong Chen. A Mathematica program for the calculation of five-body Moshinsky brackets. Computer Physics Communications, 203 (??):238–244, June 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465516300315.

Xinhe:1995:SCD

[XW95] Xu Xinhe and Xiao Wendong.

Symbolic computation for discrete event systems. Proceedings of the IEEE Conference on Decision and Control, 3:2618–2623, 1995. CODEN PCDCDZ. ISSN 0191-2216. IEEE catalog number 95CH3580-3.

Xu:1995:SCD

[XX95] Xinhe Xu and Wendong Xiao.

Symbolic computation for discrete event systems. In IEEE [IEE95e], pages 2618–2623 vol.3.

CODEN PCDCDZ. ISBN 0-7803-2685-7. ISSN 0191-2216. LCCN TJ217 .I17 1995. Four volumes. IEEE Catalog No. 95CH35803.

Yager:1992:CVF

[Yag92] Tom Yager. Cut to video: Four programs for moving presentations: The BYTE Lab looks at four video-presentation programs. BYTE Magazine, 17(12): 238–240, 242, 244, 246, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yared:1995:EBS

[Yar95] W. Yared. Electron-beam scanning of ZnS crystals: a thermal comparison of CRTs and field-emission devices. Journal of Imaging Science and Technology, 39(5):448–452, September– October 1995. CODEN JIMTE6. ISSN 8750-9237.

Yang:2004:EME

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from PostScript documents. In Gutierrez [Gut04], pages 305–311. ISBN 1-58113-827-X. LCCN ???? URL http:/

                                        /www.ocf.berkeley.edu/~mlyang/
                                        papers/MichaelYangPsmath.
                                        pdf.

Yoo:1992:LIL

[YJ92] Sung Mi Yoo and J. Javanainen. Low-intensity limit of the laser cooling of a multistate atom. Physical Review A (Atomic, Molecular, and Optical Physics), 45(5):3071–3083, March 1, 1992. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Yasutake:1995:IAP

[YK95] K. Yasutake and Y. Kagawa.

Inverse analysis of a potential problem by symbolic computational language ‘Mathematica’- charge simulation method. Journal of the Japan Society for Simulation Technology, 14(4):311– 319, December 1995. CODEN SHIMDM. ISSN 0285-9947.

Yousif:2003:CBF

[YM03] Hashim A. Yousif and Richard Melka. Computing Bessel functions of the second kind in extreme parameter regimes. Computer Physics Communications, 151(1):25–34, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Yu:1994:SSN

[Yu94] K. W. Yu. Symbolic simulation of nonlinear random resistor networks by MATHEMATICA. Computers in Physics, 8(6):735– ??, 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Ylinen:1992:CAA

[YZPD92] R. Ylinen, K. Zenger, F. Pichler, and R. M. Diaz. Computer aided analysis and design of time-varying systems. In Pichler and Moreno-Diaz [PMD92], pages 73–94. ISBN 0-387-55354-1 (New York), 3-540-55354-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA345 .I62 1991.

Zachariah:1993:MMM

[Zac93] T. M. Zachariah. Mathematical modeling and Mathematica. In Lum [Lum93], pages 437– 440. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Zachary:1998:ISP

[Zac98] Joseph L. Zachary. Introduction to scientific programming: computational problem solving using Mathematica and C. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98250-7 (hardcover). xxiv + 433 pp. LCCN QA76.6.Z324 1998.

Zhao:2002:MPD

[ZD02] Xuan Zhao and Haibao Duan. A Mathematica program for the degrees of certain Schubert varieties. Journal of Symbolic Computation, 33(4):507–517, April 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhou:1997:BPD

[Zho97] Ling Zhou. Biofilm process design using Mathematica. Thesis (M.S. in Environmental and Water Resources Engineering), Vanderbilt University, Nashville, TN, USA, 1997. xiii + 132 pp.

Zitko:2011:SMP

[Zit11] Rok Zitko. SNEG — Mathematica package for symbolic calculations with second-quantization-operator expressions. Computer Physics Communications, 182 (10):2259–2264, October 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465511001792.

Zizza:1992:APD

[Ziz92] F. Zizza. Algebraic programming and differential forms. Mathematica Journal, 2(1):88– 90, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Zizi:1993:MIE

[Ziz93] Jacqueline Zizi. Mathématiques, informatique et enseignement. Ed. Archimède, Ed. du Choix, Argenteuil, France, 1993. ISBN 2-909028-13-5. xxv + 303 pp. LCCN ????

Zizi:1997:MPC

[Ziz97a]  J. Zizi.  Mathematica pour

classes préparatoires et premiers cycles scientifiques. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62736-7. xii + 460 pp. LCCN ????

Zizi:1997:MVP

[Ziz97b] Jacqueline Zizi. Mathematica, version 2 et 3: pour classes préparatoires et DEUG scientifiques. Tome 1, Programme commun aux 1ères et 2èmes années des classes de MPSI, PCSI, PTSI, MP, PC, PSI, PT, TPC, TSI. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62736-7. xviii + 461 pp. LCCN ????

Zhang:1995:SFV

[ZL95] Qichang Zhang and A. Y. T. Leung. Studying the focal value of ordinary differential equations by using the normal form theory. Applied mathematics and mechanics = Ying yung shu hsueh ho li hsueh, 16(9):891–900, September 1995. CODEN AM-MEEQ. ISSN 0253-4827.

Zhibin:1993:TWS

[ZM93] L. Zhibin and W. Mingliang. Travelling wave solutions to the two-dimensional KdV–Burgers equation. Journal of Physics A (Mathematical and General), 26(21):6027–6631, November 7, 1993. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Zimmerman:1995:MP

[ZO95] Robert L. Zimmerman and Fredrick Iver Olness. Mathematica for Physics. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-53796-6. xxiii + 436 pp. LCCN QC20 .Z56 1995.

Zimmerman:2002:MP

[ZO02] Robert L. Zimmerman and Fredrick Iver Olness. Mathematica for physics. Addison-Wesley, Reading, MA, USA, second edition, 2002. ISBN 0-8053-8700-5. 645 pp. LCCN QC20 .Z56 2002 UCB.

Zotos:2007:PCM

[Zot07] Kostas Zotos. Performance comparison of Maple and Mathematica. Applied Mathematics and Computation, 188(2):1426– 1429, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zelniker:1994:ADS

[ZT94] Glenn Zelniker and Fred J. Taylor. Advanced Digital Signal Processing. Marcel Dekker, Inc., New York, NY, USA, 1994. ISBN 0-8247-9145-2. x + 666 pp. LCCN TK5102.9.Z45 1994.

Zypman:1993:SPH

[Zyp93] F. R. Zypman. Symbolic programming helps to teach Debye-Scherrer diffraction. Computers in Physics, 7(1):22–26, January/February 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).